### v10

# Burnside type results for fusion categories

Abstract.
In this paper, we extend a classical vanishing result of Burnside from the character tables of finite groups to the character tables of commutative fusion rings, or more generally to a certain class of abelian normalizable hypergroups. We also treat the dual vanishing result. We show that any nilpotent fusion categories satisfy both Burnside’s property and its dual. Using Drinfeld’s map, we obtain that the Grothendieck ring of any weakly-integral modular fusion category satisfies both properties. As applications, we prove new identities that hold in the Grothendieck ring of any weakly-integral fusion category satisfying the dual-Burnside’s property, thus providing new categorification criteria. In particular we improve [OY23, Theorem 4.5] as follows: A weakly integral modular fusion category of $\mathrm{FPdim}$ $md$ with $d$ square-free coprime with $m$ and $\mathrm{FPdim}(X)^{2}$ for every simple object $X$, has a pointed modular fusion subcategory of $\mathrm{FPdim}$ $d$. We also prove some new results on the perfect fusion categories.
1991 Mathematics Subject Classification:
18M20; 16T30
The first author is supported by a grant of the Ministry of Research, Innovation and Digitization, CNCS/CCCDI - UEFISCDI, project number PN-III-P4-ID-PCE-2020-0878, within PNCDI III. The second author is supported by BIMSA Start-up Research Fund and Foreign Youth Talent Program from the Ministry of Sciences and Technology of China.
1. Introduction
A classical result of Burnside in the character theory of finite groups states that any irreducible non-linear character of a finite group vanishes on at least one element of the group. This can be stated as follows: in the character table of a finite group, the row of every character of degree $\neq 1$ contains a zero entry.
More recently, the dual version of this result was also intensively studied in the literature. This dual version describes the group elements that vanish on at least one irreducible character, see [INW99] and the references therein. Burnside’s result was extended to all weakly integral fusion categories, initially in the context of modular categories in [GNN09, Appendix], and subsequently in the general framework in [Bur23].
The main goal of this paper is to develop an analogue of Burnside’s result for a certain large class of fusion rings, or even more generally, to a certain class of abelian rational normalizable hypergroups. We also consider the dual version of Burnside’s result in this more general settings.
Recall that in [Bur21] the author introduced the notion of dual of a Grothendieck ring. This was achieved based on work of Harrison [Har79] on dualizable probability groups and on [ZZ21]. More generally, this notion of dual of a fusion ring was extended to arbitrary fusion rings in [BP22]. It was noticed in [Bur21] that the dual of the Grothendieck ring of a pivotal fusion category is isomorphic to the center of the category as defined in [Shi17].
In this paper we show that both Burnside’s property and its dual are related with the ring structure of the dual of the involving Grothendieck/fusion ring. In general, the dual of a fusion ring is no longer a fusion ring, but an abelian normalizable hypergroup, see [Bur21]. The advantage of the class of hypergroups is to be closed under duality, see [Bur21, Proposition 2.3].
For a finite dimensional $\mathbb{C}$-algebra $H$, fix a (finite) $\mathbb{C}$-linear basis $\mathcal{B}$ with the unit $1\in\mathcal{B}$ and write $ab=\sum_{c\in{\mathcal{B}}}N^{c}_{ab}c$ for some scalars $N^{c}_{ab}\in\mathbb{C}$. Note that $H={\mathbb{C}[\mathcal{B}]}$, the complex linear span of $\mathcal{B}$.
Definition 1.1.
A finite dimensional algebra $H$ together with a $\mathbb{C}$-linear basis $\mathcal{B}$ is called a hypergroup if there is an involution map $(-)^{*}:{\mathcal{B}}\rightarrow{\mathcal{B}}$ such that $N^{1}_{aa^{*}}>0$ and $N^{1}_{ab}=0$ if $b\neq a^{*}$. The elements $b\in\mathcal{B}$ are called the standard elements of ${(H,\mathcal{B})}$.
We may refer to ${\mathcal{B}}$ (or $H$) only as a hypergroup, when $H$ (or ${\mathcal{B}}$) is implicitly understood. A hypergroup $\mathcal{B}$ is called:
  • symmetric if $N^{1}_{ab}=N^{1}_{ba}$,
  • normalized if $\sum_{c\in{\mathcal{B}}}N^{c}_{ab}=1$,
  • real if $N_{ab}^{c}\in\mathbb{R}$,
  • rational if $N_{ab}^{c}\in\mathbb{Q}$,
  • real non-negative if $N_{ab}^{c}\geq 0$,
  • abelian if $ab=ba$,
for all $a,b,c\in{\mathcal{B}}$. Throughout this paper we use the abbreviation (A)RN-hypergroup for an (abelian) real non-negative hypergroup.
For a normalized hypergroup, the linear map $\mu_{1}:H\rightarrow\mathbb{C}$ defined on the linear basis ${\mathcal{B}}$ by $\mu_{1}(a)=1$ for all $a\in{\mathcal{B}}$, is a morphism of algebras. It is called the augmentation map of $(H,\mathcal{B})$. The underlying algebra $H$ of a hypergroup $(H,\mathcal{B})$ is a finite dimensional *-algebra, so is semisimple. For a normalized hypergroup ${(H,\mathcal{B})}$ the central primitive idempotent $F_{1}$ corresponding to the character $\mu_{1}$ is denoted by $\lambda_{(H,\mathcal{B})}$ or simply $\lambda_{H}$ if $\mathcal{B}$ is implicitly understood. A formula for this idempotent is described in Corollary 2.8.
A RN-hypergroup is sometimes called a table algebra in the literature. Recall that a fusion ring ${\mathcal{B}}$ is a hypergroup such that $N^{a}_{bc}\in\mathbb{Z}_{\geq 0}$ and $N^{1}_{aa^{*}}=N^{1}_{a^{*}a}=1$ for all $a,b,c\in\mathcal{B}$. The first main result of this paper is the following generalization of Burnside’s result to some weakly-integral fusion rings:
Theorem 1.2.
Let ${(H,\mathcal{B})}$ be a commutative fusion ring with a $h$-integral dual. Let $a$ be a standard element such that $\mathrm{FPdim}(a)>1$. Then there is $\mu\in{\widehat{\mathcal{B}}}$ such that $\mu(a)=0$.
A fusion ring whose dual is h-integral (see Definition 3.6) is always weakly-integral ($\mathrm{FPdim}{(H,\mathcal{B})}\in\mathbb{Z}$). The converse is true for a weakly-integral fusion category (see Remark 3.7). The set $^{\widehat{\mathcal{B}}}$ is the set of characters $\mu_{j}:H\rightarrow\mathbb{C}$ (see §2 for more details). In fact we prove a more general version of Theorem 1.2, for a certain class of abelian rational normalizable hypergroups, see Theorem 4.17.
Definition 1.3.
For any hypergroup ${(H,\mathcal{B})}$, an element $a\in\mathcal{B}$ with
(1.1) $aa^{*}=N_{a,a^{*}}^{1}1\text{ (in other words, }N_{a,a^{*}}^{c}=\delta_{1,c}N_{a,a^{*}}^{1}\text{)}$
is called a grouplike element of ${(H,\mathcal{B})}$.
By Lemma 2.25, the set of grouplike elements form a group (with the multiplication structure) denoted ${G{(H,\mathcal{B})}}$.
Definition 1.4.
A hypergroup ${(H,\mathcal{B})}$ has Burnside’s vanishing property (or shortly, is Burnside) if for all $a\in\mathcal{B}$, the following are equivalent:
  1. (1)
    There is some $\mu\in{\widehat{\mathcal{B}}}$ such that $\mu(a)=0$,
  2. (2)
    the standard element $a$ is not a grouplike element.
Theorem 1.2 states that a commutative fusion ring with h-integral dual is Burnside. Non-Burnside simple integral fusion rings can be found in §12. It is easy to see that a hypergroup is Burnside if and only if the set of invertible standard elements coincides to the set of grouplike elements, see Proposition 4.3.
1.1. The dual hypergroup of a normalized hypergroup
Let ${(H,\mathcal{B})}$ be an abelian normalized hypergroup. By the abelian assumption, $^{\widehat{\mathcal{B}}}$ forms a basis for the dual $H^{*}$, which then is $\mathbb{C}[{\widehat{\mathcal{B}}}]$, and on which one can define a multiplication. On the basis $^{\widehat{\mathcal{B}}}$, the multiplication $\mu_{i}\star\mu_{j}$ is defined by declaring
(1.2) $[\mu_{i}\star\mu_{j}](a):=\mu_{i}(a)\mu_{j}(a),\;\text{for all}\;a\in\mathcal{B}$
and then extending linearly $\mu_{i}\star\mu_{j}$ on the entire ${\mathbb{C}[\mathcal{B}]}$. In this way, one obtains an algebra structure on $H^{*}$.
It follows that there are some non-zero scalars ${\widehat{p}}_{k}(i,j)\in\mathbb{C}$ such that
(1.3) $\mu_{i}\star\mu_{j}=\sum_{k=1}^{m}{\widehat{p}}_{k}(i,j)\mu_{k}.$
By [Bur21, Proposition 2.3], for any abelian normalized hypergroup $(H,{\mathcal{B}})$, the pair $(H^{*},{\widehat{\mathcal{B}}})$ is also an abelian normalized hypergroup. Moreover, the involution on $^{\widehat{\mathcal{B}}}$ is given by $\mu_{j}\mapsto\mu_{j^{\#}}$, where $\mu_{j^{\#}}\in{\widehat{\mathcal{B}}}$ is defined by $\mu_{j^{\#}}(a):=\mu_{j}(a^{*})$ for all $a\in\mathcal{B}$. We denote by $\widehat{{(H,\mathcal{B})}}:=(H^{*},{\widehat{\mathcal{B}}})$ the dual hypergroup of the abelian normalized hypergroup ${(H,\mathcal{B})}$. It is also well known that for a normalized hypergroup ${(H,\mathcal{B})}$, then $\mu_{j}(a^{*})=\overline{\mu_{j}(a)}$ for all $a\in\mathcal{B}$. We also use the notation ${\widehat{h}}_{j}:={\widehat{p}}_{1}(j,j^{\#})^{-1}$, called the order of $\mu_{j}$ in the dual hypergroup ${\widehat{(H,\mathcal{B})}}$, see Lemma 2.6.
For any RN-hypergroup ${(H,\mathcal{B})}$, we denote by ${(H,\mathcal{B})}_{ad}$ its adjoint sub-hypergroup, see §6.2. As for fusion rings, this is defined as the sub-hypergroup of ${(H,\mathcal{B})}$ generated by the element $I(1):=\sum_{i=1}^{m}h_{i}x_{i}x_{i^{*}}$, where $h_{i}:=1/N^{1}_{i,i^{*}}$.
For any RN-hypergroup ${(H,\mathcal{B})}$, the central primitive idempotent corresponding to the character $\mathrm{FPdim}:H\rightarrow\mathbb{C}$ is denoted by $\lambda_{H}$. For $H=K(\mathcal{C})$, the Grothendieck ring of a fusion category $\mathcal{C}$, this primitive idempotent corresponds to the regular element.
Definition 1.5.
A hypergroup ${(H,\mathcal{B})}$ is called normalizable if one obtains a normalized hypergroup after rescaling it. (see §2.2 for more details.)
Based on Lemma 2.2 we denote such a hypergroup by ${(H,\mathcal{B},\mu_{1})}$ where $\mu_{1}\in{\widehat{\mathcal{B}}}$.
Inspired by the notion of dualizable probability groups from [Har79], we propose the following:
Definition 1.6.
An ARN-hypergroup is called dualizable if its dual is also ARN.
1.2. Main results of the paper
We characterize Burnside dualizable ARN-hypergroups as follows:
Theorem 1.7.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup such that ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN. The following are equivalent (see Definition 2.4 for $n(\_)$):
  1. (1)
    The hypergroup ${(H,\mathcal{B})}$ is Burnside.
  2. (2)
    The following identity holds on ${\widehat{(H,\mathcal{B})}}$:
    (1.4) $\big{(}\prod_{j=1}^{m}\mu_{j}\big{)}^{2}=\frac{1}{n(\widehat{(H,{\mathcal{B}})}_{ad})}\big{(}\sum_{\mu_{j}\in{{\widehat{\mathcal{B}}}}_{ad}}{\widehat{h}}_{j}\mu_{j}\big{)}.$
Note that the right hand side of the above equation coincides to the integral $\lambda_{{}_{{\widehat{(H,\mathcal{B})}}_{ad}}}$ of the adjoint hypergroup ${\widehat{(H,\mathcal{B})}}_{ad}$.
For a fusion category $\mathcal{C}$, its Grothendieck ring is denoted by $K(\mathcal{C})$, and the set of isomorphism class representatives of simple objects of $\mathcal{C}$ by $\mathrm{Irr}(\mathcal{C}):=\{X_{1},\dots,X_{m}\}$. Let $d_{i}:=\mathrm{FPdim}(X_{i})$ be the Frobenius-Perron dimension of $X_{i}$, $i\in{\mathcal{I}}=\{1,\dots,m\}$, and $x_{i}:=[X_{i}]$ be the class of the simple object $X_{i}$ in the Grothendieck ring $K(\mathcal{C})$.
For any weakly-integral premodular category $\mathcal{C}$, ${\widehat{{K(\mathcal{C})}}}$ is both RN and rational, as demonstrated in [Bur21, Theorem 1.2]. Furthermore, for any unitary fusion category $\mathcal{C}$, ${\widehat{{K(\mathcal{C})}}}$ is also RN, see [LPW21, ENO21], and primary $3$-criterion in [HLPW24]. A normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$ is called dual-Burnside if its dual ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is Burnside (see examples from group theory in §12). By duality, one obtains the following:
Theorem 1.8.
The following are equivalent for an abelian RN-hypergroup ${(H,\mathcal{B})}$.
  1. (1)
    The hypergroup ${(H,\mathcal{B})}$ is dual-Burnside.
  2. (2)
    The following identity holds:
    (1.5) $\big{(}\prod_{i=1}^{m}{\frac{x_{i}}{d_{i}}}\big{)}^{2}=\frac{1}{n({{(H,\mathcal{B})}_{ad}})}\big{(}\sum_{x_{i}\in\mathcal{B}_{ad}}d_{i}x_{i}\big{)}$
The dual-Burnside fusion categories can be characterized by:
Theorem 1.9.
Let $\mathcal{C}$ be a fusion category with a commutative Grothendieck ring. Then ${K(\mathcal{C})}$ is dual-Burnside if and only if the following holds:
(1.6) $\big{(}\prod_{i=1}^{m}{\frac{x_{i}}{d_{i}}}\big{)}^{2}=\frac{1}{\mathrm{FPdim}({\mathcal{C}_{\mathrm{ad}}})}\big{(}\sum_{x_{i}\in{\mathcal{C}_{\mathrm{ad}}}}d_{i}x_{i}\big{)}.$
About the dual-Burnside hypergroups, we prove the following:
Theorem 1.10.
Let $(H,{\mathcal{B}})$ be an abelian rational RN-hypergroup. If $(H,{\mathcal{B}})$ is dual-Burnside then $\mathrm{FPdim}(H,{\mathcal{B}})\in\mathbb{Q}$.
As a corollary, any commutative dual-Burnside fusion ring is weakly-integral. Similar to the fusion ring settings, [GN08], the universal grading group can be defined for any ARN-hypergroup. Using this grading, we can similarly define the concept of nilpotent ARN-hypergroup, and prove the following:
Theorem 1.11.
Let ${(H,\mathcal{B})}$ be a dualizable ARN-hypergroup. Then ${(H,\mathcal{B})}$ is nilpotent if and only if its dual ${\widehat{(H,\mathcal{B})}}$ is also nilpotent. Moreover, in this case they have the same nilpotency class.
For a nilpotent ARN-hypergroup, we prove the following:
Theorem 1.12.
A nilpotent dualizable ARN-hypergroup is both Burnside and dual-Burnside.
Let $\mathcal{C}$ be a modular fusion category. Then Corollary 10.3 states that ${K(\mathcal{C})}$ is Burnside if and only if it is dual-Burnside. The proof involves the Drinfeld map of $\mathcal{C}$. Then Theorem 1.9 implies that:
Theorem 1.13.
In any modular fusion category $\mathcal{C}$ with ${K(\mathcal{C})}$ Burnside:
(1.7) $\big{(}\prod_{i=1}^{m}\frac{x_{i}}{d_{i}}\big{)}^{2}=\frac{1}{\mathrm{FPdim}({\mathcal{C}_{\mathrm{ad}}})}\big{(}\sum_{x_{i}\in\mathrm{Irr}({\mathcal{C}_{\mathrm{ad}}})}d_{i}x_{i}\big{)}$
Note that the above theorem holds for weakly-integral modular categories since their Grothendieck rings are Burnside by [Bur23, Theorem 2].
Corollary 1.14.
In any modular fusion category $\mathcal{C}$ with ${K(\mathcal{C})}$ Burnside and with the group of invertible objects $G(\mathcal{C})$ of odd order,
(1.8) $\prod_{i=1}^{m}\frac{x_{i}}{d_{i}}=\frac{1}{\dim({\mathcal{C}_{\mathrm{ad}}})}\big{(}\sum_{x_{i}\in\mathrm{Irr}({\mathcal{C}_{\mathrm{ad}}})}d_{i}x_{i}\big{)}$
For any non-negative integer $n$, let $\mathcal{V}(n)$ be the set of prime divisors of $n$. For a weakly-integral fusion category we also let $\mathcal{V}(\mathcal{C}):=\mathcal{V}(\mathrm{FPdim}(\mathcal{C}))$. From Theorem 1.9, we can also derive the following result:
Theorem 1.15.
Let $\mathcal{C}$ be a fusion category such that ${K(\mathcal{C})}$ is commutative and dual-Burnside. Then
(1.9) $\frac{({\prod_{i=1}^{m}}d_{i})^{2}}{\mathrm{FPdim}({\mathcal{C}_{\mathrm{ad}}})}\in\mathbb{Z}.$
Moreover, if additionally $\mathcal{C}$ is nilpotent then
(1.10) $\mathcal{V}({\mathcal{C}_{\mathrm{ad}}})=\bigcup_{i=1}^{m}\mathcal{V}(d_{i}^{2}).$
Note that since any nilpotent fusion category $\mathcal{C}$ is weakly-integral then $d_{i}^{2}\in\mathbb{Z}$ for all $i$, see [GN08, Theorem 3.10]. Note that in the case of an integral fusion category Equation (1.10) can be written as
(1.11) $\mathcal{V}({\mathcal{C}_{\mathrm{ad}}})=\bigcup_{i=1}^{m}\mathcal{V}(d_{i}).$
Using the above identities, we prove the following results about the $\mathrm{FPdim}$ of simple objects:
Theorem 1.16.
For any modular fusion category $\mathcal{C}$ such that ${K(\mathcal{C})}$ is dual-Burnside,
(1.12) $\frac{({\prod_{i=1}^{m}}d_{i})^{2}}{\mathrm{FPdim}({\mathcal{C}_{\mathrm{ad}}})}\in\mathbb{Z}.$
For any weakly-integral modular fusion category $\mathcal{C}$:
(1.13) $\mathcal{V}(\mathcal{C})=\mathcal{V}({\mathcal{C}_{\mathrm{pt}}})\cup\big{(}\bigcup_{i=1}^{m}\mathcal{V}(d_{i}^{2})\big{)}$
Again, recall $d_{i}^{2}\in\mathbb{Z}$ for all $i$ by [GN08, Theorem 3.10]. Moreover, if $\mathcal{C}$ is an integral modular fusion category then the above equation becomes:
(1.14) $\mathcal{V}(\mathcal{C})=\mathcal{V}({\mathcal{C}_{\mathrm{pt}}})\cup\big{(}\bigcup_{i=1}^{m}\mathcal{V}(d_{i})\big{)}$
Concrete examples on which these results apply can be found in §12.3. Next theorem improves [OY23, Theorem 4.5].
Theorem 1.17.
Let $\mathcal{C}$ be a weakly-integral modular fusion category of $\mathrm{FPdim}(\mathcal{C})=dm$ with $d$ square-free coprime with $m$ and $d_{i}^{2}$ for all $i$. Then $d\mid\mathrm{FPdim}({\mathcal{C}_{\mathrm{pt}}})$ and $\mathcal{C}$ admits a decomposition
$\mathcal{C}\simeq{\mathcal{D}}\boxtimes{\mathcal{D}}^{\prime}$
where ${\mathcal{D}}$ is a pointed non-degenerate braided fusion category of dimension $d$ and ${\mathcal{D}}^{\prime}$ a non-degenerate braided fusion category of dimension $m$.
Corollary 1.18.
Let $\mathcal{C}$ be any integral modular fusion category of $\mathrm{FPdim}(\mathcal{C})=dm$ with $d$ square-free coprime with $m$. Then $d\mid\mathrm{FPdim}({\mathcal{C}_{\mathrm{pt}}})$ and $\mathcal{C}$ admits a decomposition
$\mathcal{C}\simeq{\mathcal{D}}\boxtimes{\mathcal{D}}^{\prime}$
where ${\mathcal{D}}$ is a pointed non-degenerate braided fusion category of dimension $d$ and ${\mathcal{D}}^{\prime}$ a non-degenerate braided fusion category of dimension $m$.
In particular, the $\mathrm{FPdim}$ of a perfect integral modular fusion category does not have a powerless prime factor (Corollary 10.11). In particular, a perfect modular fusion category of even $\mathrm{FPdim}$ has its $\mathrm{FPdim}$ divisible by $4$.
Remark 1.19.
By the classification of non-pointed integral modular fusion categories of small rank in [ABPP] and by Corollary 1.18, the $\mathrm{FPdim}$ of a non-pointed integral modular fusion category does not have a powerless prime factor for every rank less than $11\times 3=33$, and every rank less than $17\times 7=119$ in the odd-dimensional case.
Integral modular fusion categories were recently intensively studied in the literature. In [CP22] it was formulated the following conjecture
Conjecture 1.20.
There is no perfect modular fusion category of odd $\mathrm{FPdim}$.
Note that the above result also applies to non-pointed simple integral modular fusion categories (since they are in particular perfect). The importance of the existence of such categories was described in [LPR23, §5], where the following two open statements are proved to be equivalent, whereas Proposition 10.14 states that they imply Conjecture 1.20.
Statement 1.21.
Every simple integral fusion category is weakly group-theoretical.
Statement 1.22.
Every simple integral modular fusion category is pointed.
We finally prove the following result:
Theorem 1.23.
A perfect modular fusion category is (dual-)Burnside if and only if it is integral.
Note that the integral modular fusion category $\mathcal{Z}(\mathrm{Rep}(G))$ is perfect if and only if the finite group $G$ is perfect and centerless (more generally, see §11.1).
Shortly, the organization of this paper is the following. In §2 we recall the basics on hypergroups and abelian rational normalizable hypergroups. In §3 we develop few basic facts about Galois symmetries of rational hypergroups. In §4 hypergroups with Burnside property are studied and Theorem 1.7 is proven. In the same section we prove Theorem 1.2 and its more general version Theorem 4.17 for hypergroups. In this section, Burnside’s property for dual hypergroups is also studied and Theorem 1.8 is proven. In §5 we extend the notion of kernels of objects from the settings of fusion categories to the settings of ARN-hypergroups. In §7 we construct the universal grading group for ARN-hypergroups. A connection between this universal grading group and the group of grouplike elements of the dual it is proven in Theorem 6.5. §7 describes the constriction of the universal grading group of an ARN-hypergroup. In §8 we define nilpotent ARN-hypergroup and prove Theorems 1.11 and 1.12. The applications to modular fusion categories are given in §10.1. In §10 we prove Proposition 10.14 showing that any of the two statements from [LPR23, §5] implies Conjecture 1.20.
Next, §11 delves into applications and supplementary results: §11.1 offers a characterization for a Drinfeld center to be perfect; §11.2 establishes some results concerning integrality and Burnside properties, proving Theorems 1.10 and 1.23, particularly highlighting their equivalence in the perfect modular case; §11.3 characterizes the near-group fusion categories which admit a modular structure.
Finally, §12 presents concrete examples and counterexamples: §12.1 introduces non-Burnside simple integral fusion rings; §12.2 classifies certain dual-Burnside $\mathrm{Rep}(G)$; §12.3 outlines the classification of fusion rings lacking modular categorification within certain families, employing skills from this paper.
Throughout of this paper, all vector spaces and linear categories are considered over the ground field $\mathbb{C}$ of complex numbers.
Acknowledgements
We thank Vicenţiu Paşol and Andrew Schopieray for fruitful discussions on Galois groups of fusion rings and hypergroups.
2. Abelian rational normalizable hypergroups
Let ${(H,\mathcal{B})}$ be a hypergroup as in Definition 1.1. A morphism of hypergroups $\psi:(H,{\mathcal{B}})\rightarrow(H^{\prime},{\mathcal{B}}^{\prime})$ is an algebra morphism $\psi:H\rightarrow H^{\prime}$ such that $\psi({\mathcal{B}})={\mathcal{B}}^{\prime}$ and $\psi(b^{*})=\psi(b)^{*}$ for all $b\in\mathcal{B}$. The morphism $\psi$ is called an isomorphism of hypergroups if it is bijective.
Let ${(H,\mathcal{B})}$ be a normalized abelian hypergroup and ${\widehat{(H,\mathcal{B})}}$ its dual as defined in the introduction. There is a formula for the dual fusion coefficients ${\widehat{p}}_{k}(i,j)$ given in [Bur21, Proposition 2.1]. For a normalized abelian hypergroup ${(H,\mathcal{B})}$, this is given by:
(2.1) $\widehat{p}_{k}({j_{1}},{j_{2}})=\frac{1}{n_{k}}\bigg{(}\sum_{a\in{\mathcal{B}}}h_{a}\mu_{j_{1}}(a)\mu_{j_{2}}(a)\mu_{k}(a^{*})\bigg{)}$
where the scalars $n_{k}$ are defined below (2.2).
Remark 2.1.
Let ${(H,\mathcal{B})}$ be an abelian normalized hypergroup. We usually let :=${\mathcal{B}}:=\{x_{i}\}_{i\in\mathcal{I}}$, with :=$\mathcal{I}:=\{1,\dots m\}$ and $x_{1}=1$, and say that ${(H,\mathcal{B})}$ has rank $m$. We can simplify the notation $N_{x_{i}x_{j}}^{x_{k}}$ by $N_{i,j}^{k}$. Let $\{x_{i}^{\circ}\}_{i\in\mathcal{I}}$be the linear dual basis of the linear basis ${\mathcal{B}}$ of $H$. Therefore $\langle x_{i}^{\circ},x_{i^{\prime}}\rangle=\delta_{i,i^{\prime}}$, for all $i,i^{\prime}\in\mathcal{I}$. It is easy to see that in this case $\widehat{E}_{i}:=x_{i}^{\circ}\in H^{*}$ is a primitive central idempotent of $H^{*}$ corresponding to the character
$\begin{array}[]{cccc}\widetilde{\omega}_{i}=\operatorname{ev}_{x_{i}}:&H^{*}&\to&\mathbb{C}\\ &f&\mapsto&f(x_{i}).\end{array}$
2.1. Function $\tau$ and associative non-degenerate bilinear form
Let $(H,\mathcal{B})$ be any abelian hypergroup. Define $h_{i}:=1/N^{1}_{i,i^{*}}$, for all $i\in\mathcal{I}$. Note that $h_{1}=1$. Then $H$ is a commutative semisimple algebra, and therefore $H\simeq\bigoplus_{j=1}^{m}\mathbb{C}F_{j}$, where $F_{j}$ are the central primitive idempotents of $H$. One can define $\tau\in H^{*}$ with $\tau(x_{i})=\delta_{i,1}$, where as above, $x_{1}=1$. Note that $\tau=x_{1}^{\circ}$ with the above notations. Moreover, in this case
$\begin{array}[]{cccc}(\;,\;)_{\tau}:&H\otimes H&\to&\mathbb{C}\\ &(a,b)&\mapsto&\tau(ab)\end{array}$
is an associative symmetric non-degenerate bilinear form. Thus, one can write
(2.2) $\tau=\sum_{j=1}^{m}\frac{1}{n_{j}}\mu_{j}.$
for some non-zero scalars $n_{j}\in\mathbb{C}^{\times}$. These scalars $n_{j}$ are called the formal codegrees of $(H,\mathcal{B})$, see [Ost15]. Note that $\sum_{j=1}^{m}\frac{1}{n_{j}}=\tau(1)=1$.
From the definition of a hypergroup it follows that $\{h_{i}x_{i}\}_{i\in\mathcal{I}}$ and $\{x_{i^{*}}\}_{i\in\mathcal{I}}$ form a pair of dual bases for $(\;,\;)_{\tau}$. Note that in this case $\{n_{j}F_{j}\}_{i\in\mathcal{I}}$ and $\{F_{j}\}_{i\in\mathcal{I}}$ form also a pair of dual basis for the bilinear form $(\;,\;)_{\tau}$. The existence of the two pairs of dual bases implies that:
(2.3) $\sum_{i=1}^{m}{h_{i}}x_{i}\otimes x_{i}^{*}=\sum_{j=1}^{m}n_{j}F_{j}\otimes F_{j}$
From here it follows that
(2.4) $F_{j}:=\frac{1}{n_{j}}\big{(}\sum_{i=1}^{m}{h_{i}}\mu_{j}(x_{i^{*}})x_{i}\big{)}$
by applying $\operatorname{id}\otimes\mu_{j}$ for the character $\mu_{j}$ corresponding to $F_{j}$. Applying $\mu_{j}\otimes\mu_{k}$ to the above, one obtains the first orthogonality relations
(2.5) $\sum_{i=1}^{m}{h_{i}}\mu_{j}(x_{i})\mu_{k}(x_{i^{*}})=\delta_{j,k}n_{j}$
Since $\mu_{j}(x_{i^{*}})=\overline{\mu_{j}(x_{i})}$, for $j=k$, it follows from here that $n_{j}>0$. By a classical argument, as for finite groups orthogonality, one obtains the second orthogonality relation:
(2.6) $\sum_{j=1}^{m}\frac{1}{n_{j}}\mu_{j}(x_{i})\mu_{j}(x_{l}^{*})=\delta_{i,l}h_{i}^{-1}$
2.2. Normalizable abelian hypergroups and their duals
Note that hypergroups can be rescaled as follows. Let $y_{i}:=\frac{x_{i}}{\alpha_{{}_{i}}}$, for some non-zero complex numbers $\alpha_{{}_{i}}$ with $\alpha_{1}=1$, $\alpha_{{}_{i^{*}}}=\overline{\alpha_{{}_{i}}}$. Let :=$\mathcal{B}^{\prime}:=\{y_{i}\}_{i\in\mathcal{I}}$. Then it is easy to see that $(H,\mathcal{B}^{\prime})$ is also a hypergroup. However note that $\tau^{\prime}(\frac{x_{i}}{\alpha_{{}_{i}}})=\delta_{i,1}$ is the same functional $\tau$ on $H$ and therefore the formal codegrees $n_{j}$ do not change after rescaling.
Recall the notion of a normalizable hypergroup from Definition 1.5. Note that any rescaled normalizable hypergroup is also normalizable.
Lemma 2.2.
Let ${(H,\mathcal{B})}$ be a hypergroup. Then ${(H,\mathcal{B})}$ is normalizable if and only if there is $\mu_{1}\in{\widehat{\mathcal{B}}}$ such that $\mu_{1}(x_{i})\neq 0$ for all $1\leq i\leq m$.
Proof.
It is easy to see that if $(H,\{\frac{x_{i}}{\alpha_{{}_{i}}}\})$ is normalized after a rescaling, then $\mu_{1}:H\rightarrow H,\;x_{i}\mapsto\alpha_{{}_{i}}$ is a morphism of algebras on $H$. Thus $\mu_{1}\in{\widehat{\mathcal{B}}}$. Conversely, one can check directly that $(H,\{\frac{x_{i}}{\mu_{1}(x_{i})}\})$ is a normalized hypergroup for any morphism $\mu_{1}\in{\widehat{\mathcal{B}}}$ such that $\mu_{1}(x_{i})\neq 0$ for all $i\in\mathcal{I}$. ∎
We denote by ${(H,\mathcal{B},\mu_{1})}$ the normalizable hypergroup obtained by normalizing ${(H,\mathcal{B})}$ by $\mu_{1}\in{\widehat{\mathcal{B}}}$. We also denote by $\overline{{(H,\mathcal{B},\mu_{1})}}$ the normalized hypergroup obtained by rescaling ${(H,\mathcal{B},\mu_{1})}$ as in Lemma 2.2.
Definition 2.3.
Recall from Subsection 1.1 that any abelian normalized hypergroup $H$ admits a dual hypergroup $\widehat{H}$. In the case of $\overline{{(H,\mathcal{B},\mu_{1})}}$, we denote this dual by ${{\widehat{{(H,\mathcal{B},\mu_{1})}}}}$.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup via $\mu_{1}\in\widehat{\mathcal{B}}$ such that $\mu_{1}(x_{i})\neq 0$ for all $1\leq i\leq m$. From Equation (1.2) it follows that the multiplication on the dual of $\widehat{{(H,\mathcal{B},\mu_{1})}}$ can be written as
(2.7) $[f\star g](\frac{x_{i}}{\mu_{1}(x_{i})})=f(\frac{x_{i}}{\mu_{1}(x_{i})})g(\frac{x_{i}}{\mu_{1}(x_{i})})$
for all $f,g\in H^{*}$. Then it is easy to see that the algebra unit of ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is $\mu_{1}$. Moreover, the involution on $^{\widehat{\mathcal{B}}}$, as given in §1.1, becomes:
(2.8) $\mu_{j}^{\#}(x_{i})=\bigg{(}\frac{\mu_{1}(x_{i})}{\overline{\mu_{1}(x_{i}})}\bigg{)}\mu_{j}(x_{i}^{*}).$
Definition 2.4.
The order of a normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$ is defined by
$n{(H,\mathcal{B},\mu_{1})}:=\sum_{i=1}^{m}h_{i}|\mu_{1}(x_{i})|^{2}.$
In the abelian case, note that
(2.9) $n{(H,\mathcal{B},\mu_{1})}=n_{1}$
by Equation (2.5).
We shortly write $n(H)$ when $\mathcal{B}$ and $\mu_{1}$ are implicitly understood. In particular, for a normalized hypergroup one may choose $\mu_{1}(x_{i})=1$ for all $1\leq i\leq m$ and therefore $n(H)=\sum_{i=1}^{m}h_{i}$. For the rest of this paper we use the notation $d_{i}:=\mu_{1}(x_{i})$ for any normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$.
We say that a normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$ is weakly-rational if its order is a rational number, i.e. $n{(H,\mathcal{B},\mu_{1})}\in\mathbb{Q}$. Also, we say that ${(H,\mathcal{B},\mu_{1})}$ is weakly-integral if its order is an integer, i.e. $n{(H,\mathcal{B},\mu_{1})}\in\mathbb{Z}$. Note that a fusion ring is weakly-rational if and only if it is weakly-integral since $\mathrm{FPdim}(R)$ is an algebraic number in this case.
Remark 2.5.
Note that the formula from Equation (2.1) can be written for abelian normalizable hypergroups as:
(2.10) $\widehat{p}_{k}({j_{1}},{j_{2}})=\frac{1}{n_{k}}\bigg{(}\sum_{x_{i}\in{\mathcal{B}}}\frac{h_{i}}{d_{i}}\mu_{j_{1}}(x_{i})\mu_{j_{2}}(x_{i})\mu_{k}(x_{i^{*}})\bigg{)}$
Lemma 2.6.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. Then with the above notations, $n{(H,\mathcal{B},\mu_{1})}=n{\widehat{{(H,\mathcal{B},\mu_{1})}}}$.
Proof.
By applying (2.10) to ${j_{1}}={j_{2}}^{\#}=j$ and $k=1$, and then (2.5), one obtains (see also [Bur21, Equation (2.10)]),
(2.11) $\widehat{p}_{1}(j,j^{\#})=\frac{n_{j}}{n{(H,\mathcal{B},\mu_{1})}}$
Let ${\widehat{h}}_{j}:=\frac{1}{\widehat{p}_{1}(j,j^{*})}$ be the order of $\mu_{j}$ in ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$. It follows that $n{\widehat{{(H,\mathcal{B},\mu_{1})}}}=\sum_{j=1}^{m}{\widehat{h}}_{j}=n{(H,\mathcal{B},\mu_{1})}(\sum_{j=1}^{m}\frac{1}{n_{j}})=n{(H,\mathcal{B},\mu_{1})}$. ∎
Next proposition holds for any normalized hypergroup, not necessarily abelian.
Proposition 2.7.
For any normalized hypergroup ${(H,\mathcal{B})}$,
(2.12) $F_{1}=\frac{1}{n{(H,\mathcal{B})}}\big{(}\sum_{i=1}^{m}h_{i^{*}}x_{i}\big{)}$
where $F_{1}$ is the idempotent corresponding to the augmentation map $\mu_{1}:H\rightarrow\mathbb{C},\;a\mapsto 1$ for all $a\in\mathcal{B}$.
Proof.
Suppose that $F_{1}=\sum_{i=1}^{m}\alpha_{{}_{i}}x_{i}$ for some $\alpha_{{}_{i}}\in\mathbb{C}$. Recall that $x_{l}F_{i}=\mu_{i}(x_{l})F_{i}$. Then, $x_{l}F_{1}=F_{1}=\sum_{i=1}^{m}\alpha_{{}_{i}}x_{l}x_{i}$. It follows that $\tau(F_{1})=\alpha_{1}=\tau(x_{l}F_{1})=\alpha_{{}_{l^{*}}}h_{l}^{-1}=\tau(F_{1}x_{l})=\alpha_{l}h_{l^{*}}^{-1}.$ Then $F_{1}=\sum_{i=1}^{m}\alpha_{i}x_{i}={\alpha_{1}}(\sum_{i=1}^{m}{h_{i^{*}}}x_{i})$ and $1=\mu_{1}(F_{1})={\alpha_{1}}(\sum_{i=1}^{m}{h_{i}})$. This implies that $\alpha_{1}=\frac{1}{n(H)}$ and therefore $\alpha_{i}=\frac{h_{i^{*}}}{n(H)}$ for all $i\in\mathcal{I}$. ∎
Corollary 2.8.
In any normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$,
(2.13) $F_{1}=\frac{1}{n{(H,\mathcal{B},\mu_{1})}}\big{(}\sum_{i=1}^{m}h_{i^{*}}d_{i^{*}}x_{i}\big{)}$
Proposition 2.9.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalized hypergroup. Then ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is also an abelian normalized hypergroup and the formal codegrees of ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ are given by
(2.14) $\widehat{n}_{i}=\frac{n{(H,\mathcal{B},\mu_{1})}}{h_{i}}.$
Proof.
As already mentioned, [Bur21, Proposition 2.3] shows that ${\widehat{(H,\mathcal{B})}}$ is also an abelian normalized hypergroup.
By Remark 2.1, $x_{i}:H^{*}\rightarrow\mathbb{C}$ are the characters of ${\widehat{(H,\mathcal{B})}}$. By definition of $\widehat{\tau}\in(H^{*})^{*}=H$, $\widehat{\tau}(\mu_{j})=\delta_{j,1}$ where $\mu_{1}$ is the unit of $H^{*}$. Thus $\widehat{\tau}=F_{1}$. Equation (2.2) for $H^{*}$ gives that $\widehat{\tau}=\sum_{i=1}^{m}\frac{1}{\widehat{n}_{i}}x_{i}$. Then Equation (2.12) finishes the proof since $n{\widehat{(H,\mathcal{B})}}=n{(H,\mathcal{B})}$, by Lemma 2.6. ∎
Corollary 2.10.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. Then the primitive central idempotent of the linear character $x_{1}\in\widehat{{\widehat{\mathcal{B}}}}=\mathcal{B}$ is given by
(2.15) $\widehat{F_{1}}=\sum_{j=1}^{m}\frac{\mu_{j}}{n_{j}}\in H^{*}.$
Proof.
Proposition 2.9 shows in its proof that $\widehat{\tau}=F_{1}$. By duality, since $\widehat{\widehat{{(H,\mathcal{B},\mu_{1})}}}=\overline{{(H,\mathcal{B},\mu_{1})}}$, it follows that $\tau=\widehat{F}_{1}$. Equation (2.2) finishes the proof. ∎
A morphism of normalizable hypergroups $\psi:(H,{\mathcal{B}},\mu_{1})\rightarrow(H^{\prime},{\mathcal{B}}^{\prime},\mu_{1}^{\prime})$ is a morphism of hypergroups with the property that $\mu^{\prime}_{1}\circ f=\mu_{1}$.
2.3. Frobenius-Perron theory for a RN-hypergroups
Frobenius-Perron theory can be defined for RN-hypergroups, in the same manner as for fusion rings, see [EGNO15, §3]. We denote by $\mathrm{FPdim}(x_{i})$ the Frobenius-Perron eigenvalue of the left multiplication operator by $x_{i}$ on $H$. As for fusion rings, by a similar argument, it can be shown that $\mathrm{FPdim}(x_{i})=\mathrm{FPdim}(x_{i^{*}})$. Indeed, it easy to see that the matrix $N_{i^{*}}$ corresponding to $x_{i^{*}}$ is the transpose of the matrix $N_{i}$ of $x_{i}$ multiplied by the scalars $\frac{h_{j}}{h_{k}}$ in the $(j,k)$-entry. More precisely, $(N_{i^{*}})_{jk}=\frac{h_{j}}{h_{k}}(N_{i})_{kj}$. Thus, the two matrices have the same set of eigenvalues, in particular the same Frobenius-Perron eigenvalue. Note that any RN-hypergroup ${(H,\mathcal{B})}$ can be normalized by $\mu_{1}=\mathrm{FPdim}$, since $\mu_{1}(x_{i})>0$ for all $1\leq i\leq m$.
Remark 2.11.
Note that any symmetric RN-hypergroup ${(H,\mathcal{B})}$ is transitive in the sense of [EGNO15, Definition 3.3.1]. Indeed the function $\tau$ from §3.1 is symmetric and therefore $\tau(x_{i}x_{j}x_{k})=\tau(x_{j}x_{k}x_{i})=\tau(x_{k}x_{i}x_{j})$. Since $\tau(x_{i}x_{j}x_{k})=h_{i}^{-1}m(x_{i^{*}},x_{j}x_{k})$ one has
(2.16) $h_{i}^{-1}m(x_{i^{*}},x_{j}x_{k})=h_{j}^{-1}m(x_{j^{*}},x_{k}x_{i})=h_{k}^{-1}m(x_{k^{*}},x_{i}x_{j}).$
which shows that ${(H,\mathcal{B})}$ is transitive.
Then it is easy to see that an analogue [EGNO15, Proposition 3.3.6] also holds for symmetric RN-hypergroups. In particular it follows that $\mathrm{FPdim}:H\rightarrow\mathbb{C}$ is the unique which takes positive values on $\mathcal{B}$.
Definition 2.12.
In this case of a RN-hypergroup ${(H,\mathcal{B})}$, one can define the Frobenius-Perron dimension as
$\mathrm{FPdim}{(H,\mathcal{B})}:=\sum_{i=1}^{m}h_{i}\mathrm{FPdim}(x_{i})^{2}.$
Lemma 2.13.
For a normalized RN-hypergroup,
$\mathrm{FPdim}{(H,\mathcal{B})}=n{(H,\mathcal{B},\mathrm{FPdim})}.$
Proof.
It is easy to see that [EGNO15, Proposition 3.3.6] also works for normalized RN-hypergroups instead of fusion rings. Therefore $\mu_{1}:H\rightarrow H,x_{i}\mapsto 1$ is the unique morphism $\phi:H\rightarrow\mathbb{C}$ with $\phi(x_{i})>0$ for all $1\leq i\leq m$. This implies that $\mathrm{FPdim}(x_{i})=1$ for all $1\leq i\leq m$ and therefore $n{(H,\mathcal{B},\mathrm{FPdim})}=\mathrm{FPdim}{(H,\mathcal{B})}$. ∎
Note that by Lemma 2.2, any RN-hypergroup is normalizable via the character $\mu_{1}=\mathrm{FPdim}$. Lemma 2.13 also shows that
$n(H,\mathcal{B},\mathrm{FPdim})=\sum_{i=1}^{m}h_{i}|\mu_{1}(x_{i})|^{2}=\sum_{i=1}^{m}h_{i}\mathrm{FPdim}(x_{i})^{2}=\mathrm{FPdim}{(H,\mathcal{B})}$
for any RN-hypergroup. For a RN-hypergroup, we sometimes write ${(H,\mathcal{B},\mathrm{FPdim})}$ instead of ${(H,\mathcal{B})}$ to specify that we normalize it by $\mathrm{FPdim}$.
Proposition 2.14.
Let ${(H,\mathcal{B},\mu_{1})}$ be a normalizable hypergroup. Then the order $n{(H,\mathcal{B},\mu_{1})}$ is invariant under rescaling. If ${(H,\mathcal{B})}$ is RN-hypergroup, then the Frobenius-Perron dimension of ${(H,\mathcal{B})}$ is also invariant under rescaling with real positive numbers.
Proof.
First we show that the order of a normalizable hypergroup is invariant under rescaling. Let $y_{i}:=\frac{x_{i}}{\alpha_{{}_{i}}}$, for some non-zero complex numbers $\alpha_{{}_{i}}$ with $\alpha_{1}=1$, $\alpha_{{}_{i^{*}}}=\overline{\alpha_{{}_{i}}}$. Consider $(H,\mathcal{B}^{\prime},\mu_{1})$ the rescaled hypergroup with :=$\mathcal{B}^{\prime}:=\{y_{i}\}_{i\in\mathcal{I}}$. Note that the order of $y_{i}$ is given by ${h^{\prime}_{i}}^{-1}=N^{1^{\prime}}_{i^{\prime}i^{{}^{\prime}*}}=\tau(x_{i^{\prime}}x_{i^{{}^{\prime}*}})=\frac{1}{|\alpha_{{}_{i}}|^{2}}h_{i}^{-1}$ and this implies that $h^{\prime}_{i}=|\alpha_{{}_{i}}|^{2}h_{i}>0.$ Therefore
$n(H,\mathcal{B}^{\prime},\mu_{1})=\sum_{i=1}^{m}h^{\prime}_{i}\mid\mu_{1}(y_{i})\mid^{2}=\sum_{i=1}^{m}h_{i}\mid\mu_{1}(x_{i})\mid^{2}=n{(H,\mathcal{B},\mu_{1})}.$
Moreover, if ${(H,\mathcal{B})}$ is a RN-hypergroup and $\alpha_{{}_{i}}=\alpha_{{}_{i^{*}}}>0$, for all $1\leq i\leq m$, it follows that
$\mathrm{FPdim}(H^{\prime},\mathcal{B}^{\prime})$ $=$ $\sum_{i=1}^{m}h^{\prime}_{i}\mathrm{FPdim}(y_{i})^{2}=$
$=$ $\sum_{i=1}^{m}\mid\alpha_{{}_{i}}\mid^{2}h_{i}(\frac{\mathrm{FPdim}(x_{i})^{2}}{\alpha_{{}_{i}}})^{2}$
$=$ $\mathrm{FPdim}{(H,\mathcal{B})}.$
Corollary 2.15.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. Then with the above notations, the dual formal codegrees are given by
(2.17) $\widehat{n}_{i}=\frac{n{(H,\mathcal{B},\mu_{1})}}{h_{i}d_{i}d_{i^{*}}}.$
Proof.
[…]
Remark 2.16.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian rational normalizable hypergroup. In order to find the central primitive idempotents of $H^{*}$, one can apply Remark 2.1 after normalization. If $x_{i}^{\circ}\in H^{*}$ are defined as above by $\langle x_{i}^{\circ},x_{i^{\prime}}\rangle=\delta_{i,i^{\prime}}$, then ${\widehat{E}}_{i}:=d_{i}x_{i}^{\circ}\in H^{*}$ are the primitive central idempotents of $H^{*}$ corresponding to the character
$\begin{array}[]{cccc}\widetilde{\omega}_{i}=\operatorname{ev}_{{\frac{x_{i}}{d_{i}}}}:&H^{*}&\to&\mathbb{C}\\ &\mu_{j}&\mapsto&\mu_{j}({\frac{x_{i}}{d_{i}}}).\end{array}$
2.4. Grouplike elements in hypergroups
In the rest of this paper all abelian RN-hypergroups are normalized by $\mu_{1}=\mathrm{FPdim}$ and therefore their duals are also considered with respect to $\mu_{1}=\mathrm{FPdim}$.
Lemma 2.17.
For any symmetric RN-hypergroup ${(H,\mathcal{B})}$, $h_{i}d_{i}^{2}\geq 1$. Moreover, $h_{i}d_{i}^{2}=1$ if and only $h_{i}x_{i}x_{i^{*}}=x_{1}.$
Proof.
One has
(2.18) $x_{i}x_{i^{*}}=\frac{1}{h_{i}}x_{1}+\sum_{k=1}^{m}N^{k}_{ii^{*}}x_{k}.$
Passing to $\mathrm{FPdim}$, $h_{i}d_{i}^{2}=1+h_{i}\big{(}\sum_{k=1}^{m}N^{k}_{ii^{*}}d_{k}\big{)}\geq 1.$ If $h_{i}d_{i}^{2}=1$ then $\sum_{k=1}^{m}N^{k}_{ii^{*}}d_{k}=1$, and therefore $N^{k}_{ii^{*}}=0$ for all $k\neq 1$. Thus $h_{i}x_{i}x_{i^{*}}=x_{1}.$ Conversely, if $h_{i}x_{i}x_{i^{*}}=x_{1}$, applying $\mathrm{FPdim}$, it follows that $h_{i}\mathrm{FPdim}(x_{i})^{2}=1$. ∎
Corollary 2.18.
In any fusion ring ${(H,\mathcal{B})}$, $x_{i}x_{i^{*}}=1$ if and only if $d_{i}=1$.
Recall the set of grouplike elements ${G{(H,\mathcal{B})}}$ from Definition 1.3.
Remark 2.19.
For a normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$ one has that $x_{i}\in G{(H,\mathcal{B})}$ if and only if
(2.19) ${\frac{x_{i}}{d_{i}}}\frac{x_{i^{*}}}{d_{i^{*}}}=1$
Indeed applying $\mu_{1}$ to Equation 1.1 one has $h_{i}d_{i}d_{i^{*}}=1$ and Equation (2.19) follows. The converse also follows by inspecting the multiplicity of $x_{1}$ in $x_{i}x_{i^{*}}$.
Lemma 2.20.
Suppose that ${(H,\mathcal{B},\mu_{1})}$ is an abelian normalizable hypergroup. Then $x_{i}\in G{(H,\mathcal{B})}$ if and only if
(2.20) $|\mu_{j}({\frac{x_{i}}{d_{i}}})|=1\;\;\text{for all}\;\;j\in\mathcal{J}.$
Proof.
If $x_{i}\in G{(H,\mathcal{B})}$ then applying $\mu_{j}$ to Equation (2.19) one obtains $|\mu_{j}({\frac{x_{i}}{d_{i}}})|^{2}=1$ since $\mu_{j}(x_{i^{*}})=\overline{\mu_{j}(x_{i})}$. Conversely, if $\mu_{j}({\frac{x_{i}}{d_{i}}}\frac{x_{i^{*}}}{d_{i^{*}}})=1$ for all $\mu_{j}$ then since ${(H,\mathcal{B})}$ is abelian it follows that ${\frac{x_{i}}{d_{i}}}\frac{x_{i^{*}}}{d_{i^{*}}}=1$ and therefore $h_{i}x_{i}x_{i^{*}}=1$. ∎
Dually one has the following:
Lemma 2.21.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. Then $\mu_{j}\in G{\widehat{(H,\mathcal{B})}}$ if and only if
(2.21) $|\mu_{j}({\frac{x_{i}}{d_{i}}})|=1\;\;\text{for all}\;\;i\in\mathcal{I}.$
Proof.
The dual version of Equation (2.19) implies that $\mu_{j}\in G{\widehat{(H,\mathcal{B})}}$ if and only if $\mu_{j}\mu_{j^{\#}}=\mu_{1}.$ Then evaluating at ${\frac{x_{i}}{d_{i}}}$ one gets the desired identity. ∎
Corollary 2.22.
In any abelian normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$ for any $\mu_{j}\in G{\widehat{(H,\mathcal{B})}}$ one has that $n_{j}=n(H,B,\mu_{1})$.
Proof.
By Lemma 2.21 one has $\mid\mu_{j}(x_{i})|=|d_{i}|$ for all $i\in\mathcal{I}$. By Equation (2.5) one has $n_{j}=\sum_{i=1}^{m}h_{i}|\mu_{j}(x_{i})|^{2}=\sum_{i=1}^{m}h_{i}|d_{i}|^{2}=n(H,B,\mu_{1})$. ∎
Lemma 2.23.
Suppose that ${(H,\mathcal{B},\mu_{1})}$ is an abelian normalizable hypergroup such that one of the following holds:
  1. (1)
    ${(H,\mathcal{B})}$ is RN and $\mu_{1}=\mathrm{FPdim}$,
  2. (2)
    ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN.
Then
(2.22) $|\mu_{j}(x_{i})|\leq|d_{i}|\;\;\text{for all}\;i\in\mathcal{I}\;\text{and}\;j\in\mathcal{J}.$
Proof.
Suppose first that ${(H,\mathcal{B})}$ is RN and $\mu_{1}=\mathrm{FPdim}$. Thus $d_{i}={\mathrm{FPdim}(x_{i})}>0$ in this case. One has $x_{i}=\sum_{j=1}^{m}\mu_{j}(x_{i})F_{j}$ and therefore $\mu_{j}(x_{i})$ is an eigenvalue of $L_{x_{i}}$. It follows that $|\mu_{j}(x_{i})|\leq d_{i}$.
Suppose now that ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN. It follows that $\mathrm{FPdim}(\mu_{j})=1$, since by Remark 2.11 in this case $\mu_{j}\mapsto 1$ is the only morphism on ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ which takes positive values on $^{\widehat{\mathcal{B}}}$. Then, the equation
(2.23) $\mu_{j}=\sum_{i=1}^{m}\mu_{j}({\frac{x_{i}}{d_{i}}}){\widehat{E}}_{i}$
implies as above that $\big{|}\mu_{j}({\frac{x_{i}}{d_{i}}})\big{|}\leq 1$. ∎
Lemma 2.24.
Suppose that ${(H,\mathcal{B},\mu_{1})}$ is an abelian normalizable hypergroup such that one of the following holds:
  1. (1)
    ${(H,\mathcal{B})}$ is RN and $\mu_{1}=\mathrm{FPdim}$
  2. (2)
    ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN.
Then $n_{j}\leq n{(H,\mathcal{B},\mu_{1})}$ for all $j\in\mathcal{J}$. Moreover, $n_{j}=n{(H,\mathcal{B},\mu_{1})}$ if and only if $\mu_{j}$ is a grouplike element.
Proof.
Lemma 2.23 and Equation (2.5) give:
$n_{j}=\sum_{i=1}^{m}h_{i}|\mu_{j}(x_{i})|^{2}\leq\sum_{i=1}^{m}h_{i}|d_{i}|^{2}=n(H,B,\mu_{1})$
and equality holds if and only if $|\mu_{j}(x_{i})|^{2}=|d_{i}|^{2}$. By Lemma 2.21 this happens if and only if $\mu_{j}\in G{(H,\mathcal{B})}$. ∎
Lemma 2.25.
In any symmetric RN-hypergroup ${(H,\mathcal{B})}$, the set $\{{\frac{x_{i}}{d_{i}}}\}$ of normalized grouplike-elements form a finite group. Moreover, the inverse of ${\frac{x_{i}}{d_{i}}}$ is given by $\frac{x_{i^{*}}}{d_{i^{*}}}$ for all $x_{i}\in{G{(H,\mathcal{B})}}$.
Proof.
If $x_{i}x_{i^{*}}=\frac{1}{h_{i}}x_{1}$ and $x_{j}x_{j^{*}}=\frac{1}{h_{j}}x_{1}$ it follows that $(x_{i}x_{j})(x_{i}x_{j})^{*}=\frac{1}{h_{i}h_{j}}x_{1}.$ Thus $(\sum_{k=1}^{m}N^{k}_{ij}x_{k})(\sum_{k=1}^{m}N^{k}_{ij}x_{k^{*}})=\frac{1}{h_{i}h_{j}}x_{1}$ and this implies that $x_{i}x_{j}=\frac{d_{j}d_{i}}{d_{k}}x_{k}$ for a certain $k\in\mathcal{J}$. Moreover, since $h_{i}=\frac{1}{d_{i}^{2}}$ and $h_{j}=\frac{1}{d_{j}^{2}}$ it follows that
$h_{k}^{-1}=m(x_{1},x_{k}x_{k}^{*})=\frac{d_{k}^{2}}{d_{i}^{2}d_{j}^{2}}m(x_{1},x_{i}x_{j}x_{i^{*}}x_{j^{*}})=\frac{d_{k}^{2}}{d_{i}^{2}d_{j}^{2}}\frac{1}{h_{i}h_{j}}=d_{k}^{2}$
which shows that $x_{k}$ is also a grouplike element ($m$ is defined in §5.1). ∎
We denote $\overline{G{(H,\mathcal{B})}}:=\{{\frac{x_{i}}{d_{i}}}\;|\;x_{i}\in G{(H,\mathcal{B})}\}$ the above group of normalized grouplike elements. Clearly $\overline{G{(H,\mathcal{B})}}=G\overline{{(H,\mathcal{B})}}$, the grouplike elements of the normalized hypergroup $\overline{{(H,\mathcal{B})}}$.
By duality one obtains the following:
Lemma 2.26.
If ${(H,\mathcal{B},\mu_{1})}$ is an abelian normalizable hypergroup and the dual ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN then ${G{\widehat{(H,\mathcal{B})}}}$ is a group.
Recall that ¯=$\overline{\mathcal{B}}=\{\frac{x_{k}}{d_{k}}\;|\;x_{k}\in{\mathcal{B}}\}$.
Lemma 2.27.
Let ${(H,\mathcal{B})}$ be a symmetric RN-hypergroup. If $x_{i}\in{G{(H,\mathcal{B})}}$ is a grouplike element then ${\frac{x_{i}}{d_{i}}}\frac{x_{j}}{d_{j}},\frac{x_{j}}{d_{j}}{\frac{x_{i}}{d_{i}}}\in\overline{\mathcal{B}}$, for any $x_{j}\in\mathcal{B}$.
Proof.
Suppose that
$x_{i}x_{j}=\sum_{k\in\mathcal{A}}N^{k}_{ij}x_{k},$
where $\mathcal{A}$ is a set such that $N^{k}_{ij}>0$. Multiplying the above equality by $x_{i^{*}}$, one obtains
$\frac{1}{h_{i}}x_{j}=\sum_{k\in\mathcal{A}}N^{k}_{ij}x_{i^{*}}x_{k},$
since $h_{i}x_{i^{*}}x_{i}=x_{1}$ as $x_{i}$ is a grouplike element. Since ${(H,\mathcal{B},\mu_{1})}$ is a RN-hypergroup, it follows that $x_{i^{*}}x_{k}=\alpha_{ik}x_{j}$, for some scalar $\alpha_{ik}\in\mathbb{C}$ and all $k\in\mathcal{A}$. Thus $\alpha_{ik}=\frac{d_{i}d_{k}}{d_{j}}$. Moreover, multiplying the last equality by $x_{i}$, one obtains that $\frac{1}{h_{i}}x_{k}=x_{i^{*}}x_{j}$, which shows that $\mathcal{A}$ has a single element. ∎
Lemma 2.28.
Suppose that ${(H,\mathcal{B})}$ is an abelian RN hypergroup and $\mu_{j}\in{\widehat{(H,\mathcal{B})}}$ a character such that $\mu_{j}(x_{m})=d_{m}\omega_{m}$ and $\mu_{j}(x_{n})=d_{n}\omega_{n}$ for some scalars $|\omega_{m}|=|\omega_{n}|=1$. Then $\mu_{j}(x_{p})=\omega_{m}\omega_{n}d_{p}$ for all constituents $x_{p}$ of $x_{m}x_{n}$.
Proof.
Straightforward from the absolute value triangle inequality. ∎
Lemma 2.29.
If ${(H,\mathcal{B})}$ is an abelian RN hypergroup then ${{G{\widehat{(H,\mathcal{B})}}}}$ is a group.
Proof.
Suppose that $\mu_{j},\mu_{k}\in{{G{(H,\mathcal{B})}}}$. By Lemma 2.21 it follows that $\mu_{j}(x_{m})=\omega_{m}\mathrm{FPdim}(x_{m})$ for some roots of unity $\omega_{m}$. Similarly $\mu_{k}(x_{m})=\eta_{m}\mathrm{FPdim}(x_{m})$ for some other root of unity $\eta_{m}$. Moreover, by lemma 2.28, since ${(H,\mathcal{B})}$ is RN it follows that all the constituents $x_{m}x_{n}$ when evaluated at $\mu_{j}$ have the scalar $\omega_{m}\omega_{n}$. On the other hand, evaluated at $\mu_{k}$ these constituents give the scalar $\eta_{m}\eta_{n}$. Thus
$(\mu_{j}\star\mu_{k})(x_{m}x_{n})$ $=$ $\sum_{p}N^{p}_{mn}\mathrm{FPdim}(x_{p})(\mu_{j}\star\mu_{k})({\frac{x_{p}}{\mathrm{FPdim}(x_{p})}})$
$=$ $\sum_{p}N^{p}_{mn}\mathrm{FPdim}(x_{p})\mu_{j}({\frac{x_{p}}{\mathrm{FPdim}(x_{p})}})\mu_{k}({\frac{x_{p}}{\mathrm{FPdim}(x_{p})}})$
$=$ $\omega_{m}\omega_{n}\eta_{m}\eta_{n}\mathrm{FPdim}(x_{m}x_{n})$
On the other hand, note that $\mu_{j}(x_{m})=\mathrm{FPdim}(x_{m})\mu_{j}({\frac{x_{m}}{\mathrm{FPdim}(x_{m})}})=\omega_{m}\mathrm{FPdim}(x_{m})$ and similarly $\mu_{k}(x_{m})=\mathrm{FPdim}(x_{m})\mu_{k}({\frac{x_{m}}{\mathrm{FPdim}(x_{m})}})=\eta_{m}\mathrm{FPdim}(x_{m}).$ Therefore
$(\mu_{j}\star\mu_{k})(x_{m})(\mu_{j}\star\mu_{k})(x_{n})=\omega_{m}\eta_{m}\mathrm{FPdim}(x_{m})\omega_{n}\eta_{n}\mathrm{FPdim}(x_{n})$
This shows that $(\mu_{j}\star\mu_{k})(x_{m}x_{n})=(\mu_{j}\star\mu_{k})(x_{m})(\mu_{j}\star\mu_{k})(x_{n})$ which shows that $(\mu_{j}\star\mu_{k})\in{\widehat{\mathcal{B}}}$. Moreover, $(\mu_{j}\star\mu_{k})({\frac{x_{p}}{\mathrm{FPdim}(x_{p})}})=\mu_{j}({\frac{x_{p}}{\mathrm{FPdim}(x_{p})}})\mu_{k}({\frac{x_{p}}{\mathrm{FPdim}(x_{p})}})=\omega_{p}\eta_{p}$ and Lemma 2.21 implies that $\mu_{j}\star\mu_{k}\in G{\widehat{(H,\mathcal{B})}}$. ∎
Proposition 2.30.
Let ${(H,\mathcal{B})}$ be an abelian RN hypergroup and $\mu\in{{G{\widehat{(H,\mathcal{B})}}}}$. It follows that $\mu\star\mu_{k}\in\widehat{B}$ for any $\mu_{k}\in\widehat{B}$.
Proof.
Since $\mu\in{{G{\widehat{(H,\mathcal{B})}}}}$ and ${(H,\mathcal{B})}$ is RN, by Lemma 2.21 one has $\mu(x_{i})=\omega_{i}d_{i}$ for some root of unity $\omega_{i}$ and for all $i\in\mathcal{I}$. Here $d_{i}=\mathrm{FPdim}(x_{i})$. Then as above it follows that $\mu(x_{k})=\omega_{i}\omega_{j}d_{k}$ for any constituent $x_{k}$ of $x_{i}x_{j}$.
One can compute directly that $\mu\star\mu_{k}$ and $\mu_{k}\star\mu$ are morphisms of algebras. Indeed, note that $(\mu\star\mu_{k})({\frac{x_{i}}{d_{i}}})=\omega_{i}\mu_{k}({\frac{x_{i}}{d_{i}}})$ and therefore
$(\mu\star\mu_{k})(x_{i}x_{j})=\sum_{l=1}^{m}N^{l}_{ij}(\mu\star\mu_{k})(x_{l})=\omega_{i}\omega_{j}\sum_{l=1}^{m}N^{l}_{ij}\mu_{k}(x_{l})=\omega_{i}\omega_{j}\mu_{k}(x_{i}x_{j}).$
On the other hand
$(\mu\star\mu_{k})(x_{i})(\mu\star\mu_{k})(x_{j})=\omega_{i}\mu_{k}(x_{i})\omega_{j}\mu_{k}(x_{j})=\omega_{i}\omega_{j}\mu_{k}(x_{i}x_{j}).$
Thus $(\mu\star\mu_{k})(x_{i}x_{j})=(\mu\star\mu_{k})(x_{i})(\mu\star\mu_{k})(x_{j})$ which shows that $\mu\star\mu_{k}\in{\widehat{\mathcal{B}}}$. ∎
In the dual situation, Lemma 2.29 and Proposition 2.30 can be written as follows:
Proposition 2.31.
If ${(H,\mathcal{B},\mu_{1})}$ is an abelian normalizable hypergroup such that ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN then $G{(H,\mathcal{B})}$ is a group. Moreover, left multiplication by $L_{{\frac{x_{i}}{d_{i}}}}$ for any $x_{i}\in{G{(H,\mathcal{B})}}$ permutes the normalized standard basis $\overline{\mathcal{B}}$.
Note that Lemmata 2.25, 2.26, 2.27, 2.29 and Propositions 2.30, 2.31 imply the following:
Corollary 2.32.
Suppose that ${(H,\mathcal{B},\mu_{1})}$ is an abelian normalizable hypergroup such that one of the following holds:
  1. (1)
    ${(H,\mathcal{B})}$ is RN and $\mu_{1}=\mathrm{FPdim}$
  2. (2)
    ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN.
Then the following statements hold:
  1. (1)
    Both $\overline{G{(H,\mathcal{B})}}$ and $G{\widehat{(H,\mathcal{B})}}$ are finite groups.
  2. (2)
    The operator given by left multiplication $L_{\frac{x_{i}}{d_{i}}}$ with $x_{i}\in G{(H,\mathcal{B})}$ permutes the normalized standard basis $\overline{\mathcal{B}}$.
  3. (3)
    The operator given by left multiplication $L_{\mu_{j}}$ with $\mu_{j}\in G{\widehat{(H,\mathcal{B})}}$ permutes the standard basis $^{{\widehat{\mathcal{B}}}}$.
2.5. Some results on $P$ and ${\widehat{P}}$.
Lemma 2.33.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalized hypergroup, and let $x_{i}\in G{(H,\mathcal{B})}$ be a grouplike element. Then
(2.24) $<\prod_{j=1}^{m}\mu_{j},\;{\frac{x_{i}}{d_{i}}}>=\pm 1.$
Proof.
By the definition of the multiplication in the dual,
$<\prod_{j=1}^{m}\mu_{j},\;{\frac{x_{i}}{d_{i}}}>=\prod_{j=1}^{m}\mu_{j}({\frac{x_{i}}{d_{i}}}).$
We group together the factors of $\mu_{j}$ and $\mu_{j\#}$ in the above product. Note that $\mu_{j}({\frac{x_{i}}{d_{i}}})\mu_{j^{\#}}({\frac{x_{i}}{d_{i}}})=\mid\mu_{j}({\frac{x_{i}}{d_{i}}})\mid^{2}=1$ by Lemma 2.20. On the other hand, if $\mu_{j}=\mu_{j^{\#}}$, then $\mu_{j}({\frac{x_{i}}{d_{i}}})\in\mathbb{R}$ and therefore $\mu_{j}({\frac{x_{i}}{d_{i}}})=\pm 1$ by the same Lemma 2.20. ∎
Definition 2.34.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup and let $x_{i}\in{G{(H,\mathcal{B})}}$ be a grouplike element. We denote by ${\mathrm{sgn}}(x_{i})$ the above value of the product $\prod_{j=1}^{m}\mu_{j}$ evaluated at ${\frac{x_{i}}{d_{i}}}$.
The dual version of the above result is the following:
Lemma 2.35.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalized hypergroup, and let $\mu_{j}\in G{\widehat{(H,\mathcal{B})}}$ be a grouplike element. Then
(2.25) $<\mu_{j},\;{\prod_{i=1}^{m}}{\frac{x_{i}}{d_{i}}}>=\pm 1.$
By duality we can also define:
Definition 2.36.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup and let $\mu_{j}\in G{\widehat{(H,\mathcal{B})}}$ be a grouplike element. We denote by ${\mathrm{sgn}}(\mu_{j})$ the above value of the product ${\prod_{i=1}^{m}}{\frac{x_{i}}{d_{i}}}$ evaluated by $\mu_{j}$.
Remark 2.37.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalized hypergroup and let $x_{i}\in{G{(H,\mathcal{B})}}$ be a grouplike element. Then $L_{{\frac{x_{i}}{d_{i}}}}$, the left multiplication by ${\frac{x_{i}}{d_{i}}}$ on $H$ induces a permutation on $\overline{\mathcal{B}}$. Because ${\frac{x_{i}}{d_{i}}}=\sum_{j=1}^{m}\mu_{j}({\frac{x_{i}}{d_{i}}})F_{j}$, the determinant $\det(L_{\frac{x_{i}}{d_{i}}})$ of the matrix of the operator $L_{{\frac{x_{i}}{d_{i}}}}$ is $\prod_{j=1}^{m}\mu_{j}({\frac{x_{i}}{d_{i}}})$. On the other hand, $L_{{\frac{x_{i}}{d_{i}}}}$ permutes the normalized basis ¯:=$\overline{\mathcal{B}}:=\{{\frac{x_{i}}{d_{i}}}\}_{i\in\mathcal{I}}$, and therefore, ${\mathrm{sgn}}(x_{i})$ is the signature of the corresponding permutation.
Remark 2.38.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup and $\mu_{j}\in G{\widehat{{(H,\mathcal{B},\mu_{1})}}}$. It follows that ${\mathrm{sgn}}(\mu_{j})=\det(L_{\mu_{j}})=\pm 1$ since $\mu_{j}$ determines a permutation of $^{\widehat{\mathcal{B}}}$.
Definition 2.39.
A symmetric hypergroup ${(H,\mathcal{B})}$ is called perfect if it has no nontrivial grouplike elements.
3. Galois symmetries for hypergroups
Let $(H,\mathcal{B})$ be an abelian rational hypergroup. For any standard element $x_{i}\in\mathcal{B}$, one can write $x_{i}=\sum_{j=1}^{m}{\alpha_{{}_{ij}}}F_{j}$, for some scalars ${\alpha_{{}_{ij}}}\in\mathbb{C}$. Note that ${\alpha_{{}_{ij}}}=\mu_{j}(x_{i})$. Let $\mathbb{K}=\mathbb{Q}({\alpha_{{}_{ij}}})$ be the field obtained by adjoining all ${\alpha_{{}_{ij}}}\in\mathbb{C}$ to $\mathbb{Q}$.
Remark 3.1.
Note that if ${\alpha_{{}_{ij}}}=\alpha_{{}_{ij^{\prime}}}$ for all $i$, then $\mu_{j}(x_{i})=\mu_{j^{\prime}}(x_{i})$, thus $\mu_{j}=\mu_{j^{\prime}}$ and therefore $j=j^{\prime}$.
3.1. Permutation of characters $\mu_{j}$
Let ${(H,\mathcal{B})}$ be any abelian rational hypergroup (i.e. $N^{k}_{ij}\in\mathbb{Q}$). For any character $\mu_{j}:H\rightarrow\mathbb{C}$ and $\sigma\in\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$, define $\sigma.\mu_{j}\in H^{*}$ as the linear function on $H$ on which the standard basis $\mathcal{B}$ is given by $[\sigma.\mu_{j}](x_{i})=\sigma(\mu_{j}(x_{i}))=\sigma({\alpha_{{}_{ij}}}).$
Lemma 3.2.
For any abelian rational hypergroup ${(H,\mathcal{B})}$, it follows that $\sigma.\mu_{j}:H\rightarrow\mathbb{C}$ is an algebra map. Thus $\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$ acts on the set $^{\widehat{\mathcal{B}}}$.
Proof.
Indeed, suppose that $x_{i_{1}}x_{i_{2}}=\sum_{k=1}^{m}N^{k}_{{i_{1}},{i_{2}}}x_{k}$. Since $N^{k}_{{i_{1}},{i_{2}}}\in\mathbb{Q}$,
$[\sigma.\mu_{j}]({x_{i_{1}}}{x_{i_{2}}})=[\sigma.\mu_{j}](\sum_{k=1}^{m}N^{k}_{{i_{1}},{i_{2}}}x_{k})=\sum_{k=1}^{m}N^{k}_{{i_{1}},{i_{2}}}[\sigma.\mu_{j}](x_{k})=\sum_{k=1}^{m}N^{k}_{{i_{1}},{i_{2}}}\sigma(\mu_{j}(x_{k})).$
On the other hand, note that
$[\sigma.\mu_{j}]({x_{i_{1}}})[\sigma.\mu_{j}]({x_{i_{2}}})$ $=$ $\sigma(\mu_{j}({x_{i_{1}}}))\sigma(\mu_{j}({x_{i_{2}}}))=\sigma(\mu_{j}({x_{i_{1}}})\mu_{j}({x_{i_{2}}}))$
$=$ $\sigma(\mu_{j}({x_{i_{1}}}{x_{i_{2}}}))=\sigma(\mu_{j}(\sum_{k=1}^{m}N^{k}_{{i_{1}},{i_{2}}}x_{k}))$
$=$ $\sum_{k=1}^{m}N^{k}_{{i_{1}},{i_{2}}}\sigma(\mu_{j}(x_{k})).$
Thus $[\sigma.\mu_{j}]({x_{i_{1}}}{x_{i_{2}}})=[\sigma.\mu_{j}]({x_{i_{1}}})[\sigma.\mu_{j}]({x_{i_{2}}})$, so that $\sigma.\mu_{j}$ is an algebra map. It is easy to see that, in this case, $\sigma.(\sigma^{\prime}.\mu_{j})=(\sigma\sigma^{\prime}).\mu_{j}.$ It follows that $\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$ acts on the set $^{\widehat{\mathcal{B}}}$ of all algebra homomorphisms $\mu_{j}:H\rightarrow\mathbb{C}$. ∎
We denote by $\mathcal{J}:=\{1,\dots,m\}$ the set of all indices for the characters $\mu_{j}:H\rightarrow\mathbb{C}$. It follows that for any $j\in\mathcal{J}$, there is an index $\tau(j)=\tau_{\sigma}(j)\in\mathcal{J}$ such that $\sigma.\mu_{j}=\mu_{\tau_{\sigma}(j)}$, and therefore
(3.1) $\sigma({\alpha_{{}_{ij}}})=\alpha_{{}_{i\tau_{\sigma}(j)}}\;\;\text{for all}\;i,j.$
It is easy to see that, in this case, $\tau_{\sigma}$ is a permutation of $\mathcal{J}$.
Corollary 3.3.
Let ${(H,\mathcal{B})}$ be an abelian rational hypergroup. Then $\mathbb{Q}\subseteq\mathbb{K}$ is a Galois extension.
Proof.
Note that Equation (3.1) shows that $\sigma(\mathbb{K})\subseteq\mathbb{K}$, for all $\sigma\in\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$. Therefore, this is a normal extension and therefore a Galois extension. ∎
Remark 3.4.
Note that the map $\rho:\mathrm{Gal}(\mathbb{K}/\mathbb{Q})\rightarrow\mathbb{S}_{m},\sigma\mapsto\tau_{\sigma}$ is injective. Indeed, if $\tau_{\sigma}=\tau_{\sigma^{\prime}}$, then $\sigma({\alpha_{{}_{ij}}})=\sigma^{\prime}({\alpha_{{}_{ij}}})$, i.e. $\sigma^{-1}\circ\sigma^{\prime}({\alpha_{{}_{ij}}})={\alpha_{{}_{ij}}}$, for all $i,j$. Thus $\sigma=\sigma^{\prime}$ on $\mathbb{K}$.
Proposition 3.5.
Let $(H,\mathcal{B})$ be an abelian rational hypergroup. For any $\sigma\in\mathrm{Gal}(\mathbb{K}/\mathbb{Q})$, with the above notations,
(3.2) $\sigma(n_{k})={n_{\tau(k)}}.$
Proof.
Applying $\sigma\in\mathrm{Gal}(\mathbb{K}/\mathbb{Q})$ to Equation (2.5), one obtains the following:
$\sum_{i=1}^{m}{h_{i}}\sigma(\alpha_{{}_{il}})\sigma(\alpha_{{}_{i^{*}k}})=\delta_{l,k}\sigma(n_{k}).$
In turn, this equation can be written as
$\sum_{i=1}^{m}{h_{i}}\alpha_{{}_{i\tau(l)}}\alpha_{{}_{i^{*}\tau(k)}}=\delta_{l,k}\sigma(n_{k}).$
On the other hand, by the same orthogonality relation:
$\sum_{i=1}^{m}{h_{i}}\alpha_{{}_{i\tau(l)}}\alpha_{{}_{i^{*}\tau(k)}}=\delta_{\tau(l),\tau(k)}{n_{\tau(k)}}.$
Therefore, for $l=k$, it follows that $\sigma(n_{k})={n_{\tau(k)}}.$
We continue with the following definition.
Definition 3.6.
A hypergroup ${(H,\mathcal{B})}$ is $h$-integral if $h_{i}:=1/N^{1}_{i,i^{*}}\in\mathbb{Z}$, for all $1\leq i\leq m$.
Remark 3.7.
In view of Equation (2.11), the Grothendieck ring of any weakly-integral fusion category $\mathcal{C}$ has a $h$-integral dual ${\widehat{{K(\mathcal{C})}}}$, see [Ost15, Theorem 2.13]. See also Remark 9.1.
Lemma 3.8.
Suppose that ${(H,\mathcal{B},\mu_{1})}$ is an abelian normalizable hypergroup with a $h$-integral dual. Then ${(H,\mathcal{B},\mu_{1})}$ is weakly-integral. Moreover, if ${(H,\mathcal{B},\mu_{1})}$ is rational, then $\widehat{h}_{k}=\widehat{h}_{\tau(k)}$, for all $1\leq k\leq m$.
Proof.
By definition,
$n{(H,\mathcal{B},\mu_{1})}=n{\widehat{{(H,\mathcal{B},\mu_{1})}}}=\sum_{j=1}^{m}{\widehat{h}}_{j}\in\mathbb{Z}.$
If ${(H,\mathcal{B},\mu_{1})}$ is rational, then Equation (3.2) can be written as $\sigma\bigg{(}\frac{n{(H,\mathcal{B},\mu_{1})}}{\widehat{h_{k}}}\bigg{)}=\frac{n{(H,\mathcal{B},\mu_{1})}}{\widehat{h_{\tau(k)}}}.$ In particular, $\widehat{h}_{k}=\widehat{h}_{\tau(k)}.$, since ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is $h$-integral. ∎
4. Burnside’s vanishing property for hypergroups
Let ${(H,\mathcal{B})}$ be any hypergroup. An element $x_{i}\in\mathcal{B}$ is called a vanishing element if there is $\mu_{j}\in{\widehat{\mathcal{B}}}$ such that $\mu_{j}(x_{i})=0$. Otherwise, $x_{i}$ is called a non-vanishing element.
Let $\mathcal{B}_{0}$ be the set of all vanishing elements of $\mathcal{B}$, and $\mathcal{B}_{1}$ be the set of all non-vanishing elements of $\mathcal{B}$. Thus $\mathcal{B}=\mathcal{B}_{1}\sqcup\mathcal{B}_{0}$. Note that if ${(H,\mathcal{B})}$ is a RN-hypergroup then ${G{(H,\mathcal{B})}}\subseteq\mathcal{B}_{1}$. Note that, in the case of an abelian hypergroup ${(H,\mathcal{B})}$, since $x_{i}=\sum_{j=1}^{m}\mu_{j}(x_{i})F_{j}$, it follows that $x_{i}$ is non-vanishing if and only if it is invertible. Thus, in this case, $\mathcal{B}_{1}$ coincides with the set of all standard elements that are invertible in $H$.
Denote ${\widehat{P}}:=\prod_{j=1}^{m}\mu_{j}$. Note that in any abelian normalizable ${(H,\mathcal{B},\mu_{1})}$ one has:
(4.1) $\prod_{j=1}^{m}\mu_{j}=\sum_{i=1}^{m}{\widehat{P}}({\frac{x_{i}}{d_{i}}}){\widehat{E}}_{i}.$
Proposition 4.1.
Let ${(H,\mathcal{B},\mu_{1})}$ be any abelian normalizable hypergroup. Then the following holds:
(4.2) $\prod_{j=1}^{m}\mu_{j}=\sum_{x_{i}\in\mathcal{B}_{1}}\det(L_{{\frac{x_{i}}{d_{i}}}}){\widehat{E}}_{i}$
where $\det(L_{{\frac{x_{i}}{d_{i}}}})$ is the determinant of the left multiplication operator by ${\frac{x_{i}}{d_{i}}}$ on $H$.
Proof.
Recall that $<{\widehat{E}}_{i},\frac{x_{j}}{d_{j}}>=\delta_{i,j}$ for all $i,j\in\mathcal{I}$. In order to prove the above equality, it is enough to evaluate both sides at each ${\frac{x_{i}}{d_{i}}}$, and prove that one obtains the same value.
If $x_{i}$ is a vanishing element, then the above left and right hand side both vanish when evaluated at ${\frac{x_{i}}{d_{i}}}$.
If $x_{i}$ is not a vanishing element then
$\big{[}\prod_{j=1}^{m}\mu_{j}\big{]}({\frac{x_{i}}{d_{i}}})=\prod_{j=1}^{m}\mu_{j}({\frac{x_{i}}{d_{i}}})=\det(L_{{\frac{x_{i}}{d_{i}}}})$
and the proof is compete. ∎
Dually, applying the above result to ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$, one obtains a decomposition of the basis ^=^${\widehat{\mathcal{B}}}={\widehat{\mathcal{B}}}_{0}\sqcup{\widehat{\mathcal{B}}}_{1}$. The above proposition implies the following:
Proposition 4.2.
Let ${(H,\mathcal{B},\mu_{1})}$ be any abelian normalizable hypergroup. Then the following holds:
(4.3) $\prod_{i=1}^{m}{\frac{x_{i}}{d_{i}}}=\sum_{\mu_{j}\in{\widehat{\mathcal{B}}}_{1}}\det(L_{\mu_{j}})F_{j},$
where $\det(L_{\mu_{j}})$ is the determinant of the left multiplication operator by $\mu_{j}$ on $H^{*}$.
Recall the definition of an abelian Burnside hypergroup from Definition 1.4. Note that the decomposition $\mathcal{B}=\mathcal{B}_{0}\sqcup\mathcal{B}_{1}$ from above implies the following:
Proposition 4.3.
An abelian hypergroup ${(H,\mathcal{B})}$ is Burnside if and only if one has ${G{(H,\mathcal{B})}}=\mathcal{B}_{1}$.
Corollary 4.4.
An abelian normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$ is Burnside if and only if the above identity becomes
(4.4) $\prod_{j=1}^{m}\mu_{j}=\sum_{x_{i}\in{G{(H,\mathcal{B})}}}{\mathrm{sgn}}(x_{i}){\widehat{E}}_{i}$
Proof.
It follows from Proposition 4.1. Recall from Definition 2.34 that ${\mathrm{sgn}}(x_{i})$ is the determinant of $L_{\frac{x_{i}}{d_{i}}}$, when $x_{i}\in{G{(H,\mathcal{B})}}$. ∎
Corollary 4.5.
An abelian normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$ is Burnside if and only if
(4.5) $(\prod_{j=1}^{m}\mu_{j})^{2}=\sum_{x_{i}\in{G{(H,\mathcal{B})}}}{\widehat{E}}_{i}.$
Corollary 4.6.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup such that ${G{(H,\mathcal{B})}}$ has odd order. Then ${(H,\mathcal{B})}$ is Burnside if and only if
(4.6) $\prod_{j=1}^{m}\mu_{j}=\sum_{x_{i}\in G(H,B)}{\widehat{E}}_{i}$
Proof.
Note that in this case ${\mathrm{sgn}}(x_{i})=1$ for all $x_{i}\in{G{(H,\mathcal{B})}}$, since each ${\frac{x_{i}}{d_{i}}}\in{G{(H,\mathcal{B})}}$ has odd order. ∎
Definition 4.7.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. We say that the hypergroup ${(H,\mathcal{B})}$ is dual-Burnside if ${\widehat{(H,\mathcal{B})}}$ is Burnside. In other words, for all $\mu_{j}\in{\widehat{\mathcal{B}}}$, the following are equivalent:
  1. (1)
    For all $x_{i}\in{\mathcal{B}}$ then $\mu_{j}(x_{i})\neq 0$,
  2. (2)
    $\mu_{j}\in G\widehat{{(H,\mathcal{B})}}$, i.e $\mu_{j}$ is a grouplike element.
The dual of Corollary 4.4 is the following:
Corollary 4.8.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. Then ${(H,\mathcal{B},\mu_{1})}$ is dual-Burnside if and only if:
(4.7) $\prod_{j=1}^{m}{\frac{x_{i}}{d_{i}}}=\sum_{\mu_{j}\in G{\widehat{(H,\mathcal{B})}}}{\mathrm{sgn}}(\mu_{j})F_{j},$
where ${\mathrm{sgn}}(\mu_{j})$ is the determinant of the left multiplication operator by $\mu_{j}$ on $H^{*}$.
Corollary 4.9.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. Then ${(H,\mathcal{B},\mu_{1})}$ is dual-Burnside if and only if :
(4.8) $(\prod_{j=1}^{m}{\frac{x_{i}}{d_{i}}})^{2}=\sum_{\mu_{j}\in G{\widehat{(H,\mathcal{B})}}}F_{j}.$
where ${\mathrm{sgn}}(\mu_{j})$ is the determinant of the left multiplication operator by $\mu_{j}$ on $H^{*}$.
Corollary 4.10.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. If ${{G{\widehat{(H,\mathcal{B})}}}}$ has odd order then ${(H,\mathcal{B})}$ is dual-Burnside if and only if
(4.9) $\prod_{j=1}^{m}{\frac{x_{i}}{d_{i}}}=\sum_{\mu_{j}\in{{G{\widehat{(H,\mathcal{B})}}}}}F_{j}.$
4.1. Inequalities for $P$ and ${\widehat{P}}$
Proposition 4.11.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup satisfying the hypothesis of Lemma 2.23. Then $|{\widehat{P}}({\frac{x_{i}}{d_{i}}})|\leq 1$. Moreover one has equality if and only if $x_{i}\in G{(H,\mathcal{B})}$.
Proof.
First suppose that ${(H,\mathcal{B})}$ is RN. Then $|\mu_{j}({\frac{x_{i}}{d_{i}}})|\leq 1$ by Frobenius-Perron theory. This implies that $|{\widehat{P}}({\frac{x_{i}}{d_{i}}})|={\prod_{j=1}^{m}}|\mu_{j}({\frac{x_{i}}{d_{i}}})|\leq 1$. Moreover one has equality if and only if $|{\frac{x_{i}}{d_{i}}}(\mu_{j})|=1$ for all $j\in\mathcal{J}$. Then Lemma 2.20 implies that $x_{i}$ is a grouplike element.
Now suppose that ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN. Then since $\mu_{j}=\sum_{i=1}^{m}\mu_{j}({\frac{x_{i}}{d_{i}}}){\widehat{E}}_{i}$ it follows that $|\mu_{j}({\frac{x_{i}}{d_{i}}})|\leq 1$ since $\mathrm{FPdim}(\mu_{j})=1$. This implies again that $|{\widehat{P}}({\frac{x_{i}}{d_{i}}})|\leq 1$. Moreover one has equality if and only if $|{\frac{x_{i}}{d_{i}}}(\mu_{j})|=1$ for all $j\in\mathcal{J}$. Then, again, Lemma 2.20 implies that $x_{i}$ is a grouplike element. ∎
Proposition 4.12.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. One has that ${\widehat{P}}^{2}$ is an idempotent if and only if the non-vanishing elements $x_{i}$ of ${(H,\mathcal{B})}$ satisfy ${\widehat{P}}({\frac{x_{i}}{d_{i}}})=\pm 1$.
Proof.
Note that Equation (4.1) one has ${\widehat{P}}^{4}={\widehat{P}}^{2}$ if and only if ${\widehat{P}}({\frac{x_{i}}{d_{i}}})^{4}={\widehat{P}}({\frac{x_{i}}{d_{i}}})^{2}$ for all $i\in\mathcal{I}$. Therefore if ${\widehat{P}}^{2}$ is idempotent then either ${\widehat{P}}({\frac{x_{i}}{d_{i}}})=0$ or ${\widehat{P}}({\frac{x_{i}}{d_{i}}})^{2}=1$ for all $i\in\mathcal{I}$. In the first case, since ${\widehat{P}}({\frac{x_{i}}{d_{i}}})={\prod_{j=1}^{m}}\mu_{j}({\frac{x_{i}}{d_{i}}})$ it follows that $x_{i}$ is a vanishing element. On the other hand if ${\widehat{P}}({\frac{x_{i}}{d_{i}}})^{2}=1$ implies ${\widehat{P}}({\frac{x_{i}}{d_{i}}})=\pm 1$. ∎
Corollary 4.13.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup satisfying the hypothesis of Lemma 2.23. Then ${\widehat{P}}^{2}$ is idempotent if and only if ${(H,\mathcal{B})}$ is Burnside.
Proof.
In this case, by Proposition 4.11, one has $|\mu_{j}({\frac{x_{i}}{d_{i}}})|\leq 1$ for all $j\in\mathcal{J}$ Thus ${\widehat{P}}({\frac{x_{i}}{d_{i}}})=\pm 1$ implies $|{\widehat{P}}({\frac{x_{i}}{d_{i}}})|=1$ and Lemma 2.20 implies that $x_{i}\in G{(H,\mathcal{B})}$.
Conversely, suppose that ${(H,\mathcal{B})}$ is Burnside. It follows from Equation (4.1) that ${\widehat{P}}=\sum_{x_{i}\in G{(H,\mathcal{B})}}{\widehat{P}}({\frac{x_{i}}{d_{i}}}){\widehat{E}}_{i}$. Note that ${\widehat{P}}({\frac{x_{i}}{d_{i}}})=\det(L_{\frac{x_{i}}{d_{i}}})$ the operator of left multiplication by ${\frac{x_{i}}{d_{i}}}$ on $H$. By Corollary 2.32 this is the determinant of a permutation matrix and therefore equal to ${\mathrm{sgn}}(x_{i})=\pm 1$. ∎
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. Define also $P={\prod_{i=1}^{m}}{\frac{x_{i}}{d_{i}}}$. One has
(4.10) $P=\sum_{j=1}^{m}\mu_{j}(P)F_{j}.$
The dual version of the last three results imply the following:
Proposition 4.14.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup satisfying the hypothesis of Lemma 2.23. Then $|\mu_{j}(P)|\leq 1$ and one has equality if and only if $\mu_{j}\in G{\widehat{(H,\mathcal{B})}}$.
Proposition 4.15.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. One has that $P^{2}$ is an idempotent if and only if the non-vanishing elements $\mu_{j}\in{\widehat{\mathcal{B}}}$ satisfy $\mu_{j}(P)=\pm 1$.
Corollary 4.16.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup satisfying the hypothesis of Lemma 2.23. One has that $P^{2}$ is an idempotent if and only if ${(H,\mathcal{B},\mu_{1})}$ is dual-Burnside.
4.2. The analogue of Burnside’s theorem
We prove the following more general version of Theorem 1.2. Its proof follows the same lines as the proof of [Bur23, Theorem 2].
Theorem 4.17.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian rational normalizable hypergroup such that ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is ${h}$-integral. Let $x_{i}\in\mathcal{B}$ be an element such that $h_{i}|d_{i}|^{2}>1$ and
(4.11) $h_{i}|\mu_{j}(x_{i})|^{2}\in\mathbb{A}$
for all $j\in\mathcal{J}$. Then there is $\mu_{j}\in{\widehat{\mathcal{B}}}$ such that $\mu_{j}(x_{i})=0$.
Proof.
For any $1\leq i\leq m$, denote $\mathcal{T}_{i}:=\mathcal{T}(x_{i})=\{j\in\mathcal{J}\;|\;\mu_{j}(x_{i})=0\}$ and ${\mathcal{D}}_{i}:=\mathcal{J}\setminus(\mathcal{T}_{i}\cup\{1\})$. One needs to show that $\mathcal{T}_{i}\neq\emptyset$, for any basis element $x_{i}\in\mathcal{B}$ satisfying the hypothesis from the theorem.
Note that the second orthogonality relation can be written as:
(4.12) $\sum_{j=1}^{m}\widehat{h}_{j}|\mu_{j}(x_{i})|^{2}=\frac{n{(H,\mathcal{B},\mu_{1})}}{h_{i}}$
Since $\widehat{h}_{1}=1$, for $j=1$, the above equation can be rewritten as:
$\frac{n{(H,\mathcal{B},\mu_{1})}}{h_{i}}=|d_{i}|^{2}+\sum_{j\in{\mathcal{D}}_{i}}{\widehat{h}_{j}}|\mu_{j}(x_{i})|^{2}$
which gives that
(4.13) $1=\frac{n{(H,\mathcal{B},\mu_{1})}}{h_{i}|d_{i}|^{2}}-\sum_{j\in{\mathcal{D}}_{i}}\frac{\widehat{h}_{j}|\mu_{j}(x_{i})|^{2}}{|d_{i}|^{2}}.$
On the other hand, note that
$n{(H,\mathcal{B},\mu_{1})}=\sum_{j=1}^{m}\widehat{h}_{j}=1+\sum_{j\in\mathcal{T}_{i}}\widehat{h}_{j}+\sum_{j\in{\mathcal{D}}_{i}}\widehat{h}_{j}.$
Therefore Equation (4.13) can be written as:
(4.14) $1=\frac{1+\sum_{j\in\mathcal{T}_{i}}\widehat{h}_{j}}{h_{i}|d_{i}|^{2}}-\big{(}\sum_{j\in{\mathcal{D}}_{i}}\frac{\widehat{h}_{j}|\mu_{j}(x_{i})|^{2}}{|d_{i}|^{2}}-\sum_{j\in{\mathcal{D}}_{i}}\frac{\widehat{h}_{j}}{h_{i}|d_{i}|^{2}}\big{)}.$
Thus, in order to finish the proof, it is enough to show that
(4.15) $\big{(}\sum_{j\in{\mathcal{D}}_{i}}\frac{\widehat{h}_{j}|\mu_{j}(x_{i})|^{2}}{|d_{i}|^{2}}-\sum_{j\in{\mathcal{D}}_{i}}\frac{\widehat{h}_{j}}{h_{i}|d_{i}|^{2}}\big{)}\geq 0,$
since then, it follows that $\frac{1+\sum\limits_{j\in\mathcal{T}_{i}}\widehat{h}_{j}}{h_{i}|d_{i}|^{2}}\geq 1$, i.e. $1+\sum\limits_{j\in\mathcal{T}_{i}}\widehat{h}_{j}\geq h_{i}|d_{i}|^{2}$. Since $h_{i}|d_{i}|^{2}>1$, it follows that $\mathcal{T}_{i}\neq\emptyset$.
The inequality from Equation (4.15) can be written as
(4.16) $\frac{1}{\sum\limits_{j\in{\mathcal{D}}_{i}}\widehat{h}_{j}}(\sum_{j\in{\mathcal{D}}_{i}}{\widehat{h}_{j}h_{i}|\mu_{j}(x_{i})|^{2}})\geq 1.$
On the other hand, the weighted AM-GM inequality gives that
(4.17) $\frac{1}{\sum\limits_{j\in{\mathcal{D}}_{i}}\widehat{h}_{j}}(\sum_{j\in{\mathcal{D}}_{i}}{\widehat{h}_{j}h_{i}|\mu_{j}(x_{i})|^{2}})\geq\prod_{j\in{\mathcal{D}}_{i}}\bigg{(}h_{i}\big{(}|\mu_{j}(x_{i})|^{2}\big{)}^{\widehat{h}_{j}}\bigg{)}^{\frac{1}{|{\mathcal{D}}_{i}|}},$
where $|{\mathcal{D}}_{i}|:=\sum_{j\in{\mathcal{D}}_{i}}\widehat{h}_{j}$. Note that Equation (3.1) implies that the set ${\mathcal{D}}_{i}$ is stable under the Galois group $\mathrm{Gal}(\mathbb{K}/\mathbb{Q})$. This in turn implies that the product
$P_{i}:=\prod_{j\in{\mathcal{D}}_{i}}(h_{i}|\mu_{j}(x_{i})|^{2})^{\widehat{h}_{j}}$
is fixed by the Galois group $\mathrm{Gal}(\mathbb{K}/\mathbb{Q})$, since ${\widehat{h}_{\tau(j)}}=\widehat{h}_{j}$ by Lemma 3.8 . It follows that $P_{i}$ is a rational number. On the other hand, each factor of $P_{i}$ is an algebraic integer (since $\widehat{h}_{j}\in\mathbb{Z}_{{}_{>0}}$), and therefore, the entire product is an integer. Since it is also positive, it must be at least $1$. ∎
Note that Theorem 1.2 follows from the above theorem, since in the case of a fusion ring, $h_{i}=1$ and $\mu_{j}(x_{i})\in\mathbb{A}$, for all $(i,j)\in{\mathcal{I}}\times{\mathcal{J}}$.
Proposition 4.18.
Let ${(H,\mathcal{B})}$ be a rational ARN-hypergroup such that $\widehat{(H,\mathcal{B})}$ is ${h}$-integral. Let $x_{i}\in\mathcal{B}$ be an invertible element in $H$ satisfying also Equation (4.11), for all $1\leq j\leq m$. Then $x_{i}\in{G{(H,\mathcal{B})}}$ is a grouplike element.
Proof.
Note that ${(H,\mathcal{B})}$ satisfies the hypothesis of Theorem 4.17. Also, by hypothesis, Equation (4.11) holds for all $1\leq j\leq m$. Suppose that $x_{i}$ is not a grouplike element, i.e. $h_{i}|d_{i}|^{2}>1$, by Lemma 2.17. In this case, by Theorem 4.17, there is $\mu_{j}\in\widehat{B}$ such that $\mu_{j}(x_{i})=0$. Thus, if $x_{i}y=1$ then $10=\mu_{j}(x_{i})\mu_{j}(y)=$ 1, which is a contradiction. ∎
Corollary 4.19.
Any abelian rational normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$ such that ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is ${h}$-integral and Equation (4.11) is satisfied for all $1\leq i,j\leq m$ is Burnside.
Proof.
By Proposition 4.3, it is enough to show that $\mathcal{B}_{1}={G{(H,\mathcal{B})}}$. Clearly, ${G{(H,\mathcal{B})}}\subseteq\mathcal{B}_{1}$. Suppose that $x_{i}\in\mathcal{B}$ is an element which is invertible in $H$. Since $x_{i}$ is not a vanishing element, it follows from Proposition 4.18 that $x_{i}\in{G{(H,\mathcal{B})}}$. ∎
Corollary 4.20.
Let ${(H,\mathcal{B})}$ be a commutative fusion ring such that $\widehat{(H,\mathcal{B})}$ is ${h}$-integral. Then ${(H,\mathcal{B})}$ is Burnside.
The following corollary recovers [Bur23, Theorem 2].
Corollary 4.21.
Any commutative Grothendieck ring of any weakly-integral fusion category is Burnside.
Proof.
It follows from above corollary by Remark 3.7. ∎
Now one can write the corresponding statement of Theorem 4.17 applied to the dual ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ of an abelian normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$. We obtain the following:
Theorem 4.22.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable ${h}$-integral hypergroup such that the dual ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is a rational hypergroup. Let $\mu_{j}\in{\widehat{\mathcal{B}}}$ be an element of ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ with $\widehat{h}_{j}>1$, and such that
(4.18) $\frac{\widehat{h}_{j}}{d_{i}^{2}}|\mu_{j}(x_{i})|^{2}\in\mathbb{A},$
for all $i\in\mathcal{I}$. Then, there is $x_{i}\in\mathcal{B}$ such that $\mu_{j}(x_{i})=0$.
Proof.
Note that the condition from Equation (4.11) applied on ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ becomes exactly Equation (4.18). ∎
5. Kernels of fusion rings, Brauer’s theorem
Let ${(H,\mathcal{B})}$ be an ARN-hypergroup. Let ${(H,\mathcal{B})}_{+}$ be the set of all elements $x\in H$ such that $x=\sum_{b\in{\mathcal{B}}}x_{b}b$ with $x_{b}\in\mathbb{R}_{\geq 0}$. A standard element $b\in\mathcal{B}$ is called a constituent of $x\in{(H,\mathcal{B})}_{+}$ if $x_{b}>0$.
Let $\psi:H\rightarrow\mathbb{C}$ be a linear character, and $E=E_{\psi}\in H$ be its primitive central idempotent. Then, since $E$ is an eigenvector for any $L_{x}$ with $x\in H_{+}$, it follows that $|\psi(x)|\leq\mathrm{FPdim}(x)$, for any object $x\in{(H,\mathcal{B})}_{+}$.
Definition 5.1.
Let $\mathcal{S}\subseteq\mathcal{B}$ be a subset closed under the involution of ${(H,\mathcal{B})}$, and such that $L:=\mathbb{C}[\mathcal{S}]$ is a subalgebra of $H$. Then ${(L,\;\mathcal{S})}$ is called a sub-hypergroup of ${(H,\mathcal{B})}$.
We mention that the notion of kernels of objects of a fusion category from [Bur16] can be extended to arbitrary ARN-hypergroups.
Definition 5.2.
Let ${(H,\mathcal{B})}$ be any ARN-hypergroup. For any $\psi\in{\widehat{\mathcal{B}}}$, define
$\ker_{{\widehat{(H,\mathcal{B})}}}(\psi):=\{x\in\mathcal{B}\;|\;\psi(x)=\mathrm{FPdim}(x)\}.$
Recall that $^{\widehat{\mathcal{B}}}$ is the set of all algebra morphisms $\mu_{j}:H\rightarrow\mathbb{C}$.
Lemma 5.3.
Let ${(H,\mathcal{B})}$ be any ARN-hypergroup and $\psi\in{\widehat{\mathcal{B}}}$. Then $\ker_{{\widehat{(H,\mathcal{B})}}}(\psi)$ is (the basis of) a sub-hypergroup of ${(H,\mathcal{B})}$.
Proof.
Let $x,y\in\ker_{{\widehat{(H,\mathcal{B})}}}(\psi)$, i.e. $\psi(x)=\mathrm{FPdim}(x)$ and $\psi(y)=\mathrm{FPdim}(y)$. Suppose that $xy=\sum_{z\in\mathcal{B}^{\prime}}N^{z}_{xy}z$, where $\mathcal{B}^{\prime}$ is the subset of $\mathcal{B}$ consisting of all standard elements for which $N^{z}_{xy}>0$.
It follows that
$\mathrm{FPdim}(x)\mathrm{FPdim}(y)=\psi(x)\psi(y)=\psi(xy)=\sum_{z\in\mathcal{B}^{\prime}}N^{z}_{xy}\psi(z).$
Using the triangle inequality for the complex absolute-value,
$\mathrm{FPdim}(x)\mathrm{FPdim}(y)$ $=$ $\mid\psi(x)\psi(y)\mid=\mid\psi(xy)\mid=\mid\sum_{z\in\mathcal{B}^{\prime}}N^{z}_{xy}\psi(z)\mid$
$\leq$ $\sum_{z\in\mathcal{B}^{\prime}}N^{z}_{xy}\mid\psi(z)\mid\leq\sum_{z\in\mathcal{B}^{\prime}}N^{z}_{xy}\mathrm{FPdim}(z)$
$=$ $\mathrm{FPdim}(x)\mathrm{FPdim}(y),$
It is easy to see that $\ker_{{\widehat{(H,\mathcal{B})}}}(\psi)$ is a sub-hypergroup of $H$. ∎
Dually, one can define the following:
Definition 5.4.
Let ${(H,\mathcal{B})}$ be an ARN-hypergroup. For any $x\in H_{+}$, one can define ${\ker_{{}_{{(H,\mathcal{B})}}}}(x)$ as the set of all standard elements $\psi\in{\widehat{\mathcal{B}}}$ such that $\psi(x)=\mathrm{FPdim}(x)$.
The dual version of Lemma 5.3 implies that ${\ker_{{}_{{(H,\mathcal{B})}}}}(x)$ is (the basis of) a sub-hypergroup of ${\widehat{(H,\mathcal{B})}}$ if ${\widehat{(H,\mathcal{B})}}$ is a RN-hypergroup.
5.1. Brauer’s theorem for ARN-hypergroup
For any hypergroup ${(H,\mathcal{B})}$, we define the bilinear function on $H$ such that $m(\sum_{i=1}^{m}\alpha_{i}x_{i},\sum_{i=1}^{m}\beta_{i}x_{i})=\sum_{i=1}^{m}\frac{\alpha_{i}\beta_{i}}{h_{i}}$. In the case of an abelian hypergroup note that Equation (2.2) implies the following:
(5.1) $m(x,y)=\tau(xy^{*})=\sum_{j=1}^{m}\frac{1}{n_{j}}\mu_{j}(x)\mu_{j}(y^{*})\;\;\text{for all}\;\;x,y\in H.$
Lemma 5.5.
Let ${(H,\mathcal{B})}$ be an ARN-hypergroup. If $x:=\sum_{s\in\mathcal{B}}p_{s}s\in{(H,\mathcal{B})}_{+}$ then
${\ker_{{}_{{(H,\mathcal{B})}}}}(x)=\bigcap_{{}_{\{s|\;p_{s}>0\}}}{\ker_{{}_{{(H,\mathcal{B})}}}}(s).$
Proof.
Note that $\mathrm{FPdim}(x)=\sum_{\{s\;|\;p_{s}>0\}}p_{s}\mathrm{FPdim}(s)$. If $\mu_{j}\in{\ker_{{}_{{(H,\mathcal{B})}}}}(x)$ then
$\mathrm{FPdim}(x)=|\mu_{j}(x)|\leq\sum_{\{s|\;p_{s}>0\}}|\mu_{j}(s)|\leq\sum_{s}p_{s}\mathrm{FPdim}(s)=\mathrm{FPdim}(x),$
and this implies that $\mu_{j}\in\bigcap_{{}_{\{s|\;p_{s}>0\}}}{\ker_{{}_{{(H,\mathcal{B})}}}}(s)$. The converse is obvious. ∎
Next theorem is a generalization of Brauer’s theorem. Although the proof is the same as the one from [Bur16], we include it here for the completeness of the paper.
For an element $x\in{(H,\mathcal{B})}_{+}$, we denote by $<x>$ the sub-hypergroup generated by $x$. By definition, it is the sub-hypergroup generated by the constituents of all powers $x^{n}$ with $n\geq 1$.
Theorem 5.6.
Let ${(H,\mathcal{B})}$ be an ARN-hypergroup and $x\in{(H,\mathcal{B})}_{+}$. Then $<x>=H$ if and only if ${\ker_{{}_{{(H,\mathcal{B})}}}}(x)=\{\mu_{1}\}$.
Proof.
Suppose that ${\ker_{{}_{{(H,\mathcal{B})}}}}(x)=\mu_{1}$. Write $x=\sum_{j=1}^{m}\mu_{j}(x)F_{j}$ and then
$x^{n}=\sum_{j=1}^{m}\mu_{j}(x)^{n}F_{j}$
Partition the set $\mathcal{J}=\mathcal{J}_{1}\sqcup\mathcal{J}_{2}\sqcup\cdots\sqcup\mathcal{J}_{s}$ such that the value $\mu_{j}(x)$ is constant on each component $\mathcal{J}_{l}$ of the partition. Denote by $\alpha_{l}:=\mu_{j}(x)$ this constant value of $\mu_{j}(x)$ with $j\in\mathcal{J}_{l}$. Note that by the hypothesis, the partition component containing $\{1\}$ is a singleton.
Also note that for any $s\in\mathcal{B}$ and any $a\in H$, then $m(s,a)=\tau(s^{*}a)$. Thus
${m(s,x^{n})}{}=\tau(s^{*}x^{n})=\sum_{j=1}^{m}\mu_{j}(s^{*})\mu_{j}(x)^{n}\tau(F_{j})=\sum_{l\in\mathcal{P}}\big{(}\sum_{j\in\mathcal{J}_{l}}\tau(F_{j})\mu_{j}(s^{*})\big{)}\alpha_{l}^{n}$
Denote
$a_{l}:=\sum_{j\in\mathcal{A}_{l}}\tau(F_{j})\mu_{j}(s^{*}).$
Note that $a_{1}=n{(H,\mathcal{B})}>0$, by Equation (2.9). If $m(s,x^{n})=0$ for all $n$, then
$\sum_{l\in\mathcal{P}}a_{l}\alpha_{l}^{n}=0$
for all $n\in\mathbb{N}$, and Vandermonde determinant gives $a_{l}=1$. This is a contradiction since $a_{1}\neq 0$.
For the converse, suppose that $<x>=H$. If $\mu_{j}\in\ker_{(H,\mathcal{B})}(x)$, then by Lemma 5.3, $\mu_{j}\in\ker(b)$, for any $b\in\mathcal{B}$. Thus $\mu_{j}=\mu_{1}$. ∎
5.2. The cointegral $\lambda_{(L,S)}$ in terms of kernels
Recall from §1 that for any RN-hypergroup ${(H,\mathcal{B})}$, we denoted by $\lambda_{H}$ the primitive central idempotent corresponding to $\mu_{1}=\mathrm{FPdim}$.
Theorem 5.7.
For any $x\in{(H,\mathcal{B})}_{+}$, define $J_{x}:={\ker_{{}_{{(H,\mathcal{B})}}}}(x)$ then
$\lambda_{<x>}=\sum_{j\in J_{x}}F_{j}$
Proof.
Suppose that $<x>={(L,\;\mathcal{S})}$. Let $G_{1},\dots,G_{s}$ be the primitive idempotents of ${(L,\;\mathcal{S})}$, and $\psi_{1},\dots,\psi_{s}$ be their associated characters. Moreover, as above one may suppose that $G_{1}=\lambda_{{(L,\;\mathcal{S})}}$, and consequently, $\psi_{1}=\mathrm{FPdim}$.
We look at the restrictions of the characters $\mu_{j}$ at ${(L,\;\mathcal{S})}$. Then, there is a surjective function $f:\{1,\dots,r\}\rightarrow\{1,\dots,s\}$ such that $\mu_{j}\big{|}_{{(L,\;\mathcal{S})}}=\psi_{f(j)}$, for all $1\leq j\leq s$.
With the above notation, it will be shown that $f^{-1}(0)=\mathcal{J}_{x}$ Indeed, since $x$ is a faithful element of ${(L,\;\mathcal{S})}$, it follows that $\ker_{{}_{(L,\;\mathcal{S})}}(x)=\{\psi_{1}\}$. On the other hand, note that $\mu_{j}(x)=\mu_{j}\big{|}_{{}_{(L,\;\mathcal{S})}}(x)=\psi_{f(j)}(x)$. Thus, $\mu_{j}\in{\ker_{{}_{{(H,\mathcal{B})}}}}(x)$ if and only if $\psi_{f(j)}\in\ker_{{}_{(L,\;\mathcal{S})}}(x)$, i.e. $f(j)=0$. ∎
6. The adjoint sub-hypergroup and grouplike elements
6.1. Notion of the center of a standard element
Let ${(H,\mathcal{B})}$ be an ARN-hypergroup. We also define the notion of center of a standard element $x_{i}\in\mathcal{B}$:
$Z_{{}_{(H,\mathcal{B})}}(x_{i})=\{\psi\in{\widehat{\mathcal{B}}}\;|\;|\psi(x_{i})|=d_{i}\}.$
If ${\widehat{(H,\mathcal{B})}}$ is a RN-hypergroup, then as in Lemma 5.3, it is easy to see that $Z_{{}_{(H,\mathcal{B})}}(x_{i})$ is a sub-hypergroup of ${\widehat{(H,\mathcal{B})}}$.
Define also the dual notion of the center,
$Z_{{}_{\widehat{(H,\mathcal{B})}}}(\psi)=\{x_{i}\in{\mathcal{B}}\;|\;|\psi(x_{i})|=d_{i}\}.$
Again, as in Lemma 5.3, it is easy to see that $Z_{{}_{\widehat{(H,\mathcal{B})}}}(\psi)$ is a sub-hypergroup of ${(H,\mathcal{B})}$ if ${(H,\mathcal{B})}$ is an ARN-hypergroup..
6.2. Adjoint sub-hypergroup
Proposition 6.1.
Let ${(H,\mathcal{B})}$ be an ARN-hypergroup and $I(1):=\sum_{i=1}^{m}h_{i}x_{i}x_{i^{*}}\in H$. Then
${\ker_{{}_{{(H,\mathcal{B})}}}}(I(1))=\bigcap_{i=1}^{m}Z_{{}_{(H,\mathcal{B})}}(x_{i})$
Proof.
Note that
$|\psi(I(1))|\leq\sum_{i=1}^{m}|h_{i}\psi(x_{i}x_{i^{*}})|=\sum_{i=1}^{m}h_{i}|\psi(x_{i})|^{2}\leq\sum_{i=1}^{m}h_{i}d_{i}^{2}=\mathrm{FPdim}{(H,\mathcal{B})}.$
So if $\psi\in{\ker_{{}_{{(H,\mathcal{B})}}}}(I(1))$, then $\psi\in\bigcap_{i=1}^{m}Z_{{}_{(H,\mathcal{B})}}(x_{i})$. The converse is immediate. ∎
Definition 6.2.
Denote by ${{(H,\mathcal{B})}_{ad}}$ the sub-hypergroup $<I(1)>$ of $H$ generated by $I(1)$. It is called the adjoint sub-hypergroup of $H$.
6.3. On the support $J_{{(L,\;\mathcal{S})}}$
Definition 6.3.
Let ${(L,\;\mathcal{S})}$ be a sub-hypergroup of an abelian hypergroup ${(H,\mathcal{B})}$. Its support is the subset $J_{{(L,\;\mathcal{S})}}$ such that
(6.1) ${\lambda_{{}_{{(L,\;\mathcal{S})}}}}=\sum\limits_{j\in{\mathcal{J}_{{}_{{(L,\;\mathcal{S})}}}}}F_{j}.$
The next result is a generalization of [Bur20, Lemma 6.4].
Proposition 6.4.
Let ${(H,\mathcal{B})}$ be an ARN-hypergroup. With the above notations,
${J_{{{(H,\mathcal{B})}_{ad}}}}=\{j\in\mathcal{J}\;|\;n_{j}=\mathrm{FPdim}{(H,\mathcal{B})}\}.$
Proof.
By Theorem 5.7, it is enough to show the following:
${\ker_{{}_{{(H,\mathcal{B})}}}}(I(1))=\{j\in\mathcal{J}\;|\;n_{j}=\mathrm{FPdim}({(H,\mathcal{B})})\}.$
Equation (2.5) gives that $n_{j}=\sum_{i=1}^{m}h_{i}|\mu_{j}(x_{i})|^{2}$, and therefore, $n_{j}=\mathrm{FPdim}{(H,\mathcal{B})}$ if and only if $|\mu_{j}(x_{i})|^{2}=d_{i}^{2}$, for all $i\in\mathcal{I}$. This is equivalent to $\mu_{j}\in Z_{{}_{(H,\mathcal{B})}}(x_{i})$, for all $i\in\mathcal{I}$. The result follows by Proposition 6.1. ∎
Theorem 6.5.
Let ${(H,\mathcal{B})}$ be an abelian RN-hypergroup. Then $\mu_{j}\in G{\widehat{(H,\mathcal{B})}}$ if and only $j\in J_{{{(H,\mathcal{B})}_{ad}}}$.
Proof.
It follows from Proposition 6.4 and Lemma 2.24.
Corollary 6.6.
Let ${(H,\mathcal{B})}$ be an ARN-hypergroup. Then
(6.2) $\lambda_{{}_{{{(H,\mathcal{B})}_{ad}}}}=\sum_{\mu_{j}\in G{\widehat{(H,\mathcal{B})}}}\widehat{F}_{j}$
Proof.
It follows from Equation (6.1) and Theorem 6.5. ∎
There is a dual version of Theorem 6.5 and its Corollary 6.6:
Proposition 6.7.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup such that ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN. Then $x_{i}\in{G{(H,\mathcal{B})}}$ if and only $i\in\mathcal{I}_{{{\widehat{(H,\mathcal{B})}}}_{ad}}$.
Dually one obtains also the following:
Corollary 6.8.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup such that ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN. Then
(6.3) $\lambda_{{}_{{{\widehat{(H,\mathcal{B})}}}_{ad}}}=\sum_{x_{i}\in{G{(H,\mathcal{B})}}}{\widehat{E}}_{i}.$
Proof of Theorem 1.7 and Theorem 1.8.
Proof.
If ${(H,\mathcal{B})}$ is Burnside then by Corollary 4.5,
$\big{(}\prod_{j=1}^{m}\mu_{j}\big{)}^{2}=\sum_{x_{i}\in G(H,B)}{\widehat{E}}_{i}.$
Corollary 6.8 implies Equation (1.4). Conversely, if Equation (1.4) holds then
$\prod_{j=1}^{m}\mu_{j}=\sum_{x_{i}\in{G{(H,\mathcal{B})}}}\pm{\widehat{E}}_{i}$
Evaluating both sides at any ${\frac{x_{i}}{d_{i}}}$, one obtains the Burnside property. ∎
By duality one obtains directly Theorem 1.8.
6.4. On the sub-hypergroup generated by $P$
Proposition 6.9.
For any two sub-hypergroups ${(L,\;\mathcal{S})},{(K,\mathcal{T})}$ of an ARN-hypergroup ${(H,\mathcal{B})}$, then $\mathcal{J}_{{}_{{(L,\;\mathcal{S})}\vee{(K,\mathcal{T})}}}={\mathcal{J}_{{}_{{(K,\mathcal{T})}}}}\cap{\mathcal{J}_{{}_{{(L,\;\mathcal{S})}}}}$.
Proof.
We prove that $\lambda_{{}_{{(L,\;\mathcal{S})}\vee{(K,\mathcal{T})}}}={\lambda_{{}_{{(L,\;\mathcal{S})}}}}{\lambda_{{}_{{(K,\mathcal{T})}}}}$. Then Equation (6.1) implies the intersection result. Note that for all $s\in\mathcal{S}$ and $t\in\mathcal{T}$, $st{\lambda_{{}_{{(L,\;\mathcal{S})}}}}{\lambda_{{}_{{(K,\mathcal{T})}}}}=d_{s}d_{t}{\lambda_{{}_{{(L,\;\mathcal{S})}}}}{\lambda_{{}_{{(K,\mathcal{T})}}}}$. Lemma 5.3 shows that for any constituent $u$ of $st$ then $u{\lambda_{{}_{{(L,\;\mathcal{S})}}}}{\lambda_{{}_{{(K,\mathcal{T})}}}}=d_{u}{\lambda_{{}_{{(L,\;\mathcal{S})}}}}{\lambda_{{}_{{(K,\mathcal{T})}}}}$ also. Thus the same relation holds for any standard element of ${(L,\;\mathcal{S})}\vee{(K,\mathcal{T})}$. ∎
Corollary 6.10.
For any two sub-hypergroups ${(L,\;\mathcal{S})},{(K,\mathcal{T})}$ of an ARN-hypergroup ${(H,\mathcal{B})}$, then ${(L,\;\mathcal{S})}\subseteq{(K,\mathcal{T})}$ if and only if ${\mathcal{J}_{{}_{{(K,\mathcal{T})}}}}\subseteq{\mathcal{J}_{{}_{{(L,\;\mathcal{S})}}}}$.
Proof.
If ${(L,\;\mathcal{S})}\subseteq{(K,\mathcal{T})}$ then ${\lambda_{{}_{{(L,\;\mathcal{S})}}}}{\lambda_{{}_{{(K,\mathcal{T})}}}}={\lambda_{{}_{{(K,\mathcal{T})}}}}$ and therefore equation (6.1) implies ${\mathcal{J}_{{}_{{(K,\mathcal{T})}}}}\subseteq{\mathcal{J}_{{}_{{(L,\;\mathcal{S})}}}}$. The converse follows from Proposition 6.9. ∎
Corollary 6.11.
For any two elements $P,Q\in{(H,\mathcal{B})}_{+}$ of an ARN-hypergroup ${(H,\mathcal{B})}$, then $<P>\subseteq<Q>$ if and only if ${\ker_{{}_{{(H,\mathcal{B})}}}}(Q)\subseteq{\ker_{{}_{{(H,\mathcal{B})}}}}(P)$.
Recall $P:=\prod_{i=1}^{m}{\frac{x_{i}}{d_{i}}}$ for any ${(H,\mathcal{B},\mu_{1})}$.
Proposition 6.12.
In any abelian RN-hypergroup ${(H,\mathcal{B})}$ one has
$<P^{2}>={{(H,\mathcal{B})}_{ad}}.$
Proof.
One can prove first that ${\ker_{{}_{{(H,\mathcal{B})}}}}(P^{2})={\ker_{{}_{{(H,\mathcal{B})}}}}(I(1))$. Indeed $\big{|}\psi(P)\big{|}=\prod_{i=1}^{m}\big{|}\psi({\frac{x_{i}}{d_{i}}})\big{|}\leq 1$. Thus, if $\psi\in{\ker_{{}_{{(H,\mathcal{B})}}}}(P)$, then $\big{|}\psi({\frac{x_{i}}{d_{i}}})\big{|}=1$, i.e $\psi\in{\ker_{{}_{{(H,\mathcal{B})}}}}(I(1))$, by Proposition 6.1. Conversely, if $\psi\in{\ker_{{}_{{(H,\mathcal{B})}}}}(I(1))$ then $\psi({\frac{x_{i}}{d_{i}}}\frac{x_{i^{*}}}{d_{i^{*}}})=\big{|}\psi({\frac{x_{i}}{d_{i}}})\big{|}^{2}=1$. Thus, if $i\neq i^{*}$ then grouping together $x_{i}$ and $x_{i^{*}}$ in $P$, one obtains $\psi({\frac{x_{i}}{d_{i}}})\psi(\frac{x_{{i^{*}}}}{{d_{i^{*}}}})=\big{|}\psi({\frac{x_{i}}{d_{i}}})^{2}\big{|}=1$. Thus,
$\psi(P)=\prod_{i={i^{*}}}\psi({\frac{x_{i}}{d_{i}}}).$
Note that if $i=i^{*}$ then $\psi({\frac{x_{i}}{d_{i}}})=\pm 1$, since in this case $\psi(x_{i})=\psi(x_{i^{*}})\in\mathbb{R}$. Thus, $\psi(P)=\pm 1$ and $\psi(P^{2})=1$ for all $\psi\in{\ker_{{}_{{(H,\mathcal{B})}}}}(I(1))$.
It follows that $\lambda_{<P^{2}>}=\lambda_{<I(1)>}$ by Theorem 5.7, and so $<P^{2}>={{(H,\mathcal{B})}_{ad}}$. ∎
Corollary 6.13.
In any ARN-hypergroup ${(H,\mathcal{B})}$ one has
${{(H,\mathcal{B})}_{ad}}\subseteq<P>.$
Corollary 6.14.
Let ${(H,\mathcal{B})}$ be an abelian RN hypergroup. Then the following holds:
  1. (1)
    ${{(H,\mathcal{B})}_{ad}}=\mathbb{C}$ if and only if ${\widehat{(H,\mathcal{B})}}$ is pointed.
  2. (2)
    ${{(H,\mathcal{B})}_{ad}}={(H,\mathcal{B})}$ if and only if ${\widehat{(H,\mathcal{B})}}$ is perfect.
Remark 6.15.
By §7, ${{(H,\mathcal{B})}_{ad}}={(H,\mathcal{B})}$ if and only if the universal grading group is trivial, if and only if there is no non-trivial grading (see Corollary 7.7).
Corollary 6.16.
A simple non-pointed abelian RN-hypergroup ${(H,\mathcal{B})}$ has a perfect dual.
Proof.
By simplicity ${{(H,\mathcal{B})}_{ad}}={(H,\mathcal{B})}$, and therefore ${\widehat{(H,\mathcal{B})}}$ is perfect. ∎
Corollary 6.17.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup with ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ RN. Then the following holds:
  1. (1)
    ${\widehat{(H,\mathcal{B})}}_{ad}=\mathbb{C}$ if and only if ${(H,\mathcal{B})}$ is pointed,
  2. (2)
    ${\widehat{(H,\mathcal{B})}}_{ad}={\widehat{(H,\mathcal{B})}}$ if and only if ${(H,\mathcal{B})}$ is perfect.
7. On the universal grading of hypergroups
We remark that the universal grading construction for fusion rings from [GN08, §2] works word for word in the ARN-hypergroup settings. For the completeness of the paper, we sketch its main steps below.
7.0.1. Based modules are completely reducible
Definition 7.1.
Given an ARN-hypergroup ${(H,\mathcal{B})}$, we define a based ${(H,\mathcal{B})}$-module as a pair $(M,\mathcal{M})$ where $M$ is a $H$-module with $\mathcal{M}$ a finite free $H$-basis for $M$, and if $x_{i}m_{j}=\sum_{k}t^{k}_{ij}m_{k}$, then $t^{k}_{ij}=t^{k^{*}}_{i^{*}j}$, for all $m_{k}\in\mathcal{M}$ and all $x_{i}\in\mathcal{B}$.
Note that the above condition is equivalent to $(L_{x_{i}^{*}})^{t}=L_{x_{i}}$, where $L_{x_{i}}$ is the matrix of the left multiplication operator by $x_{i}$ on $M$.
Then one can define a inner product on $M$ such that for all $m_{i},m_{j}\in\mathcal{M}$
(7.1) $(m_{i},m_{j})=\delta_{i,j}$
This inner product can be used to prove that any based ${(H,\mathcal{B})}$-modules is completely reducible, as in [Ost03a, Lemma 2.1].
Let ${(H,\mathcal{B})}$ be a RN-hypergroup and $x,y\in{(H,\mathcal{B})}_{+}$. We say that $x$ contains $y$ (or $y$ is a constituent of $x$) if the difference $x-y$ is in ${(H,\mathcal{B})}_{+}$.
Definition 7.2.
Let ${(H,\mathcal{B})}$ be a symmetric RN-hypergroup, $G$ be a finite group. Suppose that $\mathcal{B}=\sqcup_{g\in G}{\mathcal{B}}_{g}$ is a partition of $\mathcal{B}$, and let $H_{g}:=\mathbb{C}[\mathcal{B}_{g}]$. The decomposition ${(H,\mathcal{B})}=\bigoplus_{g\in G}(H_{g},\mathcal{B}_{g})$ is called a grading of ${(H,\mathcal{B})}$ by $G$ if $H_{g}H_{h}\subseteq H_{gh}$, for all $g,h\in G$. The grading is called faithful if $\mathcal{B}_{g}\neq\emptyset$, for all $g\in G$.
Let ${(H,\mathcal{B})}=\bigoplus_{g\in G}(H_{g},\mathcal{B}_{g})$ be a grading of ${(H,\mathcal{B})}$, and define $R_{g}:=\sum_{x_{i}\in\mathcal{B}_{g}}h_{i}d_{i}x_{i}$, for all $g\in G$. The analogue of [ENO05, Proposition 8.20] holds in the settings of ARN-hypergroup:
(7.2) $\mathrm{FPdim}(R_{g})=\frac{\mathrm{FPdim}{(H,\mathcal{B})}}{|G|}.$
Indeed, consider $R:=\sum_{g\in G}R_{g}$, and note that $\frac{1}{\mathrm{FPdim}{(H,\mathcal{B})}}R=F_{1}$, the idempotent corresponding to $\mathrm{FPdim}$ in Corollary 2.8. Thus, $RR_{h}=\mathrm{FPdim}(R_{h})R$. Since $R_{g}R_{h}\subseteq R_{gh}$, one obtains from the previous equality that $R_{g}R_{h}=\mathrm{FPdim}(R_{h})R_{gh}$, for all $g,h\in G$. Passing to $\mathrm{FPdim}$, one get that $\mathrm{FPdim}(R_{gh})=\mathrm{FPdim}(R_{g})$, for all $g,h\in G$.
7.1. Constructing the universal grading
Proposition 7.3.
Let ${(H,\mathcal{B})}$ be a RN-hypergroup. As above, define ${I(1)}:=\sum\limits_{a\in\mathcal{B}}h_{a}aa^{*}$. With the above notations:
  1. (1)
    $I(1)$ is central in $H$.
  2. (2)
    Every based left $H_{ad}$-submodule $M$ of $H$ is also a $H_{ad}$-subbimodule of $H$.
  3. (3)
    A based $H_{ad}$-subbimodule $M$ of $H$ is indecomposable if and only if it is indecomposable as $H_{ad}$-bimodule.
Proof.
  1. (1)
    The proof of this fact is the same as in [GN08].
  2. (2)
    Since $m{I(1)}^{n}={I(1)}^{n}m$, for all $m\in M$, this shows that both the left and the right $H_{ad}$-submodules have the same standard constituents.
  3. (3)
    Clear. ∎
Let $H=\oplus_{a\in A}H_{a}$ be a decomposition of $H$ as a $H_{ad}$-bimodule. This also corresponds to a partition $\mathcal{B}=\sqcup_{a\in A}\mathcal{B}_{a}$. One may suppose that there is an element $1\in A$ such that $H_{1}=H_{ad}$.
7.1.1. Definition of $a^{*}$ for any $a\in A$
Note that if $H_{a}$ is an indecomposable $H_{ad}$-bimodule then, $(H_{a})^{*}$ is also an indecomposable $H_{ad}$-bimodule. Therefore, we may suppose that $(H_{a})^{*}=H_{a^{*}}$, for some $a^{*}\in A$. Next lemma has the same proof as [GN08, Lemma 3.4].
Lemma 7.4.
If $x_{a},y_{a}\in\mathcal{B}_{a}$ then, $x_{a}y_{a}^{*}\in H_{ad}$.
Proof.
Note that
$M:=\{y_{a}\in{\mathcal{B}}_{a}\;|\;m(y_{a},x_{a}{I(1)}^{n})>0,\;\;\text{for some}\;\;n\geq 0\}$
is a $H_{ad}$-submodule of $H_{a}$. Since $H_{a}$ is indecomposable, it follows that $M=H_{a}$. It follows that $m(y_{a}^{*},x_{a}{I(1)}^{n})>0$, which implies that $x_{a}y_{a}^{*}$ has a constituent in $H_{1}$. This gives that $x_{a}y_{a}^{*}$ is contained in $x_{a}x_{a}^{*}{I(1)}^{n}\in R_{1}.$
Theorem 7.5.
There is a well-defined binary operation on $A$ such that $a\star b=c$ if and only if $x_{a}\star y_{b}\in H_{c}$, for all $(x_{a},y_{b})\in\mathcal{B}_{a}\times{\mathcal{B}}_{b}$. It defines a group structure on $A$, with $a^{*}$ the inverse of $a$, and $1$ the unit of $A$.
Proof.
As for the proof of [GN08, Theorem 3.5]. ∎
Definition 7.6.
Above group $A$ is called the universal grading group of ${(H,\mathcal{B})}$. It is denoted $U_{H}$ when the standard basis $\mathcal{B}$ is implicitly understood.
Corollary 7.7.
Every ARN-hypergroup ${(H,\mathcal{B})}$ has a canonical faithful grading by the group $U(H)$. Any other faithful grading of ${(H,\mathcal{B})}$ by a group $G$ is determined by a surjective homomorphism $\pi:U(H)\rightarrow G$.
Proof.
Let ${(H,\mathcal{B})}=\oplus_{g\in G}(S_{g},\mathcal{D}_{g})$ be another grading. It follows that $xx^{*}\in\mathcal{D}_{1}$, for any $x\in\mathcal{D}_{g}$, and therefore, $\mathcal{B}_{ad}\subseteq\mathcal{D}_{1}$. Moreover, each $S_{g}$ is a ${H_{ad}}$-subbimodule of $H$. Since $H_{a}$ is an indecomposable module, then $H_{a}\subseteq S_{\pi(a)}$, for some well-defined $\pi(a)\in G$. It follows that $\pi:U(H)\rightarrow G$ is an epimorphism of groups, since we started with a grading. ∎
7.2. On the dual universal grading group and invertible
In this subsection, we use the notation $H$ for a hypergroup ${(H,\mathcal{B})}$, and $\widehat{H}$ for its dual ${\widehat{(H,\mathcal{B})}}$. The goal of this subsection is to prove the following:
Theorem 7.8.
Let $H$ be an abelian RN-hypergroup. Then,
$U_{H}\simeq{G(\widehat{H})}$
Dually one has the following:
Theorem 7.9.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup such that ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN. Then
$U_{\widehat{H}}\simeq G(H)$
as abelian groups.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup satisfying any of the two hypothesis from Lemma 2.23. Recall that this means that al least one of the following holds:
  1. (1)
    ${(H,\mathcal{B})}$ is RN and $\mu_{1}=\mathrm{FPdim}$
  2. (2)
    ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN.
By Corollary 2.32, for any such hypergroup, if $\mu_{j}\in{G(\widehat{H})}$ then $\mu_{j}^{n}=\mu_{1}$, for some $n\geq 0$. Thus, for all $i$, $\mu_{j}(x_{i})=\omega_{ij}d_{i}$ for some complex number $|\omega_{ij}|=1$.
Definition 7.10.
Let ${(H,\mathcal{B})}$ be a hypergroup satisfying any of the two hypothesis from Lemma 2.23. For any standard element $x_{i}\in\mathcal{B}$, we define a morphism $\omega_{i}:{G(\widehat{H})}\rightarrow\mathbb{C},\mu_{j}\mapsto\omega_{ij}=\mu_{j}({\frac{x_{i}}{d_{i}}})$.
Next lemma shows that $\omega_{i}$ is a character of ${G(\widehat{H})}$.
Lemma 7.11.
Let ${(H,\mathcal{B})}$ be a hypergroup satisfying any of the two hypothesis from Lemma 2.23. For any $x_{i}$, then $\omega_{i}$ is a linear character of ${G(\widehat{H})}$.
Proof.
Since ${\frac{x_{i}}{d_{i}}}:\widehat{H}\rightarrow\mathbb{C}$ are algebra characters, then their restrictions ${\frac{x_{i}}{d_{i}}}\big{|}_{{G(\widehat{H})}}$ are also algebra characters. Note that ${\frac{x_{i}}{d_{i}}}\big{|}_{{G(\widehat{H})}}=\omega_{i}$. ∎
Let ${(H,\mathcal{B})}$ be a satisfying any of the two hypothesis from Lemma 2.23. Given $\psi\in\widehat{{G(\widehat{H})}}$, let
(7.3) $\mathcal{B}_{\psi}:=\{x_{i}\in\mathcal{B}\;\mid\;\omega_{i}=\psi\}=\{x_{i}\;|\;\mu_{j}({\frac{x_{i}}{d_{i}}})=\psi(\mu_{j}),\;\text{for all}\;\mu_{j}\in{G(\widehat{H})}\}$
Note that the above definition can be rewritten as
$x_{i}\in\mathcal{B}_{\psi}\iff{\frac{x_{i}}{d_{i}}}\big{|}_{G(\widehat{H})}=\psi$
Proposition 7.12.
Let ${(H,\mathcal{B})}$ be an abelian RN-hypergroup. Then, Equation (7.3) defines a grading on ${(H,\mathcal{B})}$.
Proof.
It reduces to show that if $x_{i}\in\mathcal{B}_{\psi}$ and $x_{k}\in\mathcal{B}_{\phi}$, then $x_{i}x_{k}$ has all constituents in $\mathcal{B}_{\psi\phi}$. Note that $\mu_{j}({\frac{x_{i}}{d_{i}}})=\psi(\mu_{j})$ and $\mu_{j}(\frac{x_{k}}{d_{k}})=\phi(\mu_{j})$, for all $\mu_{j}\in{G(\widehat{H})}$. It follows that
$\psi\phi(\mu_{j})=\psi(\mu_{j})\phi(\mu_{j})=\mu_{j}({\frac{x_{i}}{d_{i}}})\mu_{j}(\frac{x_{k}}{d_{k}})=\mu_{j}({\frac{x_{i}}{d_{i}}}\frac{x_{k}}{d_{k}}).$
Now, since $|\psi\phi(\mu_{j})|=1$ and ${\frac{x_{i}}{d_{i}}}\frac{x_{k}}{d_{k}}=\frac{1}{d_{i}d_{k}}\sum_{l=1}^{m}d_{l}N^{l}_{ik}\frac{x_{l}}{d_{l}}$, the triangle inequality implies that $\mu_{j}(\frac{x_{l}}{d_{l}})=\mu_{j}(\frac{x_{i}}{d_{i}})\mu_{j}(\frac{x_{k}}{d_{k}})$, for all $l$ with $N^{l}_{ik}>0$. Indeed, note that
$1=|\psi\phi(\mu_{j})|=|\mu_{j}({\frac{x_{i}}{d_{i}}}\frac{x_{k}}{d_{k}})|\leq\frac{1}{d_{i}d_{k}}\sum_{l=1}^{m}d_{l}N^{l}_{ik}|\mu_{j}(\frac{x_{l}}{d_{l}})|\leq\frac{1}{d_{i}d_{k}}\sum_{l=1}^{m}d_{l}N^{l}_{ik}=1,$
and this forces $\mu_{j}(\frac{x_{l}}{d_{l}})=\mu_{j}(\frac{x_{i}}{d_{i}})\mu_{j}(\frac{x_{k}}{d_{k}})$. ∎
7.3. Computations of the regular components
For any grading component $\mathcal{B}_{\psi}$, we define (as above) the regular part of the grading:
$R_{\psi}=\sum_{x_{i}\in\mathcal{B}_{\psi}}h_{i}d_{i}x_{i}.$
Lemma 7.13.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian hypergroup satisfying any of the two hypothesis from Lemma 2.23. With the above notations, for all $\mu_{j}\in G{\widehat{(H,\mathcal{B})}}$,
(7.4) $F_{j}=\frac{1}{n(H)}\big{(}\sum_{\psi\in\widehat{{G(\widehat{H})}}}\psi(\mu_{j})R_{\psi^{-1}}\big{)}.$
Proof.
For any $j$ such that $\mu_{j}\in{G(\widehat{H})}$,
$F_{j}$ $=$ $\frac{1}{n_{j}}\big{(}\sum_{i=1}^{m}h_{i}\mu_{j}(x_{i^{*}})x_{i}\big{)}=\frac{1}{n_{j}}\big{(}\sum_{\psi\in\widehat{{G(\widehat{H})}}}\sum_{x_{i}\in\mathcal{B}_{\psi}}\big{(}h_{i}\mu_{j}(x_{i^{*}})x_{i}\big{)}=$
$=$ $\frac{1}{n_{j}}\big{(}\sum_{\psi\in\widehat{{G(\widehat{H})}}}\psi^{-1}(\mu_{j})R_{\psi}\big{)}=\frac{1}{n(H)}\big{(}\sum_{\psi\in\widehat{{G(\widehat{H})}}}\psi(\mu_{j})R_{\psi^{-1}}\big{)}.$
If $\mu_{j}\in G{\widehat{(H,\mathcal{B})}}$ then $n_{j}=n(H)$ by Corrolary 2.22. ∎
Corollary 7.14.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian hypergroup satisfying any of the two hypothesis from Lemma 2.23. With above notations, for any linear character $\psi$ of ${G(\widehat{H})}$, the set $\mathcal{B}_{\psi}$ is not empty.
Proof.
Previous corollary shows that the linear span of $<F_{j}\;|\;\mu_{j}\in{G(\widehat{H})}>$ lies inside the span of $<\mathcal{B}_{\psi}\;|\;\mathcal{B}_{\psi}\neq\emptyset>$. By counting the dimension of these vector spaces, it follows that $\mathcal{B}_{\psi}$ is not empty, for any $\psi$ in the dual. ∎
From Equation (7.4), it follows immediately:
Lemma 7.15.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian hypergroup satisfying any of the two hypothesis from Lemma 2.23. Then
$R_{\psi}=\frac{1}{|{G(\widehat{H})}|}\big{(}\sum_{\mu_{j}\in{G(\widehat{H})}}\psi(\mu_{j})F_{j}\big{)}.$
Proof of Theorem 7.8
By Lemma 7.15, if $\mu_{j}=\mu_{1}$, then
$R_{1}=\frac{1}{|{G(\widehat{H})}|}\big{(}\sum_{\mu_{j}\in{G(\widehat{H})}}F_{j}\big{)}=\lambda_{{H_{ad}}}.$
Corollary 6.10 proves that $\mathcal{B}_{1}=\mathcal{B}_{ad}$. Therefore the above grading
$\mathcal{B}=\bigsqcup_{\psi\in\widehat{G(\widehat{H})}}{\mathcal{B}}_{\psi}$
coincides with the universal grading of ${(H,\mathcal{B})}$. Thus $U_{H}={G(\widehat{H})}$. By duality, $U_{\widehat{H}}=G(H)$ if ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is abelian.
7.4. On the perp of the adjoint subring
Recall from Definition 1.6 that an abelian RN-hypergroup is called dualizable if ${\widehat{(H,\mathcal{B})}}$ is also RN.
For any sub-hypergroup ${(L,\;\mathcal{S})}$ of an abelian dualizable hypergroup ${(H,\mathcal{B})}$, define
$\mathcal{S}^{\perp}:=\{\mu_{j}\;|\;\mu_{j}(s)=\mathrm{FPdim}(s)\;\text{for all}\;s\in\mathcal{S}\}.$
With the above notations, $\mathcal{S}^{\perp}=\bigcap_{s\in\mathcal{S}}\ker_{(H,\mathcal{B})}(s)$. Moreover, [Har79, Proposition 2.11] implies that $\mathcal{S}^{\perp}$ is a sub-hypergroup of ${\widehat{(H,\mathcal{B})}}$, and $(\mathcal{S}^{\perp})^{\perp}=\mathcal{S}$.
Corollary 7.16.
Let $H$ be an abelian dualizable RN-hypergroup. With the above notations,
${G(\widehat{H})}^{\perp}=H_{ad}\text{ and }G(H)^{\perp}={(\widehat{H})_{ad}}.$
Proof.
Consider the trivial character $\psi_{1}=\mu_{1}\big{|}_{G(\widehat{H})}$. It follows that
${H_{ad}}=B_{\psi_{1}}$ $=$ $\{x_{i}\;|\;\mu_{j}({\frac{x_{i}}{d_{i}}})=1\;\text{for all}\;\mu_{j}\in{G(\widehat{H})}\}$
$=$ $\bigcap_{\mu_{j}\in{G(\widehat{H})}}\ker_{\widehat{H}}(\mu_{j})={G(\widehat{H})}^{\perp}.$
By duality, one obtains the other equality. ∎
7.5. Quotients of hypergroups
In this subsection, we define the quotient hypergroup of a RN-hypergroup ${(H,\mathcal{B})}$ by a sub-hypergroup ${(L,\;\mathcal{S})}$. This is the analogue of the quotient construction from [Har79], in the case of probability groups. Recall that a probability group is a symmetric normalized RN-hypergroup.
One can define the following equivalence relation on $\mathcal{B}$. For two elements $a,b\in\mathcal{B}$, say $a\sim_{\mathcal{S}}b$ if and only if there are $s_{1},s_{2}\in{\mathcal{S}},x\in{\mathcal{B}}$ such that $m(x,as_{1})>0$ and $m(x,s_{2}b)>0$.
Let $[a]_{\mathcal{S}}$ denote the equivalence class of any $a\in\mathcal{B}$, with respect to $\sim_{\mathcal{S}}$. Define $H//{\mathcal{S}}$ as the set of all these equivalence classes of $\sim_{\mathcal{S}}$. Recall that $\lambda_{\mathcal{S}}$ is the primitive central idempotent of ${\mathcal{S}}$ corresponding to $\mu_{1}=\mathrm{FPdim}$.
For an element $a\in\mathcal{B}$, one can see that $[a]_{\mathcal{S}}=[b]_{\mathcal{S}}$ if and only if $\lambda_{\mathcal{S}}a\lambda_{\mathcal{S}}=\lambda_{\mathcal{S}}b\lambda_{\mathcal{S}}$. Therefore there is a set bijection
$\begin{array}[]{cccc}\phi:&\mathbb{C}[H//{\mathcal{S}}]&\to&\lambda_{\mathcal{S}}H\lambda_{\mathcal{S}}\\ &[a]_{\mathcal{S}}&\mapsto&\lambda_{\mathcal{S}}a\lambda_{\mathcal{S}}\end{array}$
Then, it easy to verify that $H//{\mathcal{S}}$ becomes a ARN-hypergroup, with the multiplication inherited from $\lambda_{\mathcal{S}}H\lambda_{\mathcal{S}}$, via the above isomorphism. We denote by $\overline{m}({[c]},\;[a][b])$ the multiplicity structure of $H//{\mathcal{S}}$. Therefore,
$[a][b]=\sum_{[c]\in H//{\mathcal{S}}}\overline{m}({[c]},\;[a][b])[c].$
We write shortly $[a]$ instead of $[a]_{\mathcal{S}}$ when no confusion is possible. Following [Har79], if ${(H,\mathcal{B})}$ is abelian, one can show that
(7.5) $\overline{m}({[c]},\;[a][b])=\sum_{w\in[c]}m(w,ab).$
It was proven in [Har79, Proposition 2.11] that if ${(H,\mathcal{B})}$ is an abelian dualizable probability group then,
$\begin{array}[]{cccc}\alpha:&{\mathcal{S}}^{\perp}&\to&\widehat{H//{\mathcal{S}}}\\ &\psi&\mapsto&\alpha(\psi)\end{array},$
with $\alpha(\psi)([a]_{\mathcal{S}}):=\psi(a)$, is an isomorphism of probability groups. It is clear that the same results remains true for dualizable ARN-hypergroups.
7.6. Applying Harrison’s results
Let ${(H,\mathcal{B})}$ be a dualizable ARN-hypergroup. In this subsection, we also use the short notation $H$ (or $\mathcal{B}$) for this hypergroup, if $\mathcal{B}$ (or $H$) are implicitly understood. For any sub-hypergroup ${(L,S)}$ of an ANR-hypergroup ${(H,\mathcal{B})}$, note that [Har79, Proposition 2.11] implies that there is an isomorphism of hypergroups:
$\widehat{H}//\mathcal{S}^{\perp}\simeq\widehat{\mathcal{S}},\;\;[\mu]_{\mathcal{S}^{\perp}}\mapsto\mu\big{|}_{L}.$
In particular, for $\mathcal{S}=H_{ad}$,
(7.6) $\widehat{H}//G(\widehat{H})\simeq\widehat{H_{ad}}$
Dually, for ${\mathcal{S}}={(\widehat{H})_{ad}}$, one obtains that
(7.7) $H//G(H)\simeq\widehat{{(\widehat{H})_{ad}}}$
For $S=G(H)$, then $\widehat{H}//{(\widehat{H})_{ad}}\simeq\widehat{G(H)}$, and dually, for ${\mathcal{S}}={G(\widehat{H})}$, then $H//H_{ad}\simeq\widehat{{G(\widehat{H})}}.$
8. On lower and upper central series of hypergroups
Let $H$ be a RN-hypergroup. Following [GN08], let $H^{(0)}=H$, $H^{(1)}={H_{ad}}$, and $H^{(n)}=H^{(n-1)}_{ad}$, for all $n\geq 1$.
Definition 8.1.
The non-increasing sequence
$H=H^{(0)}\supseteq H^{(1)}\supseteq\dots\supseteq H^{(n)}\supseteq\dots$
will be called the upper central series.
Definition 8.2.
Let ${(L,\;\mathcal{S})}$ be a sub-hypergroup of an abelian RN-hypergroup ${(H,\mathcal{B})}$. We define the commutator ${(L,\;\mathcal{S})}^{{co}}$ of ${(L,\;\mathcal{S})}$ in ${(H,\mathcal{B})}$ as the sub-hypergroup of ${(H,\mathcal{B})}$ generated by all standard elements $x\in\mathcal{B}$ such that $xx^{*}\in\mathcal{{\mathcal{S}}}$.
As in [GN08, Remark 4.9], it follows that the linear span of ${\mathcal{S}}^{co}$ is already a sub-hypergroup of $H$. Moreover, [GN08, Lemma 4.15] works as well in the settings of ARN-hypergroups. Thus,
(8.1) $(\mathcal{S}^{{co}})_{ad}\subseteq\mathcal{S}\subseteq(\mathcal{S}_{ad})^{{co}}$
for any sub-hypergroup $\mathcal{S}$ of $\mathcal{B}$. By applying $\perp$ to (8.1), one obtains
(8.2) $\big{(}(\mathcal{S}^{{co}})_{ad}\big{)}^{\perp}\supseteq\mathcal{S}^{\perp}\supseteq\big{(}(\mathcal{S}_{ad})^{{co}}\big{)}^{\perp}.$
For an ARN-hypergroup ${(H,\mathcal{B})}$, define $H_{(0)}=\mathbb{C}$ and $H_{(n)}=(H_{(n-1)})^{{co}}$, for all $n\geq 1$. Then, $H_{(1)}=\mathbb{C}^{{co}}=G({(H,\mathcal{B})})=H_{pt}$.
Definition 8.3.
Let $H$ be an ARN-hypergroup. The non-decreasing sequence
$\mathbb{C}=H_{(0)}\subseteq H_{(1)}\subseteq\dots\subseteq H_{(n)}\subseteq\dots$
will be called the lower central series of $H$.
Similarly to [GN08, Theorem 4.16], one can prove that $H^{(n)}=\mathbb{C}$ if and only if $H_{(n)}=H$, where $H$ is an ARN-hypergroup.
Definition 8.4.
An ARN-hypergroup is called nilpotent when $H^{(n)}=\mathbb{C}$, for some $n\in\mathbb{N}$.
Lemma 8.5.
Let ${(H,\mathcal{B})}$ be a dualizable ARN-hypergroup. For any sub-hypergroup ${(L,\;\mathcal{S})}$,
(8.3) $(\mathcal{S}^{\perp})_{ad}\subseteq(\mathcal{S}^{co})^{\perp}.$
Proof.
It reduces to show that $\mu_{j}\mu_{j^{\#}}\in({\mathcal{S}}^{co})^{\perp}$, for all $\mu_{j}\in{\mathcal{S}}^{\perp}$.
Let $x\in{\mathcal{S}}^{co}$ be a standard element. It follows that $xx^{*}\in{\mathcal{S}}$, and therefore, $\mu_{j}(x){\mu_{j^{\#}}}(x)=\mu_{j}(xx^{*})=\mathrm{FPdim}(x)^{2}$, for all $\mu_{j}\in{\mathcal{S}}^{\perp}$. This implies
$[\mu_{j}\star\mu_{j^{\#}}]({\frac{x}{\mathrm{FPdim}(x)}})$ $=$ $\mu_{j}({\frac{x}{\mathrm{FPdim}(x)}})\mu_{j^{\#}}({\frac{x}{\mathrm{FPdim}(x)}})$
$=$ $\frac{1}{\mathrm{FPdim}(x)^{2}}\mu_{j}(x)\mu_{j^{\#}}(x)=1,$
which shows that $\mu_{j}\star{\mu_{j^{\#}}}\in({\mathcal{S}}^{co})^{\perp}$. Thus $({\mathcal{S}}^{\perp})_{ad}\subseteq(S^{co})^{\perp}$. ∎
Proposition 8.6.
Let ${(H,\mathcal{B})}$ be a dualizable ARN-hypergroup. With the above notations, for all $n\geq 0$,
(8.4) $\widehat{H}^{(n)}\subseteq H_{(n)}^{\perp}$
and
(8.5) $H^{(n)}\subseteq(\widehat{H}_{(n)})^{\perp}.$
Proof.
We will prove the first inclusion by induction on $n$. If $n=0$, then
$\widehat{H}^{(0)}=H=\mathbb{C}^{\perp}=H_{(0)}^{\perp}.$
Now, suppose that ${\widehat{H}^{(n)}}\subseteq{H_{(n)}}^{\perp}$. Then
$\widehat{H}^{(n+1)}=({\widehat{H}^{(n)}})_{ad}\subseteq({H_{(n)}}^{\perp})_{ad}\subseteq({H_{(n)}}^{co})^{\perp}=H_{(n+1)}^{\perp}.$
By duality, one obtains the second inclusion. ∎
Observe that the equality holds for $n=1$ also, because
$\widehat{H}^{(1)}={(\widehat{H})_{ad}}=G(H)^{\perp}=H_{(1)}^{\perp}.$
8.1. Proof of Theorem 1.11
Proof.
If ${(H,\mathcal{B})}$ is nilpotent then $H_{(n)}=H$, for some $n$. Then, $\widehat{H}^{(n)}=\mathbb{C}$ by Equation (8.4), and therefore ${\widehat{(H,\mathcal{B})}}$ is nilpotent. Idem for the converse with Equation (8.5).
8.2. Burnside property for nilpotent ARN-hypergroups
Let ${(H,\mathcal{B})}$ be an abelian RN-hypergroup. Recall that we call $x_{i}\in\mathcal{B}$ a vanishing element if there is $\mu_{j}\in{\widehat{\mathcal{B}}}$ such that $\mu_{j}(x_{i})=0$.
Lemma 8.7.
Let ${(H,\mathcal{B})}$ be an abelian RN-hypergroup and ${(L,\;\mathcal{S})}$ be a sub-hypergroup of ${(H,\mathcal{B})}$. Consider a standard element $x_{i}\in{\mathcal{B}}$ such that $[x_{i}]$ is a vanishing element in $H//{\mathcal{S}}$. Then, $x_{i}$ is also vanishing element in ${(H,\mathcal{B})}$.
Proof.
Recall that $H//{\mathcal{S}}\simeq\lambda_{\mathcal{S}}H\lambda_{\mathcal{S}}$ as hypergroups, with $[x]\mapsto\lambda_{\mathcal{S}}x\lambda_{\mathcal{S}}$. Moreover, the hypergroup $\widehat{H//{\mathcal{S}}}$ can be identified with a subring of $\widehat{H}$. If $\lambda_{\mathcal{S}}=\sum_{j\in\mathcal{J}_{\mathcal{S}}}F_{j}$, then $\lambda_{\mathcal{S}}H\lambda_{\mathcal{S}}=\oplus_{j\in\mathcal{J}_{\mathcal{S}}}\mathbb{C}F_{j}$, as a subalgebra of $H$. Thus, the characters $\mu_{j}$, with $j\in\mathcal{J}_{\mathcal{S}}$, have distinct restriction to $\lambda_{\mathcal{S}}H\lambda_{\mathcal{S}}$, and cover all the characters of $\lambda_{\mathcal{S}}H\lambda_{\mathcal{S}}$. So if $[x_{i}]$ is a vanishing element in $H//{\mathcal{S}}$, then $\lambda_{\mathcal{S}}x_{i}\lambda_{\mathcal{S}}$ is a vanishing element in $\lambda_{\mathcal{S}}H\lambda_{\mathcal{S}}$, which in turn implies that $x_{i}$ is a vanishing element in $H$. ∎
Lemma 8.8.
Let ${(H,\mathcal{B})}$ be an ARN-hypergroup. A grouplike element $g\in{G{(H,\mathcal{B})}}$ is a constituent of $x_{i}x_{i^{*}}$ if and only if $gx_{i}=\mathrm{FPdim}(g)x_{i}$. The set of all grouplike elements that are constituents of $x_{i}x_{i^{*}}$ form a subgroup of ${G{(H,\mathcal{B})}}$.
Proof.
Equation (5.1) gives that
$m(g,x_{i}x_{i}^{*})=\tau(x_{i}x_{i^{*}}g^{*})=\tau(x_{i^{*}}g^{*}x_{i})=m(x_{i^{*}},x_{i^{*}}g^{*})=m(x_{i},gx_{i}).$
Thus, by Lemma 2.27 one has $g^{-1}x_{i}=\mathrm{FPdim}(g^{-1})x_{i}$ if $m(g,x_{i}x_{i}^{*})>0$. The second assertion is a straightforward consequence of this. ∎
Proof of Theorem 1.12 It is enough to prove the following:
Proposition 8.9.
A nilpotent dualizable ARN-hypergroup ${(H,\mathcal{B})}$ is Burnside.
Proof.
We will proceed by induction on the nilpotency class of ${(H,\mathcal{B})}$.
If ${(H,\mathcal{B})}$ is of nilpotency class $1$, then ${(H,\mathcal{B})}$ is pointed and we are done.
Now, suppose that ${(H,\mathcal{B})}$ is nilpotent of class $n$ and that the result holds for the nilpotency classes less than $n$. We will show that any standard element $x_{i}\in\mathcal{B}$ is either a vanishing or a grouplike element. Assume that $x_{i}\notin G(H)$. We will show that $x_{i}$ is a vanishing element.
Consider $[x_{i}]\in H//G(H)$. Since $H//G(H)\simeq{(\widehat{H})_{ad}}$, it is nilpotent, with nilpotency class strictly less that of $H$, by Theorem 1.11. By induction hypothesis, we know that $[x_{i}]$ is either a grouplike element or a vanishing element in $H//G(H)$.
If $[x_{i}]$ is a vanishing element, then Lemma 8.7 shows that $x_{i}$ is a vanishing element.
Now, if $[x_{i}]$ is a grouplike element of $H//G(H)$, then
$\lambda_{G(H)}h_{[x_{i}]}x_{i}x_{i}^{*}\lambda_{G(H)}=\lambda_{G(H)},$
which implies that $h_{[x_{i}]}x_{i}x_{i}^{*}\in G(H)$. Thus, $h_{[x_{i}]}x_{i}x_{i}^{*}=\sum_{g\in G_{1}}g$, for a subgroup $G_{1}\subset G$. Since $x_{i}$ is not a grouplike element in $H$, $G_{1}\neq\{1\}$. Then, there is $\mu_{j}$ such that $\mu_{j}(\Lambda_{G_{1}})=0$ where $\Lambda_{G_{1}}=\frac{1}{|G_{1}|}\big{(}\sum_{x\in G_{1}}x\big{)}$ is the integral of $G_{1}$. This implies that $\mu_{j}(x_{i}x_{i}^{*})=0$. Thus $\mu_{j}(x_{i})=0$, which means that $x_{i}$ is a vanishing element. ∎
Now, since $H$ and $\widehat{H}$ are simultaneously nilpotent, Theorem 1.12 follows.
9. Applications to fusion categories
Let $\mathcal{C}$ be a pivotal fusion category with a commutative Grothendieck ring. As in §1, we denote the set of isomorphism classes of simple objects of $\mathcal{C}$ by $\mathrm{Irr}(\mathcal{C}):=\{X_{1},\dots,X_{m}\}$ and let :=$\mathcal{I}:=\{1,\dots,m\}$. Let also $d_{i}:=\dim(X_{i})$ be the quantum dimension of $X_{i}$ for all $i\in\mathcal{I}$. For any simple object $X_{i}\in\mathrm{Irr}(\mathcal{C})$ denote by $x_{i}:=[X_{i}]$ the class of $X_{i}$ in the Grothendieck ring $K(\mathcal{C})$ of $\mathcal{C}$.
Let ${\mathcal{Z}(\mathcal{C})}$ be the Drinfeld center of the fusion category $\mathcal{C}$. The forgetful functor $F:{\mathcal{Z}(\mathcal{C})}\rightarrow\mathcal{C}$ admits a right adjoint functor $R:\mathcal{C}\rightarrow{\mathcal{Z}(\mathcal{C})}$. It is known that $A:=R(\mathbf{1})$ has the structure of a central commutative algebra in $\mathcal{Z}(\mathcal{C})$, where $\mathbf{1}$ is the unit object of $A$. The vector space $\mathrm{{CE}(\mathcal{C})}:=\mathrm{{Hom}}_{\mathcal{C}}(\mathbf{1},A)$ is called the set of central elements. The space ${\mathrm{CF}(\mathcal{C})}:=\mathrm{{Hom}}_{\mathcal{C}}(A,\mathbf{1})$ is called the space of class functions of $\mathcal{C}$. Recall [Shi17] that ${\mathrm{CF}(\mathcal{C})}\simeq K(\mathcal{C})$, the complexification of the Grothendieck ring $K(\mathcal{C})$ of $\mathcal{C}$.
For any simple object $X_{i}$ of $\mathcal{C}$ we denote by $\chi_{i}:=\mathrm{ch}(X_{i})\in{\mathrm{CF}(\mathcal{C})}$ its associated character, [Shi17]. The central element space $\mathrm{{CE}(\mathcal{C})}$ has a basis of primitive orthogonal idempotents $E_{i}$ such that $<\chi_{i},E_{j}>=\delta_{i,j}d_{i}$ for all $1\leq i,j\leq m$.
As explained in [Shi17, Theorem 3.8] the adjunction between $F$ and $R$ gives a canonical isomorphism of algebras
(9.1) ${\mathrm{CF}(\mathcal{C})}\xrightarrow{\cong}\mathrm{End}_{{\mathcal{Z}(\mathcal{C})}}(R(\mathbf{1})),$
where ${\mathrm{CF}(\mathcal{C})}$ is the algebra of class functions on $\mathcal{C}$. Since ${\mathcal{Z}(\mathcal{C})}$ is also fusion category we can write $R(\mathbf{1})=\bigoplus_{j=1}^{m}\mathcal{C}^{j}$ as a direct sum of simple objects in ${\mathcal{Z}(\mathcal{C})}$. Recall that $\mathcal{C}^{j}$ are called conjugacy classes for $\mathcal{C}$. The above isomorphism also gives a canonical bijection between the set of central primitive idempotents $\{F_{j}\}_{j=1}^{m}$ and the set of conjugacy classes $\{\mathcal{C}^{j}\}$. We denote as above $\mathcal{J}:=\{1,\dots m\}$.
For any fusion category $\mathcal{C}$, by abuse of notation, the symmetric RN-hypergroup ${(H,\mathcal{B},\mu_{1})}:=(K(\mathcal{C}),\mathrm{Irr}(\mathcal{C}),\mathrm{FPdim})$ is also denoted by $K(\mathcal{C})$. We denote by ${\widehat{{K(\mathcal{C})}}}$ its dual hypergroup.
Recall also from [GN08] the notion of a universal group grading of a fusion category $\mathcal{C}$ and its adjoint subcategory ${\mathcal{C}_{\mathrm{ad}}}$.
Remark 9.1.
By [Bur20, Equation (4.8)], it follows that for any spherical fusion category, $\dim(\mathcal{C}^{j})=\frac{\dim(\mathcal{C})}{n_{j}}$. See also [Ost15, Theorem 2.13] for a related statement. Then Equation (2.11) implies that
(9.2) $\widehat{h}_{j}=\dim(\mathcal{C}^{j})$
in the dual hypergroup ${\widehat{{K(\mathcal{C})}}}$. In particular, for a weakly-integral fusion category $\mathcal{C}$, by [ENO05, Proposition 8.27], $\dim(\mathcal{C}^{j})\in\mathbb{Z}$, and therefore ${\widehat{{K(\mathcal{C})}}}$ is $h$-integral.
Corollary 9.2.
Let $\mathcal{C}$ be a a fusion category with a Burnside commutative Grothendieck ring and let ${\widehat{{K(\mathcal{C})}}}$ be its dual. Then
(9.3) $\prod_{j=1}^{m}\mu_{j}=\sum_{X_{i}\in{\mathcal{C}_{\mathrm{pt}}}}{\mathrm{sgn}}(x_{i}){\widehat{E}}_{i}$
where ${\mathrm{sgn}}(x_{i})$ is the determinant of the permutation matrix determined by $L_{{\frac{x_{i}}{d_{i}}}}$ on $K(\mathcal{C})$.
Proof.
Since $K(\mathcal{C})$ is a Burnside hypergroup the result follows from Corollary 4.4. ∎
Note that the dual hypergroup ${\widehat{{K(\mathcal{C})}}}$ is denoted by $\widehat{{\mathrm{CF}(\mathcal{C})}}$ in [Bur21]. Theorem 3.4 from the same paper implies that for any pseudo-unitary fusion category there is a canonical isomorphism of hypergroups
(9.4) $\alpha:\widehat{{\mathrm{CF}(\mathcal{C})}}\rightarrow\mathrm{{CE}(\mathcal{C})},\;\mu_{j}\mapsto\frac{C_{j}}{\dim(\mathcal{C}^{j})}.$
For a pivotal fusion category, Shimizu has defined $C_{j}:={\mathcal{F}_{\lambda}}^{-1}(F_{j})\in\mathrm{{CE}(\mathcal{C})}$ to be the conjugacy class sums corresponding to the conjugacy class $\mathcal{C}^{j}$. Here $\lambda\in{\mathrm{CF}(\mathcal{C})}$ is a cointegral of $\mathcal{C}$ such that $\langle\lambda,u\rangle=1$, see [Shi17, §5]. Recall also the Fourier transform of $\mathcal{C}$ associated to $\lambda$ is the linear map
(9.5) $\mathcal{F}_{\lambda}:\mathrm{{CE}(\mathcal{C})}\rightarrow{\mathrm{CF}(\mathcal{C})}\;\;\text{given by}\;\;a\mapsto\lambda\leftharpoonup\mathcal{S}(a)$
Since any weakly-integral fusion category is pseudo-unitary, by applying the canonical morphism $\alpha$ one obtains the following:
Corollary 9.3.
For any weakly-internal fusion category $\mathcal{C}$ the following identity holds in $\mathrm{{CE}(\mathcal{C})}$:
(9.6) $\prod_{j=1}^{m}C_{j}=\sum_{X_{i}\in{\mathcal{C}_{\mathrm{pt}}}}{\mathrm{sgn}}(x_{i})E_{i}$
where $E_{i}\in\mathrm{{CE}(\mathcal{C})}$ is the primitive central idempotent of $X_{i}\in{\mathcal{C}_{\mathrm{pt}}}$.
For a spherical fusion category note the following:
Proposition 9.4.
Let $\mathcal{C}$ be a spherical fusion category with the Grothendieck ring ${K(\mathcal{C})}$ commutative. Let also $\mu_{j}\in{\widehat{{K(\mathcal{C})}}}$ a character of ${K(\mathcal{C})}$. Suppose that ${\widehat{{K(\mathcal{C})}}}$ is a RN-hypergroup.
Then $\mu_{j}$ is a grouplike element of ${\widehat{{K(\mathcal{C})}}}$ if and only if $\dim(\mathcal{C}^{j})=1$.
Proof.
By Lemma 2.17, $\mu_{j}$ is a grouplike element if and only if ${\widehat{h}}_{j}d_{j}^{2}=1$. Since ${\widehat{{K(\mathcal{C})}}}$ is normalized it follows that $d_{j}=1$ for all $j\in\mathcal{J}$. Thus $\mu_{j}$ is a grouplike element if and only if ${\widehat{h}}_{j}=1$. On the other hand, it follows from Equation (9.2) that ${\widehat{h}}_{j}=1$ if and only if $\mathcal{C}^{j}$ is a one dimensional conjugacy class of $\mathcal{C}$. ∎
Suppose now that $\mathcal{C}$ is a pivotal fusion category and ${\mathcal{D}}\subseteq\mathcal{C}$ is a fusion subcategory. There is a subset $\mathcal{J}_{\mathcal{D}}\subseteq\mathcal{J}$ such that
(9.7) $\lambda_{{\mathcal{D}}}=\sum_{j\in\mathcal{J}_{\mathcal{D}}}F_{j}$
since $\lambda_{\mathcal{D}}$ is an idempotent element inside ${\mathrm{CF}(\mathcal{C})}$. Note that this set is denoted by $\mathcal{L}_{\mathcal{D}}$ in [Bur20, §4.2]. Also, this is the corresponding notion of the support from Definition 6.3.
Note that [Bur20, Lemma 4.6 ] implies
(9.8) $J_{{}_{{\mathcal{C}_{\mathrm{ad}}}}}=\{j\in\mathcal{J}\;|\;\dim(\mathcal{C}^{j})=1\},$
for any fusion category with a commutative Grothendieck ring ${K(\mathcal{C})}$.
Proposition 9.5.
Let $\mathcal{C}$ be a fusion category with a commutative Grothendieck ring. Then ${\widehat{{K(\mathcal{C})}}}$ is Burnside if and only if:
(9.9) $\prod_{i=1}^{m}{\frac{x_{i}}{d_{i}}}=\sum_{j\in J_{{\mathcal{C}_{\mathrm{ad}}}}}{\mathrm{sgn}}(\mu_{j})F_{j}$
Proof.
It follows from Corollary 4.8 and Theorem (6.5). ∎
Proof of Theorem 1.9
Proof.
Suppose that the dual ${\widehat{{K(\mathcal{C})}}}$ is Burnside. Squaring Equation (9.9) it follows that
$\big{(}\prod_{i=1}^{m}{\frac{x_{i}}{d_{i}}}\big{)}^{2}=\sum_{j\in J_{{}_{{\mathcal{C}_{\mathrm{ad}}}}}}F_{j}=\lambda_{{\mathcal{C}_{\mathrm{ad}}}}=\frac{1}{\dim({\mathcal{C}_{\mathrm{ad}}})}\big{(}\sum_{x_{i}\in{\mathcal{C}_{\mathrm{ad}}}}d_{i}x_{i}\big{)}.$
The converse also follows immediately from Proposition 9.5. ∎
Let us provide a dual statement in the group case:
Theorem 9.6.
For any finite nilpotent group $G$ the dual $\widehat{K(G)}$ is Burnside and therefore
(9.10) $\left(\prod_{x_{i}\in\mathrm{Irr}(G)}{\frac{x_{i}}{d_{i}}}\right)^{2}=\frac{|Z(G)|}{|G|}\left(\sum_{x_{i}\in\mathrm{Irr}(G/Z(G))}d_{i}x_{i}\right).$
Proof.
Let $\mathcal{C}=\mathrm{Rep}(G)$. It is easy to verify that $\widehat{K(G)}$ is a real-non negative hypergroup. For, example see [Bur21, Equation (3.12)]. Then Theorem B of [INW99] says that if $G$ is a nilpotent group then ${K(G)}$ is dual-Burnside. ∎
Theorem 1.12 implies the following:
Corollary 9.7.
Suppose that $\mathcal{C}$ is a nilpotent fusion category. Then ${K(\mathcal{C})}$ and ${\widehat{{K(\mathcal{C})}}}$ are both Burnside.
10. Premodular categories
Recall that a premodular category is a braided and spherical fusion category. For the rest of this section suppose that $\mathcal{C}$ is a pseudo-unitary premodular category. By [Shi17, Example 6.14] there is $\mathbb{C}$-algebra map $\mathrm{f}_{Q}:{{\mathrm{CF}(\mathcal{C})}}\rightarrow\mathrm{{CE}(\mathcal{C})}$ given by the following formula:
(10.1) $\mathrm{f}_{Q}(\chi_{i})=\sum_{i^{\prime}=1}^{m}\frac{s_{ii^{\prime}}}{d_{i^{\prime}}}E_{i^{\prime}}.$
where $S=(s_{ij})$ is the $S$-matrix of $\mathcal{C}$ and $E_{i^{\prime}}$ are the primitive central idempotents of $\mathrm{{CE}(\mathcal{C})}$ as defined in the previous section.
As in [Bur22, §4] there is a braided function $M:\mathcal{I}\rightarrow\mathcal{J}$ such that if
$\mathrm{f}_{Q}(F_{j})=\sum_{i\in\mathcal{B}_{j}}E_{i}$
[…]
Recall from [BB15] the notion of cosets of a fusion category with respect to a fusion subcategory. By [Bur22, Theorem 4.10] two simple objects $X_{i},X_{i^{\prime}}$ of a pseudo-unitary premodular fusion category $\mathcal{C}$ are in the same coset with respect to $\mathcal{Z}_{2}(\mathcal{C})$ if and only if $M(i)=M(i^{\prime})$. For any $j\in\mathcal{J}_{2}$ we denote by $\mathcal{R}_{j}:=\{[X_{i}]\;|\;M(i)=j\}$. The aforementioned theorem implies also that $\mathcal{R}_{j}$ with $j\in\mathcal{J}_{2}$ are exactly the right cosets of $\mathcal{C}$ with respect to $\mathcal{C}^{\prime}:=\mathcal{Z}_{2}(\mathcal{C})$. Denote also ${\mathrm{R}}_{j}:=\sum_{[X_{i}]\in\mathcal{R}_{j}}d_{i}\chi_{i}\in{\mathrm{CF}(\mathcal{C})}$ the regular part of their characters. By [Bur22, Equation (4.20)],
(10.2) $\dim({\mathrm{R}}_{j})=\dim({\mathcal{C}^{\prime}})\dim(\mathcal{C}^{j}),\;\text{for all}\;j\in\mathcal{J}_{2}.$
Proposition 10.1.
Suppose that $\mathcal{C}$ is a weakly-integral premodular category such that $\mathcal{Z}_{2}(\mathcal{C})\subseteq{\mathcal{C}_{\mathrm{pt}}}$ and it acts freely on the set $\mathrm{Irr}(\mathcal{C})$ of isomorphism classes of simple objects of $\mathcal{C}$. Then for any $j\in\mathcal{J}_{2}$ such that $\dim(\mathcal{C}^{j})>1$ there is an $x_{i}$ such that $\mu_{j}(x_{i})\neq 0$.
Proof.
Note that in ${K(\mathcal{C})}$ one has $h_{i}=1$ for all $i$, and therefore ${K(\mathcal{C})}$ is $h$-integral. If $\mathcal{C}$ is weakly-integral, the fact that the dual ${\widehat{{K(\mathcal{C})}}}$ is rational RN follows from [Bur21, Equation (4.5)]. Thus one may try to apply Theorem 4.22 for ${(H,\mathcal{B},\mu_{1})}={K(\mathcal{C})}$. In order to do be able to do this, one needs to verify that the additional condition (4.18) is satisfied for all $1\leq i\leq m$, i.e. $\dim(\mathcal{C}_{j})\frac{|{\alpha_{{}_{ij}}}|^{2}}{d_{i}^{2}}\in\mathbb{A}$ for all $i\in\mathcal{I}$. In [Bur22, Lemma 4.2] it was shown that
(10.3) $\frac{\alpha_{{}_{iM(i^{\prime})}}}{d_{i}}=\frac{s_{{}_{ii^{\prime}}}}{d_{i}d_{i^{\prime}}}=\frac{\alpha_{{}_{i^{\prime}M(i)}}}{d_{i^{\prime}}}.$
for all $1\leq i,i^{\prime}\leq m$. Fix an index $i_{j}\in\mathcal{I}$ such that $M(i_{j})=j$. It follows that
(10.4) ${\alpha_{{}_{ij}}}=\frac{d_{i}}{d_{i_{j}}}\alpha_{i_{j}M(i)}.$
Therefore
[…]
Since $\mathcal{C}$ is spherical one has $d_{i}=\overline{d_{i}}$ and therefore $|d_{i}|^{2}=d_{i}^{2}$. Thus
(10.6) $\dim(\mathcal{C}_{j})\frac{|{\alpha_{{}_{ij}}}|^{2}}{d_{i}^{2}}=\frac{\dim(\mathcal{C}_{j})}{d_{i_{j}}^{2}}\mid\alpha_{i_{j}M(i)}\mid^{2}$
We will show that $\dim(\mathcal{C}^{j})=d_{i_{j}}^{2}$ and then the proof follows since $\alpha_{i_{j}M(i)}\in\mathbb{A}$. Indeed, applying Equation (10.2), then $\dim(\mathcal{C}^{j})=\frac{\dim(R_{j})}{\dim(\mathcal{C}^{\prime})}=d_{i_{j}}^{2}$ since the action of $\mathcal{C}^{\prime}$ is free on $\mathrm{Irr}(\mathcal{C})$. ∎
10.1. Modular category case
In this section let $\mathcal{C}$ be a weakly-integral modular fusion category. We apply the results of previous section to $K(\mathcal{C})$ and prove Theorem 1.13 and Theorem 1.16. Recall the definition of $\mathcal{J}_{{}_{\mathcal{D}}}$ from Equation (9.7).
Composing Drinfeld’s map $\mathrm{f}_{Q}:{\mathrm{CF}(\mathcal{C})}\rightarrow\mathrm{{CE}(\mathcal{C})}$ with the inverse of the natural isomorphism $\alpha:\mathrm{{CE}(\mathcal{C})}\rightarrow\widehat{{\mathrm{CF}(\mathcal{C})}}$ from Equation (9.4) one obtains an algebra isomorphism ${\widehat{\mathrm{f}_{Q}}}:{\mathrm{CF}(\mathcal{C})}\rightarrow\widehat{{\mathrm{CF}(\mathcal{C})}}$. This can also be written as an algebra isomorphism ${\widehat{\mathrm{f}_{Q}}}:{K(\mathcal{C})}\rightarrow{\widehat{{K(\mathcal{C})}}}$. It was shown in [Bur21, Theorem 4.1] that in fact this is an isomorphism of normalized hypergroup ${\widehat{\mathrm{f}_{Q}}}:\overline{{K(\mathcal{C})}}\rightarrow{\widehat{{K(\mathcal{C})}}}.$ Recall that $\overline{{K(\mathcal{C})}}$ denotes the normalized version of ${K(\mathcal{C})}$.
Remark 10.2.
This isomorphism establishes a canonical bijection between the sets $\mathcal{I}$ and $\mathcal{J}$ which is given by $\mathrm{f}_{Q}(F_{i})=E_{i}$ for all $i\in\mathcal{I}$.
Through the rest of the paper we identify $\mathcal{I}$ and $\mathcal{J}$ via the above bijection.
Corollary 10.3.
Let $\mathcal{C}$ be a modular fusion category. Then $K(\mathcal{C})$ is Burnside if and only if it is dual-Burnside.
Proof.
As mentioned above ${\widehat{\mathrm{f}_{Q}}}:\overline{{K(\mathcal{C})}}\rightarrow{\widehat{{K(\mathcal{C})}}}$ is an isomorphism of normalized hypergroup, where $\overline{{K(\mathcal{C})}}$ denotes the normalized version of ${K(\mathcal{C})}$. Therefore $\overline{{K(\mathcal{C})}}$ (and thus ${K(\mathcal{C})}$) is Burnside if and only if ${\widehat{{K(\mathcal{C})}}}$ is also Burnside.∎
Lemma 10.4.
Let $\mathcal{C}$ be a modular fusion category. Then with the above notations,
(10.7) $X_{i}\in\mathrm{Irr}({\mathcal{C}_{\mathrm{pt}}})\iff i\in{J_{{}_{{\mathcal{C}_{\mathrm{ad}}}}}}$
Proof.
The result follows since $({\mathcal{C}_{\mathrm{ad}}})^{\prime}={\mathcal{C}_{\mathrm{pt}}}$. Then $X_{i}\in\mathrm{Irr}(({\mathcal{C}_{\mathrm{ad}}})^{\prime})$ if and only if $i\in J_{\mathcal{C}_{\mathrm{ad}}}$ by the second item of [Bur20, Theorem 6.2]. ∎
Theorem 10.5.
In any modular fusion category $\mathcal{C}$ with ${K(\mathcal{C})}$ Burnside the following identity holds:
(10.8) $\prod_{i=1}^{m}\frac{x_{i}}{d_{i}}=\sum_{j\in{J_{{}_{{\mathcal{C}_{\mathrm{ad}}}}}}}{\mathrm{sgn}}(x_{j})F_{j}$
Proof.
Since $K(\mathcal{C})$ is Burnside we may apply ${\widehat{\mathrm{f}_{Q}}}^{-1}$ to Equation (9.3). Under the above bijection between $\mathcal{I}$ and $\mathcal{J}$ one has ${\widehat{\mathrm{f}_{Q}}}({\frac{x_{i}}{d_{i}}})=\mu_{i}$, see [Bur20, Theorem 6.1]. Thus the above Equation (9.3) becomes
$\prod_{i=1}^{m}{\frac{x_{i}}{d_{i}}}=\sum_{X_{i}\in{\mathcal{C}_{\mathrm{pt}}}}{\mathrm{sgn}}(x_{i})F_{i}$
and Lemma 10.4 finishes the proof. ∎
Proof of Theorem 1.13
Proof.
Squaring Equation (10.8) one obtains that $\big{(}\prod_{i=1}^{m}\frac{x_{i}}{d_{i}}\big{)}^{2}=\sum_{j\in{J_{{}_{{\mathcal{C}_{\mathrm{ad}}}}}}}F_{j}$. Note that $\lambda_{{\mathcal{C}_{\mathrm{ad}}}}=\sum_{j\in{J_{{}_{{\mathcal{C}_{\mathrm{ad}}}}}}}F_{j}$ by the definition of ${J_{{}_{{\mathcal{C}_{\mathrm{ad}}}}}}$.∎
Note that the above theorem holds for weakly-integral modular categories since their Grothendieck rings are Burnside.
Proof of Corollary 1.14
Proof.
In this case every invertible element $x_{j}$ has odd order since this order divides $|{G(\mathcal{C})}|$. Thus every ${\mathrm{sgn}}(x_{j})=1$ for any invertible object $X_{j}\in{\mathcal{C}_{\mathrm{pt}}}$. ∎
Proposition 10.6.
Let $\mathcal{C}$ be a fusion category such that the dual ${\widehat{{K(\mathcal{C})}}}$ is a Burnside hypergroup. Then
$\frac{({\prod_{i=1}^{m}}d_{i})^{2}}{\mathrm{FPdim}({\mathcal{C}_{\mathrm{ad}}})}\in\mathbb{Z}.$
Proof.
For any category $\mathcal{C}$ define $P_{\mathcal{C}}:=\prod_{i=1}^{m}d_{i}$. Suppose that $\big{(}\prod_{i=1}^{m}x_{i}\big{)}^{2}=\sum_{i=1}^{m}N_{i}x_{i}$ with $N_{i}\in\mathbb{Z}_{\geq 0}$. Then equalizing coefficients in Equation (1.7) one obtains the following $\frac{N_{i}}{P_{\mathcal{C}}^{2}}=\frac{d_{i}}{\mathrm{FPdim}({\mathcal{C}_{\mathrm{ad}}})}$. For $i=1$ it follows that $n_{1}\mathrm{FPdim}({\mathcal{C}_{\mathrm{ad}}})=P_{\mathcal{C}}^{2}$ which proves the divisibility. ∎
Proof of Theorem 1.15.
Proof.
By [GN08, Theorem 3.10] $d_{i}^{2}\in\mathbb{Z}$ for all $i$ since $\mathcal{C}$ is weakly-integral. Note that since ${\widehat{{K(\mathcal{C})}}}$ is Burnside the first divisibility result follows from Proposition 10.6. Moreover, if $\mathcal{C}$ is nilpotent then [GN08, Corollary 5.3] implies $d_{i}^{2}\mid{\mathrm{FPdim}({\mathcal{C}_{\mathrm{ad}}})}$. Thus $\mathcal{V}({\mathcal{C}_{\mathrm{ad}}})=\bigcup_{i=1}^{m}\mathcal{V}(d_{i}^{2})$. ∎
Remark 10.7.
In the case of an integral fusion category Equation (1.10) can be written as
$\mathcal{V}({\mathcal{C}_{\mathrm{ad}}})=\bigcup_{i=1}^{m}\mathcal{V}(d_{i}).$
Proof of Theorem 1.16
Proof.
The first divisibility follows from Proposition (10.6). Suppose now that $\mathcal{C}$ is a weakly-integral modular fusion category. By [ENO05, Proposition 8.27] one has that $\mathrm{FPdim}({\mathcal{C}_{\mathrm{ad}}})$ is an integer. By [GN08, Theorem 3.10] $d_{i}^{2}\in\mathbb{Z}$ for all $i$. The second item follows since for a modular category one has $\mathrm{FPdim}(\mathcal{C})=\mathrm{FPdim}({\mathcal{C}_{\mathrm{ad}}})\mathrm{FPdim}({\mathcal{C}_{\mathrm{pt}}})$. Thus $\mathcal{V}(\mathcal{C})=\mathcal{V}({\mathcal{C}_{\mathrm{ad}}})\cup\mathcal{V}({\mathcal{C}_{\mathrm{pt}}})$. On the other hand, from the first item, $\mathcal{V}({\mathcal{C}_{\mathrm{ad}}})\subseteq\bigcup_{i=1}^{m}\mathcal{V}(d_{i}^{{2}})$ which proves the inclusion $\mathcal{V}(\mathcal{C})\subseteq\mathcal{V}({\mathcal{C}_{\mathrm{pt}}})\cup\big{(}\bigcup_{i=1}^{m}\mathcal{V}(d_{i}^{{2}})\big{)}$.
Conversely, $\mathcal{V}({\mathcal{C}_{\mathrm{pt}}})\cup\big{(}\bigcup_{i=1}^{m}\mathcal{V}(d_{i}^{2})\big{)}\subseteq\mathcal{V}(\mathcal{C})$ since $\mathrm{FPdim}({\mathcal{C}_{\mathrm{pt}}})\mid\mathrm{FPdim}(\mathcal{C})$ and $d_{i}^{{2}}\mid\mathrm{FPdim}(\mathcal{C})$ for all $1\leq i\leq m$. ∎
Proof of Theorem 1.17
Proof.
Since $g.c.d(d_{i}^{2},\;d)=1$ it follows from Equation (1.14) that $d\mid\mathrm{FPdim}({\mathcal{C}_{\mathrm{pt}}})$. Let now ${\mathcal{D}}\subseteq{\mathcal{C}_{\mathrm{pt}}}$ a pointed fusion subcategory such that $\mathrm{FPdim}({\mathcal{D}})=d$. Since $\mathrm{FPdim}({\mathcal{D}}^{\prime})=N$ and $\mathrm{g.c.d}(d,N)=1$ it follows that ${\mathcal{D}}\cap{\mathcal{D}}^{\prime}=\mathrm{Vec}$ and therefore ${\mathcal{D}}$ is non-degenerate. Thus $\mathcal{C}\simeq{\mathcal{D}}\boxtimes{\mathcal{D}}^{\prime}$ and ${\mathcal{D}}^{\prime}$ is non-degenerate as well. ∎
Remark 10.8.
Note that the above Corollary improves [OY23, Theorem 4.5] since the authors also assume the existence of a Tannakian subcategory ${\mathcal{E}}=\mathrm{Rep}(G)\subseteq\mathcal{C}$ such that $\mathcal{C}^{0}_{G}\simeq\mathcal{C}(\mathbb{Z}_{d},q)\boxtimes\mathcal{A}$ for some non-degenerate braided fusion category $\mathcal{A}$.
Proof of Corollary 1.18:
Proof.
Immediate from Theorem 1.17, we just need to check that $g.c.d(d_{i}^{2},\;d)=1$, but it is clear since $d_{i}^{2}$ divides $\mathrm{FPdim}(\mathcal{C})$ and $d$ is a square-free factor. ∎
Lemma 10.9.
With the above notations, for an integral perfect modular category, $\mathcal{V}(\mathcal{C})=\bigcup_{i=1}^{m}\mathcal{V}(d_{i})$ and
$d_{i}^{2}\mid\mathrm{FPdim}(\mathcal{C})\mid P_{\mathcal{C}}^{2}$
Proof.
Since ${\mathcal{C}_{\mathrm{pt}}}$ is trivial then ${\mathcal{C}_{\mathrm{ad}}}=\mathcal{C}$, see Remark 11.13. Thus $\mathcal{V}(\mathcal{C})=\mathcal{V}({\mathcal{C}_{\mathrm{ad}}})=\bigcup_{i=1}^{m}\mathcal{V}(d_{i})$ by Theorem 1.16. Since $d_{i}^{2}\mid\mathrm{FPdim}(\mathcal{C})$ the result follows. ∎
Lemma 10.9 extends trivially to the weakly-integral case because:
Proposition 10.10.
A weakly-integral perfect modular fusion category $\mathcal{C}$ is integral.
Proof.
By the proof of Lemma 10.9 then ${\mathcal{C}_{\mathrm{ad}}}=\mathcal{C}$. But [EGNO15, Exercice 9.6.12] states that the adjoint subcategory of a weakly-integral fusion category is integral. ∎
Corollary 10.11.
Let $\mathcal{C}$ be an integral perfect modular fusion category. If $p$ is a prime number dividing $\mathrm{FPdim}(\mathcal{C})$ then $p^{2}\mid\mathrm{FPdim}(\mathcal{C})$.
Corollary 10.12.
Any integral modular category $\mathcal{C}$ such that there is a prime $p$ whose exponent in $\mathrm{FPdim}(\mathcal{C})$ is $1$ (i.e $m\neq 1$) it is not perfect.
Corollary 10.13.
If $\mathcal{C}$ is a perfect, integral, even dimensional modular fusion category then $4\mid\mathrm{FPdim}(\mathcal{C})$.
10.2. On the Statements (1.21) and (1.22) and Conjecture (1.20)
In [LPR23, Theorem 5.6] it was shown that the statements (1.21) and (1.22) are equivalent. Next we will show that any of these statements implies Conjecture 1.20 formulated in [CP22].
Proposition 10.14.
Existence of a perfect odd dimensional modular fusion category implies the existence of a non-pointed simple integral modular fusion category.
Proof.
Suppose that there exists a perfect modular odd-dimensional fusion category $\mathcal{C}$.
Note first that $\mathcal{C}$ has no pointed subcategories since it is perfect. It also has no symmetric subcategories, since then these subcategories are of the type $\mathrm{Rep}(G)$ with $G$ a perfect group. But then $|G|$ is divisible by $4$ and this is a contradiction since the $\mathrm{FPdim}(\mathcal{C})$ is odd.
On the other hand, by a dimension argument, $\mathcal{C}$ has at least one minimal subcategory under the inclusion (and hence a simple subcategory), say ${\mathcal{D}}$. If ${\mathcal{D}}$ is a modular fusion category then ${\mathcal{D}}$ is an example of a non-pointed simple integral modular fusion category.
If ${\mathcal{D}}$ is not modular then ${\mathcal{D}}^{\prime}\cap{\mathcal{D}}$ is a non-trivial symmetric fusion subcategory which as explained above, it is not possible. ∎
Recall that in [CP22] the authors have also shown that Conjecture 1.20 is equivalent to either of the following:
Conjecture 10.15.
Odd-dimensional fusion categories are solvable.
Conjecture 10.16.
Odd-dimensional modular fusion categories are solvable.
11. Applications and extra results
11.1. Perfect Drinfeld center
We will characterize the perfect fusion categories with a perfect center. That will provide a large class of perfect modular fusion categories.
Lemma 11.1.
Let $\mathcal{C}$ be a fusion category. Let $F:\mathcal{Z}(\mathcal{C})\to\mathcal{C}$ be the forgetful functor. Let $F_{1}:{\mathcal{Z}(\mathcal{C})}_{\mathrm{pt}}\rightarrow{\mathcal{C}_{\mathrm{pt}}}$ be the induced group homomorphism. Let $G$ be the universal grading group of $\mathcal{C}$. Then,
  1. (1)
    $\ker F_{1}$ is isomorphic to $\widehat{G_{ab}}$, with $G_{ab}:=G/G^{\prime}$ the abelianization of $G$,
  2. (2)
    $G$ is perfect if and only if $\ker F_{1}$ is trivial (i.e. $F(Z)=1$ implies $Z=1$).
Proof.
By definition of the center $\mathcal{Z}(\mathcal{C})$, see [EGNO15, Definition 7.13.1], $F(Z)=1$ if and only if $Z=(1,\gamma)$ with $\gamma\in\mathrm{Aut}_{\otimes}(\operatorname{id}_{\mathcal{C}})$. So $\mathrm{Aut}_{\otimes}(\operatorname{id}_{\mathcal{C}})$ is isomorphic to $\ker F_{1}$. Now, [GN08, Proposition 3.9] states that $\mathrm{Aut}_{\otimes}(\operatorname{id}_{\mathcal{C}})$ is isomorphic to $\widehat{G_{ab}}$. In particular, $\ker F_{1}$ is trivial if and only if $G^{\prime}=G$ (i.e. $G$ is perfect). ∎
Note that Lemma 11.1 covers [DNV15, Lemma 2.1].
Proposition 11.2.
Let $\mathcal{C}$ be a perfect fusion category. Then its universal grading group is perfect if and only if $\mathcal{Z}(\mathcal{C})$ is perfect.
Proof.
Let $Z$ be a simple object of $\mathcal{Z}(\mathcal{C})$ with $\mathrm{FPdim}(Z)=1$. The forgetful functor $F:\mathcal{Z}(\mathcal{C})\to\mathcal{C}$ is a tensor functor, so $\mathrm{FPdim}(F(Z))=1$, but $\mathcal{C}$ is perfect, so $F(Z)=1$. The result follows by Lemma 11.1. ∎
Corollary 11.3.
Let $\mathcal{C}$ be a braided fusion category. Then $\mathcal{Z}(\mathcal{C})$ is perfect if and only if $\mathcal{C}$ is perfect with a perfect universal grading group.
Proof.
Immediate by Proposition 11.2 because $\mathcal{C}$ is isomorphic to a fusion subcategory of $\mathcal{Z}(\mathcal{C})$. ∎
Lemma 11.4.
The universal grading group of $\mathrm{Rep}(G)$ is isomorphic to the center $Z(G)$ of the finite group $G$.
Proof.
Immediate by Theorem 7.8, because a grouplike element in the dual hypergroup of the Grothendieck ring of $\mathrm{Rep}(G)$ corresponds to a column of the character table of $G$ with squared norm $|G|$, so to a conjugacy class of size one, thus to a central element. ∎
Corollary 11.5.
Let $G$ be a finite group. Then $\mathcal{Z}(\mathrm{Rep}(G))$ is perfect if and only if $G$ is perfect with a trivial center $Z(G)$.
Proof.
Immediate by Corollary 11.3 and Lemma 11.4, because $\mathrm{Rep}(G)$ is perfect if and only if $G$ is perfect, and (as an abelian group) $Z(G)$ is perfect if and only if it is trivial. ∎
Remark 11.6.
Note that Corollary 11.5 can be proved alternatively using the fact (see [EGNO15, §8.5]) that the simple objects of $\mathcal{Z}(\mathrm{Rep}(G))$ are in bijection with the pairs $(C,V)$ where $C$ is a finite conjugacy class in $G$, and $V$ is an irreducible finite dimensional representation of the centralizer of $g$ in $C$, and the Frobenius-Perron dimension of the object corresponding to a pair $(C,V)$ is $|C|dim_{\mathbb{C}}(V)$.
Corollary 11.7.
Let $G$ be a non-abelian finite simple group. Then $\mathcal{Z}(\mathrm{Rep}(G))$ is perfect.
Proof.
Immediate by Corollary 11.5, because a non-abelian finite simple group is perfect with a trivial center. ∎
By Corollary 11.7, $\mathcal{Z}(\mathrm{Rep}(A_{5}))$ is a perfect integral modular fusion category of $\mathrm{FPdim}$ $360060^{2}=$ 3600. Using Remark 11.6 and GAP, it is of rank $22$ and type
$[[1,1],[3,2],[4,1],[5,1],[12,10],[15,4],[20,3]].$
Question 11.8.
Is there a perfect integral modular fusion category of rank less than $22$? of $\mathrm{FPdim}$ less than $3600$?
Recall that a Hopf algebra $A$ is called perfect if $\mathrm{Rep}(A)$ is perfect, i.e. $G(A^{*})$ is a trivial group; that a finite dimensional semisimple Hopf algebra $A$ (over $\mathbb{C}$) is factorizable if and only if $\mathrm{Rep}(A)$ is modular; that $\mathcal{Z}(\mathrm{Rep}(G))=\mathrm{Rep}(D(G))$ where the Hopf algebra $D(G)$ is the Drinfeld double of $G$. So, for any non-abelian finite simple group $G$ then $D(G)$ is a finite dimensional semisimple factorizable perfect Hopf algebra (over $\mathbb{C}$) of dimension $|G|^{2}$.
Question 11.9.
Is there a finite dimensional semisimple factorizable perfect Hopf algebra (over $\mathbb{C}$) of dimension less than $\dim(D(A_{5}))=3600$?
Lemma 11.10.
Let $A$ be a semisimple factorizable Hopf algebra. If $A^{*}$ is perfect then so is $A$.
Proof.
By [Sch01, Theorem 2.3(b)] there is an isomorphism $G(A^{*})\simeq G(A)\cap Z(A)$. If $G(A)=\{1\}$ then clearly $G(A^{*})=\{1\}$. ∎
[…]
11.2. Burnside and integrality properties
Remark 11.11.
Let $\mathcal{C}$ be a fusion category with a commutative Grothendieck ring $K(\mathcal{C})$. By Theorem 7.8, the dual of $K(\mathcal{C})$ is perfect if and only if its universal grading group is trivial.
Remark 11.12.
The Grothendieck ring of a modular fusion category is perfect if and only if its dual is perfect (as it is isomorphic to its dual).
Remark 11.13.
By Remarks 11.11 and 11.12, a modular fusion category $\mathcal{C}$ is perfect if and only if its universal grading group is trivial (i.e. $\mathcal{C}_{ad}=\mathcal{C}$).
Proof of Theorem 1.10
Proof.
We use the Galois action on $\widehat{\mathcal{B}}$ from Lemma 3.2. Let $\mu_{1}$ be the $\mathrm{FPdim}$ character. By Proposition 3.5, it follows that any Galois conjugate of $\mu_{1}$, say $\mu_{j}=\sigma.\mu_{1}$ has for formal codegree $n_{j}=\sigma(\mathrm{FPdim}(H,{\mathcal{B}}))$. On the other hand, since $\mu_{j}=\sigma.\mu_{1}$ is also a non-vanishing character, it should be a grouplike element, by the definition of a dual-Burnside hypergroup. Thus, its formal codegree should also be $\mathrm{FPdim}(H,{\mathcal{B}})$, which implies that $\sigma(\mathrm{FPdim}(H,{\mathcal{B}}))=\mathrm{FPdim}(H,{\mathcal{B}})$, for all $\sigma$ in the Galois group. Then, we are done since this implies that $\mathrm{FPdim}(H,{\mathcal{B}})$ is a rational number. ∎
The following result is partially due to Andrew Schopieray, see [Sch23].
Corollary 11.14.
Let ${(H,\mathcal{B},\mu_{1})}$ be rational abelian normalizable hypergroup. If the dual ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is perfect and Burnside then ${(H,\mathcal{B},\mu_{1})}$ is integral, i.e $d_{i}\in\mathbb{Q}$.
Proof.
The dual ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is perfect, so as for the above proof of Theorem 1.10, $\sigma.\mu_{1}=\mu_{1}$ for all $\sigma\in\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$. Thus $\sigma(d_{i})=d_{i}$, which shows that $d_{i}\in\mathbb{Q}$. ∎
Corollary 11.15.
A commutative dual-Burnside fusion ring with a trivial universal grading group is integral.
Proof.
It follows from Corollary 11.14 and Remark 11.11. ∎
Recall Remark 6.15 about a trivial universal grading group.
Corollary 11.16.
A fusion category with a dual-Burnside commutative Grothendieck ring and a trivial universal grading group is integral.
Proof.
Immediate from Corollary 11.15. ∎
Proof of Theorem 1.23
Proof.
Recall that a modular fusion category is Burnside if and only if it is dual-Burnside (Corollary 10.3). Now, by Remarks 11.11 and 11.12, the universal grading group is trivial, so by Corollary 11.15, it must be integral. The converse follows from [GNN09, Theorem 6.1]. ∎
Remark 11.17.
By Theorem 1.23, a simple non-integral modular fusion category is not (dual-)Burnside. Plenty of modular fusion categories of Lie type (Verlinde categories) are simple non-integral, so is the one from [Sch22, Theorem 1], and the center of Extended-Haagerup, so none is (dual-)Burnside.
11.3. Near-group modular fusion categories
Let $G$ be a finite abelian group and $m$ be a non-negative integer. Let $K(G,m)$ be the fusion ring with basis $G\cup\{\rho\}$ and fusion rules:
$\rho^{2}=\sum_{g\in G}g+m\rho\text{ and }g\rho=\rho g=\rho,$
for all $g\in G$. A fusion category whose Grothendieck ring is $K(G,m)$ is called near-group (or also Tambara-Yamagami when $m=0$), see for example [Izu17].
Proposition 11.18.
There is no modular fusion category whose Grothendieck ring is $K(G,m)$, with $G$ non-trivial and $m>0$.
Proof.
Let $\mathcal{C}$ be such a category and $R=K(\mathcal{C})$ be its Grothendieck ring. Since $\rho^{2}=m\rho+\sum_{g\in G}g$, then $\mathrm{FPdim}(\rho)^{2}=m\mathrm{FPdim}(\rho)+|G|$. Let $x_{1}>0$ and $x_{2}<0$ be the two solutions of the equation $x^{2}-mx-|G|=0$, then $\mathrm{FPdim}(\rho)=x_{1}$.
Any non-trivial character $\eta\in\widehat{G}$ can be extended to a character of $R$ (denoted also by $\eta$) by declaring $\eta(\rho)=0$. On the other hand, the trivial character $\eta=\epsilon$ ($\epsilon(g)=1$, for all $g\in G$) has two extensions on $R$: $\psi_{+}(\rho)=x_{1}$ and $\psi_{-}(\rho)=x_{2}$.
It follows that $\psi_{+}=\mathrm{FPdim}$. Thus
^=${\widehat{\mathcal{B}}}=\{\eta|\;\eta\in\widehat{G},\;\eta\neq\epsilon\}\cup\{\psi_{\pm}\}$
The multiplication on $\widehat{R}$ is given by
$\eta\star\psi_{-}=\eta,\;\eta\star\eta^{\prime}=\eta\eta^{\prime}\in\widehat{G},\;$
Moreover, it is easy to compute that:
$\psi_{-}^{2}=\frac{x_{2}^{2}+|G|}{x_{1}^{2}+|G|}\psi_{+}+\frac{x_{1}^{2}-x_{2}^{2}}{x_{1}^{2}+|G|}\psi_{-}.$
Thus $R$ has RN-dual since $x_{1}^{2}=mx_{1}+|G|>mx_{2}+|G|=x_{2}^{2}$.
Note that $R$ is a self dual hypergroup since $\mathcal{C}$ is modular. In particular, Drinfeld’s map $\mathrm{f}_{Q}:R\rightarrow\widehat{R}$, since an isomorphism of hypergroups, sends grouplike elements of $R$ to grouplike elements of $\widehat{R}$. Since $G(\widehat{R})=1$ and $G(R)=G$ is non-trivial, this shows that $\mathrm{f}_{Q}$ is not bijective, which is a contradiction. ∎
Proposition 11.19.
Let $G$ be a finite group and $\mathcal{C}$ be a modular category whose Grothendieck ring ${K(\mathcal{C})}$ is $K(G,0)$. Then $G\simeq C_{2}$.
Proof.
As above, if $m=0$ then $\psi_{-}^{2}=\mathrm{FPdim}$ and therefore ${\widehat{{K(\mathcal{C})}}}$ has two grouplike elements $\psi_{+}=\mathrm{FPdim}$ and $\psi_{-}$. Note that Drinfeld’s map $\mathrm{f}_{Q}:{K(\mathcal{C})}\rightarrow{\widehat{{K(\mathcal{C})}}}$ send bijectively grouplike elements of ${K(\mathcal{C})}$ into grouplike elements of ${\widehat{{K(\mathcal{C})}}}$. Since all $g\in G$ are grouplike elements of ${K(\mathcal{C})}$ it follows $|G|=2$. ∎
Remark 11.20.
Ising’s category is the only example of a modular category whose Grothendieck ring is of type $K(C_{2},0)$, see [DGNO07, Appendix B].
Corollary 11.21.
A modular fusion category is near-group if and only if its Grothendieck ring is $K(G,m)$ with $(G,m)=(C_{1},0),(C_{1},1),(C_{2},0)$.
Proof.
Immediate by Propositions 11.18, 11.19, Remark 11.20 and [Ost03b]. ∎
Remark 11.22.
Note that modular (generalized) Tambara-Yamagami categories were classified before, see [Nat13, Lemma 5.3 and Theorem 5.4], [Tho11] and [Sie00].
12. Concrete examples
This section focus on providing some concrete examples of fusion rings/categories to which the main results of the paper apply as categorification criteria, and also some counter-examples showing that some results cannot be generalized.
12.1. Burnside property
Recall that a fusion ring is called Burnside if for every basis element then its fusion matrix has norm $1$ if and only if its determinant is nonzero (in other words, grouplike is equivalent to invertible, see Proposition 4.3). Recall that William Burnside proved that the Grothendieck ring of $\mathrm{Rep}(G)$ is Burnside, for every finite group $G$, and that first author’s [Bur23, Theorem 2] generalizes this result to every weakly-integral fusion category (over $\mathbb{C}$).
A fusion ring $R$ is called $\alpha$-Frobenius if for all basis element $x$ then $\frac{\mathrm{FPdim}(R)^{\alpha}}{\mathrm{FPdim}(x)}$ is an algebraic integer. Kaplansky’s 6th conjecture states that the Grothendieck ring of a complex fusion category is $1$-Frobenius. Under the following bounds:
$\begin{array}[]{c|c|c|c|c|c|c|c|c}\text{rank}(R)&\leq 5&6&7&8&9&10&11&12\\ \hline\cr\mathrm{FPdim}(R)<&1000000&150000&15000&10000&4000&2000&500&200\end{array}$
there are exactly $192$ non-pointed simple integral $1$-Frobenius fusion rings with $\mathrm{FPdim}\neq p^{a}q^{b},pqr$ (the last restriction is motivated by [ENO11]). This is the current state of a classification initiated in [LPW21]. Among them, we found only $4$ non-Burnside ones. They have the type of $\mathrm{Rep}(A_{7})$, but different fusion data.
  • Rank: $9$,
  • FPdim: $2520=2^{3}\times 3^{2}\times 5\times 7$,
  • Type: $[[1,1],[6,1],[10,2],[14,2],[15,1],[21,1],[35,1]]$,
  • Fusion data 1:
    $\begin{smallmatrix}1&0&0&0&0&0&0&0&0\\ 0&1&0&0&0&0&0&0&0\\ 0&0&1&0&0&0&0&0&0\\ 0&0&0&1&0&0&0&0&0\\ 0&0&0&0&1&0&0&0&0\\ 0&0&0&0&0&1&0&0&0\\ 0&0&0&0&0&0&1&0&0\\ 0&0&0&0&0&0&0&1&0\\ 0&0&0&0&0&0&0&0&1\end{smallmatrix},\ \begin{smallmatrix}0&1&0&0&0&0&0&0&0\\ 1&0&0&0&0&1&0&1&0\\ 0&0&0&1&0&0&1&0&1\\ 0&0&1&0&0&0&1&0&1\\ 0&0&0&0&1&2&0&2&0\\ 0&1&0&0&2&0&1&0&1\\ 0&0&1&1&0&1&0&1&1\\ 0&1&0&0&2&0&1&2&1\\ 0&0&1&1&0&1&1&1&4\end{smallmatrix},\ \begin{smallmatrix}0&0&1&0&0&0&0&0&0\\ 0&0&0&1&0&0&1&0&1\\ 0&1&0&1&1&1&0&1&1\\ 1&0&0&0&1&1&1&1&1\\ 0&0&1&1&0&0&1&0&3\\ 0&0&1&1&0&1&1&1&2\\ 0&1&1&0&1&1&1&1&2\\ 0&0&1&1&0&1&1&1&4\\ 0&1&1&1&3&2&2&4&4\end{smallmatrix},\ \begin{smallmatrix}0&0&0&1&0&0&0&0&0\\ 0&0&1&0&0&0&1&0&1\\ 1&0&0&0&1&1&1&1&1\\ 0&1&1&0&1&1&0&1&1\\ 0&0&1&1&0&0&1&0&3\\ 0&0&1&1&0&1&1&1&2\\ 0&1&0&1&1&1&1&1&2\\ 0&0&1&1&0&1&1&1&4\\ 0&1&1&1&3&2&2&4&4\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&1&0&0&0&0\\ 0&0&0&0&1&2&0&2&0\\ 0&0&1&1&0&0&1&0&3\\ 0&0&1&1&0&0&1&0&3\\ 1&1&0&0&3&3&0&5&0\\ 0&2&0&0&3&0&2&2&2\\ 0&0&1&1&0&2&1&2&3\\ 0&2&0&0&5&2&2&4&2\\ 0&0&3&3&0&2&3&2&9\end{smallmatrix},$
    $\begin{smallmatrix}0&0&0&0&0&1&0&0&0\\ 0&1&0&0&2&0&1&0&1\\ 0&0&1&1&0&1&1&1&2\\ 0&0&1&1&0&1&1&1&2\\ 0&2&0&0&3&0&2&2&2\\ 1&0&1&1&0&3&0&3&2\\ 0&1&1&1&2&0&2&1&3\\ 0&0&1&1&2&3&1&4&3\\ 0&1&2&2&2&2&3&3&8\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&0&0&1&0&0\\ 0&0&1&1&0&1&0&1&1\\ 0&1&1&0&1&1&1&1&2\\ 0&1&0&1&1&1&1&1&2\\ 0&0&1&1&0&2&1&2&3\\ 0&1&1&1&2&0&2&1&3\\ 1&0&1&1&1&2&1&2&3\\ 0&1&1&1&2&1&2&2&5\\ 0&1&2&2&3&3&3&5&7\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&0&0&0&1&0\\ 0&1&0&0&2&0&1&2&1\\ 0&0&1&1&0&1&1&1&4\\ 0&0&1&1&0&1&1&1&4\\ 0&2&0&0&5&2&2&4&2\\ 0&0&1&1&2&3&1&4&3\\ 0&1&1&1&2&1&2&2&5\\ 1&2&1&1&4&4&2&6&4\\ 0&1&4&4&2&3&5&4&12\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&0&0&0&0&1\\ 0&0&1&1&0&1&1&1&4\\ 0&1&1&1&3&2&2&4&4\\ 0&1&1&1&3&2&2&4&4\\ 0&0&3&3&0&2&3&2&9\\ 0&1&2&2&2&2&3&3&8\\ 0&1&2&2&3&3&3&5&7\\ 0&1&4&4&2&3&5&4&12\\ 1&4&4&4&9&8&7&12&15\end{smallmatrix}$
  • Fusion data 2:
    $\begin{smallmatrix}1&0&0&0&0&0&0&0&0\\ 0&1&0&0&0&0&0&0&0\\ 0&0&1&0&0&0&0&0&0\\ 0&0&0&1&0&0&0&0&0\\ 0&0&0&0&1&0&0&0&0\\ 0&0&0&0&0&1&0&0&0\\ 0&0&0&0&0&0&1&0&0\\ 0&0&0&0&0&0&0&1&0\\ 0&0&0&0&0&0&0&0&1\end{smallmatrix},\ \begin{smallmatrix}0&1&0&0&0&0&0&0&0\\ 1&0&0&0&0&1&0&1&0\\ 0&0&0&1&0&0&1&0&1\\ 0&0&1&0&0&0&1&0&1\\ 0&0&0&0&1&2&0&2&0\\ 0&1&0&0&2&0&1&0&1\\ 0&0&1&1&0&1&0&1&1\\ 0&1&0&0&2&0&1&2&1\\ 0&0&1&1&0&1&1&1&4\end{smallmatrix},\ \begin{smallmatrix}0&0&1&0&0&0&0&0&0\\ 0&0&0&1&0&0&1&0&1\\ 0&1&1&2&1&1&1&1&0\\ 1&0&1&1&1&1&2&1&0\\ 0&0&1&1&0&0&1&0&3\\ 0&0&1&1&0&1&1&1&2\\ 0&1&2&1&1&1&2&1&1\\ 0&0&1&1&0&1&1&1&4\\ 0&1&0&0&3&2&1&4&5\end{smallmatrix},\ \begin{smallmatrix}0&0&0&1&0&0&0&0&0\\ 0&0&1&0&0&0&1&0&1\\ 1&0&1&1&1&1&2&1&0\\ 0&1&2&1&1&1&1&1&0\\ 0&0&1&1&0&0&1&0&3\\ 0&0&1&1&0&1&1&1&2\\ 0&1&1&2&1&1&2&1&1\\ 0&0&1&1&0&1&1&1&4\\ 0&1&0&0&3&2&1&4&5\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&1&0&0&0&0\\ 0&0&0&0&1&2&0&2&0\\ 0&0&1&1&0&0&1&0&3\\ 0&0&1&1&0&0&1&0&3\\ 1&1&0&0&3&3&0&5&0\\ 0&2&0&0&3&0&2&2&2\\ 0&0&1&1&0&2&1&2&3\\ 0&2&0&0&5&2&2&4&2\\ 0&0&3&3&0&2&3&2&9\end{smallmatrix},$
    $\begin{smallmatrix}0&0&0&0&0&1&0&0&0\\ 0&1&0&0&2&0&1&0&1\\ 0&0&1&1&0&1&1&1&2\\ 0&0&1&1&0&1&1&1&2\\ 0&2&0&0&3&0&2&2&2\\ 1&0&1&1&0&3&0&3&2\\ 0&1&1&1&2&0&2&1&3\\ 0&0&1&1&2&3&1&4&3\\ 0&1&2&2&2&2&3&3&8\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&0&0&1&0&0\\ 0&0&1&1&0&1&0&1&1\\ 0&1&2&1&1&1&2&1&1\\ 0&1&1&2&1&1&2&1&1\\ 0&0&1&1&0&2&1&2&3\\ 0&1&1&1&2&0&2&1&3\\ 1&0&2&2&1&2&2&2&2\\ 0&1&1&1&2&1&2&2&5\\ 0&1&1&1&3&3&2&5&8\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&0&0&0&1&0\\ 0&1&0&0&2&0&1&2&1\\ 0&0&1&1&0&1&1&1&4\\ 0&0&1&1&0&1&1&1&4\\ 0&2&0&0&5&2&2&4&2\\ 0&0&1&1&2&3&1&4&3\\ 0&1&1&1&2&1&2&2&5\\ 1&2&1&1&4&4&2&6&4\\ 0&1&4&4&2&3&5&4&12\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&0&0&0&0&1\\ 0&0&1&1&0&1&1&1&4\\ 0&1&0&0&3&2&1&4&5\\ 0&1&0&0&3&2&1&4&5\\ 0&0&3&3&0&2&3&2&9\\ 0&1&2&2&2&2&3&3&8\\ 0&1&1&1&3&3&2&5&8\\ 0&1&4&4&2&3&5&4&12\\ 1&4&5&5&9&8&8&12&14\end{smallmatrix}$
  • Fusion data 3:
    $\begin{smallmatrix}1&0&0&0&0&0&0&0&0\\ 0&1&0&0&0&0&0&0&0\\ 0&0&1&0&0&0&0&0&0\\ 0&0&0&1&0&0&0&0&0\\ 0&0&0&0&1&0&0&0&0\\ 0&0&0&0&0&1&0&0&0\\ 0&0&0&0&0&0&1&0&0\\ 0&0&0&0&0&0&0&1&0\\ 0&0&0&0&0&0&0&0&1\end{smallmatrix},\ \begin{smallmatrix}0&1&0&0&0&0&0&0&0\\ 1&1&0&0&0&1&1&0&0\\ 0&0&0&1&0&0&1&0&1\\ 0&0&1&0&0&0&1&0&1\\ 0&0&0&0&2&0&0&1&1\\ 0&1&0&0&0&3&1&1&0\\ 0&1&1&1&0&1&1&0&1\\ 0&0&0&0&1&1&0&3&1\\ 0&0&1&1&1&0&1&1&4\end{smallmatrix},\ \begin{smallmatrix}0&0&1&0&0&0&0&0&0\\ 0&0&0&1&0&0&1&0&1\\ 0&1&0&1&0&1&0&0&2\\ 1&0&0&0&0&1&1&0&2\\ 0&0&0&0&1&1&0&2&2\\ 0&0&1&1&1&0&1&1&2\\ 0&1&1&0&0&1&1&0&3\\ 0&0&0&0&2&1&0&3&3\\ 0&1&2&2&2&2&3&3&4\end{smallmatrix},\ \begin{smallmatrix}0&0&0&1&0&0&0&0&0\\ 0&0&1&0&0&0&1&0&1\\ 1&0&0&0&0&1&1&0&2\\ 0&1&1&0&0&1&0&0&2\\ 0&0&0&0&1&1&0&2&2\\ 0&0&1&1&1&0&1&1&2\\ 0&1&0&1&0&1&1&0&3\\ 0&0&0&0&2&1&0&3&3\\ 0&1&2&2&2&2&3&3&4\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&1&0&0&0&0\\ 0&0&0&0&2&0&0&1&1\\ 0&0&0&0&1&1&0&2&2\\ 0&0&0&0&1&1&0&2&2\\ 1&2&1&1&1&1&2&0&3\\ 0&0&1&1&1&0&1&2&3\\ 0&0&0&0&2&1&0&3&3\\ 0&1&2&2&0&2&3&0&5\\ 0&1&2&2&3&3&3&5&6\end{smallmatrix},$
    $\begin{smallmatrix}0&0&0&0&0&1&0&0&0\\ 0&1&0&0&0&3&1&1&0\\ 0&0&1&1&1&0&1&1&2\\ 0&0&1&1&1&0&1&1&2\\ 0&0&1&1&1&0&1&2&3\\ 1&3&0&0&0&6&2&3&0\\ 0&1&1&1&1&2&2&2&2\\ 0&1&1&1&2&3&2&3&3\\ 0&0&2&2&3&0&2&3&9\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&0&0&1&0&0\\ 0&1&1&1&0&1&1&0&1\\ 0&1&1&0&0&1&1&0&3\\ 0&1&0&1&0&1&1&0&3\\ 0&0&0&0&2&1&0&3&3\\ 0&1&1&1&1&2&2&2&2\\ 1&1&1&1&0&2&2&0&4\\ 0&0&0&0&3&2&0&5&4\\ 0&1&3&3&3&2&4&4&7\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&0&0&0&1&0\\ 0&0&0&0&1&1&0&3&1\\ 0&0&0&0&2&1&0&3&3\\ 0&0&0&0&2&1&0&3&3\\ 0&1&2&2&0&2&3&0&5\\ 0&1&1&1&2&3&2&3&3\\ 0&0&0&0&3&2&0&5&4\\ 1&3&3&3&0&3&5&0&7\\ 0&1&3&3&5&3&4&7&10\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&0&0&0&0&1\\ 0&0&1&1&1&0&1&1&4\\ 0&1&2&2&2&2&3&3&4\\ 0&1&2&2&2&2&3&3&4\\ 0&1&2&2&3&3&3&5&6\\ 0&0&2&2&3&0&2&3&9\\ 0&1&3&3&3&2&4&4&7\\ 0&1&3&3&5&3&4&7&10\\ 1&4&4&4&6&9&7&10&17\end{smallmatrix}$
  • Fusion data 4:
    $\begin{smallmatrix}1&0&0&0&0&0&0&0&0\\ 0&1&0&0&0&0&0&0&0\\ 0&0&1&0&0&0&0&0&0\\ 0&0&0&1&0&0&0&0&0\\ 0&0&0&0&1&0&0&0&0\\ 0&0&0&0&0&1&0&0&0\\ 0&0&0&0&0&0&1&0&0\\ 0&0&0&0&0&0&0&1&0\\ 0&0&0&0&0&0&0&0&1\end{smallmatrix},\ \begin{smallmatrix}0&1&0&0&0&0&0&0&0\\ 1&0&0&0&0&1&0&1&0\\ 0&0&0&1&0&0&1&0&1\\ 0&0&1&0&0&0&1&0&1\\ 0&0&0&0&1&2&0&2&0\\ 0&1&0&0&2&0&1&0&1\\ 0&0&1&1&0&1&0&1&1\\ 0&1&0&0&2&0&1&2&1\\ 0&0&1&1&0&1&1&1&4\end{smallmatrix},\ \begin{smallmatrix}0&0&1&0&0&0&0&0&0\\ 0&0&0&1&0&0&1&0&1\\ 1&0&0&2&1&1&2&1&0\\ 0&1&2&1&1&1&1&1&0\\ 0&0&1&1&0&0&1&0&3\\ 0&0&1&1&0&1&1&1&2\\ 0&1&2&1&1&1&2&1&1\\ 0&0&1&1&0&1&1&1&4\\ 0&1&0&0&3&2&1&4&5\end{smallmatrix},\ \begin{smallmatrix}0&0&0&1&0&0&0&0&0\\ 0&0&1&0&0&0&1&0&1\\ 0&1&2&1&1&1&1&1&0\\ 1&0&1&1&1&1&2&1&0\\ 0&0&1&1&0&0&1&0&3\\ 0&0&1&1&0&1&1&1&2\\ 0&1&1&2&1&1&2&1&1\\ 0&0&1&1&0&1&1&1&4\\ 0&1&0&0&3&2&1&4&5\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&1&0&0&0&0\\ 0&0&0&0&1&2&0&2&0\\ 0&0&1&1&0&0&1&0&3\\ 0&0&1&1&0&0&1&0&3\\ 1&1&0&0&3&3&0&5&0\\ 0&2&0&0&3&0&2&2&2\\ 0&0&1&1&0&2&1&2&3\\ 0&2&0&0&5&2&2&4&2\\ 0&0&3&3&0&2&3&2&9\end{smallmatrix},$
    $\begin{smallmatrix}0&0&0&0&0&1&0&0&0\\ 0&1&0&0&2&0&1&0&1\\ 0&0&1&1&0&1&1&1&2\\ 0&0&1&1&0&1&1&1&2\\ 0&2&0&0&3&0&2&2&2\\ 1&0&1&1&0&3&0&3&2\\ 0&1&1&1&2&0&2&1&3\\ 0&0&1&1&2&3&1&4&3\\ 0&1&2&2&2&2&3&3&8\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&0&0&1&0&0\\ 0&0&1&1&0&1&0&1&1\\ 0&1&2&1&1&1&2&1&1\\ 0&1&1&2&1&1&2&1&1\\ 0&0&1&1&0&2&1&2&3\\ 0&1&1&1&2&0&2&1&3\\ 1&0&2&2&1&2&2&2&2\\ 0&1&1&1&2&1&2&2&5\\ 0&1&1&1&3&3&2&5&8\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&0&0&0&1&0\\ 0&1&0&0&2&0&1&2&1\\ 0&0&1&1&0&1&1&1&4\\ 0&0&1&1&0&1&1&1&4\\ 0&2&0&0&5&2&2&4&2\\ 0&0&1&1&2&3&1&4&3\\ 0&1&1&1&2&1&2&2&5\\ 1&2&1&1&4&4&2&6&4\\ 0&1&4&4&2&3&5&4&12\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&0&0&0&0&1\\ 0&0&1&1&0&1&1&1&4\\ 0&1&0&0&3&2&1&4&5\\ 0&1&0&0&3&2&1&4&5\\ 0&0&3&3&0&2&3&2&9\\ 0&1&2&2&2&2&3&3&8\\ 0&1&1&1&3&3&2&5&8\\ 0&1&4&4&2&3&5&4&12\\ 1&4&5&5&9&8&8&12&14\end{smallmatrix}$
The determinant of the second fusion matrix (of each fusion data above) is ±$\pm 36$ whereas its norm is $6$, proving that these integral fusion ring are non-Burnside, and so cannot be categorified.
Similar classification were done in the non-1-Frobenius case, and unexpectedly, most (not all) of the ones we found are non-Burnside (see one example below), so without categorification, which can be seen as an evidence for Kaplansky 6th conjecture.
  • Rank: $7$,
  • FPdim: $798=2\times 3\times 7\times 19$,
  • Type: $[[1,1],[7,1],[8,1],[9,3],[21,1]]$,
  • Fusion data:
    $\begin{smallmatrix}1&0&0&0&0&0&0\\ 0&1&0&0&0&0&0\\ 0&0&1&0&0&0&0\\ 0&0&0&1&0&0&0\\ 0&0&0&0&1&0&0\\ 0&0&0&0&0&1&0\\ 0&0&0&0&0&0&1\end{smallmatrix},\ \begin{smallmatrix}0&1&0&0&0&0&0\\ 1&0&0&1&1&1&1\\ 0&0&1&1&1&1&1\\ 0&1&1&1&1&1&1\\ 0&1&1&1&1&1&1\\ 0&1&1&1&1&1&1\\ 0&1&1&1&1&1&5\end{smallmatrix},\ \begin{smallmatrix}0&0&1&0&0&0&0\\ 0&0&1&1&1&1&1\\ 1&1&1&1&1&1&1\\ 0&1&1&2&1&1&1\\ 0&1&1&1&2&1&1\\ 0&1&1&1&1&2&1\\ 0&1&1&1&1&1&6\end{smallmatrix},\ \begin{smallmatrix}0&0&0&1&0&0&0\\ 0&1&1&1&1&1&1\\ 0&1&1&2&1&1&1\\ 1&1&2&1&1&2&1\\ 0&1&1&1&2&2&1\\ 0&1&1&2&2&1&1\\ 0&1&1&1&1&1&7\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&1&0&0\\ 0&1&1&1&1&1&1\\ 0&1&1&1&2&1&1\\ 0&1&1&1&2&2&1\\ 1&1&2&2&1&1&1\\ 0&1&1&2&1&2&1\\ 0&1&1&1&1&1&7\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&0&1&0\\ 0&1&1&1&1&1&1\\ 0&1&1&1&1&2&1\\ 0&1&1&2&2&1&1\\ 0&1&1&2&1&2&1\\ 1&1&2&1&2&1&1\\ 0&1&1&1&1&1&7\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&0&0&1\\ 0&1&1&1&1&1&5\\ 0&1&1&1&1&1&6\\ 0&1&1&1&1&1&7\\ 0&1&1&1&1&1&7\\ 0&1&1&1&1&1&7\\ 1&5&6&7&7&7&8\end{smallmatrix}$
The determinant of the third fusion matrix is $16$ whereas its norm is $8$, proving that this integral fusion ring is non-Burnside, and so cannot be categorified.
12.2. Dual-Burnside property
We made the abbreviation (A)RN for (abelian) real non-negative. Note that [LPW21, Corollary 8.5] means that the Grothendieck ring, if commutative, of a unitary fusion category, has a RN dual (hypergroup).
Question 12.1.
Is the dual of the Grothendieck ring of a complex fusion category always RN?
Recall that a commutative fusion ring $F$ with RN dual is called dual-Burnside when a column of its character table has a zero entry if and only if the squared norm of the column (i.e. the corresponding formal codegree) has value strictly less than $\mathrm{FPdim}(F)$, or alternatively, when $P^{2}$ is an idempotent, where
$P=\prod_{i\in I}\frac{x_{i}}{\mathrm{FPdim}(x_{i})}$
and $\{x_{i}\}_{i\in I}$ is the set of basis elements (Corollary 4.16). We warn the reader that for a non-abelian group $G$, the product of all its elements depends on the ordering, in fact, the set of all possible products forms a $G^{\prime}$-coset, see [DH82].
Recall that [INW99, Theorem B] states that for all finite nilpotent group $G$ then the Grothendieck ring of $\mathrm{Rep}(G)$ is dual-Burnside. It is extended to every nilpotent dualizable ARN-hypergroup by Theorem 1.12. So every commutative nilpotent fusion ring with RN dual is both Burnside and dual-Burnside. Note that all the commutative nilpotent fusion rings we checked have RN dual.
Question 12.2.
Is there a commutative nilpotent fusion ring without RN dual?
Let us call a finite group $G$ dual-Burnside if $K(\mathrm{Rep}(G))$ is dual-Burnside. So a nilpotent finite group is dual-Burnside. There are exactly $144$ finite groups of order less than $32$, and among them, exactly $30$ ones are non-nilpotent, and among them, only $\mathrm{SL}(2,3)$ is dual-Burnside (so the $29$ other ones, e.g. $S_{3}$ and $D_{5}$, are not dual-Burnside).
Observe that a finite group $G$ is centerless if and only if $K(\mathrm{Rep}(G))$ has a perfect dual (see Definition 2.39, Theorem 7.8 and Lemma 11.4). So a centerless finite group is dual-Burnside if and only if every non-$\mathrm{FPdim}$ column of its character table has a zero entry. We checked that every centerless and dual-Burnside finite group $G$ of order $|G|\leq 1000$ is almost simple (i.e. $S\subseteq G\subseteq\mathrm{Aut}(S)$ with $S$ non-abelian simple); but $G=A_{5}\times A_{5}$, of order $3600$, is not almost simple. A finite non-abelian simple group is centerless, so by [Pal] and its answers, all the non-alternating finite simple groups are dual-Burnside, except the Mathieu groups $M_{22}$ and $M_{24}$. The alternating group $A_{n}$, $5\leq n\leq 19$, is dual-Burnside if and only if $n\not\in\{7,11,13,15,16,18,19\}$. Recall that a finite group $G$ is simple if and only if $K(\mathrm{Rep}(G))$ is simple.
12.3. Modular categories
Let us provide some applications of Theorem 1.16. There are exactly $43$ different types for a $\frac{1}{2}$-Frobenius integral fusion ring up to rank $11$ (see [ABPP]), and among them, exactly $7$ ones (below) are excluded from modular categorification by Theorem 1.18.
$[1,1,1,1,2,2],[1,1,1,1,2,2,2,2],[1,1,1,1,2,2,2,2,2],[1,1,1,1,2,2,2,4,4],$
$[1,1,1,1,1,1,1,1,1,3],[1,1,1,1,2,2,2,2,2,2],[1,1,1,1,2,2,2,4,4,4,4].$
For example, let us provide all the fusion data for the first type above:
  • Rank: $6$,
  • FPdim: $12=2^{2}\times 3$,
  • Type: $[1,1,1,1,2,2]$,
  • Fusion data 1:
    $\begin{smallmatrix}1&0&0&0&0&0\\ 0&1&0&0&0&0\\ 0&0&1&0&0&0\\ 0&0&0&1&0&0\\ 0&0&0&0&1&0\\ 0&0&0&0&0&1\end{smallmatrix},\ \begin{smallmatrix}0&1&0&0&0&0\\ 1&0&0&0&0&0\\ 0&0&0&1&0&0\\ 0&0&1&0&0&0\\ 0&0&0&0&0&1\\ 0&0&0&0&1&0\end{smallmatrix},\ \begin{smallmatrix}0&0&1&0&0&0\\ 0&0&0&1&0&0\\ 1&0&0&0&0&0\\ 0&1&0&0&0&0\\ 0&0&0&0&0&1\\ 0&0&0&0&1&0\end{smallmatrix},\ \begin{smallmatrix}0&0&0&1&0&0\\ 0&0&1&0&0&0\\ 0&1&0&0&0&0\\ 1&0&0&0&0&0\\ 0&0&0&0&1&0\\ 0&0&0&0&0&1\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&1&0\\ 0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 0&0&0&0&1&0\\ 1&0&0&1&0&1\\ 0&1&1&0&1&0\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&0&1\\ 0&0&0&0&1&0\\ 0&0&0&0&1&0\\ 0&0&0&0&0&1\\ 0&1&1&0&1&0\\ 1&0&0&1&0&1\end{smallmatrix}$
  • Fusion data 2:
    $\begin{smallmatrix}1&0&0&0&0&0\\ 0&1&0&0&0&0\\ 0&0&1&0&0&0\\ 0&0&0&1&0&0\\ 0&0&0&0&1&0\\ 0&0&0&0&0&1\end{smallmatrix},\ \begin{smallmatrix}0&1&0&0&0&0\\ 1&0&0&0&0&0\\ 0&0&0&1&0&0\\ 0&0&1&0&0&0\\ 0&0&0&0&1&0\\ 0&0&0&0&0&1\end{smallmatrix},\ \begin{smallmatrix}0&0&1&0&0&0\\ 0&0&0&1&0&0\\ 0&1&0&0&0&0\\ 1&0&0&0&0&0\\ 0&0&0&0&0&1\\ 0&0&0&0&1&0\end{smallmatrix},\ \begin{smallmatrix}0&0&0&1&0&0\\ 0&0&1&0&0&0\\ 1&0&0&0&0&0\\ 0&1&0&0&0&0\\ 0&0&0&0&0&1\\ 0&0&0&0&1&0\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&1&0\\ 0&0&0&0&1&0\\ 0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 1&1&0&0&0&1\\ 0&0&1&1&1&0\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 0&0&0&0&1&0\\ 0&0&0&0&1&0\\ 0&0&1&1&1&0\\ 1&1&0&0&0&1\end{smallmatrix}$
  • Fusion data 3:
    $\begin{smallmatrix}1&0&0&0&0&0\\ 0&1&0&0&0&0\\ 0&0&1&0&0&0\\ 0&0&0&1&0&0\\ 0&0&0&0&1&0\\ 0&0&0&0&0&1\end{smallmatrix},\ \begin{smallmatrix}0&1&0&0&0&0\\ 1&0&0&0&0&0\\ 0&0&0&1&0&0\\ 0&0&1&0&0&0\\ 0&0&0&0&1&0\\ 0&0&0&0&0&1\end{smallmatrix},\ \begin{smallmatrix}0&0&1&0&0&0\\ 0&0&0&1&0&0\\ 1&0&0&0&0&0\\ 0&1&0&0&0&0\\ 0&0&0&0&1&0\\ 0&0&0&0&0&1\end{smallmatrix},\ \begin{smallmatrix}0&0&0&1&0&0\\ 0&0&1&0&0&0\\ 0&1&0&0&0&0\\ 1&0&0&0&0&0\\ 0&0&0&0&1&0\\ 0&0&0&0&0&1\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&1&0\\ 0&0&0&0&1&0\\ 0&0&0&0&1&0\\ 0&0&0&0&1&0\\ 0&0&0&0&0&2\\ 1&1&1&1&0&0\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 1&1&1&1&0&0\\ 0&0&0&0&2&0\end{smallmatrix}$
  • Fusion data 4:
    $\begin{smallmatrix}1&0&0&0&0&0\\ 0&1&0&0&0&0\\ 0&0&1&0&0&0\\ 0&0&0&1&0&0\\ 0&0&0&0&1&0\\ 0&0&0&0&0&1\end{smallmatrix},\ \begin{smallmatrix}0&1&0&0&0&0\\ 1&0&0&0&0&0\\ 0&0&0&1&0&0\\ 0&0&1&0&0&0\\ 0&0&0&0&1&0\\ 0&0&0&0&0&1\end{smallmatrix},\ \begin{smallmatrix}0&0&1&0&0&0\\ 0&0&0&1&0&0\\ 0&1&0&0&0&0\\ 1&0&0&0&0&0\\ 0&0&0&0&1&0\\ 0&0&0&0&0&1\end{smallmatrix},\ \begin{smallmatrix}0&0&0&1&0&0\\ 0&0&1&0&0&0\\ 1&0&0&0&0&0\\ 0&1&0&0&0&0\\ 0&0&0&0&1&0\\ 0&0&0&0&0&1\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&1&0\\ 0&0&0&0&1&0\\ 0&0&0&0&1&0\\ 0&0&0&0&1&0\\ 0&0&0&0&0&2\\ 1&1&1&1&0&0\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 1&1&1&1&0&0\\ 0&0&0&0&2&0\end{smallmatrix}$
They are excluded by Theorem 1.18 from modular categorification, because $3$ is a powerless factor of the $\mathrm{FPdim}$, whereas $3$ does not divide $4$, the number of grouplike elements. Note that the above fusion data $3$ and $4$ can be generalized as follows. Let $G$ be a group of order $n^{2}$, and let $K$ be a group of order $m+1$. Consider the following $\frac{1}{2}$-Frobenius integral fusion ring:
  • Rank: $n^{2}+m$,
  • FPdim: $n^{2}(m+1)$,
  • Type: $[[1,n^{2}],[n,m]]$,
  • Basis: =$\mathcal{B}=\{x_{g}\}_{g\in G}\cup\{\rho_{k}\}_{k\in K\setminus\{e\}}$,
  • Fusion data:
    • $\rho_{e}:=\frac{1}{n}\sum_{g\in G}x_{g}$,
    • $x_{g}x_{h}=x_{gh}$, for all $g,h\in G$,
    • $\rho_{k}\rho_{l}=n\rho_{kl}$, for all $k,l\in K$,
    • $x_{g}\rho_{k}=\rho_{k}x_{g}=\rho_{k}$, for all $(g,k)\in G\times K$,
It is excluded by Theorem 1.16 (1.13) from modular categorification if $n>1$ and $m+1$ has a prime factor not dividing $n$.
Availability of data and materials
The datasets generated during and/or analyzed during the current study are available from the second author on reasonable request.
Declaration
Conflict of interests: The authors declare that they have no conflict of interest.