### v11

# Burnside type results for fusion categories

Abstract.
In this paper, we extend a classical vanishing result of Burnside from the character tables of finite groups to the character tables of commutative fusion rings, or more generally to a certain class of abelian normalizable hypergroups. We also treat the dual vanishing result. We show that any nilpotent unitary fusion categories satisfy both Burnside’s property and its dual. Using Drinfeld’s map, we obtain that the Grothendieck ring of any weakly-integral modular fusion category satisfies both properties. As applications, we prove new identities that hold in the Grothendieck ring of any weakly-integral fusion category satisfying the dual-Burnside’s property, thus providing new categorification criteria. In particular we improve [OY23, Theorem 4.5] as follows: A weakly integral modular fusion category of $\mathrm{FPdim}$ $md$ with $d$ square-free coprime with $m$ and $\mathrm{FPdim}(X)^{2}$ for every simple object $X$, has a pointed modular fusion subcategory of $\mathrm{FPdim}$ $d$. We also present new results on perfect modular fusion categories, including a Cauchy-type theorem.
1991 Mathematics Subject Classification:
18M20; 20N20; 20C15; 18N25; 16T20; 16T30
The first author is supported by a grant of the Ministry of Research, Innovation and Digitization, CNCS/CCCDI - UEFISCDI, project number PN-III-P4-ID-PCE-2020-0878, within PNCDI III. The second author is supported by BIMSA Start-up Research Fund, Foreign Youth Talent Program from the Ministry of Sciences and Technology of China and National Natural Science Foundation of China (NSFC, Grant no. 12471031)
1. Introduction
A classical result of Burnside in the character theory of finite groups states that any irreducible non-linear character of a finite group vanishes on at least one element of the group. This can be stated as follows: in the character table of a finite group, the row of every character of degree 1$\neq$ 1≠ 1 contains a zero entry.
More recently, the dual version of this result was also intensively studied in the literature. This dual version describes the group elements that vanish on at least one irreducible character, see [INW99] and the references therein. Burnside’s result was extended to every weakly integral fusion category with a commutative Grothendieck ring, initially in the context of modular categories in [GNN09, Appendix], and subsequently in the general framework in [Bur23].
The main goal of this paper is to develop an analogue of Burnside’s result for a certain large class of fusion rings, or even more generally, to a certain class of abelian rational normalizable hypergroups. We also consider the dual version of Burnside’s result in this more general settings.
Recall that in [Bur21] the author introduced the notion of dual of a Grothendieck ring. This was achieved based on work of Harrison [Har79] on dualizable probability groups and on [ZZ21]. More generally, this notion of dual of a fusion ring was extended to arbitrary fusion rings in [BP25]. It was noticed in [Bur21] that the dual of the Grothendieck ring of a pivotal fusion category is isomorphic to the center of the category as defined in [Shi17].
In this paper we show that both Burnside’s property and its dual are related with the ring structure of the dual of the involving Grothendieck/fusion ring. In general, the dual of a fusion ring is no longer a fusion ring, but an abelian normalizable hypergroup, see [Bur21]. The advantage of the class of hypergroups is to be closed under duality, see [Bur21, Proposition 2.3].
For a finite dimensional $\mathbb{C}$-algebra $H$, fix a (finite) $\mathbb{C}$-linear basis $\mathcal{B}$ with the unit $1\in\mathcal{B}$ and write $ab=\sum_{c\in{\mathcal{B}}}N^{c}_{ab}c$ for some scalars $N^{c}_{ab}\in\mathbb{C}$. Note that $H={\mathbb{C}[\mathcal{B}]}$, the complex linear span of $\mathcal{B}$.
Definition 1.1.
A finite dimensional algebra $H$ together with a $\mathbb{C}$-linear basis $\mathcal{B}$ is called a hypergroup if there is an involution map $(-)^{*}:{\mathcal{B}}\rightarrow{\mathcal{B}}$ such that $N^{1}_{aa^{*}}>0$ and $N^{1}_{ab}=0$ if $b\neq a^{*}$. The elements $b\in\mathcal{B}$ are called the standard elements of ${(H,\mathcal{B})}$.
We may refer to ${\mathcal{B}}$ (or $H$) only as a hypergroup, when $H$ (or ${\mathcal{B}}$) is implicitly understood. A hypergroup $\mathcal{B}$ is called:
  • symmetric if $N^{1}_{ab}=N^{1}_{ba}$,
  • normalized if $\sum_{c\in{\mathcal{B}}}N^{c}_{ab}=1$,
  • real if $N_{ab}^{c}\in\mathbb{R}$,
  • rational if $N_{ab}^{c}\in\mathbb{Q}$,
  • real non-negative if $N_{ab}^{c}\geq 0$,
  • abelian if $ab=ba$,
for all $a,b,c\in{\mathcal{B}}$. Throughout this paper we use the abbreviation (A)RN-hypergroup for an (abelian) real non-negative hypergroup.
For a normalized hypergroup, the linear map $\mu_{1}:H\rightarrow\mathbb{C}$ defined on the linear basis ${\mathcal{B}}$ by $\mu_{1}(a)=1$ for all $a\in{\mathcal{B}}$, is a morphism of algebras. It is called the augmentation map of $(H,\mathcal{B})$. The underlying algebra $H$ of a hypergroup $(H,\mathcal{B})$ is a finite dimensional *-algebra, so is semisimple. For a normalized hypergroup ${(H,\mathcal{B})}$ the central primitive idempotent $F_{1}$ corresponding to the character $\mu_{1}$ is denoted by $\lambda_{(H,\mathcal{B})}$ or simply $\lambda_{H}$ if $\mathcal{B}$ is implicitly understood. A formula for this idempotent is described in Corollary 2.10.
A RN-hypergroup is sometimes called a table algebra in the literature. Recall that a fusion ring ${\mathcal{B}}$ is a hypergroup such that $N^{a}_{bc}\in\mathbb{Z}_{\geq 0}$ and $N^{1}_{aa^{*}}=N^{1}_{a^{*}a}=1$ for all $a,b,c\in\mathcal{B}$. The first main result of this paper is the following generalization of Burnside’s result to some weakly-integral fusion rings:
Theorem 1.2.
Let ${(H,\mathcal{B})}$ be a commutative fusion ring with a $h$-integral dual. Let $a$ be a standard element such that $\mathrm{FPdim}(a)>1$. Then there is $\mu\in{\widehat{\mathcal{B}}}$ such that $\mu(a)=0$.
A fusion ring whose dual is h-integral (see Definition 3.6) is always weakly-integral ($\mathrm{FPdim}{(H,\mathcal{B})}\in\mathbb{Z}$). The converse is true for a weakly-integral fusion category (see Remark 3.7). The set $^{\widehat{\mathcal{B}}}$ is the set of characters $\mu_{j}:H\rightarrow\mathbb{C}$ (see §2 for more details). In fact we prove a more general version of Theorem 1.2, for a certain class of abelian rational normalizable hypergroups, see Theorem 4.17.
Definition 1.3.
For any hypergroup ${(H,\mathcal{B})}$, an element $a\in\mathcal{B}$ with
(1.1) $aa^{*}=N_{a,a^{*}}^{1}1\text{ (or equivalently, }h_{a}aa^{*}=1\text{, with }h_{a}:=1/N_{a,a^{*}}^{1})$
is called a grouplike element of ${(H,\mathcal{B})}$. In other words, $N_{a,a^{*}}^{c}=\delta_{1,c}N_{a,a^{*}}^{1}$.
By Lemma 2.32, the set of grouplike elements form a group (with the multiplication structure) denoted ${G{(H,\mathcal{B})}}$.
Definition 1.4.
A hypergroup ${(H,\mathcal{B})}$ has Burnside’s vanishing property (or shortly, is Burnside) if for all $a\in\mathcal{B}$, the following are equivalent:
  1. (1)
    There is some $\mu\in{\widehat{\mathcal{B}}}$ such that $\mu(a)=0$,
  2. (2)
    the standard element $a$ is not a grouplike element.
Theorem 1.2 states that a commutative fusion ring with h-integral dual is Burnside. Non-Burnside simple integral fusion rings can be found in §12. It is easy to see that a hypergroup is Burnside if and only if the set of invertible standard elements coincides to the set of grouplike elements, see Proposition 4.3.
1.1. The dual hypergroup of a normalized hypergroup
Let ${(H,\mathcal{B})}$ be an abelian normalized hypergroup. By the abelian assumption, $^{\widehat{\mathcal{B}}}$ forms a basis for the dual $H^{*}$, which then is $\mathbb{C}[{\widehat{\mathcal{B}}}]$, and on which we can define a multiplication. On the basis $^{\widehat{\mathcal{B}}}$, the multiplication $\mu_{i}\star\mu_{j}$ is defined by declaring
(1.2) $[\mu_{i}\star\mu_{j}](a):=\mu_{i}(a)\mu_{j}(a),\;\text{for all}\;a\in\mathcal{B}$
and then extending linearly $\mu_{i}\star\mu_{j}$ on the entire ${\mathbb{C}[\mathcal{B}]}$. In this way, we obtain an algebra structure on $H^{*}$.
It follows that there are some non-zero scalars ${\widehat{p}}_{k}(i,j)\in\mathbb{C}$ such that
(1.3) $\mu_{i}\star\mu_{j}=\sum_{k=1}^{m}{\widehat{p}}_{k}(i,j)\mu_{k}.$
By [Bur21, Proposition 2.3], for any abelian normalized hypergroup $(H,{\mathcal{B}})$, the pair $(H^{*},{\widehat{\mathcal{B}}})$ is also an abelian normalized hypergroup. Moreover, the involution on $^{\widehat{\mathcal{B}}}$ is given by $\mu_{j}\mapsto\mu_{j^{\#}}$, where $\mu_{j^{\#}}\in{\widehat{\mathcal{B}}}$ is defined by $\mu_{j^{\#}}(a):=\mu_{j}(a^{*})$ for all $a\in\mathcal{B}$. We denote by $\widehat{{(H,\mathcal{B})}}:=(H^{*},{\widehat{\mathcal{B}}})$ the dual hypergroup of the abelian normalized hypergroup ${(H,\mathcal{B})}$. It is also well known that for a normalized hypergroup ${(H,\mathcal{B})}$, then $\mu_{j}(a^{*})=\overline{\mu_{j}(a)}$ for all $a\in\mathcal{B}$. We also use the notation ${\widehat{h}}_{j}:={\widehat{p}}_{1}(j,j^{\#})^{-1}$, called the order of $\mu_{j}$ in the dual hypergroup $^{\widehat{(H,\mathcal{B})}}$, see Lemma 2.8.
For any RN-hypergroup ${(H,\mathcal{B})}$, we denote by ${(H,\mathcal{B})}_{ad}$ its adjoint sub-hypergroup, see §6.2. As for fusion rings, this is defined as the sub-hypergroup of ${(H,\mathcal{B})}$ generated by the element $I(1):=\sum_{i\in\mathcal{I}}h_{i}x_{i}x_{i^{*}}$, where $h_{i}:=1/N^{1}_{i,i^{*}}$.
For any RN-hypergroup ${(H,\mathcal{B})}$, the central primitive idempotent corresponding to the character $\mathrm{FPdim}:H\rightarrow\mathbb{C}$ is denoted by $\lambda_{H}$. For $H=K(\mathcal{C})$, the Grothendieck ring of a fusion category $\mathcal{C}$, this primitive idempotent corresponds to the regular element.
Definition 1.5.
A hypergroup ${(H,\mathcal{B})}$ is called normalizable if we obtain a normalized hypergroup after rescaling it. (see §2.2 for more details.)
Based on Lemma 2.3 we denote such a hypergroup by ${(H,\mathcal{B},\mu_{1})}$ where $\mu_{1}\in{\widehat{\mathcal{B}}}$.
Inspired by the notion of dualizable probability groups from [Har79], we propose the following:
Definition 1.6.
An ARN-hypergroup is called dualizable if its dual is also ARN.
1.2. Main results of the paper
Theorem 1.7.
Let $(H,B,\mu_{1})$ be an abelian normalizable hypergroup such that $G{(H,\mathcal{B})}$ is a finite group. Then $(H,{\mathcal{B}})$ is Burnside if and only if
$(\prod_{j=1}^{m}\mu_{j})^{2}=\frac{|G{(H,\mathcal{B})}|}{n{(H,\mathcal{B},\mu_{1})}}\big{(}\sum_{j\in\mathcal{I}_{G{(H,\mathcal{B})}}}{\widehat{h}}_{j}\mu_{j}\big{)}$
See Definition 2.5 for $n(\_)$). We characterize Burnside dualizable ARN-hypergroups as follows:
Theorem 1.8.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup such that $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN. Then $(H,{\mathcal{B}})$ is Burnside if and only if the following identity holds on $^{\widehat{(H,\mathcal{B})}}$:
(1.4) $\big{(}\prod_{j=1}^{m}\mu_{j}\big{)}^{2}=\frac{1}{n(\widehat{(H,{\mathcal{B}})}_{ad})}\big{(}\sum_{\mu_{j}\in{{\widehat{\mathcal{B}}}}_{ad}}{\widehat{h}}_{j}\mu_{j}\big{)}.$
Note that the RHS of the equation above corresponds to the integral $\lambda_{{}_{{\widehat{(H,\mathcal{B})}}_{ad}}}$ of the adjoint hypergroup ${\widehat{(H,\mathcal{B})}}_{ad}$.
Remark 1.9.
If the abelian normalizable hypergroup ${(H,\mathcal{B})}$ is either RN or dual RN, then according to Lemmas 2.32 and 2.36, the group $G{(H,\mathcal{B})}$ must be finite. As a result, we can apply Theorem 1.7. Specifically, by comparing this with Theorem 1.8, we can infer that in the dual RN case, $n(\widehat{(H,{\mathcal{B}})}_{ad})|G{(H,\mathcal{B})}|=n{(H,\mathcal{B})}$. By duality, we also have $n((H,{\mathcal{B}})_{ad})|G{\widehat{(H,\mathcal{B})}}|=n{(H,\mathcal{B})}$ in the RN case. These equalities can also be obtained from Harrison’s Equations (7.7) and (7.6), although in the more restrictive dualizable ARN case.
For a fusion category $\mathcal{C}$, its Grothendieck ring is denoted by $K(\mathcal{C})$, and the set of isomorphism class representatives of simple objects of $\mathcal{C}$ by $\mathrm{Irr}(\mathcal{C}):=\{X_{1},\dots,X_{m}\}$. Let $d_{i}:=\mathrm{FPdim}(X_{i})$ be the Frobenius-Perron dimension of $X_{i}$, $i\in{\mathcal{I}}=\{1,\dots,m\}$, and $x_{i}:=[X_{i}]$ be the class of the simple object $X_{i}$ in the Grothendieck ring $K(\mathcal{C})$.
For any weakly-integral premodular category $\mathcal{C}$, $^{\widehat{{K(\mathcal{C})}}}$ is both RN and rational, as demonstrated in [Bur21, Theorem 1.2]. Furthermore, for any unitary fusion category $\mathcal{C}$, $^{\widehat{{K(\mathcal{C})}}}$ is also RN, see [LPW21, ENO25], and primary 333-criterion in [HLPW24]. A normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$ is called dual-Burnside if its dual $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is Burnside (see examples from group theory in §12). The dual version of Theorem 1.7, is the following:
Theorem 1.10.
Let $(H,B,\mu_{1})$ be an abelian normalizable hypergroup such that $G{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is a finite group. Then $(H,{\mathcal{B}})$ is dual-Burnside if and only if
$(\prod_{i=1}^{m}{\frac{x_{i}}{d_{i}}})^{2}=\frac{|G{\widehat{{(H,\mathcal{B},\mu_{1})}}}|}{n{(H,\mathcal{B},\mu_{1})}}\big{(}\sum_{i\in\mathcal{I}_{G{\widehat{{(H,\mathcal{B},\mu_{1})}}}}}h_{i}d_{i^{*}}x_{i}\big{)}$
On the other hand, the dual version of Theorem 1.8 is:
Theorem 1.11.
The following are equivalent for an ARN-hypergroup ${(H,\mathcal{B})}$.
  1. (1)
    The hypergroup ${(H,\mathcal{B})}$ is dual-Burnside.
  2. (2)
    The following identity holds:
    (1.5) $\big{(}\prod_{i=1}^{m}{\frac{x_{i}}{d_{i}}}\big{)}^{2}=\frac{1}{n({{(H,\mathcal{B})}_{ad}})}\big{(}\sum_{x_{i}\in\mathcal{B}_{ad}}h_{i}d_{i^{*}}x_{i}\big{)}$
The dual-Burnside fusion categories can be characterized by:
Theorem 1.12.
Let $\mathcal{C}$ be a fusion category with a commutative Grothendieck ring. Then ${K(\mathcal{C})}$ is dual-Burnside if and only if the following holds:
(1.6) $\big{(}\prod_{i=1}^{m}{\frac{x_{i}}{d_{i}}}\big{)}^{2}=\frac{1}{\mathrm{FPdim}({\mathcal{C}_{\mathrm{ad}}})}\big{(}\sum_{x_{i}\in{\mathcal{C}_{\mathrm{ad}}}}d_{i}x_{i}\big{)}.$
About the dual-Burnside hypergroups, we prove the following:
Theorem 1.13.
Let $(H,{\mathcal{B}})$ be an abelian rational RN-hypergroup. If $(H,{\mathcal{B}})$ is dual-Burnside then $\mathrm{FPdim}(H,{\mathcal{B}})\in\mathbb{Q}$.
As a corollary, any commutative dual-Burnside fusion ring is weakly-integral. Similar to the fusion ring settings, [GN08], the universal grading group can be defined for any ARN-hypergroup. Using this grading, we can similarly define the concept of nilpotent ARN-hypergroup, and prove the following:
Theorem 1.14.
Let ${(H,\mathcal{B})}$ be a dualizable ARN-hypergroup. Then ${(H,\mathcal{B})}$ is nilpotent if and only if its dual $^{\widehat{(H,\mathcal{B})}}$ is also nilpotent. Moreover, in this case they have the same nilpotency class.
For a nilpotent ARN-hypergroup, we prove the following:
Theorem 1.15.
A nilpotent dualizable ARN-hypergroup is both Burnside and dual-Burnside.
Let $\mathcal{C}$ be a modular fusion category. Then Corollary 10.3 states that ${K(\mathcal{C})}$ is Burnside if and only if it is dual-Burnside. The proof involves the Drinfeld map of $\mathcal{C}$. Then Theorem 1.12 implies that:
Theorem 1.16.
In any modular fusion category $\mathcal{C}$ with ${K(\mathcal{C})}$ Burnside:
(1.7) $\big{(}\prod_{i=1}^{m}\frac{x_{i}}{d_{i}}\big{)}^{2}=\frac{1}{\mathrm{FPdim}({\mathcal{C}_{\mathrm{ad}}})}\big{(}\sum_{x_{i}\in\mathrm{Irr}({\mathcal{C}_{\mathrm{ad}}})}d_{i}x_{i}\big{)}$
Note that the above theorem holds for weakly-integral modular categories since their Grothendieck rings are Burnside by [GNN09, Appendix] or [Bur23, Theorem 2].
Corollary 1.17.
In any modular fusion category $\mathcal{C}$ with ${K(\mathcal{C})}$ Burnside and with the group of invertible objects $G(\mathcal{C})$ of odd order,
(1.8) $\prod_{i=1}^{m}\frac{x_{i}}{d_{i}}=\frac{1}{\dim({\mathcal{C}_{\mathrm{ad}}})}\big{(}\sum_{x_{i}\in\mathrm{Irr}({\mathcal{C}_{\mathrm{ad}}})}d_{i}x_{i}\big{)}$
For any non-negative integer $n$, let $\mathcal{V}(n)$ be the set of prime divisors of $n$. For a weakly-integral fusion category we also let $\mathcal{V}(\mathcal{C}):=\mathcal{V}(\mathrm{FPdim}(\mathcal{C}))$. From Theorem 1.12, we can also derive the following result:
Theorem 1.18.
Let $\mathcal{C}$ be a fusion category such that ${K(\mathcal{C})}$ is commutative and dual-Burnside. Then
(1.9) $\frac{({\prod_{i=1}^{m}}d_{i})^{2}}{\mathrm{FPdim}({\mathcal{C}_{\mathrm{ad}}})}\in\mathbb{Z}.$
Moreover, if additionally $\mathcal{C}$ is nilpotent then
(1.10) $\mathcal{V}({\mathcal{C}_{\mathrm{ad}}})=\bigcup_{i=1}^{m}\mathcal{V}(d_{i}^{2}).$
Note that since any nilpotent fusion category $\mathcal{C}$ is weakly-integral then $d_{i}^{2}\in\mathbb{Z}$ for all $i$, see [GN08, Theorem 3.10]. Note that in the case of an integral fusion category Equation (1.10) can be written as
(1.11) $\mathcal{V}({\mathcal{C}_{\mathrm{ad}}})=\bigcup_{i=1}^{m}\mathcal{V}(d_{i}).$
Using the above identities, we prove the following results about the $\mathrm{FPdim}$ of simple objects:
Theorem 1.19.
For any modular fusion category $\mathcal{C}$ such that ${K(\mathcal{C})}$ is dual-Burnside,
(1.12) $\frac{({\prod_{i=1}^{m}}d_{i})^{2}}{\mathrm{FPdim}({\mathcal{C}_{\mathrm{ad}}})}\in\mathbb{Z}.$
For any weakly-integral modular fusion category $\mathcal{C}$:
(1.13) $\mathcal{V}(\mathcal{C})=\mathcal{V}({\mathcal{C}_{\mathrm{pt}}})\cup\big{(}\bigcup_{i=1}^{m}\mathcal{V}(d_{i}^{2})\big{)}$
Again, recall $d_{i}^{2}\in\mathbb{Z}$ for all $i$ by [GN08, Theorem 3.10]. Moreover, if $\mathcal{C}$ is an integral modular fusion category then the above equation becomes:
(1.14) $\mathcal{V}(\mathcal{C})=\mathcal{V}({\mathcal{C}_{\mathrm{pt}}})\cup\big{(}\bigcup_{i=1}^{m}\mathcal{V}(d_{i})\big{)}$
In particular,
Remark 1.20 (Cauchy-type theorem).
For every perfect integral modular fusion category,
(1.15) $\mathcal{V}(\mathcal{C})=\bigcup_{i=1}^{m}\mathcal{V}(d_{i}),$
thus, for all prime $p$ dividing $\mathrm{FPdim}(\mathcal{C})$, there is a simple object $X$ in $\mathcal{C}$ such that $p$ divides $\mathrm{FPdim}(X)$; which can be interpreted as a Cauchy-type theorem. Consequently, $p^{2}$ divides $\mathrm{FPdim}(\mathcal{C})$ since $\mathrm{FPdim}(X)^{2}$ divides $\mathrm{FPdim}(\mathcal{C})$ by [EGNO15, Proposition 8.14.6].
Concrete examples where these results apply can be found in §12.3. Next theorem improves [OY23, Theorem 4.5].
Theorem 1.21.
Let $\mathcal{C}$ be a weakly-integral modular fusion category of $\mathrm{FPdim}=dm$ with $d$ square-free coprime with $m$ and $d_{i}^{2}$ for all $i$. Then $d\mid\mathrm{FPdim}({\mathcal{C}_{\mathrm{pt}}})$ and $\mathcal{C}$ admits a decomposition
$\mathcal{C}\simeq{\mathcal{D}}\boxtimes{\mathcal{D}}^{\prime}$
where ${\mathcal{D}}$ is a pointed modular fusion category of dimension $d$ and ${\mathcal{D}}^{\prime}$ a modular fusion category of dimension $m$.
Corollary 1.22.
Let $\mathcal{C}$ be any integral modular fusion category of $\mathrm{FPdim}=dm$ with $d$ square-free coprime with $m$. Then $d\mid\mathrm{FPdim}({\mathcal{C}_{\mathrm{pt}}})$ and $\mathcal{C}$ admits a decomposition
$\mathcal{C}\simeq{\mathcal{D}}\boxtimes{\mathcal{D}}^{\prime}$
where ${\mathcal{D}}$ is a pointed modular fusion category of dimension $d$ and ${\mathcal{D}}^{\prime}$ a modular fusion category of dimension $m$.
In particular, the $\mathrm{FPdim}$ of a perfect integral modular fusion category does not have a powerless prime factor (Corollary 10.10). In particular, a perfect modular fusion category of even $\mathrm{FPdim}$ has its $\mathrm{FPdim}$ divisible by 444.
Remark 1.23.
By the classification of non-pointed integral modular fusion categories of small rank in [ABPP] and by Corollary 1.22, the $\mathrm{FPdim}$ of a non-pointed integral modular fusion category does not have a powerless prime factor for every rank less than 11×3=3311$\times$ 3=3311 × 3 = 33, and every rank less than 17×7=11917$\times$ 7=11917 × 7 = 119 in the odd-dimensional case.
Integral modular fusion categories were recently intensively studied in the literature. The following conjecture is formulated in [CP22]:
Conjecture 1.24.
There is no non-trivial perfect modular fusion category of odd $\mathrm{FPdim}$.
This conjecture was checked for ranks less than 252525 in [ABPP] or [CGP23]. Note that the above result also applies to non-pointed simple integral modular fusion categories (since they are in particular perfect). The importance of the existence of such categories was described in [LPR23, §5], where the following two open statements are proved to be equivalent, whereas Proposition 10.13 states that they imply Conjecture 1.24.
Statement 1.25.
Every simple integral fusion category is weakly group-theoretical.
Statement 1.26.
Every simple integral modular fusion category is pointed.
We finally prove the following result:
Theorem 1.27.
A perfect modular fusion category is (dual-)Burnside if and only if it is integral.
Note that the integral modular fusion category $\mathcal{Z}(\mathrm{Rep}(G))$ is perfect if and only if the finite group $G$ is perfect and centerless (more generally, see §11.1).
Shortly, the organization of this paper is the following. In §2 we recall the basics on hypergroups and abelian rational normalizable hypergroups. In §3 we develop few basic facts about Galois symmetries of rational hypergroups. In §4 hypergroups with Burnside property are studied. In the same section, we prove Theorem 1.2 and its more general version Theorem 4.17 for hypergroups. In this section, Burnside’s property for dual hypergroups is also studied. In §5 we extend the notion of kernels of objects from the settings of fusion categories to the settings of ARN-hypergroups. In §6 and §7, we construct the universal grading group for ARN-hypergroups. A connection between this universal grading group and the group of grouplike elements of the dual is proven in Theorem 6.5. Theorems 1.8 and 1.11 are proven. In §8 we define nilpotent ARN-hypergroup and prove Theorems 1.14 and 1.15. In §9 we prove Theorem 1.12. In §10, the applications to modular fusion categories are given (e.g. the proofs from Theorem 1.16 to Corollary 1.22), and we prove Proposition 10.13 showing that any of the two statements from [LPR23, §5] implies Conjecture 1.24.
Next, §11 delves into applications and supplementary results: §11.1 offers a characterization for a Drinfeld center to be perfect; §11.2 establishes some results concerning integrality and Burnside properties, proving Theorems 1.13 and 1.27, particularly highlighting their equivalence in the perfect modular case; §11.3 characterizes the near-group fusion categories which admit a modular structure.
Finally, §12 presents concrete examples and counterexamples: §12.1 introduces non-Burnside simple integral fusion rings; §12.2 classifies certain dual-Burnside $\mathrm{Rep}(G)$; §12.3 provides certain families of fusion rings lacking modular categorification by Theorems 1.19 or 1.22.
Throughout of this paper, all vector spaces and linear categories are considered over the ground field $\mathbb{C}$ of complex numbers.
Acknowledgements
We thank Vicenţiu Paşol and Andrew Schopieray for fruitful discussions on Galois groups of fusion rings and hypergroups.
2. Abelian rational normalizable hypergroups
Let ${(H,\mathcal{B})}$ be a hypergroup as in Definition 1.1. A morphism of hypergroups $\psi:(H,{\mathcal{B}})\rightarrow(H^{\prime},{\mathcal{B}}^{\prime})$ is an algebra morphism $\psi:H\rightarrow H^{\prime}$ such that $\psi({\mathcal{B}})={\mathcal{B}}^{\prime}$ and $\psi(b^{*})=\psi(b)^{*}$ for all $b\in\mathcal{B}$. The morphism $\psi$ is called an isomorphism of hypergroups if it is bijective.
Let ${(H,\mathcal{B})}$ be a normalized abelian hypergroup and $^{\widehat{(H,\mathcal{B})}}$ its dual as defined in the introduction. There is a formula for the dual fusion coefficients ${\widehat{p}}_{k}(i,j)$ given in [Bur21, Proposition 2.1] by:
(2.1) $\widehat{p}_{k}(i,j)=\frac{1}{n_{k}}\bigg{(}\sum_{a\in{\mathcal{B}}}h_{a}\mu_{i}(a)\mu_{j}(a)\mu_{k}(a^{*})\bigg{)}$
where the scalars $n_{k}$ are defined below (2.2), and $h_{a}:=1/N^{1}_{a,a^{*}}$ (then $h_{1}=1$).
Remark 2.1.
Let ${(H,\mathcal{B})}$ be an abelian normalized hypergroup. We usually let :=${\mathcal{B}}:=\{x_{i}\}_{i\in\mathcal{I}}$, with :=$\mathcal{I}:=\{1,\dots m\}$ and $x_{1}=1$, and say that ${(H,\mathcal{B})}$ has rank $m$. We can simplify the notation $N_{x_{i}x_{j}}^{x_{k}}$ by $N_{i,j}^{k}$, and $h_{x_{i}}$ by $h_{i}$.
Remark 2.2.
Let $\{x_{i}^{\circ}\}_{i\in\mathcal{I}}$ be the linear dual basis of the linear basis ${\mathcal{B}}$ of $H$. Therefore $x_{i}^{\circ}(x_{j})=\delta_{i,j}$, for all $i,j\in\mathcal{I}$. It is easy to see that in this case $\widehat{F}_{i}:=x_{i}^{\circ}\in H^{*}$ is a primitive central idempotent of $H^{*}$ corresponding to the character
$\begin{array}[]{cccc}\widetilde{\omega}_{i}=\operatorname{ev}_{x_{i}}:&H^{*}&\to&\mathbb{C}\\ &f&\mapsto&f(x_{i}).\end{array}$
2.1. Function $\tau$ and associative non-degenerate bilinear form
Let $(H,\mathcal{B})$ be any abelian hypergroup. The finite dimensional *-algebra $H$ (so semismple) is commutative, therefore $H\simeq\bigoplus_{i=1}^{m}\mathbb{C}F_{i}$, where $(F_{i})$ are the central primitive idempotents of $H$. We can define $\tau\in H^{*}$ with $\tau(x_{i})=\delta_{i,1}$, where as above, $x_{1}=1$. Note that $\tau=x_{1}^{\circ}$ with the above notations. Moreover, in this case
$\begin{array}[]{cccc}(\;,\;)_{\tau}:&H\times H&\to&\mathbb{C}\\ &(a,b)&\mapsto&\tau(ab)\end{array}$
is an associative symmetric non-degenerate bilinear form. Thus, $\tau(F_{i})$ must be nonzero for all $i$. The nonzero scalars $n_{i}:=\tau(F_{i})^{-1}$ are called the formal codegrees of $(H,\mathcal{B})$, see [Ost15]. We can write
(2.2) $\tau=\sum_{i=1}^{m}\tau(F_{i})\mu_{i}=\sum_{i=1}^{m}\frac{1}{n_{i}}\mu_{i}.$
Note that $\sum_{i\in\mathcal{I}}\frac{1}{n_{i}}=\tau(1)=1$.
From the definition of a hypergroup it follows that $\{h_{i}x_{i}\}_{i\in\mathcal{I}}$ and $\{x_{i^{*}}\}_{i\in\mathcal{I}}$ form a pair of dual bases for $(\;,\;)_{\tau}$. Note that in this case $\{n_{j}F_{j}\}_{j\in\mathcal{I}}$ and $\{F_{j}\}_{j\in\mathcal{I}}$ form also a pair of dual basis for the bilinear form $(\;,\;)_{\tau}$. The existence of the two pairs of dual bases implies that:
(2.3) $\sum_{i\in\mathcal{I}}{h_{i}}x_{i}\otimes x_{i}^{*}=\sum_{j\in\mathcal{I}}n_{j}F_{j}\otimes F_{j}$
By applying $\operatorname{id}\otimes\mu_{j}$, for the character $\mu_{j}$ corresponding to $F_{j}$, it follows that
(2.4) $F_{j}=\frac{1}{n_{j}}\sum_{i\in\mathcal{I}}{h_{i}}\mu_{j}(x_{i^{*}})x_{i}$
Applying $\mu_{j}\otimes\mu_{k}$ to the above, we obtain the first orthogonality relation
(2.5) $\sum_{i\in\mathcal{I}}{h_{i}}\mu_{j}(x_{i})\mu_{k}(x_{i^{*}})=\delta_{j,k}n_{j}.$
Since $\mu_{j}(x_{i^{*}})=\overline{\mu_{j}(x_{i})}$, for $j=k$, it follows from here that $n_{j}>0$. By a classical argument (see, for example, the proof of [LPR23, Theorem 2.4]), as for finite groups orthogonality, we obtain the second orthogonality relation:
(2.6) $\sum_{j\in\mathcal{I}}\frac{1}{n_{j}}\mu_{j}(x_{i})\mu_{j}(x_{l^{*}})=\delta_{i,l}h_{i}^{-1}.$
2.2. Normalizable abelian hypergroups and their duals
Note that hypergroups can be rescaled as follows. Let $y_{i}:=\frac{x_{i}}{\alpha_{{}_{i}}}$, for some non-zero complex numbers $\alpha_{{}_{i}}$ with $\alpha_{1}=1$, $\alpha_{{}_{i^{*}}}=\overline{\alpha_{{}_{i}}}$. Let :=$\mathcal{B}^{\prime}:=\{y_{i}\}_{i\in\mathcal{I}}$. Then, $(H,\mathcal{B}^{\prime})$ is also a hypergroup, with coefficients rescaled as $\frac{\alpha_{k}}{\alpha_{i}\alpha_{j}}N_{i,j}^{k}$; and then $h_{i}:=(N_{i,i^{*}}^{1})^{-1}$ rescaled as $|\alpha_{i}|^{2}h_{i}$. Consider $\tau^{\prime}$ defined by $\tau^{\prime}(y_{i}):=\delta_{i,1}$, then $\tau^{\prime}(x_{i})=\alpha_{i}\delta_{i,1}=\delta_{i,1}$, because $\alpha_{1}=1$. Therefore $\tau^{\prime}=\tau$ and the formal codegrees $(n_{j})$ are invariant by rescaling.
Recall the notion of a normalizable hypergroup from Definition 1.5. Note that any rescaled normalizable hypergroup is also normalizable.
Lemma 2.3.
An hypergroup ${(H,\mathcal{B})}$ is normalizable if and only if there is $\mu_{1}\in{\widehat{\mathcal{B}}}$ such that $\mu_{1}(x_{i})$ nonzero, for all $i\in\mathcal{I}$.
Proof.
The rescaling $(H,\{\frac{x_{i}}{\alpha_{{}_{i}}}\})$ is normalized if and only if $\sum_{i\in\mathcal{I}}\frac{\alpha_{k}}{\alpha_{i}\alpha_{j}}N_{i,j}^{k}=1$, if and only if $\alpha_{i}\alpha_{j}=\sum_{i\in\mathcal{I}}N_{i,j}^{k}\alpha_{k}$, if and only if $\mu_{1}:x_{i}\mapsto\alpha_{i}$ is an algebra morphism (i.e. an element of $^{\widehat{\mathcal{B}}}$) with $\mu_{1}(x_{i})$ nonzero, for all $i\in\mathcal{I}$. ∎
The normalizable hypergroup ${(H,\mathcal{B})}$ together with $\mu_{1}$ as in Lemma 2.3 is denoted ${(H,\mathcal{B},\mu_{1})}$. The normalized hypergroup $(H,\{\frac{x_{i}}{\mu_{1}(x_{i})}\})$ is denoted $\overline{{(H,\mathcal{B},\mu_{1})}}$.
Definition 2.4.
Recall from §1.1 that any abelian normalized hypergroup $H$ admits a dual hypergroup $H^{*}$. In the case of $\overline{{(H,\mathcal{B},\mu_{1})}}$, we denote this dual by $^{{\widehat{{(H,\mathcal{B},\mu_{1})}}}}$.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. From Equation (1.2), it follows that the multiplication on the dual $\widehat{{(H,\mathcal{B},\mu_{1})}}$ can be written as
(2.7) $[f\star g]\left(\frac{x_{i}}{\mu_{1}(x_{i})}\right)=f\left(\frac{x_{i}}{\mu_{1}(x_{i})}\right)g\left(\frac{x_{i}}{\mu_{1}(x_{i})}\right)$
for all $f,g\in H^{*}$. Then it is easy to see that the algebra unit of $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is $\mu_{1}$. Moreover, the involution on $^{\widehat{\mathcal{B}}}$, as given in §1.1, becomes:
(2.8) $\mu_{j^{\#}}(\frac{x_{i}}{\mu_{1}(x_{i})})=\mu_{j}(\frac{x_{i^{*}}}{\overline{\mu_{1}(x_{i}})}).$
Definition 2.5.
The order of a normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$ is defined by
$n{(H,\mathcal{B},\mu_{1})}:=\sum_{i\in\mathcal{I}}h_{i}|\mu_{1}(x_{i})|^{2}.$
In the abelian case, the following holds by Equation (2.5):
(2.9) $n{(H,\mathcal{B},\mu_{1})}=n_{1}$
Proposition 2.6.
Let ${(H,\mathcal{B},\mu_{1})}$ be a normalizable hypergroup. Then the order $n{(H,\mathcal{B},\mu_{1})}$ is invariant under rescaling.
Proof.
Let $x^{\prime}_{i}:=\frac{x_{i}}{\alpha_{{}_{i}}}$, for some non-zero complex numbers $\alpha_{{}_{i}}$ with $\alpha_{1}=1$, $\alpha_{{}_{i^{*}}}=\overline{\alpha_{{}_{i}}}$. Consider $(H,\mathcal{B}^{\prime},\mu_{1})$ the rescaled hypergroup with :=$\mathcal{B}^{\prime}:=\{x^{\prime}_{i}\}_{i\in\mathcal{I}}$. Recall that the order of $x^{\prime}_{i}$ is $h^{\prime}_{i}=|\alpha_{{}_{i}}|^{2}h_{i}>0.$ Therefore
$n(H,\mathcal{B}^{\prime},\mu_{1})=\sum_{i\in\mathcal{I}}h^{\prime}_{i}\mid\mu_{1}(x^{\prime}_{i})\mid^{2}=\sum_{i\in\mathcal{I}}h_{i}\mid\mu_{1}(x_{i})\mid^{2}=n{(H,\mathcal{B},\mu_{1})}.\qed$
We shortly write $n(H)$ when $\mathcal{B}$ and $\mu_{1}$ are implicitly understood. In particular, for a normalized hypergroup we may choose $\mu_{1}(x_{i})=1$ for all $1\leq i\leq m$ and therefore $n(H)=\sum_{i\in\mathcal{I}}h_{i}$. For the rest of this paper we use the notation $d_{i}:=\mu_{1}(x_{i})$ for any normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$.
We say that a normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$ is weakly-rational if its order is a rational number, i.e. $n{(H,\mathcal{B},\mu_{1})}\in\mathbb{Q}$. Also, we say that ${(H,\mathcal{B},\mu_{1})}$ is weakly-integral if its order is an integer, i.e. $n{(H,\mathcal{B},\mu_{1})}\in\mathbb{Z}$. Note that a fusion ring is weakly-rational if and only if it is weakly-integral since $\mathrm{FPdim}(R)$ is an algebraic integer in this case.
Remark 2.7.
Note that the formula from Equation (2.1) can be written for abelian normalizable hypergroups as follows (using the fact that $h_{l}$ rescales as $|d_{l}|^{2}h_{l}$):
(2.10) $\widehat{p}_{k}(i,j)=\frac{1}{n_{k}}\bigg{(}\sum_{x_{l}\in{\mathcal{B}}}\frac{h_{l}}{d_{l}}\mu_{i}(x_{l})\mu_{j}(x_{l})\mu_{k}(x_{l^{*}})\bigg{)}$
Lemma 2.8.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. Then with the above notations, $n{(H,\mathcal{B},\mu_{1})}=n{\widehat{{(H,\mathcal{B},\mu_{1})}}}$.
Proof.
By applying (2.10) to $i=j^{\#}$ and $k=1$, and then (2.5), we obtain (see also [Bur21, Equation (2.10)]),
(2.11) $\widehat{p}_{1}(i,i^{\#})=\frac{n_{i}}{n{(H,\mathcal{B},\mu_{1})}}$
Let ${\widehat{h}}_{i}:=\frac{1}{\widehat{p}_{1}(i,i^{\#})}$ be the order of $\mu_{i}$ in $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$. It follows that $n{\widehat{{(H,\mathcal{B},\mu_{1})}}}=\sum_{i\in\mathcal{I}}{\widehat{h}}_{i}=n{(H,\mathcal{B},\mu_{1})}(\sum_{i\in\mathcal{I}}\frac{1}{n_{i}})=n{(H,\mathcal{B},\mu_{1})}$. ∎
The next proposition holds for any normalized hypergroup, not necessarily abelian.
Proposition 2.9.
For any normalized hypergroup ${(H,\mathcal{B})}$,
(2.12) $F_{1}=\frac{1}{n{(H,\mathcal{B})}}\big{(}\sum_{i\in\mathcal{I}}h_{i^{*}}x_{i}\big{)}$
where $F_{1}$ is the idempotent corresponding to the augmentation map $\mu_{1}:H\rightarrow\mathbb{C},\;a\mapsto 1$ for all $a\in\mathcal{B}$.
Proof.
Suppose that $F_{1}=\sum_{i\in\mathcal{I}}\alpha_{{}_{i}}x_{i}$ for some $\alpha_{{}_{i}}\in\mathbb{C}$. Recall that $x_{l}F_{i}=\mu_{i}(x_{l})F_{i}$. Then, $x_{l}F_{1}=F_{1}=\sum_{i\in\mathcal{I}}\alpha_{{}_{i}}x_{l}x_{i}$. It follows that $\tau(F_{1})=\alpha_{1}=\tau(x_{l}F_{1})=\alpha_{l^{*}}h_{l}^{-1}.$ Thus $\alpha_{l^{*}}=h_{l}\alpha_{1}$, so $\alpha_{l}=h_{l^{*}}\alpha_{1}$, $F_{1}=\sum_{i\in\mathcal{I}}\alpha_{i}x_{i}={\alpha_{1}}(\sum_{i\in\mathcal{I}}{h_{i^{*}}}x_{i})$ and $1=\mu_{1}(F_{1})={\alpha_{1}}(\sum_{i\in\mathcal{I}}{h_{i}})$. This implies that $\alpha_{1}=\frac{1}{n(H)}$ and therefore $\alpha_{i}=\frac{h_{i^{*}}}{n(H)}$ for all $i\in\mathcal{I}$. ∎
Corollary 2.10.
In any normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$,
(2.13) $F_{1}=\frac{1}{n{(H,\mathcal{B},\mu_{1})}}\big{(}\sum_{i\in\mathcal{I}}h_{i^{*}}d_{i^{*}}x_{i}\big{)}$
Remark 2.11.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian (rational) normalizable hypergroup. In order to find the central primitive idempotents of $H^{*}$, we can apply Remark 2.2 after normalization. If $x_{i}^{\circ}\in H^{*}$ are defined as above by $\langle x_{i}^{\circ},x_{i^{\prime}}\rangle=\delta_{i,i^{\prime}}$, then $\widehat{F}_{i}:=d_{i}x_{i}^{\circ}\in H^{*}$ are the primitive central idempotents of $H^{*}$ corresponding to the character
$\begin{array}[]{cccc}\widetilde{\omega}_{i}=\operatorname{ev}_{\overline{x}_{i}}:&H^{*}&\to&\mathbb{C}\\ &\mu_{j}&\mapsto&\mu_{j}(\overline{x}_{i}),\end{array}$
where $\overline{x}_{i}=\frac{x_{i}}{d_{i}}$. By identifying $H^{**}$ with $H$, we get that $\operatorname{ev}_{\overline{x}_{i}}=\overline{x}_{i}$.
Recall that $\overline{x}_{1}=x_{1}$.
Proposition 2.12.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. Then $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is an abelian normalized hypergroup whose formal codegrees are
(2.14) $\widehat{n}_{i}=\frac{n{(H,\mathcal{B},\mu_{1})}}{h_{i}|d_{i}|^{2}}.$
Proof.
As already mentioned, [Bur21, Proposition 2.3] shows that $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is an abelian normalized hypergroup.
Let $\overline{x}_{i}=\frac{x_{i}}{d_{i}}$ be the normalization of $x_{i}$. By Remark 2.11, $\overline{x}_{i}:H^{*}\rightarrow\mathbb{C}$ are the characters of $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$. By definition of $\widehat{\tau}\in(H^{*})^{*}=H$, $\widehat{\tau}(\mu_{j})=\delta_{j,1}$ as $\mu_{1}$ is the unit of $H^{*}$. Thus $\widehat{\tau}=F_{1}$. Equation (2.2) for $H^{*}$ gives that $\widehat{\tau}=\sum_{i\in\mathcal{I}}\frac{1}{\widehat{n}_{i}}\overline{x}_{i}$. So, by Equation (2.13),
$\frac{1}{n{(H,\mathcal{B},\mu_{1})}}\big{(}\sum_{i\in\mathcal{I}}h_{i^{*}}d_{i^{*}}x_{i}\big{)}=F_{1}=\widehat{\tau}=\sum_{i\in\mathcal{I}}\frac{1}{\widehat{n}_{i}}\frac{x_{i}}{d_{i}}.$
The result follows by $h_{i^{*}}=(N_{i^{*},i}^{1})^{-1}=(N_{i,i^{*}}^{1})^{-1}=h_{i}$, in the abelian case. ∎
Corollary 2.13.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. Then the primitive central idempotent of the linear character $x_{1}\in\widehat{\widehat{\mathcal{B}}}=\mathcal{B}$ is given by
(2.15) $\widehat{F_{1}}=\sum_{j\in\mathcal{I}}\frac{\mu_{j}}{n_{j}}\in H^{*}.$
Proof.
Proposition 2.12 shows in its proof that $\widehat{\tau}=F_{1}$. By duality, since $\widehat{\widehat{{(H,\mathcal{B},\mu_{1})}}}=\overline{{(H,\mathcal{B},\mu_{1})}}$, it follows that $\tau=\widehat{F}_{1}$. Equation (2.2) finishes the proof. ∎
A morphism of normalizable hypergroups $\psi:(H,{\mathcal{B}},\mu_{1})\rightarrow(H^{\prime},{\mathcal{B}}^{\prime},\mu_{1}^{\prime})$ is a morphism of hypergroups with the property that $\mu^{\prime}_{1}\circ\psi=\mu_{1}$.
2.3. Frobenius-Perron theory for a RN-hypergroups
Frobenius-Perron theory can be defined for RN-hypergroups, in the same manner as for fusion rings, see [EGNO15, §3]. We denote by $\mathrm{FPdim}(x_{i})$ the Frobenius-Perron eigenvalue of the left multiplication operator by $x_{i}$ on $H$, i.e. the ($\ell^{2}$) matrix norm of $N_{i}=(N_{i,j}^{k})_{j,k}$. Recall that a hypergroup is called symmetric when $N_{a,b}^{1}=N_{b,a}^{1}$, for all $a,b\in\mathcal{B}$, but $N_{a,b}^{1}=\delta_{a^{*},b}h_{a}^{-1}$, so symmetric means that $h_{a}=h_{a^{*}}$, for all $a\in\mathcal{B}$
Lemma 2.14.
An abelian hypergroup is symmetric.
Proof.
By definition and abelian assumption, $h_{i}^{-1}:=N_{i,i^{*}}^{1}=N_{i^{*},i}^{1}=h_{i^{*}}^{-1}$. ∎
Lemma 2.15.
If $N_{i,j}^{k}$ is nonzero then
(2.16) $\frac{h_{i^{*}}h_{j^{*}}h_{k}}{h_{i}h_{j}h_{k^{*}}}=1.$
Proof.
The associativity of $H$ reformulates as $\sum_{s}N_{i,j}^{s}N_{s,k}^{t}=\sum_{s}N_{j,k}^{s}N_{i,s}^{t}$, for all $i,j,k,t$. If $t=1$, we get that $\sum_{s}N_{i,j}^{s}N_{s,k}^{1}=\sum_{s}N_{j,k}^{s}N_{i,s}^{1}$. But recall that $N_{a,b}^{1}=\delta_{a^{*},b}h_{a}^{-1}$. It follows that $N_{i,j}^{k^{*}}h_{k^{*}}^{-1}=N_{j,k}^{i^{*}}h_{i}^{-1}$, in other words, $N_{i,j}^{k}=\frac{h_{k}}{h_{i}}N_{j,k^{*}}^{i^{*}}$. By applying this last equality, we get that $N_{j,k^{*}}^{i^{*}}=\frac{h_{i^{*}}}{h_{j}}N_{k^{*},i}^{j^{*}}$, and $N_{k^{*},i}^{j^{*}}=\frac{h_{j^{*}}}{h_{k^{*}}}N_{i,j}^{k}$. Thus
(2.17) $N_{i,j}^{k}=\frac{h_{k}}{h_{i}}N_{j,k^{*}}^{i^{*}}=\frac{h_{i^{*}}h_{k}}{h_{i}h_{j}}N_{k^{*},i}^{j^{*}}=\frac{h_{i^{*}}h_{j^{*}}h_{k}}{h_{i}h_{j}h_{k^{*}}}N_{i,j}^{k}$
The result follows. ∎
Theorem 2.16.
A (finite-dimensional) hypergroup is symmetric, i.e. $h_{i^{*}}=h_{i}$, for all $i\in\mathcal{I}$. In particular, the function $\tau$ from §2.1 is symmetric.
Proof.
Introduce the notation $r_{i}:=\frac{h_{i^{*}}}{h_{i}}$. Lemma 2.15 reformulates as:
$N_{i,j}^{k}\neq 0\Rightarrow r_{i}r_{j}=r_{k}.$
We aim to demonstrate that $r_{i}$ is a root of unity for every $i\in\mathcal{I}$. First, select $k_{2}$ such that $N_{i,i}^{k_{2}}$ is nonzero, which implies $r_{i}^{2}=r_{k_{2}}$. Next, choose $k_{3}$ such that $N_{i,k_{2}}^{k_{3}}$ is nonzero, leading to $r_{i}r_{k_{2}}=r_{k_{3}}$ and thus $r_{i}^{3}=r_{k_{3}}$. Continuing this process, we find that the sequence $\{r_{i}^{n}\}$ for $n\in\mathbb{N}$ matches the set $\{r_{j}\}$ for $j\in\mathcal{J}$, with $\mathcal{J}$ being a subset of $\mathcal{I}$. Due to the finite dimensionality of the hypergroup, $\mathcal{I}$ is finite, which implies that $\mathcal{J}$ is also finite. Consequently, there exist integers $m$ and $n$ with $m>n$ such that $r_{i}^{m}=r_{i}^{n}$. Thus, $r_{i}^{s}=1$ where $s=m-n>0$, confirming that $r_{i}$ is a root of unity. Additionally, the axioms of the hypergroup state that $N_{a,a^{*}}^{1}$ is positive for all $a\in\mathcal{B}$; hence $h_{i}$ is positive for all $i\in\mathcal{I}$. Therefore, $r_{i}=\frac{h_{i^{*}}}{h_{i}}$ is a positive root of unity, which must be 111. ∎
In the rest of the paper, we assume that the involution on $\mathcal{B}$ extends into an anti-involution on $H$, i.e. $N_{i,j}^{k}=N_{j^{*},i^{*}}^{k^{*}}$, for all $i,j,k$.
Proposition 2.17 (Frobenius Reciprocity).
The following equalities hold:
(2.18) $h_{k}^{-1}N_{i,j}^{k}=h_{j}^{-1}N_{i^{*},k}^{j}=h_{i}^{-1}N_{j,k^{*}}^{i^{*}}=h_{k}^{-1}N_{j^{*},i^{*}}^{k^{*}}=h_{j}^{-1}N_{k^{*},i}^{j^{*}}=h_{i}^{-1}N_{k,j^{*}}^{i}.$
Proof.
By (2.17), $N_{i,j}^{k}=\frac{h_{i^{*}}h_{k}}{h_{i}h_{j}}N_{k^{*},i}^{j^{*}}$, then by anti-involution, $N_{k^{*},i}^{j^{*}}=N_{i^{*},k}^{j}$, and by (2.16), $\frac{h_{i^{*}}h_{k}}{h_{i}h_{j}}=\frac{h_{k^{*}}}{h_{j^{*}}}$. But $h_{k^{*}}=h_{k}$ and $h_{j^{*}}=h_{j}$ by Theorem 2.16. The first equality follows. The rest is similar. ∎
Corollary 2.18.
For every $x_{i}\in\mathcal{B}$ then $\mathrm{FPdim}(x_{i})=\mathrm{FPdim}(x_{i^{*}})$.
Proof.
Let $N_{i}$ be the matrix $(N_{i,j}^{k})_{j,k}$. By the first equality of (2.18), $N_{i,j}^{k}=\frac{h_{k}}{h_{j}}N_{i^{*},k}^{j}$, so $N_{i}=DN_{i^{*}}^{T}D^{-1}$, with $D={\rm diag}(h_{i})$ and $()^{T}$ the matrix transpose. Thus
$\mathrm{FPdim}(x_{i^{*}})=\|N_{i^{*}}\|=\|N_{i^{*}}^{T}\|=\|D^{-1}N_{i}D\|=\|N_{i}\|=\mathrm{FPdim}(x_{i}).$
The second last equality hold because $D^{-1}N_{i}D$ have the same eigenvalues than $N_{i}$, because if $N_{i}v=\lambda v$ then $D^{-1}N_{i}Dw=\lambda w$, with $w=D^{-1}v$. ∎
Since we are dealing with the finite-dimensional case, by Theorem 2.16, every hypergroup is symmetric. Therefore, we can omit this assumption in the rest of the paper, as it is automatically satisfied.
Remark 2.19.
Note that any hypergroup ${(H,\mathcal{B})}$ is transitive in the sense of [EGNO15, Definition 3.3.1], i.e. $\forall i,j\in\mathcal{I}$ there are $k_{1},k_{2}\in\mathcal{I}$ such that $N_{i,k_{1}}^{j}$ and $N_{k_{2},i}^{j}$ are nonzero. Indeed, take $k_{1},k_{2}$ such that $N_{i^{*},j}^{k_{1}}$ and $N_{j,i^{*}}^{k_{2}}$ are nonzero, the result follows by Frobenius reciprocity (2.18).
Observe that an analogue of [EGNO15, Proposition 3.3.6] holds for RN-hypergroups as it requires Frobenius-Perron [EGNO15, Theorem 3.2.1] involving a RN-matrix (i.e. with real nonnegative entries). In particular, $\mathrm{FPdim}:H\rightarrow\mathbb{C}$ is the unique algebra morphism which takes positive values on $\mathcal{B}$.
Definition 2.20.
The Frobenius-Perron dimension of a RN-hypergroup ${(H,\mathcal{B})}$ is
$\mathrm{FPdim}{(H,\mathcal{B})}:=\sum_{i\in\mathcal{I}}h_{i}\mathrm{FPdim}(x_{i})^{2}.$
By Lemma 2.3 and Remark 2.19, any RN-hypergroup is normalizable via the linear character $\mu_{1}=\mathrm{FPdim}$. By Definition 2.5 and positivity of $\mathrm{FPdim}$,
$\mathrm{FPdim}{(H,\mathcal{B})}=n{(H,\mathcal{B},\mathrm{FPdim})}.$
Moreover, in the normalized case, by uniqueness in Remark 2.19, $\mathrm{FPdim}$ must be the augmentation map, i.e. $\mathrm{FPdim}(x_{i})=1$ for all $i\in\mathcal{I}$.
Recall from Proposition 2.6 that for a normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$, the order $n{(H,\mathcal{B},\mu_{1})}$ is invariant under rescaling.
Proposition 2.21.
If ${(H,\mathcal{B})}$ is a RN-hypergroup, then the Frobenius-Perron dimension of ${(H,\mathcal{B})}$ is invariant under rescaling with real positive numbers.
Proof.
If ${(H,\mathcal{B})}$ is a RN-hypergroup and $\alpha_{{}_{i}}=\alpha_{{}_{i^{*}}}>0$, for all $i\in\mathcal{I}$, then $(H^{\prime},\mathcal{B}^{\prime})=(H,\{\frac{x_{i}}{\alpha_{i}}\})$ is RN and
$\mathrm{FPdim}(H^{\prime},\mathcal{B}^{\prime})$ $=$ $\sum_{i\in\mathcal{I}}h^{\prime}_{i}\mathrm{FPdim}(\frac{x_{i}}{\alpha_{i}})^{2}=$
$=$ $\sum_{i\in\mathcal{I}}|\alpha_{{}_{i}}|^{2}h_{i}\frac{\mathrm{FPdim}(x_{i})^{2}}{\alpha_{{}_{i}}^{2}}$
$=$ $\mathrm{FPdim}{(H,\mathcal{B})}.\qed$
2.4. Fourier transform
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. Define
(2.19) $\mathcal{F}:{(H,\mathcal{B})}\rightarrow{\widehat{(H,\mathcal{B})}},\;x_{i}\mapsto\frac{n{(H,\mathcal{B},\mu_{1})}}{h_{i^{*}}}x_{i^{*}}^{\circ}.$
It is clear that $\mathcal{F}$ is a linear isomorphism. Additionally, as noted in Remark 2.11, we have $\widehat{F}_{i}:=d_{i}x_{i}^{\circ}$. Therefore:
(2.20) $\mathcal{F}(x_{i})=\frac{n{(H,\mathcal{B},\mu_{1})}}{d_{i^{*}}h_{i^{*}}}\widehat{F}_{i^{*}}.$
For the sake of brevity, we define $|H|:=n{(H,\mathcal{B},\mu_{1})}$.
Lemma 2.22.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. With the above notations, for all $x,y\in H$:
$\langle\mathcal{F}(y),x\rangle=|H|\tau(xy).$
Proof.
It suffices to verify the above identity for $(x,y)=(x_{i},x_{j})$, for all $i,j$. We have:
[…]
since $h_{i}=h_{i^{*}}$ (see Lemma 2.14), and $\delta_{i,j^{*}}h_{i}^{-1}=\tau(x_{i}x_{j})$, because $h_{i}^{-1}=N_{i,i^{*}}^{1}$ and $\tau=x_{1}^{\circ}$. ∎
Proposition 2.23.
Using the above notations, we have that
(2.21) $\mathcal{F}(F_{k})=\frac{|H|}{n_{k}}\mu_{k}.$
Proof.
By taking $(x,y)=(F_{j},F_{k})$ in Lemma 2.22, we obtain:
$\langle\mathcal{F}(F_{k}),F_{j}\rangle=|H|\tau(F_{j}F_{k})=\delta_{j,k}|H|\tau(F_{j}).$
From Equation (2.4), we have:
$\tau(F_{j})=\frac{1}{n_{j}}\big{(}\sum_{i\in\mathcal{I}}h_{i}\mu_{j}(x_{i^{*}})\tau(x_{i})\big{)}=\frac{1}{n_{j}}.$
Thus
$\langle\mathcal{F}(F_{k}),F_{j}\rangle=\delta_{j,k}\frac{|H|}{n_{j}},$
which shows that $\mathcal{F}(F_{k})=\frac{|H|}{n_{k}}\mu_{k}$, since $\mu_{i}(F_{j})=\delta_{i,j}$. ∎
2.5. Grouplike elements in hypergroups
In the rest of this paper, all abelian RN-hypergroups are normalized by $\mu_{1}=\mathrm{FPdim}$, and therefore their duals are also considered with respect to $\mu_{1}=\mathrm{FPdim}$. Recall that $d_{i}:=\mu_{1}(x_{i})=\mathrm{FPdim}(x_{i})$.
Lemma 2.24.
For any RN-hypergroup ${(H,\mathcal{B})}$, $h_{i}d_{i}^{2}\geq 1$. Moreover, $h_{i}d_{i}^{2}=1$ if and only $h_{i}x_{i}x_{i^{*}}=x_{1}.$
Proof.
The following equality holds
(2.22) $x_{i}x_{i^{*}}=\frac{1}{h_{i}}x_{1}+\sum_{k=2}^{m}N^{k}_{ii^{*}}x_{k}.$
Passing to $\mathrm{FPdim}$, $h_{i}d_{i}^{2}=1+h_{i}\big{(}\sum_{k=2}^{m}N^{k}_{ii^{*}}d_{k}\big{)}\geq 1.$ If $h_{i}d_{i}^{2}=1$ then $\sum_{k=2}^{m}N^{k}_{ii^{*}}d_{k}=0$, and therefore $N^{k}_{ii^{*}}=0$ for all $k\neq 1$. Thus $h_{i}x_{i}x_{i^{*}}=x_{1}.$ Conversely, if $h_{i}x_{i}x_{i^{*}}=x_{1}$, applying $\mathrm{FPdim}$, it follows that $h_{i}d_{i}^{2}=1$. ∎
Corollary 2.25.
In any fusion ring ${(H,\mathcal{B})}$, $x_{i}x_{i^{*}}=x_{1}$ if and only if $d_{i}=1$.
Recall the set of grouplike elements ${G{(H,\mathcal{B})}}$ from Definition 1.3.
Remark 2.26.
For a normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$, we establish that $x_{i}\in G{(H,\mathcal{B})}$ if and only if:
(2.23) $\frac{x_{i}x_{i^{*}}}{d_{i}d_{i^{*}}}=x_{1}.$
Applying $\mu_{1}$ to Equation (1.1), we find that $h_{i}d_{i}d_{i^{*}}=1$, from which Equation (2.23) follows. The converse is similar.
Lemma 2.27.
Suppose that ${(H,\mathcal{B},\mu_{1})}$ is an abelian normalizable hypergroup. Then $x_{i}\in G{(H,\mathcal{B})}$ if and only if
(2.24) $|\mu_{j}({\frac{x_{i}}{d_{i}}})|=1\;\;\text{for all}\;\;j\in\mathcal{I}.$
Proof.
If $x_{i}\in G{(H,\mathcal{B})}$ then applying $\mu_{j}$ to Equation (2.23) results in $|\mu_{j}({\frac{x_{i}}{d_{i}}})|^{2}=1$, since $\mu_{j}(x_{i^{*}})=\overline{\mu_{j}(x_{i})}$. Conversely, if $\mu_{j}({\frac{x_{i}}{d_{i}}}\frac{x_{i^{*}}}{d_{i^{*}}})=1$ for all $\mu_{j}$, then ${\frac{x_{i}}{d_{i}}}\frac{x_{i^{*}}}{d_{i^{*}}}=x_{1}$, since ${(H,\mathcal{B})}$ is abelian, and therefore $x_{i}\in G{(H,\mathcal{B})}$ by Remark 2.26. ∎
Dually we have the following:
Lemma 2.28.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. Then $\mu_{j}\in G{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ if and only if
(2.25) $|\mu_{j}({\frac{x_{i}}{d_{i}}})|=1\;\;\text{for all}\;\;i\in\mathcal{I}.$
Proof.
The dual version of Equation (2.23) implies that $\mu_{j}\in G{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ if and only if $\mu_{j}\mu_{j^{\#}}=\mu_{1}.$ Evaluating at ${\frac{x_{i}}{d_{i}}}$ results in the desired identity. ∎
Corollary 2.29.
In any abelian normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$, for any $\mu_{j}\in G{\widehat{{(H,\mathcal{B},\mu_{1})}}}$, it holds that $n_{j}=n(H,B,\mu_{1})$.
Proof.
By Lemma 2.28, we find that $\mid\mu_{j}(x_{i})|=|d_{i}|$ for all $i\in\mathcal{I}$. From Equation (2.5), we deduce that $n_{j}=\sum_{i\in\mathcal{I}}h_{i}|\mu_{j}(x_{i})|^{2}=\sum_{i\in\mathcal{I}}h_{i}|d_{i}|^{2}=n(H,B,\mu_{1})$. ∎
Lemma 2.30.
Suppose that ${(H,\mathcal{B},\mu_{1})}$ is an abelian normalizable hypergroup such that one of the following holds:
  1. (1)
    ${(H,\mathcal{B})}$ is RN and $\mu_{1}=\mathrm{FPdim}$,
  2. (2)
    $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN.
Then
(2.26) $|\mu_{j}(x_{i})|\leq|\mu_{1}(x_{i})|=|d_{i}|\;\;\text{, for all}\;i,j\in\mathcal{I}.$
Proof.
Suppose first that ${(H,\mathcal{B})}$ is RN and $\mu_{1}=\mathrm{FPdim}$. Thus $d_{i}={\mathrm{FPdim}(x_{i})}>0$. But $x_{i}=\sum_{j\in\mathcal{I}}\mu_{j}(x_{i})F_{j}$, and therefore $x_{i}F_{j}=\mu_{j}(x_{i})F_{j}$ meaning that $\mu_{j}(x_{i})$ is an eigenvalue of $L_{x_{i}}$. It follows that $|\mu_{j}(x_{i})|\leq d_{i}$, by definition of $\mathrm{FPdim}$.
Now, suppose that $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN. It follows that $\mathrm{FPdim}(\mu_{j})=1$, since by Remark 2.19, in this case, $\mu_{j}\mapsto 1$ is the only algebra morphism on $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ which takes positive values on $^{\widehat{\mathcal{B}}}$. Then, the equation
(2.27) $\mu_{j}=\sum_{i\in\mathcal{I}}\mu_{j}({\frac{x_{i}}{d_{i}}})\widehat{F}_{i}$
implies as above that $\big{|}\mu_{j}({\frac{x_{i}}{d_{i}}})\big{|}\leq 1$. ∎
Lemma 2.31.
Under the same assumptions as Lemma 2.30, we have $n_{j}\leq n{(H,\mathcal{B},\mu_{1})}$, for all $j\in\mathcal{I}$. Moreover, $n_{j}=n{(H,\mathcal{B},\mu_{1})}$ if and only if $\mu_{j}$ is a grouplike element.
Proof.
Lemma 2.30 and Equation (2.5) give:
$n_{j}=\sum_{i\in\mathcal{I}}h_{i}|\mu_{j}(x_{i})|^{2}\leq\sum_{i\in\mathcal{I}}h_{i}|d_{i}|^{2}=n(H,B,\mu_{1}),$
and the equality holds if and only if $|\mu_{j}(x_{i})|^{2}=|d_{i}|^{2}$, if and only if $\mu_{j}\in G{\widehat{{(H,\mathcal{B},\mu_{1})}}}$, by Lemma 2.28. ∎
Lemma 2.32.
In any RN-hypergroup ${(H,\mathcal{B})}$, the set $\{{\frac{x_{i}}{d_{i}}}\}$ of normalized grouplike-elements forms a finite group. Moreover, the inverse of ${\frac{x_{i}}{d_{i}}}$ is given by $\frac{x_{i^{*}}}{d_{i^{*}}}$ for all $x_{i}\in{G{(H,\mathcal{B})}}$.
Proof.
If $x_{i}x_{i^{*}}=\frac{1}{h_{i}}x_{1}$ and $x_{j}x_{j^{*}}=\frac{1}{h_{j}}x_{1}$ then $(x_{i}x_{j})(x_{i}x_{j})^{*}=\frac{1}{h_{i}h_{j}}x_{1}.$ Thus $(\sum_{k=1}^{m}N^{k}_{ij}x_{k})(\sum_{k=1}^{m}N^{k}_{ij}x_{k^{*}})=\frac{1}{h_{i}h_{j}}x_{1},$ so that there is a unique $k\in\mathcal{I}$ such that $N_{i,j}^{k}$ is nonzero (if it were not the case, then by the axioms of hypergroups coupled with the RN assumption, the left-hand side of the aforementioned equation would decompose into several components). Consequently, we have $(N^{k}_{ij}x_{k})(N^{k}_{ij}x_{k^{*}})=\frac{1}{h_{i}h_{j}}x_{1}$, and furthermore, $x_{i}x_{j}=N^{k}_{ij}x_{k}$. However, applying $\mathrm{FPdim}$ yields $N^{k}_{ij}=(d_{i}d_{j})/d_{k}$, so $\frac{x_{i}}{d_{i}}\frac{x_{j}}{d_{j}}=\frac{x_{k}}{d_{k}}$, and by the definition of a grouplike element, we have $h_{i}^{-1}=d_{i}^{2}$. Synthesizing all these equalities, we deduce that $((d_{i}d_{j})/d_{k})^{2}x_{k}x_{k^{*}}=(d_{i}d_{j})^{2}x_{1}$, leading to the conclusion that $x_{k}x_{k^{*}}=d_{k}^{2}x_{1}=h_{k}^{-1}x_{1}$. There remain to prove that if $x_{i}\in{G{(H,\mathcal{B})}}$ then so is $x_{i^{*}}$. By Lemma 2.24, $x_{i^{*}}$ is grouplike if and only if $h_{i^{*}}d_{i^{*}}^{2}=1$. But $h_{i^{*}}=h_{i}$ by Theorem 2.16, and $d_{i^{*}}=d_{i}$ by Corollary 2.18, so $h_{i^{*}}d_{i^{*}}^{2}=h_{i}d_{i}^{2}=1$, because $x_{i}$ is grouplike. Finally, by above, there is $k$ such that $\frac{x_{i}}{d_{i}}\frac{x_{i^{*}}}{d_{i^{*}}}=\frac{x_{k}}{d_{k}}$, and by the hypergroup axioms, $k$ must be 111. ∎
We denote $\overline{G{(H,\mathcal{B})}}:=\{{\frac{x_{i}}{d_{i}}}\;|\;x_{i}\in G{(H,\mathcal{B})}\}$ the above group of normalized grouplike elements. Clearly $\overline{G{(H,\mathcal{B})}}=G\overline{{(H,\mathcal{B})}}$, the grouplike elements of the normalized hypergroup $\overline{{(H,\mathcal{B})}}$.
By duality we obtain the following:
Lemma 2.33.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup such that its dual $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN. Then $^{G{\widehat{(H,\mathcal{B})}}}$ is a group.
Recall that ¯=$\overline{\mathcal{B}}=\{\frac{x_{k}}{d_{k}}\;|\;x_{k}\in{\mathcal{B}}\}$.
Lemma 2.34.
Let ${(H,\mathcal{B})}$ be a RN-hypergroup. Let $x_{i}\in{G{(H,\mathcal{B})}}$, then ${\frac{x_{i}}{d_{i}}}\frac{x_{j}}{d_{j}},\frac{x_{j}}{d_{j}}{\frac{x_{i}}{d_{i}}}\in\overline{\mathcal{B}}$, for any $x_{j}\in\mathcal{B}$.
Proof.
Suppose that
$x_{i}x_{j}=\sum_{k\in\mathcal{A}}N^{k}_{ij}x_{k},$
where $\mathcal{A}$ is a set such that $N^{k}_{ij}>0$. Multiplying the above equality by $x_{i^{*}}$, we obtain
$\frac{1}{h_{i}}x_{j}=\sum_{k\in\mathcal{A}}N^{k}_{ij}x_{i^{*}}x_{k},$
since $h_{i^{*}}x_{i^{*}}x_{i}=x_{1}$ as $x_{i^{*}}$ is also grouplike by Lemma 2.32, and $h_{i^{*}}=h_{i}$ by Theorem 2.16. By RN assumption, $x_{i^{*}}x_{k}=\alpha_{ik}x_{j}$, for some scalar $\alpha_{ik}>0$ and all $k\in\mathcal{A}$. Thus $\alpha_{ik}=\frac{d_{i}d_{k}}{d_{j}}$. Multiplying the second last equality by $x_{i}$, we get that $h_{i}^{-1}x_{k}=\alpha_{ik}x_{i}x_{j}$, which reformulates as ${\frac{x_{i}}{d_{i}}}\frac{x_{j}}{d_{j}}=\frac{x_{k}}{d_{k}}\in\overline{\mathcal{B}}$. Idem for $\frac{x_{j}}{d_{j}}{\frac{x_{i}}{d_{i}}}$ using right multiplications. ∎
Lemma 2.35.
Suppose that ${(H,\mathcal{B})}$ is an abelian RN hypergroup and $\mu_{j}\in\widehat{\mathcal{B}}$ a character such that $\mu_{j}(x_{m})=d_{m}\omega_{m}$ and $\mu_{j}(x_{n})=d_{n}\omega_{n}$ for some scalars $|\omega_{m}|=|\omega_{n}|=1$. Then $\mu_{j}(x_{p})=\omega_{m}\omega_{n}d_{p}$ for all constituents $x_{p}$ of $x_{m}x_{n}$.
Proof.
Apply the absolute value triangle inequality and Lemma 2.30. ∎
Lemma 2.36.
If ${(H,\mathcal{B})}$ is an abelian RN hypergroup then $^{{G{\widehat{(H,\mathcal{B})}}}}$ is a group.
Proof.
Suppose that $\mu_{j},\mu_{k}\in{{G{\widehat{(H,\mathcal{B})}}}}$. By Lemma 2.28, $\mu_{j}(x_{m})=\omega_{m}d_{m}$ and $\mu_{k}(x_{m})=\eta_{m}d_{m}$, for some roots of unity $\omega_{m}$ and $\eta_{m}$. Now,
$(\mu_{j}\star\mu_{k})(x_{m}x_{n})=\sum_{p}N^{p}_{m,n}d_{p}(\mu_{j}\star\mu_{k})({\frac{x_{p}}{d_{p}}}),$
but by Equation (2.7),
$(\mu_{j}\star\mu_{k})({\frac{x_{p}}{d_{p}}})=\mu_{j}({\frac{x_{p}}{d_{p}}})\mu_{k}({\frac{x_{p}}{d_{p}}}),$
and by Lemma 2.35, $\mu_{j}(x_{p})=\omega_{m}\omega_{n}d_{p}$ and $\mu_{k}(x_{p})=\eta_{m}\eta_{n}d_{p}$, when $N^{p}_{m,n}$ is nonzero. Thus
$(\mu_{j}\star\mu_{k})(x_{m}x_{n})=\sum_{p}N^{p}_{m,n}d_{p}\omega_{m}\omega_{n}\eta_{m}\eta_{n}=\omega_{m}\omega_{n}\eta_{m}\eta_{n}d_{m}d_{n}$
On the other hand,
$(\mu_{j}\star\mu_{k})(x_{m})=d_{m}(\mu_{j}\star\mu_{k})({\frac{x_{m}}{d_{m}}})=d_{m}\mu_{j}({\frac{x_{m}}{d_{m}}})\mu_{k}({\frac{x_{m}}{d_{m}}})=d_{m}\omega_{m}\eta_{m}.$
Idem, $(\mu_{j}\star\mu_{k})(x_{n})=d_{n}\omega_{n}\eta_{n}$. So
$(\mu_{j}\star\mu_{k})(x_{m})(\mu_{j}\star\mu_{k})(x_{n})=d_{m}\omega_{m}\eta_{m}d_{n}\omega_{n}\eta_{n}=(\mu_{j}\star\mu_{k})(x_{m}x_{n}),$
by above, which shows that $(\mu_{j}\star\mu_{k})\in{\widehat{\mathcal{B}}}$.
Moreover, $(\mu_{j}\star\mu_{k})({\frac{x_{p}}{d_{p}}})=\mu_{j}({\frac{x_{p}}{d_{p}}})\mu_{k}({\frac{x_{p}}{d_{p}}})=\omega_{p}\eta_{p}$, which by Lemma 2.28, implies that $\mu_{j}\star\mu_{k}\in G{\widehat{(H,\mathcal{B})}}$. ∎
Proposition 2.37.
Let ${(H,\mathcal{B})}$ be an abelian RN hypergroup and $\mu\in{{G{\widehat{(H,\mathcal{B})}}}}$. Then, $\mu\star\mu_{k}\in\widehat{B}$, for any $\mu_{k}\in\widehat{B}$.
Proof.
Since $\mu\in{{G{\widehat{(H,\mathcal{B})}}}}$ and ${(H,\mathcal{B})}$ is RN, by Lemma 2.28, $\mu(x_{i})=\omega_{i}d_{i}$ for some root of unity $\omega_{i}$ and for all $i\in\mathcal{I}$. By Lemma 2.35, $\mu(x_{l})=\omega_{i}\omega_{j}d_{l}$ for any constituent $x_{l}$ of $x_{i}x_{j}$, therefore
$(\mu\star\mu_{k})(x_{i}x_{j})=\sum_{l=1}^{m}N^{l}_{ij}(\mu\star\mu_{k})(x_{l})=\omega_{i}\omega_{j}\sum_{l=1}^{m}N^{l}_{ij}\mu_{k}(x_{l})=\omega_{i}\omega_{j}\mu_{k}(x_{i}x_{j}).$
On the other hand
$(\mu\star\mu_{k})(x_{i})(\mu\star\mu_{k})(x_{j})=\omega_{i}\mu_{k}(x_{i})\omega_{j}\mu_{k}(x_{j})=\omega_{i}\omega_{j}\mu_{k}(x_{i}x_{j}).$
Thus $(\mu\star\mu_{k})(x_{i}x_{j})=(\mu\star\mu_{k})(x_{i})(\mu\star\mu_{k})(x_{j})$ which shows that $\mu\star\mu_{k}\in{\widehat{\mathcal{B}}}$. ∎
By duality, Lemma 2.36 and Proposition 2.37 reformulate as follows:
Proposition 2.38.
If ${(H,\mathcal{B},\mu_{1})}$ is an abelian normalizable hypergroup such that $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN, then $\overline{G{(H,\mathcal{B})}}$ is a group. Moreover, the left multiplication $L_{{\frac{x_{i}}{d_{i}}}}$, for any $x_{i}\in{G{(H,\mathcal{B})}}$, permutes the normalized basis $\overline{\mathcal{B}}$.
Lemmas 2.32, 2.33, 2.34, 2.36 and Propositions 2.37, 2.38 imply the following:
Corollary 2.39.
Suppose that ${(H,\mathcal{B},\mu_{1})}$ is an abelian normalizable hypergroup such that one of the following holds:
  1. (1)
    ${(H,\mathcal{B})}$ is RN and $\mu_{1}=\mathrm{FPdim}$
  2. (2)
    $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN.
Then, the following statements hold:
  1. (1)
    Both $\overline{G{(H,\mathcal{B})}}$ and $G{\widehat{(H,\mathcal{B})}}$ are finite groups.
  2. (2)
    The left multiplication operator $L_{\frac{x_{i}}{d_{i}}}$, with $x_{i}\in G{(H,\mathcal{B})}$, permutes the normalized basis $\overline{\mathcal{B}}$.
  3. (3)
    The left multiplication operator $L_{\mu_{i}}$, with $\mu_{i}\in G{\widehat{(H,\mathcal{B})}}$, permutes the basis $^{{\widehat{\mathcal{B}}}}$.
2.6. Some results on $P$ and $^{\widehat{P}}$.
They denote $\prod_{i\in\mathcal{I}}{\frac{x_{i}}{d_{i}}}$ and $\prod_{j\in\mathcal{I}}\mu_{j}$, respectively.
Lemma 2.40.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup, and let $x_{i}\in G{(H,\mathcal{B})}$ be a grouplike element. Then
(2.28) $(\prod_{j\in\mathcal{I}}\mu_{j})({\frac{x_{i}}{d_{i}}})=\pm 1.$
Proof.
By Equation (2.7), defining the multiplication in the dual,
$(\prod_{j\in\mathcal{I}}\mu_{j})({\frac{x_{i}}{d_{i}}})=\prod_{j\in\mathcal{I}}\mu_{j}({\frac{x_{i}}{d_{i}}}).$
By Equation (2.8), defining the involution in the dual, and then Lemma 2.27,
$\mu_{j}({\frac{x_{i}}{d_{i}}})\mu_{j^{\#}}({\frac{x_{i}}{d_{i}}})=|\mu_{j}({\frac{x_{i}}{d_{i}}})|^{2}=1.$
We group together the factors $\mu_{j}$ and $\mu_{j\#}$ in the product above. Now, if $\mu_{j}=\mu_{j^{\#}}$, then $\mu_{j}({\frac{x_{i}}{d_{i}}})\in\mathbb{R}$, and therefore, $\mu_{j}({\frac{x_{i}}{d_{i}}})=\pm 1$ by Lemma 2.27. ∎
Definition 2.41.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup, and let $x_{i}\in{G{(H,\mathcal{B})}}$ be a grouplike element. Then $(\prod_{j\in\mathcal{I}}\mu_{j})({\frac{x_{i}}{d_{i}}})$ will be denoted ${\mathrm{sgn}}(x_{i})$.
The dual version of the above result is the following:
Lemma 2.42.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup, and let $\mu_{j}\in G{\widehat{(H,\mathcal{B})}}$ be a grouplike element. Then
(2.29) $\mu_{j}(\prod_{i\in\mathcal{I}}{\frac{x_{i}}{d_{i}}})=\pm 1.$
By duality, we can also define:
Definition 2.43.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup, and let $\mu_{j}\in G{\widehat{(H,\mathcal{B})}}$ be a grouplike element. Then $\mu_{j}(\prod_{i\in\mathcal{I}}{\frac{x_{i}}{d_{i}}})$ will be denoted ${\mathrm{sgn}}(\mu_{j})$.
Proposition 2.44.
Following the assumptions of Corollary 2.39, let $x_{i}\in{G{(H,\mathcal{B})}}$ and $\mu_{j}\in G{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ be grouplike elements. Then
${\mathrm{sgn}}(x_{i})=\det(L_{\frac{x_{i}}{d_{i}}})\text{ and }{\mathrm{sgn}}(\mu_{j})=\det(L_{\mu_{j}})$
are the signature of the permutations of $\overline{\mathcal{B}}$ by $L_{{\frac{x_{i}}{d_{i}}}}$, and of $\widehat{\mathcal{B}}$ by $L_{\mu_{j}}$, respectively.
Proof.
By Corollary 2.39, the left multiplication operator $L_{{\frac{x_{i}}{d_{i}}}}$ permutes $\overline{\mathcal{B}}$, but the signature of this permutation is the determinant $\det(L_{\frac{x_{i}}{d_{i}}})$. Now ${\frac{x_{i}}{d_{i}}}=\sum_{j\in\mathcal{I}}\mu_{j}({\frac{x_{i}}{d_{i}}})F_{j}$, so $\det(L_{\frac{x_{i}}{d_{i}}})=\prod_{j\in\mathcal{I}}\mu_{j}({\frac{x_{i}}{d_{i}}})$. Idem for ${\mathrm{sgn}}(\mu_{j})$. ∎
Definition 2.45.
A hypergroup ${(H,\mathcal{B})}$ is called perfect if it has no nontrivial grouplike elements.
3. Galois symmetries for hypergroups
Let $(H,\mathcal{B})$ be an abelian rational hypergroup (i.e. $N^{k}_{ij}\in\mathbb{Q}$). For any standard element $x_{i}\in\mathcal{B}$, we can write $x_{i}=\sum_{j\in\mathcal{I}}{\alpha_{{}_{ij}}}F_{j}$, with ${\alpha_{{}_{ij}}}=\mu_{j}(x_{i})\in\mathbb{C}$. Let $\mathbb{K}=\mathbb{Q}({\alpha_{{}_{ij}}})$ be the field obtained by adjoining all ${\alpha_{{}_{ij}}}$ to $\mathbb{Q}$.
Remark 3.1.
Note that if ${\alpha_{{}_{ij}}}=\alpha_{{}_{ij^{\prime}}}$ for all $i$, then $\mu_{j}(x_{i})=\mu_{j^{\prime}}(x_{i})$, thus $\mu_{j}=\mu_{j^{\prime}}$ and therefore $j=j^{\prime}$.
3.1. Permutation of characters $\mu_{j}$
Let ${(H,\mathcal{B})}$ be any abelian rational hypergroup. For any character $\mu_{j}:H\rightarrow\mathbb{C}$ and $\sigma\in\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$, define $\sigma.\mu_{j}\in H^{*}$ as the linear function on $H$ such that $[\sigma.\mu_{j}](x_{i})=\sigma(\mu_{j}(x_{i}))=\sigma({\alpha_{{}_{ij}}})$, for all $x_{i}\in{\mathcal{B}}$.
Lemma 3.2.
For any abelian rational hypergroup ${(H,\mathcal{B})}$, the function $\sigma.\mu_{j}$ is an algebra map. Thus $\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$ acts on the set $^{\widehat{\mathcal{B}}}$.
Proof.
Suppose that $x_{i_{1}}x_{i_{2}}=\sum_{k\in\mathcal{I}}N^{k}_{{i_{1}},{i_{2}}}x_{k}$. Since $N^{k}_{{i_{1}},{i_{2}}}\in\mathbb{Q}$,
$[\sigma.\mu_{j}]({x_{i_{1}}}{x_{i_{2}}})=[\sigma.\mu_{j}](\sum_{k\in\mathcal{I}}N^{k}_{{i_{1}},{i_{2}}}x_{k})=\sum_{k\in\mathcal{I}}N^{k}_{{i_{1}},{i_{2}}}[\sigma.\mu_{j}](x_{k})=\sum_{k\in\mathcal{I}}N^{k}_{{i_{1}},{i_{2}}}\sigma(\mu_{j}(x_{k})).$
On the other hand,
$[\sigma.\mu_{j}]({x_{i_{1}}})[\sigma.\mu_{j}]({x_{i_{2}}})$ $=$ $\sigma(\mu_{j}({x_{i_{1}}}))\sigma(\mu_{j}({x_{i_{2}}}))=\sigma(\mu_{j}({x_{i_{1}}})\mu_{j}({x_{i_{2}}}))$
$=$ $\sigma(\mu_{j}({x_{i_{1}}}{x_{i_{2}}}))=\sigma(\mu_{j}(\sum_{k\in\mathcal{I}}N^{k}_{{i_{1}},{i_{2}}}x_{k}))$
$=$ $\sum_{k\in\mathcal{I}}N^{k}_{{i_{1}},{i_{2}}}\sigma(\mu_{j}(x_{k})).$
Thus $[\sigma.\mu_{j}]({x_{i_{1}}}{x_{i_{2}}})=[\sigma.\mu_{j}]({x_{i_{1}}})[\sigma.\mu_{j}]({x_{i_{2}}})$, so that $\sigma.\mu_{j}$ is an algebra map. Now,
$[\sigma.(\sigma^{\prime}.\mu_{j})](x_{i})=\sigma([\sigma^{\prime}.\mu_{j}](x_{i}))=\sigma(\sigma^{\prime}(\mu_{j}(x_{i})))=(\sigma\sigma^{\prime})(\mu_{j}(x_{i}))=[\sigma\sigma^{\prime}.\mu_{j}](x_{i})$
It follows that $\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$ acts on the set $^{\widehat{\mathcal{B}}}$. ∎
Thus, there is a permutation $\tau_{\sigma}$ of ${\mathcal{I}}$ such that $\sigma.\mu_{j}=\mu_{\tau_{\sigma}(j)}$, and therefore
(3.1) $\sigma({\alpha_{{}_{ij}}})=\alpha_{{}_{i\tau_{\sigma}(j)}}\;\;\text{for all}\;i,j\in{\mathcal{I}}.$
Corollary 3.3.
Let ${(H,\mathcal{B})}$ be an abelian rational hypergroup. Then $\mathbb{Q}\subseteq\mathbb{K}$ is a Galois extension.
Proof.
Equation (3.1) shows that $\sigma(\mathbb{K})\subseteq\mathbb{K}$, for all $\sigma\in\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$. Therefore, this is a normal extension and consequently a Galois extension (as ${\mathrm{c}har}(\mathbb{Q})=0$). ∎
Remark 3.4.
The map $\rho:\mathrm{Gal}(\mathbb{K}/\mathbb{Q})\rightarrow\mathbb{S}_{\mathcal{I}},\sigma\mapsto\tau_{\sigma}$ is injective. Indeed, if $\tau_{\sigma}=\tau_{\sigma^{\prime}}$, then $\sigma({\alpha_{{}_{ij}}})=\sigma^{\prime}({\alpha_{{}_{ij}}})$, i.e. $\sigma^{-1}\circ\sigma^{\prime}({\alpha_{{}_{ij}}})={\alpha_{{}_{ij}}}$, for all $i,j$. Thus $\sigma=\sigma^{\prime}$.
Proposition 3.5.
Let $(H,\mathcal{B})$ be an abelian rational hypergroup. For any $\sigma\in\mathrm{Gal}(\mathbb{K}/\mathbb{Q})$, let $\tau=\tau_{\sigma}$, with the above notations,
(3.2) $\sigma(n_{j})={n_{\tau(j)}}.$
Proof.
Applying $\sigma\in\mathrm{Gal}(\mathbb{K}/\mathbb{Q})$ to the orthogonality relation (2.5), we obtain:
$\sum_{i\in\mathcal{I}}h_{i}\mu_{\tau(j)}(x_{i})\mu_{\tau(k)}(x_{i^{*}})=\delta_{j,k}\sigma(n_{j}),$
On the other hand, by the same orthogonality relation:
$\sum_{i\in\mathcal{I}}h_{i}\mu_{\tau(j)}(x_{i})\mu_{\tau(k)}(x_{i^{*}})=\delta_{\tau(j),\tau(k)}{n_{\tau(j)}}.$
Therefore, with $j=k$, we get that $\sigma(n_{j})={n_{\tau(j)}}$. ∎
Definition 3.6.
A hypergroup ${(H,\mathcal{B})}$ is $h$-integral if $h_{i}:=1/N^{1}_{i,i^{*}}\in\mathbb{Z}$, for all $i\in\mathcal{I}$.
Remark 3.7.
Let ${(H,\mathcal{B},\mu_{1})}$ be the Grothendieck ring of any weakly-integral fusion category $\mathcal{C}$. By [Ost15, Theorem 2.13] and Equation (2.11), $\widehat{h}_{i}=\frac{n{(H,\mathcal{B},\mu_{1})}}{n_{i}}$ is an integer, for all $i\in\mathcal{I}$. Thus, the dual $^{\widehat{{K(\mathcal{C})}}}$ is $h$-integral. See also Remark 9.1.
Lemma 3.8.
Suppose that ${(H,\mathcal{B},\mu_{1})}$ is an abelian normalizable hypergroup with a $h$-integral dual. Then it is weakly-integral. Moreover, if it is rational, then $\widehat{h}_{k}=\widehat{h}_{\tau(k)}$, for all $k\in\mathcal{I}$.
Proof.
By Lemma 2.8 and Definition 2.5,
$n{(H,\mathcal{B},\mu_{1})}=n{\widehat{{(H,\mathcal{B},\mu_{1})}}}=\sum_{i\in\mathcal{I}}{\widehat{h}}_{i}\in\mathbb{Z}.$
If ${(H,\mathcal{B},\mu_{1})}$ is rational, then Equation (3.2) can be written as $\sigma\big{(}\frac{n{(H,\mathcal{B},\mu_{1})}}{\widehat{h}_{i}}\big{)}=\frac{n{(H,\mathcal{B},\mu_{1})}}{\widehat{h}_{\tau(i)}}.$ In particular, $\widehat{h}_{k}=\widehat{h}_{\tau(i)}$, since $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is $h$-integral. ∎
4. Burnside’s vanishing property for hypergroups
Let ${(H,\mathcal{B})}$ be any hypergroup. An element $x_{i}\in\mathcal{B}$ is called a vanishing element if there is $\mu_{j}\in{\widehat{\mathcal{B}}}$ such that $\mu_{j}(x_{i})=0$. Otherwise, $x_{i}$ is called a non-vanishing element.
Let $\mathcal{B}_{0}$ be the set of all vanishing elements of $\mathcal{B}$, and $\mathcal{B}_{1}$ be the set of all non-vanishing elements of $\mathcal{B}$. Thus $\mathcal{B}=\mathcal{B}_{1}\sqcup\mathcal{B}_{0}$. By the proof of Lemma 2.27, if ${(H,\mathcal{B})}$ is a RN-hypergroup then ${G{(H,\mathcal{B})}}\subseteq\mathcal{B}_{1}$. In the abelian case, since $x_{i}=\sum_{j\in\mathcal{I}}\mu_{j}(x_{i})F_{j}$, it follows that $x_{i}$ is non-vanishing if and only if it is invertible in $H$ (i.e. $\det(L_{x_{i}})$ nonzero). Thus, in this case, $\mathcal{B}_{1}$ coincides with the set of all standard elements that are invertible in $H$.
In any abelian normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$:
(4.1) ${\widehat{P}}:=\prod_{j\in\mathcal{I}}\mu_{j}=\sum_{i\in\mathcal{I}}{\widehat{P}}({\frac{x_{i}}{d_{i}}})\widehat{F}_{i}.$
Proposition 4.1.
Let ${(H,\mathcal{B},\mu_{1})}$ be any abelian normalizable hypergroup. Then:
(4.2) $\prod_{j\in\mathcal{I}}\mu_{j}=\sum_{x_{i}\in\mathcal{B}_{1}}\det(L_{{\frac{x_{i}}{d_{i}}}})\widehat{F}_{i}$
where $\det(L_{{\frac{x_{i}}{d_{i}}}})$ is the determinant of the left multiplication operator by ${\frac{x_{i}}{d_{i}}}$ on $H$.
Proof.
Recall that $\widehat{F}_{i}(\frac{x_{j}}{d_{j}})=\delta_{i,j}$, and $\frac{x_{j}}{d_{j}}=\sum_{i\in\mathcal{I}}\mu_{i}(\frac{x_{j}}{d_{j}})F_{i}$. Thus, for all $j\in\mathcal{I}$,
$\left[\sum_{i\in\mathcal{I}}\det(L_{{\frac{x_{i}}{d_{i}}}})\widehat{F}_{i}\right](\frac{x_{j}}{d_{j}})=\det(L_{\frac{x_{j}}{d_{j}}})=\prod_{i\in\mathcal{I}}\mu_{i}(\frac{x_{j}}{d_{j}})=\left[\prod_{i\in\mathcal{I}}\mu_{i}\right](\frac{x_{j}}{d_{j}})$
It follows that
$\prod_{i\in\mathcal{I}}\mu_{i}=\sum_{i\in\mathcal{I}}\det(L_{{\frac{x_{i}}{d_{i}}}})\widehat{F}_{i}=\sum_{x_{i}\in\mathcal{B}_{1}}\det(L_{{\frac{x_{i}}{d_{i}}}})\widehat{F}_{i}.$
Indeed, $x_{i}\in{\mathcal{B}_{1}}$ if and only if $\det(L_{{\frac{x_{i}}{d_{i}}}})$ is nonzero, justifying the last equality.∎
Dually, applying the above result to $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$, we obtain a decomposition of the basis ^=^${\widehat{\mathcal{B}}}={\widehat{\mathcal{B}}}_{0}\sqcup{\widehat{\mathcal{B}}}_{1}$. Proposition 4.1 implies the following:
Proposition 4.2.
Let ${(H,\mathcal{B},\mu_{1})}$ be any abelian normalizable hypergroup. Then:
(4.3) $\prod_{i\in\mathcal{I}}{\frac{x_{i}}{d_{i}}}=\sum_{\mu_{j}\in{\widehat{\mathcal{B}}}_{1}}\det(L_{\mu_{j}})F_{j},$
where $\det(L_{\mu_{j}})$ is the determinant of the left multiplication operator by $\mu_{j}$ on $H^{*}$.
Recall the definition of an abelian Burnside hypergroup from Definition 1.4. Note that the decomposition $\mathcal{B}=\mathcal{B}_{0}\sqcup\mathcal{B}_{1}$ from above implies the following:
Proposition 4.3.
An abelian hypergroup ${(H,\mathcal{B})}$ is Burnside if and only if we have ${G{(H,\mathcal{B})}}=\mathcal{B}_{1}$.
Corollary 4.4.
An abelian normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$ is Burnside if and only if the following equality holds:
(4.4) $\prod_{j\in\mathcal{I}}\mu_{j}=\sum_{x_{i}\in{G{(H,\mathcal{B})}}}{\mathrm{sgn}}(x_{i})\widehat{F}_{i},$
where the notation ${\mathrm{sgn}}(x_{i})$ was introduced in Definition 2.41.
Proof.
Immediate from Propositions 4.1 and 4.3. ∎
Corollary 4.5.
An abelian normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$ is Burnside if and only if
(4.5) $(\prod_{j\in\mathcal{I}}\mu_{j})^{2}=\sum_{x_{i}\in{G{(H,\mathcal{B})}}}\widehat{F}_{i}.$
Corollary 4.6.
An abelian normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$, such that ${G{(H,\mathcal{B})}}$ has odd order, is Burnside if and only if
(4.6) $\prod_{j\in\mathcal{I}}\mu_{j}=\sum_{x_{i}\in G(H,B)}\widehat{F}_{i}.$
Proof.
Every ${\frac{x_{i}}{d_{i}}}\in{G{(H,\mathcal{B})}}$ has odd order, so ${\mathrm{sgn}}(x_{i})=1$. ∎
Definition 4.7.
An abelian normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$ is called dual-Burnside if $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is Burnside; in other words, if for all $\mu_{j}\in{\widehat{\mathcal{B}}}$, the following are equivalent:
  1. (1)
    For all $x_{i}\in{\mathcal{B}}$ then $\mu_{j}(x_{i})\neq 0$,
  2. (2)
    $\mu_{j}\in G\widehat{{(H,\mathcal{B})}}$, i.e $\mu_{j}$ is a grouplike element.
This can be shortened to $G\widehat{{(H,\mathcal{B})}}={\widehat{\mathcal{B}}}_{1}$.
Here are the dual of Corollaries 4.4, 4.5 and 4.6:
Corollary 4.8.
An abelian normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$ is dual-Burnside if and only if:
(4.7) $\prod_{i\in\mathcal{I}}{\frac{x_{i}}{d_{i}}}=\sum_{\mu_{j}\in G{\widehat{(H,\mathcal{B})}}}{\mathrm{sgn}}(\mu_{j})F_{j}.$
Corollary 4.9.
An abelian normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$ is dual-Burnside if and only if :
(4.8) $(\prod_{i\in\mathcal{I}}{\frac{x_{i}}{d_{i}}})^{2}=\sum_{\mu_{j}\in G{\widehat{(H,\mathcal{B})}}}F_{j}.$
Corollary 4.10.
An abelian normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$, with $^{{G{\widehat{(H,\mathcal{B})}}}}$ of odd order, is dual-Burnside if and only if
(4.9) $\prod_{i\in\mathcal{I}}{\frac{x_{i}}{d_{i}}}=\sum_{\mu_{j}\in{{G{\widehat{(H,\mathcal{B})}}}}}F_{j}.$
4.1. Inequalities for $P$ and $^{\widehat{P}}$
Recall that $P:=\prod_{i\in\mathcal{I}}\frac{x_{i}}{d_{i}}$ and ${\widehat{P}}:=\prod_{j\in\mathcal{I}}\mu_{j}$.
Proposition 4.11.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup satisfying the hypothesis of Lemma 2.30. Then $|{\widehat{P}}({\frac{x_{i}}{d_{i}}})|\leq 1$. Moreover, the equality holds if and only if $x_{i}\in G{(H,\mathcal{B})}$.
Proof.
First, suppose that ${(H,\mathcal{B})}$ is RN. Then $|\mu_{j}({\frac{x_{i}}{d_{i}}})|\leq 1$ by Frobenius-Perron theory. This implies that $|{\widehat{P}}({\frac{x_{i}}{d_{i}}})|=\prod_{j\in\mathcal{I}}|\mu_{j}({\frac{x_{i}}{d_{i}}})|\leq 1$. Moreover, the equality holds if and only if $|{\frac{x_{i}}{d_{i}}}(\mu_{j})|=1$ for all $j\in\mathcal{I}$, if and only if $x_{i}$ is grouplike, by Lemma 2.27.
Next, suppose that $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN. Since $\mu_{j}=\sum_{i\in\mathcal{I}}\mu_{j}({\frac{x_{i}}{d_{i}}})\widehat{F}_{i}$, then $\{\mu_{j}({\frac{x_{i}}{d_{i}}})\}_{i\in\mathcal{I}}$ are the eigenvalues of $L_{\mu_{j}}$. By Frobenius-Perron theory, $|\mu_{j}({\frac{x_{i}}{d_{i}}})|\leq\mathrm{FPdim}(\mu_{j})=1$, as $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is normalized. Thus, $|{\widehat{P}}({\frac{x_{i}}{d_{i}}})|\leq 1$, again. The rest is as before. ∎
Proposition 4.12.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. Then, ${\widehat{P}}^{2}$ is an idempotent if and only if ${\widehat{P}}({\frac{x_{i}}{d_{i}}})=\pm 1$, for all $x_{i}\in\mathcal{B}_{1}$.
Proof.
By Equation (4.1), ${\widehat{P}}^{4}={\widehat{P}}^{2}$ if and only if ${\widehat{P}}({\frac{x_{i}}{d_{i}}})^{4}={\widehat{P}}({\frac{x_{i}}{d_{i}}})^{2}$ for all $i\in\mathcal{I}$. Therefore, ${\widehat{P}}^{2}$ is idempotent if and only if ${\widehat{P}}({\frac{x_{i}}{d_{i}}})=0$ or ${\widehat{P}}({\frac{x_{i}}{d_{i}}})^{2}=1$, for all $i\in\mathcal{I}$. Since ${\widehat{P}}({\frac{x_{i}}{d_{i}}})=\prod_{j\in\mathcal{I}}\mu_{j}({\frac{x_{i}}{d_{i}}})$, the second case occurs if and only if $x_{i}\in\mathcal{B}_{1}$, if and only if ${\widehat{P}}({\frac{x_{i}}{d_{i}}})=\pm 1$. ∎
Corollary 4.13.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup satisfying the hypothesis of Lemma 2.30. Then, ${\widehat{P}}^{2}$ is an idempotent if and only if ${(H,\mathcal{B},\mu_{1})}$ is Burnside.
Proof.
By Proposition 4.12, ${\widehat{P}}^{2}$ is an idempotent if and only if ${\widehat{P}}({\frac{x_{i}}{d_{i}}})=\pm 1$, for all $x_{i}\in\mathcal{B}_{1}$. But, by assumption and Proposition 4.11, $|{\widehat{P}}({\frac{x_{i}}{d_{i}}})|=1$ if and only if $x_{i}\in G{(H,\mathcal{B})}$. Thus, ${\widehat{P}}^{2}$ idempotent implies that ${\mathcal{B}_{1}}\subset G{(H,\mathcal{B})}$. But $G{(H,\mathcal{B})}\subset\mathcal{B}_{1}$, so the equality holds, i.e. ${(H,\mathcal{B})}$ is Burnside.
Conversely, if ${(H,\mathcal{B})}$ is Burnside, then ${\widehat{P}}^{2}=\sum_{x_{i}\in G{(H,\mathcal{B})}}\widehat{F}_{i}$, by Corollary 4.5, which is an idempotent. ∎
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup, and $P:=\prod_{i\in\mathcal{I}}{\frac{x_{i}}{d_{i}}}$. Then
(4.10) $P=\sum_{j\in\mathcal{I}}\mu_{j}(P)F_{j}.$
Here are the dual version of the last three results:
Proposition 4.14.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup satisfying the hypothesis of Lemma 2.30. Then $|\mu_{j}(P)|\leq 1$. Moreover, the equality holds if and only if $\mu_{j}\in G{\widehat{(H,\mathcal{B})}}$.
Proposition 4.15.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. Then, $P^{2}$ is an idempotent if and only if $\mu_{j}(P)=\pm 1$, for all $\mu_{j}\in\widehat{\mathcal{B}}_{1}$.
Corollary 4.16.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup satisfying the hypothesis of Lemma 2.30. Then, $P^{2}$ is an idempotent if and only if ${(H,\mathcal{B},\mu_{1})}$ is dual-Burnside.
4.2. The analogue of Burnside’s theorem
We will prove the following generalization of Theorem 1.2, using the same approach as in [Bur23, Theorem 2].
Theorem 4.17.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian rational normalizable hypergroup with ${h}$-integral dual. Let $x_{i}\in\mathcal{B}$. If $h_{i}|d_{i}|^{2}>1$, and
(4.11) $h_{i}|\mu_{j}(x_{i})|^{2}\in\mathbb{A}$
for all $j\in\mathcal{I}$. Then $x_{i}\in\mathcal{B}_{0}$
Proof.
For any $i\in\mathcal{I}$, denote $\mathcal{T}_{i}:=\{j\in\mathcal{I}\;|\;\mu_{j}(x_{i})=0\}$ and ${\mathcal{D}}_{i}:=\mathcal{I}\setminus(\mathcal{T}_{i}\cup\{1\})$. We need to show that $\mathcal{T}_{i}\neq\emptyset$, for any $x_{i}\in\mathcal{B}$ satisfying the theorem’s hypothesis.
The second orthogonality relation (2.6) and Equation (2.11) implies
(4.12) $\sum_{j\in\mathcal{I}}\widehat{h}_{j}|\mu_{j}(x_{i})|^{2}=\frac{n{(H,\mathcal{B},\mu_{1})}}{h_{i}}.$
Since $\widehat{h}_{1}=1$, the above equation can be written as:
$\frac{n{(H,\mathcal{B},\mu_{1})}}{h_{i}}=|d_{i}|^{2}+\sum_{j\in{\mathcal{D}}_{i}}{\widehat{h}_{j}}|\mu_{j}(x_{i})|^{2}$
which gives that
(4.13) $1=\frac{n{(H,\mathcal{B},\mu_{1})}}{h_{i}|d_{i}|^{2}}-\sum_{j\in{\mathcal{D}}_{i}}\frac{\widehat{h}_{j}|\mu_{j}(x_{i})|^{2}}{|d_{i}|^{2}}.$
On the other hand,
$n{(H,\mathcal{B},\mu_{1})}=\sum_{j\in\mathcal{I}}\widehat{h}_{j}=1+\sum_{j\in\mathcal{T}_{i}}\widehat{h}_{j}+\sum_{j\in{\mathcal{D}}_{i}}\widehat{h}_{j}.$
Therefore, Equation (4.13) can be written as
(4.14) $1=\frac{1+\sum_{j\in\mathcal{T}_{i}}\widehat{h}_{j}}{h_{i}|d_{i}|^{2}}-\big{(}\sum_{j\in{\mathcal{D}}_{i}}\frac{\widehat{h}_{j}|\mu_{j}(x_{i})|^{2}}{|d_{i}|^{2}}-\sum_{j\in{\mathcal{D}}_{i}}\frac{\widehat{h}_{j}}{h_{i}|d_{i}|^{2}}\big{)}.$
Thus, in order to finish the proof, it is enough to show that
(4.15) $\big{(}\sum_{j\in{\mathcal{D}}_{i}}\frac{\widehat{h}_{j}|\mu_{j}(x_{i})|^{2}}{|d_{i}|^{2}}-\sum_{j\in{\mathcal{D}}_{i}}\frac{\widehat{h}_{j}}{h_{i}|d_{i}|^{2}}\big{)}\geq 0,$
since then, $\frac{1+\sum_{j\in\mathcal{T}_{i}}\widehat{h}_{j}}{h_{i}|d_{i}|^{2}}\geq 1$, i.e. $1+\sum\limits_{j\in\mathcal{T}_{i}}\widehat{h}_{j}\geq h_{i}|d_{i}|^{2}$; and $h_{i}|d_{i}|^{2}>1$ implies $\mathcal{T}_{i}\neq\emptyset$.
The inequality (4.15) can be written as
(4.16) $(\sum_{j\in{\mathcal{D}}_{i}}\widehat{h}_{j})^{-1}(\sum_{j\in{\mathcal{D}}_{i}}{\widehat{h}_{j}h_{i}|\mu_{j}(x_{i})|^{2}})\geq 1.$
On the other hand, the weighted AM-GM inequality gives that
(4.17) $(\sum_{j\in{\mathcal{D}}_{i}}\widehat{h}_{j})^{-1}(\sum_{j\in{\mathcal{D}}_{i}}{\widehat{h}_{j}h_{i}|\mu_{j}(x_{i})|^{2}})\geq\bigg{(}\prod_{j\in{\mathcal{D}}_{i}}\big{(}h_{i}|\mu_{j}(x_{i})|^{2}\big{)}^{\widehat{h}_{j}}\bigg{)}^{(\sum_{j\in{\mathcal{D}}_{i}}\widehat{h}_{j})^{-1}}.$
By rational assumption, $h_{i}$ is rational and $\mu_{j}(x_{i})$ is an algebraic number. Equation (3.1) implies that the set ${\mathcal{D}}_{i}$ is stable under the action of $\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$. It follows that the product
$P_{i}:=\prod_{j\in{\mathcal{D}}_{i}}(h_{i}|\mu_{j}(x_{i})|^{2})^{\widehat{h}_{j}}$
is fixed by the action of $\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$, since ${\widehat{h}_{\tau(j)}}=\widehat{h}_{j}$ by Lemma 3.8. Thus $P_{i}$ is a rational number. On the other hand, each factor of $P_{i}$ is an algebraic integer (since $\widehat{h}_{j}\in\mathbb{Z}_{{}_{>0}}$), and therefore, the entire product is an integer. Since it is also positive, it must be at least 111. ∎
Note that Theorem 1.2 follows from the above theorem, since in the case of a fusion ring, $h_{i}=1$, and $\mu_{j}(x_{i})\in\mathbb{A}$, for all $i,j\in{\mathcal{I}}$.
Proposition 4.18.
Let ${(H,\mathcal{B})}$ be a rational ARN-hypergroup with $h$-integral dual. Let $x_{i}\in\mathcal{B}$ be an invertible element in $H$ satisfying also Equation (4.11), for all $j\in\mathcal{I}$. Then $x_{i}\in{G{(H,\mathcal{B})}}$.
Proof.
Suppose that $x_{i}$ is not a grouplike element, i.e. $h_{i}|d_{i}|^{2}>1$, by Lemma 2.24. Then, all the assumptions of Theorem 4.17 are satisfied, so that, $x_{i}\in\mathcal{B}_{0}$, contradiction with $x_{i}$ invertible.∎
Corollary 4.19.
Any rational ARN-hypergroup ${(H,\mathcal{B})}$ with $h$-integral dual, such that Equation (4.11) holds for all $i,j\in\mathcal{I}$, is Burnside.
Proof.
By Proposition 4.3, it is enough to show that ${\mathcal{B}_{1}}\subseteq{G{(H,\mathcal{B})}}$. Let $x_{i}\in\mathcal{B}_{1}$, thus $x_{i}$ is invertible, so by Proposition 4.18, $x_{i}\in{G{(H,\mathcal{B})}}$. ∎
Corollary 4.20.
Any commutative fusion ring with $h$-integral dual is Burnside.
The following corollary recovers [Bur23, Theorem 2].
Corollary 4.21.
A weakly-integral fusion category with a commutative Grothendieck ring is Burnside.
Proof.
It follows from Corollary 4.20 by Remark 3.7. ∎
We can write a kind of dual version of Theorem 4.17:
Theorem 4.22.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable $h$-integral hypergroup with a rational dual. Let $\mu_{j}\in{\widehat{\mathcal{B}}}$ such that $\widehat{h}_{j}>1$ and
(4.18) $\widehat{h}_{j}\left|\mu_{j}(\frac{x_{i}}{d_{i}})\right|^{2}\in\mathbb{A},$
for all $i\in\mathcal{I}$. Then, $\mathcal{B}_{0}$ is non-empty.
Proof.
The condition from Equation (4.11) applied on $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ becomes exactly Equation (4.18). So we can apply Theorem 4.17 to $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$. Thus $\mu_{j}\in\widehat{\mathcal{B}}_{0}$, meaning the existence of $x_{i}\in\mathcal{B}$ such that $\mu_{j}(\frac{x_{i}}{d_{i}})=0$, so $x_{i}\in\mathcal{B}_{0}$. ∎
5. Kernels of fusion rings, Brauer’s theorem
Let ${(H,\mathcal{B})}$ be an ARN-hypergroup. Let ${(H,\mathcal{B})}_{+}$ be the set of all elements $x\in H$ such that $x=\sum_{b\in{\mathcal{B}}}x_{b}b$ with $x_{b}\in\mathbb{R}_{\geq 0}$. A standard element $b\in\mathcal{B}$ is called a constituent of $x\in{(H,\mathcal{B})}_{+}$ if $x_{b}>0$. Recall that $^{\widehat{\mathcal{B}}}$ is the set of all algebra morphisms $\mu_{j}:H\rightarrow\mathbb{C}$. Let $\psi\in{\widehat{\mathcal{B}}}$. For all $x\in{(H,\mathcal{B})}_{+}$,
$|\psi(x)|=|\psi(\sum_{b\in{\mathcal{B}}}x_{b}b)|\leq\sum_{b\in{\mathcal{B}}}x_{b}|\psi(b)|\leq\sum_{b\in{\mathcal{B}}}x_{b}\mathrm{FPdim}(b)=\mathrm{FPdim}(x),$
by Frobenius-Perron theory.
Definition 5.1.
Let $\mathcal{S}\subseteq\mathcal{B}$ be a subset such that
  • $\mathcal{S}$ is closed under the involution of ${(H,\mathcal{B})}$,
  • $L:=\mathbb{C}[\mathcal{S}]$ is a subalgebra of $H$.
Then ${(L,\mathcal{S})}$ is called a sub-hypergroup of ${(H,\mathcal{B})}$.
The notion of kernel of object of a fusion category from [Bur16] can be extended to arbitrary ARN-hypergroups.
Definition 5.2.
Let ${(H,\mathcal{B})}$ be any ARN-hypergroup. For any $\psi\in{\widehat{\mathcal{B}}}$, define
$\ker_{{\widehat{(H,\mathcal{B})}}}(\psi):=\{x\in\mathcal{B}\;|\;\psi(x)=\mathrm{FPdim}(x)\}.$
Lemma 5.3.
Let ${(H,\mathcal{B})}$ be any ARN-hypergroup and $\psi\in{\widehat{\mathcal{B}}}$. Then $\ker_{{\widehat{(H,\mathcal{B})}}}(\psi)$ is (the basis of) a sub-hypergroup of ${(H,\mathcal{B})}$.
Proof.
Let $x,y\in\ker_{{\widehat{(H,\mathcal{B})}}}(\psi)$, i.e. $\psi(x)=\mathrm{FPdim}(x)$ and $\psi(y)=\mathrm{FPdim}(y)$. Suppose that $xy=\sum_{z\in\mathcal{B}^{\prime}}N^{z}_{xy}z$, where $\mathcal{B}^{\prime}$ is the subset of $\mathcal{B}$ consisting of all standard elements for which $N^{z}_{xy}>0$.
It follows that
$\mathrm{FPdim}(x)\mathrm{FPdim}(y)=\psi(x)\psi(y)=\psi(xy)=\sum_{z\in\mathcal{B}^{\prime}}N^{z}_{xy}\psi(z).$
Using the triangle inequality for the complex absolute-value,
$\mathrm{FPdim}(x)\mathrm{FPdim}(y)$ $=$ $\left|\sum_{z\in\mathcal{B}^{\prime}}N^{z}_{xy}\psi(z)\right|$
$\leq$ $\sum_{z\in\mathcal{B}^{\prime}}N^{z}_{xy}|\psi(z)|\leq\sum_{z\in\mathcal{B}^{\prime}}N^{z}_{xy}\mathrm{FPdim}(z)$
$=$ $\mathrm{FPdim}(x)\mathrm{FPdim}(y).$
It is easy to deduce that $\ker_{{\widehat{(H,\mathcal{B})}}}(\psi)$ is a sub-hypergroup of $H$. ∎
Dually, we can define the following:
Definition 5.4.
Let ${(H,\mathcal{B})}$ be an ARN-hypergroup. Let $x\in H_{+}$. Define
${\ker_{{}_{{(H,\mathcal{B})}}}}(x):=\{\psi\in{\widehat{\mathcal{B}}}\;|\;\psi(x)=\mathrm{FPdim}(x)\}.$
The dual version of Lemma 5.3 implies that ${\ker_{{}_{{(H,\mathcal{B})}}}}(x)$ is (the basis of) a sub-hypergroup of $^{\widehat{(H,\mathcal{B})}}$ if $^{\widehat{(H,\mathcal{B})}}$ is a RN-hypergroup.
5.1. Brauer’s theorem for ARN-hypergroup
For any hypergroup ${(H,\mathcal{B})}$, we define the bilinear function $m$ on $H$ by
$m(\sum_{i\in\mathcal{I}}\alpha_{i}x_{i},\sum_{i\in\mathcal{I}}\beta_{i}x_{i})=\sum_{i\in\mathcal{I}}\frac{\alpha_{i}\beta_{i}}{h_{i}}.$
In the abelian case, Equation (2.2) implies
(5.1) $m(x,y)=\tau(xy^{*})=\sum_{j\in\mathcal{I}}\frac{1}{n_{j}}\mu_{j}(x)\mu_{j}(y^{*}),\;\;\text{for all}\;\;x,y\in H.$
Lemma 5.5.
Let ${(H,\mathcal{B})}$ be an ARN-hypergroup. If $x:=\sum_{s\in\mathcal{B}}p_{s}s\in{(H,\mathcal{B})}_{+}$ then
${\ker_{{}_{{(H,\mathcal{B})}}}}(x)=\bigcap_{{}_{\{s\in\mathcal{B}|\;p_{s}>0\}}}{\ker_{{}_{{(H,\mathcal{B})}}}}(s).$
Proof.
Note that $\mathrm{FPdim}(x)=\sum_{\{s\in\mathcal{B}\;|\;p_{s}>0\}}p_{s}\mathrm{FPdim}(s)$. Let $\mu_{j}\in\widehat{B}$ then
$|\mu_{j}(x)|\leq\sum_{\{s\in\mathcal{B}|\;p_{s}>0\}}p_{s}|\mu_{j}(s)|\leq\sum_{s}p_{s}\mathrm{FPdim}(s)=\mathrm{FPdim}(x).$
Thus, $\mu_{j}\in{\ker_{{}_{{(H,\mathcal{B})}}}}(x)$ if and only if $\mu_{j}\in\bigcap_{{}_{\{s\in\mathcal{B}|\;p_{s}>0\}}}{\ker_{{}_{{(H,\mathcal{B})}}}}(s)$. ∎
The following theorem is a generalization of Brauer’s theorem. Although the proof is the same as that in [Bur16], we include it here for the sake of completeness.
For $x\in{(H,\mathcal{B})}_{+}$, we denote by $\langle x\rangle$ the sub-hypergroup generated by $x$, i.e. whose basis is the set of constituents of all powers $x^{n}$, with $n\geq 1$.
Theorem 5.6.
Let ${(H,\mathcal{B})}$ be an ARN-hypergroup, and $x\in{(H,\mathcal{B})}_{+}$. Then $\langle x\rangle=H$ if and only if ${\ker_{{}_{{(H,\mathcal{B})}}}}(x)=\{\mu_{1}\}$.
Proof.
Partition the set $\mathcal{I}=\bigsqcup_{l\in\mathcal{P}}\mathcal{I}_{l}$ such that the values $\mu_{j}(x)$ are constant on each component $\mathcal{I}_{l}$ of the partition, but distinct otherwise (i.e. for all $l,l^{\prime}\in\mathcal{P}$, $j\in\mathcal{I}_{l}$, $j^{\prime}\in\mathcal{I}_{l^{\prime}}$, $\mu_{j}(x)=\mu_{j^{\prime}}(x)$ if and only if $l=l^{\prime}$). Let $s\in\mathcal{B}$. By Equation (5.1),
${m(x^{n},s)}{}=\sum_{j\in\mathcal{I}}\frac{1}{n_{j}}\mu_{j}(x)^{n}\mu_{j}(s^{*})=\sum_{l\in\mathcal{P}}\left(\sum_{j\in\mathcal{I}_{l}}\frac{1}{n_{j}}\mu_{j}(s^{*})\right)\alpha_{l}^{n}=\sum_{l\in\mathcal{P}}a_{l}\alpha_{l}^{n}$
with
$a_{l}:=\sum_{j\in\mathcal{I}_{l}}\frac{1}{n_{j}}\mu_{j}(s^{*}),\text{ and }\alpha_{l}:=\mu_{j}(x)\text{ when }j\in\mathcal{I}_{l}.$
Note that $s\not\in\langle x\rangle$ if and only if $m(x^{n},s)=0$ for all $n$, if and only if $Mv=0$, with $M$ be the matrix $(\alpha_{l}^{n})$ and $v$ be the vector $(a_{l})$. But $M$ is invertible because its determinant is nonzero, because all $\alpha_{l}$ are distinct (by Vandermonde determinant formula). Thus, $s\in\langle x\rangle$ if and only if there is $l\in\mathcal{P}$ such that $a_{l}$ is nonzero.
[…]
Finally, $\mu_{j}\in\ker_{(H,\mathcal{B})}(x)$ if and only if $x\in\ker_{{\widehat{(H,\mathcal{B})}}}(\mu_{j})$, if and only if $\langle x\rangle\subset\ker_{{\widehat{(H,\mathcal{B})}}}(\mu_{j})$ by Lemma 5.3. So if $\langle x\rangle=H$ then $\ker_{{\widehat{(H,\mathcal{B})}}}(\mu_{j})=H$, i.e $\mu_{j}=\mu_{1}$. ∎
5.2. The cointegral $\lambda_{(L,\mathcal{S})}$ in terms of kernels
Recall from §1 that for any ARN-hypergroup ${(H,\mathcal{B})}$, we denoted by $\lambda_{H}$ the primitive central idempotent $F_{1}$ corresponding to $\mu_{1}=\mathrm{FPdim}$. Let ${(L,\mathcal{S})}$ be a sub-hypergroup of ${(H,\mathcal{B})}$, and let $\lambda_{(L,\mathcal{S})}$ be the idempotent in $L\subset H$ corresponding to $\mu_{1}\big{|}_{(L,\mathcal{S})}$. By Equation (2.4), we get that:
(5.2) $\lambda_{(L,\mathcal{S})}=\frac{1}{n{(L,\mathcal{S})}}\sum_{x_{i}\in\mathcal{S}}h_{i}d_{i}x_{i}.$
In particular, $\lambda_{(L,\mathcal{S})}=\lambda_{(K,\mathcal{T})}$ if and only if ${(L,\mathcal{S})}={(K,\mathcal{T})}$.
Theorem 5.7.
For any $x\in{(H,\mathcal{B})}_{+}$, let $\mathcal{I}_{x}$ be the index set of ${\ker_{{}_{{(H,\mathcal{B})}}}}(x)$. Then
$\lambda_{\langle x\rangle}=\sum_{j\in\mathcal{I}_{x}}F_{j}$
Proof.
Suppose that $\langle x\rangle={(L,\mathcal{S})}$. Let $G_{1},\dots,G_{s}$ be the primitive central idempotents of ${(L,\mathcal{S})}$, and $\psi_{1},\dots,\psi_{s}$ be their associated characters. As above, let us take $G_{1}=\lambda_{{(L,\mathcal{S})}}$, and consequently, $\psi_{1}=\mathrm{FPdim}$.
We look at the restrictions of the characters $\mu_{j}$ at ${(L,\mathcal{S})}$. Then, there is a surjective function $f:\mathcal{I}\rightarrow\{1,\dots,s\}$ such that $\mu_{j}\big{|}_{{(L,\mathcal{S})}}=\psi_{f(j)}$, for all $j\in\mathcal{I}$.
With the above notation, we are reduced to show that $f^{-1}(\{1\})=\mathcal{I}_{x}$ By Theorem 5.6, since $\langle x\rangle={(L,\mathcal{S})}$, then $\ker_{{}_{(L,\mathcal{S})}}(x)=\{\psi_{1}\}$. On the other hand,
$\mu_{j}(x)=\mu_{j}\big{|}_{{}_{(L,\mathcal{S})}}(x)=\psi_{f(j)}(x).$
Thus, $\mu_{j}\in{\ker_{{}_{{(H,\mathcal{B})}}}}(x)$ if and only if $\psi_{f(j)}\in\ker_{{}_{(L,\mathcal{S})}}(x)$, i.e. $f(j)=1$. ∎
6. The adjoint sub-hypergroup and grouplike elements
6.1. Notion of the center of a standard element
Let ${(H,\mathcal{B})}$ be an ARN-hypergroup. We also define the notion of center of a standard element $x_{i}\in\mathcal{B}$:
$Z_{{}_{(H,\mathcal{B})}}(x_{i})=\{\psi\in{\widehat{\mathcal{B}}}\;|\;|\psi(x_{i})|=d_{i}\}.$
Define also the dual notion of the center,
$Z_{{}_{\widehat{(H,\mathcal{B})}}}(\psi)=\{x_{i}\in{\mathcal{B}}\;|\;|\psi(x_{i})|=d_{i}\}.$
As in Lemma 5.3, it is easy to see that $Z_{{}_{\widehat{(H,\mathcal{B})}}}(\psi)$ is a sub-hypergroup of ${(H,\mathcal{B})}$. Idem for $Z_{{}_{(H,\mathcal{B})}}(x_{i})$, if $^{\widehat{(H,\mathcal{B})}}$ is a RN-hypergroup.
6.2. Adjoint sub-hypergroup
Proposition 6.1.
Let ${(H,\mathcal{B})}$ be an ARN-hypergroup and $I(1):=\sum_{i\in\mathcal{I}}h_{i}x_{i}x_{i^{*}}\in H$. Then
${\ker_{{}_{{(H,\mathcal{B})}}}}(I(1))=\bigcap_{i\in\mathcal{I}}Z_{{}_{(H,\mathcal{B})}}(x_{i})$
Proof.
Note that
$\psi(I(1))=\sum_{i\in\mathcal{I}}h_{i}\psi(x_{i}x_{i^{*}})=\sum_{i\in\mathcal{I}}h_{i}|\psi(x_{i})|^{2}\leq\sum_{i\in\mathcal{I}}h_{i}d_{i}^{2}=\mathrm{FPdim}{(H,\mathcal{B})}.$
So if $\psi\in{\ker_{{}_{{(H,\mathcal{B})}}}}(I(1))$, then $\psi\in\bigcap_{i\in\mathcal{I}}Z_{{}_{(H,\mathcal{B})}}(x_{i})$. The converse is immediate. ∎
Definition 6.2.
Denote by ${{(H,\mathcal{B})}_{ad}}$ the sub-hypergroup $\langle I(1)\rangle$ of $H$ generated by $I(1)$. It is called the adjoint sub-hypergroup of $H$.
6.3. On the support ${\mathcal{I}_{{}_{{(L,\mathcal{S})}}}}$
Definition 6.3.
Let ${(L,\mathcal{S})}$ be a sub-hypergroup of an abelian hypergroup ${(H,\mathcal{B})}$. Its support is the subset ${\mathcal{I}_{{}_{{(L,\mathcal{S})}}}}\subset\mathcal{I}$ such that
(6.1) ${\lambda_{{}_{{(L,\mathcal{S})}}}}=\sum_{j\in{\mathcal{I}_{{}_{{(L,\mathcal{S})}}}}}F_{j}.$
In particular, by Equation (5.2), ${\mathcal{I}_{{}_{{(L,\mathcal{S})}}}}={\mathcal{I}_{{}_{{(K,\mathcal{T})}}}}$ if and only if ${(L,\mathcal{S})}={(K,\mathcal{T})}$.
The next result generalizes [Bur20, Lemma 6.4].
Proposition 6.4.
Let ${(H,\mathcal{B})}$ be an ARN-hypergroup. With the above notations,
${{\mathcal{I}}_{{{(H,\mathcal{B})}_{ad}}}}=\{j\in\mathcal{I}\;|\;n_{j}=\mathrm{FPdim}{(H,\mathcal{B})}\}.$
Proof.
By Theorem 5.7 and Definition 6.2, ${{\mathcal{I}}_{{{(H,\mathcal{B})}_{ad}}}}={\mathcal{I}}_{I(1)}$, the index set of ${\ker_{{}_{{(H,\mathcal{B})}}}}(I(1))$. Thus,
${{\mathcal{I}}_{{{(H,\mathcal{B})}_{ad}}}}=\{j\in\mathcal{I}\;|\;\mu_{j}(I(1))=\mathrm{FPdim}(I(1))\}.$
But $\mathrm{FPdim}(I(1))=\mathrm{FPdim}{(H,\mathcal{B})}$, and $\mu_{j}(I(1))=n_{j}$ by Equation (2.5). ∎
Theorem 6.5.
Let ${(H,\mathcal{B})}$ be an ARN-hypergroup. Then $\mu_{j}\in G{\widehat{(H,\mathcal{B})}}$ if and only $j\in{{\mathcal{I}}_{{{(H,\mathcal{B})}_{ad}}}}$.
Proof.
It follows from Proposition 6.4 and Lemma 2.31. ∎
Corollary 6.6.
Let ${(H,\mathcal{B})}$ be an ARN-hypergroup. Then
(6.2) $\lambda_{{}_{{{(H,\mathcal{B})}_{ad}}}}=\sum_{\mu_{j}\in G{\widehat{(H,\mathcal{B})}}}F_{j}$
Proof.
It follows from Equation (6.1) and Theorem 6.5. ∎
Here are dual versions of Theorem 6.5 and Corollary 6.6:
Theorem 6.7.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup such that $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN. Then $x_{i}\in{G{(H,\mathcal{B})}}$ if and only $i\in\mathcal{I}_{{{\widehat{(H,\mathcal{B})}}}_{ad}}$.
Corollary 6.8.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup such that $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN. Then
(6.3) $\lambda_{{}_{{{\widehat{(H,\mathcal{B})}}}_{ad}}}=\sum_{x_{i}\in{G{(H,\mathcal{B})}}}\widehat{F}_{i}.$
6.4. Proof of Theorems 1.7 and 1.10
Recall Definition 6.3, which defines the support $\mathcal{I}_{{(L,\mathcal{S})}}\subset\mathcal{I}$ of a sub-hypergroup ${(L,\mathcal{S})}$ within an abelian hypergroup ${(H,\mathcal{B})}$. For the sake of brevity, we define $|L|:=n(L,\mathcal{S},\mu_{1})$.
Theorem 6.9.
Let ${(L,\mathcal{S})}$ be a sub-hypergroup of an abelian normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$. With the above notations, we have:
(6.4) $\sum\limits_{x_{i}\in\mathcal{S}}\widehat{F}_{i}=\frac{|L|}{|H|}\big{(}\sum\limits_{j\in\mathcal{I}_{{(L,\mathcal{S})}}}{\widehat{h}}_{j}\mu_{j}\big{)}.$
Proof.
Applying $\mathcal{F}$ to Equation (6.1) we have that
[…]
On the other hand, from Equation (2.13) we have ${\lambda_{{}_{{(L,\mathcal{S})}}}}=\frac{1}{|L|}\big{(}\sum\limits_{x_{i}\in\mathcal{S}}h_{i^{*}}d_{i^{*}}x_{i}\big{)}$ and therefore:
[…]
Comparing (6.5) and (6.6), we obtain:
$\sum\limits_{x_{i}\in\mathcal{S}}\widehat{F}_{i}=|L|\big{(}\sum\limits_{j\in\mathcal{I}_{(L,\mathcal{S})}}\frac{\mu_{j}}{n_{j}}\big{)}.$
According to Equation (2.11), we have ${\widehat{h}}_{j}=\frac{|H|}{n_{j}}$, thus concluding the proof. ∎
Proof of Theorem 1.7: By applying Corollary 4.5, we get that
$\big{(}\prod_{j\in\mathcal{I}}\mu_{j}\big{)}^{2}=\sum_{x_{i}\in{G{(H,\mathcal{B})}}}\widehat{F}_{i}.$
On the other hand, by utilizing Theorem 6.9 with $\mathcal{S}=G{(H,\mathcal{B})}$, we arrive at the desired conclusion.
Proof of Theorem 1.10: To derive the result through duality from Theorem 1.7, we need to consider the normalized basis $(\frac{x_{i}}{d_{i}})_{i\in\mathcal{I}}$. Recall that by definition, $\langle x_{i}x_{i^{*}},x_{1}\rangle^{-1}=h_{i}$. Therefore, $\langle\frac{x_{i}}{d_{i}}\frac{x_{i^{*}}}{d_{i^{*}}},x_{1}\rangle^{-1}=d_{i}d_{i^{*}}h_{i}$. Consequently, $d_{i}d_{i^{*}}h_{i}\frac{x_{i}}{d_{i}}=d_{i^{*}}h_{i}x_{i}$, which precisely matches the summand on the RHS.
6.5. Proof of Theorems 1.8 and 1.11.
Proof.
By Corollary 4.5, ${(H,\mathcal{B})}$ is Burnside if and only if
$\big{(}\prod_{j\in\mathcal{I}}\mu_{j}\big{)}^{2}=\sum_{x_{i}\in G(H,B)}\widehat{F}_{i},$
if and only if Equation (1.4) holds, by Corollary 6.8, (the dual version of) Equation (2.12) and Lemma 2.14.∎
By duality, we obtain directly Theorem 1.11.
6.6. On the sub-hypergroup generated by $P$
Lemma 6.10.
For any sub-hypergroup ${(L,\mathcal{S})}$ of an ARN-hypergroup ${(H,\mathcal{B})}$, then ${\mathcal{I}}_{{(L,\mathcal{S})}}=\bigcap_{s\in\mathcal{S}}{\mathcal{I}}_{s}$. Thus, $s\in\mathcal{S}$ if and only if $s\lambda_{{(L,\mathcal{S})}}=d_{s}\lambda_{{(L,\mathcal{S})}}$.
Proof.
Note that ${(L,\mathcal{S})}=\langle x\rangle$, with $x=\sum_{s\in\mathcal{S}}s\in{(H,\mathcal{B})}_{+}$. Recall that ${\mathcal{I}}_{x}$ is the index set of ${\ker_{{}_{{(H,\mathcal{B})}}}}(x)$. By Theorem 5.7, ${\mathcal{I}_{x}}={\mathcal{I}_{\langle x\rangle}}$, and by Lemma 5.5, ${\mathcal{I}_{x}}=\bigcap_{s\in\mathcal{S}}{\mathcal{I}_{s}}$. The first sentence follows. Recall that $x=\sum_{j\in\mathcal{I}}\mu_{j}(x)F_{j}$, thus $x\lambda_{{(L,\mathcal{S})}}=\sum_{j\in\mathcal{I}_{{(L,\mathcal{S})}}}\mu_{j}(x)F_{j}$. But the first sentence means that $j\in{\mathcal{I}}_{{(L,\mathcal{S})}}$ if and only if $\mu_{j}(s)=\mathrm{FPdim}(s)$, for all $s\in\mathcal{S}$. So if $s\in\mathcal{S}$ then $s\lambda_{{(L,\mathcal{S})}}=d_{s}\lambda_{{(L,\mathcal{S})}}$. Regarding the converse, we deduce by positivity in Equation (5.2) that $sS\subseteq S$. Specifically, since $s=sx_{1}\in sS$, it follows that $s\in S$. ∎
In particular, for an ARN-hypergroup with $\mu_{1}=\mathrm{FPdim}$, then ${\mathcal{I}_{{}_{{(L,\mathcal{S})}}}}=\{j\in{\mathcal{I}}\text{ such that }\mu_{j}|_{(L,\mathcal{S})}=\mathrm{FPdim}\}$. Let ${(L,\mathcal{S})},{(K,\mathcal{T})}$ be two sub-hypergroups of an ARN-hypergroup ${(H,\mathcal{B})}$. Let ${(L,\mathcal{S})}\vee{(K,\mathcal{T})}$ be the sub-hypergroup generated by ${(L,\mathcal{S})}$ and ${(K,\mathcal{T})}$.
Proposition 6.11.
With above notations, $\mathcal{I}_{{}_{{(L,\mathcal{S})}\vee{(K,\mathcal{T})}}}={\mathcal{I}_{{}_{{(K,\mathcal{T})}}}}\cap{\mathcal{I}_{{}_{{(L,\mathcal{S})}}}}$. In other words, $\lambda_{{(L,\mathcal{S})}\vee{(K,\mathcal{T})}}=\lambda_{{(L,\mathcal{S})}}\lambda_{{(K,\mathcal{T})}}$.
Proof.
Let ${\mathcal{S}}\vee{\mathcal{T}}$ denotes the basis of ${(L,\mathcal{S})}\vee{(K,\mathcal{T})}$. By Lemmas 6.10 and 5.3,
${\mathcal{I}_{{}_{{(L,\mathcal{S})}}}}\cap{\mathcal{I}_{{}_{{(K,\mathcal{T})}}}}=\bigcap_{a\in{\mathcal{S}}\cup{\mathcal{T}}}{\mathcal{I}}_{a}$ =$=\{j\in{\mathcal{I}}\ |\ \mu_{j}\in\ker_{{(H,\mathcal{B})}}(a),\ \forall a\in{\mathcal{S}}\cup{\mathcal{T}}\}$
=$=\{j\in{\mathcal{I}}\ |\ {\mathcal{S}}\cup{\mathcal{T}}\subset\ker_{{\widehat{(H,\mathcal{B})}}}(\mu_{j})\}$
=$=\{j\in{\mathcal{I}}\ |\ {\mathcal{S}}\vee{\mathcal{T}}\subset\ker_{{\widehat{(H,\mathcal{B})}}}(\mu_{j})\}$
=$=\{j\in{\mathcal{I}}\ |\ \mu_{j}\in\ker_{{(H,\mathcal{B})}}(a),\ \forall a\in{\mathcal{S}}\vee{\mathcal{T}}\}$
$=\bigcap_{a\in{\mathcal{S}}\vee{\mathcal{T}}}{\mathcal{I}}_{a}=\mathcal{I}_{{}_{{(L,\mathcal{S})}\vee{(K,\mathcal{T})}}}.\qed$
Proposition 6.12.
For any two sub-hypergroups ${(L,\mathcal{S})},{(K,\mathcal{T})}$ of an ARN-hypergroup ${(H,\mathcal{B})}$, then ${(L,\mathcal{S})}\subseteq{(K,\mathcal{T})}$ if and only if ${\mathcal{I}_{{}_{{(K,\mathcal{T})}}}}\subseteq{\mathcal{I}_{{}_{{(L,\mathcal{S})}}}}$.
Proof.
By Lemma 6.10, if ${(L,\mathcal{S})}\subseteq{(K,\mathcal{T})}$ then ${\mathcal{I}_{{}_{{(K,\mathcal{T})}}}}\subseteq{\mathcal{I}_{{}_{{(L,\mathcal{S})}}}}$. By Proposition 6.11, if ${\mathcal{I}_{{}_{{(K,\mathcal{T})}}}}\subseteq{\mathcal{I}_{{}_{{(L,\mathcal{S})}}}}$, then ${\lambda_{{}_{{(L,\mathcal{S})}}}}{\lambda_{{}_{{(K,\mathcal{T})}}}}={\lambda_{{}_{{(K,\mathcal{T})}}}}$. Pick $s\in\mathcal{S}$, then
$s{\lambda_{{}_{{(K,\mathcal{T})}}}}=s{\lambda_{{}_{{(L,\mathcal{S})}}}}{\lambda_{{}_{{(K,\mathcal{T})}}}}=d_{s}{\lambda_{{}_{{(L,\mathcal{S})}}}}{\lambda_{{}_{{(K,\mathcal{T})}}}}=d_{s}{\lambda_{{}_{{(K,\mathcal{T})}}}},$
thus $s\in\mathcal{T}$, by Lemma 6.10. ∎
Corollary 6.13.
For any two elements $P,Q\in{(H,\mathcal{B})}_{+}$ of an ARN-hypergroup ${(H,\mathcal{B})}$, then $\langle P\rangle\subseteq\langle Q\rangle$ if and only if ${\ker_{{}_{{(H,\mathcal{B})}}}}(Q)\subseteq{\ker_{{}_{{(H,\mathcal{B})}}}}(P)$.
For any ${(H,\mathcal{B},\mu_{1})}$, recall that $P:=\prod_{i\in\mathcal{I}}{\frac{x_{i}}{d_{i}}}$.
Proposition 6.14.
For any abelian RN-hypergroup ${(H,\mathcal{B})}$ then
$\langle P^{2}\rangle={{(H,\mathcal{B})}_{ad}}.$
Proof.
By Definition 6.2, ${{(H,\mathcal{B})}_{ad}}:=\langle I(1)\rangle$, so we are reduced to show that $\mathcal{I}_{\langle P^{2}\rangle}=\mathcal{I}_{\langle I(1)\rangle}$, so (by Theorem 5.7) that $\mathcal{I}_{P^{2}}=\mathcal{I}_{I(1)}$ i.e. ${\ker_{{}_{{(H,\mathcal{B})}}}}(P^{2})={\ker_{{}_{{(H,\mathcal{B})}}}}(I(1))$.
Note that $\big{|}\psi(P)\big{|}=\prod_{i\in\mathcal{I}}\big{|}\psi({\frac{x_{i}}{d_{i}}})\big{|}\leq 1$. Thus, if $\psi\in{\ker_{{}_{{(H,\mathcal{B})}}}}(P)$, then $\big{|}\psi({\frac{x_{i}}{d_{i}}})\big{|}=1$, i.e $\psi\in{\ker_{{}_{{(H,\mathcal{B})}}}}(I(1))$, by Proposition 6.1. Thus, ${\ker_{{}_{{(H,\mathcal{B})}}}}(P)\subseteq{\ker_{{}_{{(H,\mathcal{B})}}}}(I(1))$, and the same argument shows that ${\ker_{{}_{{(H,\mathcal{B})}}}}(P^{2})\subseteq{\ker_{{}_{{(H,\mathcal{B})}}}}(I(1))$. Conversely, if $\psi\in{\ker_{{}_{{(H,\mathcal{B})}}}}(I(1))$ then $\psi({\frac{x_{i}}{d_{i}}}\frac{x_{i^{*}}}{d_{i^{*}}})=\big{|}\psi({\frac{x_{i}}{d_{i}}})\big{|}^{2}=1$. Thus, if $i\neq i^{*}$ then, grouping together $x_{i}$ and $x_{i^{*}}$ in $P$, we obtain $\psi({\frac{x_{i}}{d_{i}}})\psi(\frac{x_{{i^{*}}}}{{d_{i^{*}}}})=\big{|}\psi({\frac{x_{i}}{d_{i}}})^{2}\big{|}=1$. Thus,
$\psi(P)=\prod_{i={i^{*}}}\psi({\frac{x_{i}}{d_{i}}}).$
Note that if $i=i^{*}$, then $\psi({\frac{x_{i}}{d_{i}}})=\pm 1$, since in this case $\psi(x_{i})=\psi(x_{i^{*}})\in\mathbb{R}$. Thus, if $\psi\in{\ker_{{}_{{(H,\mathcal{B})}}}}(I(1))$, then $\psi(P)=\pm 1$ and $\psi(P^{2})=1$, i.e. $\psi\in{\ker_{{}_{{(H,\mathcal{B})}}}}(P^{2})$. ∎
Corollary 6.15.
For any ARN-hypergroup ${(H,\mathcal{B})}$, then ${{(H,\mathcal{B})}_{ad}}\subseteq\langle P\rangle$.
Corollary 6.16.
Let ${(H,\mathcal{B})}$ be an ARN hypergroup. The following holds:
  1. (1)
    ${{(H,\mathcal{B})}_{ad}}=\mathbb{C}$ if and only if ${(H,\mathcal{B})}$ is pointed,
  2. (2)
    ${{(H,\mathcal{B})}_{ad}}=\mathbb{C}$ if and only if $^{\widehat{(H,\mathcal{B})}}$ is pointed,
  3. (3)
    ${{(H,\mathcal{B})}_{ad}}={(H,\mathcal{B})}$ if and only if $^{\widehat{(H,\mathcal{B})}}$ is perfect.
Proof.
First, recall that ${{(H,\mathcal{B})}_{ad}}=\langle I(1)\rangle$, so ${{(H,\mathcal{B})}_{ad}}=\mathbb{C}$ if and only if $I(1)=n{(H,\mathcal{B})}x_{1}$, if and only if ${(H,\mathcal{B})}$ is pointed, as $I(1)=\sum_{i\in\mathcal{I}}h_{i}x_{i}x_{i^{*}}$. Next, by Equation (5.2), ${{(H,\mathcal{B})}_{ad}}=\mathbb{C}$ if and only if $\lambda_{{{(H,\mathcal{B})}_{ad}}}=\lambda_{\mathbb{C}}=x_{1}=\operatorname{id}=\sum_{i\in\mathcal{I}}F_{i},$ if and only if $G{\widehat{(H,\mathcal{B})}}=\widehat{\mathcal{B}}$, by Equation 6.2, meaning that $^{\widehat{(H,\mathcal{B})}}$ is pointed. Similary, ${{(H,\mathcal{B})}_{ad}}={(H,\mathcal{B})}$ if and only if $\lambda_{{{(H,\mathcal{B})}_{ad}}}=F_{1}$, if and only if $G{\widehat{(H,\mathcal{B})}}=\{\mu_{1}\}$, meaning that $^{\widehat{(H,\mathcal{B})}}$ is perfect. ∎
Remark 6.17.
By §7, ${{(H,\mathcal{B})}_{ad}}={(H,\mathcal{B})}$ if and only if the universal grading group is trivial, if and only if there is no non-trivial grading (see Corollary 7.7).
Corollary 6.18.
A simple non-pointed ARN-hypergroup ${(H,\mathcal{B})}$ has a perfect dual.
Proof.
By simplicity ${{(H,\mathcal{B})}_{ad}}=\mathbb{C}$ or ${(H,\mathcal{B})}$. By non-pointed assumption and Corollary 6.16, ${{(H,\mathcal{B})}_{ad}}\neq\mathbb{C}$, therefore ${{(H,\mathcal{B})}_{ad}}={(H,\mathcal{B})}$, and so $^{\widehat{(H,\mathcal{B})}}$ is perfect by Corollary 6.16. ∎
Corollary 6.19.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup with $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ RN. Then the following holds:
  1. (1)
    ${\widehat{(H,\mathcal{B})}}_{ad}=\mathbb{C}$ if and only if $^{\widehat{(H,\mathcal{B})}}$ is pointed,
  2. (2)
    ${\widehat{(H,\mathcal{B})}}_{ad}=\mathbb{C}$ if and only if ${(H,\mathcal{B})}$ is pointed,
  3. (3)
    $^{\widehat{(H,\mathcal{B})}}_{ad}={\widehat{(H,\mathcal{B})}}$ if and only if ${(H,\mathcal{B})}$ is perfect.
7. On the universal grading of hypergroups
The universal grading construction for fusion rings from [GN08, §2] works word for word in the ARN-hypergroup settings. For the completeness of the paper, we sketch its main steps below.
7.1. Based modules are completely reducible
Definition 7.1.
Given an ARN-hypergroup ${(H,\mathcal{B})}$, we define a based left ${(H,\mathcal{B})}$-module as a pair $(M,\mathcal{M})$ where $M$ is a $H$-module with $\mathcal{M}$ a finite free $H$-basis for $M$, i.e. for all $x_{i}\in\mathcal{B}$ and all $m_{j}\in\mathcal{M}$, then $x_{i}m_{j}=\sum_{k}t^{k}_{ij}m_{k}$; and moreover, for all $i,j,k$ then $t^{k}_{ij}$ is zero if and only if $t^{j}_{i^{*}k}$ is zero.
The last condition above means that $(x_{i}m_{j},m_{k})$ vanishes if and only if $(m_{j},x_{i^{*}}m_{k})$ vanishes, for the inner product on $M$ such that for all $m_{i},m_{j}\in\mathcal{M}$
(7.1) $(m_{i},m_{j})=\delta_{i,j}.$
It is easy to deduce that any based ${(H,\mathcal{B})}$-modules is completely reducible, as in [Ost03a, Lemma 2.1]. Indeed, if $N\subset M$ and $HN\subset N$ then $(N,N^{\perp})=\{0\}=(HN,N^{\perp})=(N,HN^{\perp})$. Thus, $HN^{\perp}\subset N^{\perp}$.
Let ${(H,\mathcal{B})}$ be a RN-hypergroup and $x,y\in{(H,\mathcal{B})}_{+}$. We say that $x$ contains $y$ (or $y$ is a constituent of $x$) if the difference $x-y$ is in ${(H,\mathcal{B})}_{+}$.
Definition 7.2.
Let ${(H,\mathcal{B})}$ be a RN-hypergroup, $G$ be a finite group. Suppose that $\mathcal{B}=\sqcup_{g\in G}{\mathcal{B}}_{g}$ is a partition of $\mathcal{B}$, and let $H_{g}:=\mathbb{C}[\mathcal{B}_{g}]$. The decomposition ${(H,\mathcal{B})}=\bigoplus_{g\in G}(H_{g},\mathcal{B}_{g})$ is called a grading of ${(H,\mathcal{B})}$ by $G$ if $H_{g}H_{h}\subseteq H_{gh}$, for all $g,h\in G$. The grading is called faithful if $\mathcal{B}_{g}\neq\emptyset$, for all $g\in G$.
Please note that a faithful grading group of an abelian hypergroup must also be abelian. Let ${(H,\mathcal{B})}=\bigoplus_{g\in G}(H_{g},\mathcal{B}_{g})$ be a faithful grading of ${(H,\mathcal{B})}$, and define $R_{g}:=\sum_{x_{i}\in\mathcal{B}_{g}}h_{i}d_{i}x_{i}$, for all $g\in G$. The analogue of [ENO05, Proposition 8.20] holds in the settings of ARN-hypergroup:
(7.2) $\mathrm{FPdim}(R_{g})=\frac{\mathrm{FPdim}{(H,\mathcal{B})}}{|G|}.$
Indeed, consider $R:=\sum_{g\in G}R_{g}$, and note that $\frac{1}{\mathrm{FPdim}{(H,\mathcal{B})}}R=F_{1}$, the idempotent corresponding to $\mathrm{FPdim}$ in Corollary 2.10. Thus, $RR_{h}=\mathrm{FPdim}(R_{h})R$. Since $H_{g}H_{h}\subseteq H_{gh}$, we obtain from the previous equality that $R_{g}R_{h}=\mathrm{FPdim}(R_{h})R_{gh}$, for all $g,h\in G$. Passing to $\mathrm{FPdim}$, we get that $\mathrm{FPdim}(R_{g})\mathrm{FPdim}(R_{h})=\mathrm{FPdim}(R_{h})\mathrm{FPdim}(R_{gh})$; but $\mathrm{FPdim}(R_{h})\neq 0$ by faithfulness, so $\mathrm{FPdim}(R_{gh})=\mathrm{FPdim}(R_{g})$, for all $g,h\in G$.
7.2. Constructing the universal grading
Proposition 7.3.
Let ${(H,\mathcal{B})}$ be a RN-hypergroup. For all $x\in H$, define $I(x):=\sum_{j\in\mathcal{I}}h_{j}x_{j}xx_{j^{*}}$. With the above notations:
  1. (1)
    $I(x)$ is central in $H$.
  2. (2)
    Every based left $H_{ad}$-submodule $M$ of $H$ is also a $H_{ad}$-subbimodule of $H$.
  3. (3)
    A based $H_{ad}$-subbimodule $M$ of $H$ is indecomposable if and only if it is indecomposable as left $H_{ad}$-module.
Proof.
  1. (1)
    For all $i\in\mathcal{I}$, by Frobenius reciprocity (2.18):
    $x_{i}I(x)=\sum_{j\in\mathcal{I}}h_{j}x_{i}x_{j}xx_{j^{*}}=\sum_{j,k\in\mathcal{I}}h_{j}N_{i,j}^{k}x_{k}xx_{j^{*}}=\sum_{j,k\in\mathcal{I}}h_{k}N_{k^{*},i}^{j^{*}}x_{k}xx_{j^{*}}$
    $=\sum_{k\in\mathcal{I}}h_{k}x_{k}x(\sum_{j\in\mathcal{I}}N_{k^{*},i}^{j^{*}}x_{j^{*}})=\sum_{k\in\mathcal{I}}h_{k}x_{k}xx_{k^{*}}x_{i}=I(x)x_{i}.$
  2. (2)
    Recall that $H_{ad}:=\langle I(1)\rangle$. By (1), $m{I(1)}^{n}={I(1)}^{n}m$, thus by RN-assumption, $mH_{ad}=H_{ad}m\subset M$, for all $m\in M$, meaning that $M$ is also a based right $H_{ad}$-submodule.
  3. (3)
    Clear. ∎
Let $H=\oplus_{a\in A}H_{a}$ be a decomposition (into indecomposable) of $H$ as a $H_{ad}$-bimodule. This also corresponds to a partition $\mathcal{B}=\sqcup_{a\in A}\mathcal{B}_{a}$. Let 111 be an element of $A$ such that $H_{1}=H_{ad}$.
7.2.1. Definition of $a^{*}$ for any $a\in A$
Since $H_{a}$ is an indecomposable $H_{ad}$-bimodule then $(H_{a})^{*}$ is also an indecomposable $H_{ad}$-bimodule. Let $a^{*}$ be an element of $A$ such that $(H_{a})^{*}=H_{a^{*}}$. Next lemma is as [GN08, Lemma 3.4].
Lemma 7.4.
If $x_{a},y_{a}\in\mathcal{B}_{a}$ then $x_{a}y_{a}^{*}\in H_{ad}$.
Proof.
Note that $M:=H_{a}\cap(H_{ad}y_{a})$ is a $H_{ad}$-submodule of $H_{a}$ (containing $y_{a}$). But $H_{a}$ is indecomposable (so irreducible by complete reducibility). It follows that $M=H_{a}$, so $H_{a}\subset H_{ad}y_{a}$. Thus $x_{a}\in H_{ad}y_{a}$ and $x_{a}y_{a}^{*}\in H_{ad}y_{a}y_{a}^{*}\subset H_{ad}$. ∎
Theorem 7.5.
For all $a,b\in A$, there is $c\in A$ such that $H_{a}H_{b}\subset H_{c}$. This induces a group structure on $A$ given by $ab:=c$. Moreover, $a^{*}$ is the inverse of $a$, and 111 the unit of $A$.
Proof.
As for the proof of [GN08, Theorem 3.5]. ∎
Definition 7.6.
Above group $A$ is called the universal grading group of ${(H,\mathcal{B})}$. It is denoted $U_{H}$ when the standard basis $\mathcal{B}$ is implicitly understood.
Corollary 7.7.
Every ARN-hypergroup ${(H,\mathcal{B})}$ has a canonical faithful grading by the group $U(H)$. Any other faithful grading of ${(H,\mathcal{B})}$ by a group $G$ is determined by a surjective group morphism $\pi:U_{H}\rightarrow G$.
Proof.
Let ${(H,\mathcal{B})}=\oplus_{g\in G}(S_{g},\mathcal{D}_{g})$ be another faithful grading. It follows that $xx^{*}\in S_{1}$, for any $x\in\mathcal{D}_{g}$, but $xx^{*}\in H_{ad}$ therefore, $\mathcal{B}_{ad}\subseteq\mathcal{D}_{1}$. Moreover, each $S_{g}$ is a ${H_{ad}}$-subbimodule of $H$. Since $H_{a}$ is an indecomposable module, then $H_{a}\subseteq S_{\pi(a)}$, for some well-defined $\pi(a)\in G$. It follows that $\pi:U_{H}\rightarrow G$ is a surjective group morphism. Indeed, for all $a,b\in A$, then $H_{a}\subset S_{\pi(a)}$ and $H_{b}\subset S_{\pi(b)}$, thus $H_{a}H_{b}\subset S_{\pi(a)}S_{\pi(b)}\subset S_{\pi(a)\pi(b)}$. But $H_{a}H_{b}\subset H_{ab}$, thus $H_{a}H_{b}\subset H_{ab}\cap S_{\pi(a)\pi(b)}\neq\emptyset$, so $H_{ab}\subset S_{\pi(a)\pi(b)}$, meaning that $\pi(ab)=\pi(a)\pi(b)$. Finally, the surjectivity comes from the faithfulness. ∎
7.3. On the dual universal grading group and invertible
In this subsection, we use the notation $H$ for a hypergroup ${(H,\mathcal{B})}$, and $\widehat{H}$ for its dual $^{\widehat{(H,\mathcal{B})}}$. The goal of this subsection is to prepare for the proof (§7.4) of the following:
Theorem 7.8.
Let $H$ be an ARN-hypergroup. Then,
$U(H)\simeq{G(\widehat{H})}$
Dually, we have the following:
Theorem 7.9.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup such that $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN. Then
$U(\widehat{H})\simeq G(H).$
Recall from Lemma 2.28 that for any abelian normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$ then $\mu_{j}\in{G(\widehat{H})}$ if and only if $|\omega_{ij}|=1$, for all $i\in\mathcal{I}$, with $\omega_{ij}:=\mu_{j}(\frac{x_{i}}{d_{i}})$.
Definition 7.10.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. For any $x_{i}\in\mathcal{B}$, we define the linear character $\omega_{i}:={\frac{x_{i}}{d_{i}}}\big{|}_{{G(\widehat{H})}}$, the restriction to ${G(\widehat{H})}$ of the linear character $^{\frac{x_{i}}{d_{i}}}:\widehat{H}\rightarrow\mathbb{C}$, i.e. $\omega_{i}(\mu_{j})=\omega_{ij}$.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. Given $\psi\in\widehat{{G(\widehat{H})}}$, let
(7.3) $\mathcal{B}_{\psi}:=\{x_{i}\in\mathcal{B}\;\mid\;\omega_{i}=\psi\}=\{x_{i}\;|\;\mu_{j}({\frac{x_{i}}{d_{i}}})=\psi(\mu_{j}),\;\text{for all}\;\mu_{j}\in{G(\widehat{H})}\}.$
Proposition 7.11.
Let ${(H,\mathcal{B})}$ be an ARN-hypergroup. Then $(\mathcal{B}_{\psi})_{\psi\in\widehat{{G(\widehat{H})}}}$ defines a grading on ${(H,\mathcal{B})}$.
Proof.
For any $\mu_{j}\in{G(\widehat{H})}$, $\psi,\phi\in\widehat{{G(\widehat{H})}}$, $x_{i}\in B_{\psi}$ and $x_{k}\in B_{\phi}$, then
$\psi\phi(\mu_{j})=\psi(\mu_{j})\phi(\mu_{j})=\mu_{j}({\frac{x_{i}}{d_{i}}})\mu_{j}(\frac{x_{k}}{d_{k}})=\mu_{j}({\frac{x_{i}}{d_{i}}}\frac{x_{k}}{d_{k}}).$
By the triangle inequality:
$1=|\psi\phi(\mu_{j})|=|\mu_{j}({\frac{x_{i}}{d_{i}}}\frac{x_{k}}{d_{k}})|\leq\frac{1}{d_{i}d_{k}}\sum_{l\in\mathcal{I}}d_{l}N^{l}_{ik}|\mu_{j}(\frac{x_{l}}{d_{l}})|\leq\frac{1}{d_{i}d_{k}}\sum_{l\in\mathcal{I}}d_{l}N^{l}_{ik}=1,$
which forces $\mu_{j}(\frac{x_{l}}{d_{l}})=\mu_{j}(\frac{x_{i}}{d_{i}})\mu_{j}(\frac{x_{k}}{d_{k}})$ if $N^{l}_{ik}$ is nonzero, meaning that if $x_{i}\in\mathcal{B}_{\psi}$ and $x_{k}\in\mathcal{B}_{\phi}$, then $x_{i}x_{k}$ has all constituents in $\mathcal{B}_{\psi\phi}$. ∎
7.4. Computation of the regular components
For any grading component $\mathcal{B}_{\psi}$, we define (as above):
$R_{\psi}=\sum_{x_{i}\in\mathcal{B}_{\psi}}h_{i}d_{i}x_{i}.$
Lemma 7.12.
Let ${(H,\mathcal{B})}$ be an ARN-hypergroup. For all $\mu_{j}\in{G(\widehat{H})}$,
(7.4) $F_{j}=\frac{1}{n(H)}\sum_{\psi\in\widehat{{G(\widehat{H})}}}\psi(\mu_{j})R_{\psi^{-1}}.$
Proof.
For all $\mu_{j}\in{G(\widehat{H})}$, by Equation (2.4):
$F_{j}$ $=$ $\frac{1}{n_{j}}\sum_{i\in\mathcal{I}}h_{i}\mu_{j}(x_{i^{*}})x_{i}=\frac{1}{n_{j}}\sum_{\psi\in\widehat{{G(\widehat{H})}}}\sum_{x_{i}\in\mathcal{B}_{\psi}}h_{i}\mu_{j}(x_{i^{*}})x_{i}=$
$=$ $\frac{1}{n_{j}}\sum_{\psi\in\widehat{{G(\widehat{H})}}}\psi^{-1}(\mu_{j})R_{\psi}=\frac{1}{n_{j}}\sum_{\psi\in\widehat{{G(\widehat{H})}}}\psi(\mu_{j})R_{\psi^{-1}}.$
Finally, if $\mu_{j}\in{G(\widehat{H})}$ then $n_{j}=n(H)$ by Corollary 2.29. ∎
Corollary 7.13.
Let ${(H,\mathcal{B})}$ be an ARN-hypergroup. The set $\mathcal{B}_{\psi}$ is not empty.
Proof.
Lemma 7.12 shows that the linear span of $\{F_{j}\;|\;\mu_{j}\in{G(\widehat{H})}\}$ lies inside the span of $\{\mathcal{R}_{\psi}\;|\;\psi\in\widehat{{G(\widehat{H})}}\text{ with }\mathcal{B}_{\psi}\neq\emptyset\}$. But $\widehat{{G(\widehat{H})}}\simeq{G(\widehat{H})}$ as finite abelian group, so have same order. Thus, an argument involving the dimension of these vector spaces implies that $\mathcal{B}_{\psi}$ is not empty, for any $\psi\in\widehat{{G(\widehat{H})}}$. ∎
By Equation (7.4) and the second orthogonality relation (2.6) applied on the (hyper)group ${G(\widehat{H})}$, it follows that:
Lemma 7.14.
Let ${(H,\mathcal{B})}$ be an ARN-hypergroup. Then
$R_{\psi}=\frac{n(H)}{|{G(\widehat{H})}|}\big{(}\sum_{\mu_{j}\in{G(\widehat{H})}}\psi(\mu_{j})F_{j}\big{)}.$
Proof of Theorem 7.8 and Theorem 7.9
By Lemma 7.14, and Corollary 6.6, if $\psi=1$, then
$R_{1}=\frac{n(H)}{|{G(\widehat{H})}|}\big{(}\sum_{\mu_{j}\in{G(\widehat{H})}}F_{j}\big{)}=\frac{n(H)}{|{G(\widehat{H})}|}\lambda_{{H_{ad}}}.$
Proposition 6.12 proves that $\mathcal{B}_{1}=\mathcal{B}_{ad}$. Therefore the above grading
$\mathcal{B}=\bigsqcup_{\psi\in\widehat{G(\widehat{H})}}{\mathcal{B}}_{\psi}$
coincides with the universal grading of ${(H,\mathcal{B})}$. Thus $U_{H}={G(\widehat{H})}$. By duality, $U_{\widehat{H}}=G(H)$ if $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN.
7.5. On the perp of the adjoint sub-hypergroup
Recall from Definition 1.6 that an ARN-hypergroup is called dualizable if $^{\widehat{(H,\mathcal{B})}}$ is also (A)RN.
For any sub-hypergroup ${(L,\mathcal{S})}$ of a dualizable ARN-hypergroup ${(H,\mathcal{B})}$, define
$\mathcal{S}^{\perp}:=\{\mu_{j}\;|\;\mu_{j}(s)=\mathrm{FPdim}(s)\;\text{for all}\;s\in\mathcal{S}\}.$
With the above notations, $\mathcal{S}^{\perp}=\bigcap_{s\in\mathcal{S}}\ker_{(H,\mathcal{B})}(s)$. Moreover, [Har79, Proposition 2.11] implies that $\mathcal{S}^{\perp}$ is a sub-hypergroup of $^{\widehat{(H,\mathcal{B})}}$, and $(\mathcal{S}^{\perp})^{\perp}=\mathcal{S}$.
Corollary 7.15.
Let $H$ be a dualizable ARN-hypergroup. With the above notations,
${G(\widehat{H})}^{\perp}=H_{ad}\text{ and }G(H)^{\perp}={(\widehat{H})_{ad}}.$
Thus, by $(\mathcal{S}^{\perp})^{\perp}=\mathcal{S}$,
${G(\widehat{H})}=(H_{ad})^{\perp}\text{ and }G(H)={(\widehat{H})_{ad}}^{\perp}.$
Proof.
Consider the trivial character $\psi_{1}=x_{1}\big{|}_{G(\widehat{H})}$. It follows that
${H_{ad}}=B_{\psi_{1}}$ $=$ $\{x_{i}\;|\;\mu_{j}({\frac{x_{i}}{d_{i}}})=1\;\text{for all}\;\mu_{j}\in{G(\widehat{H})}\}$
$=$ $\bigcap_{\mu_{j}\in{G(\widehat{H})}}\ker_{\widehat{H}}(\mu_{j})={G(\widehat{H})}^{\perp}.$
By duality, we obtain the second equality. ∎
7.6. Quotients of hypergroups
In this subsection, we define the quotient hypergroup of a RN-hypergroup ${(H,\mathcal{B})}$ by a sub-hypergroup ${(L,\mathcal{S})}$. This is the analogue of the quotient construction from [Har79], in the case of probability groups. Recall that a probability group is a normalized RN-hypergroup.
We can define the following equivalence relation on $\mathcal{B}$. For two elements $a,b\in\mathcal{B}$, say $a\sim_{\mathcal{S}}b$ if and only if there are $s_{1},s_{2}\in{\mathcal{S}},x\in{\mathcal{B}}$ such that $m(x,as_{1})>0$ and $m(x,s_{2}b)>0$.
Let $[a]_{\mathcal{S}}$ denote the equivalence class of any $a\in\mathcal{B}$, with respect to $\sim_{\mathcal{S}}$. Define $H//{\mathcal{S}}$ as the set of all these equivalence classes of $\sim_{\mathcal{S}}$. Recall that $\lambda_{\mathcal{S}}$ is the central idempotent of ${\mathcal{S}}$ corresponding to $\mu_{1}=\mathrm{FPdim}$.
For two elements $a,b\in\mathcal{B}$, we can see that $[a]_{\mathcal{S}}=[b]_{\mathcal{S}}$ if and only if $\lambda_{\mathcal{S}}a\lambda_{\mathcal{S}}=\lambda_{\mathcal{S}}b\lambda_{\mathcal{S}}$. Therefore there is a set bijection
$\begin{array}[]{cccc}\phi:&\mathbb{C}[H//{\mathcal{S}}]&\to&\lambda_{\mathcal{S}}H\lambda_{\mathcal{S}}\\ &[a]_{\mathcal{S}}&\mapsto&\lambda_{\mathcal{S}}a\lambda_{\mathcal{S}}\end{array}$
Then, it is easy to verify that $H//{\mathcal{S}}$ becomes a RN-hypergroup, with the multiplication inherited from $\lambda_{\mathcal{S}}H\lambda_{\mathcal{S}}$, via the above isomorphism. We denote by $\overline{m}({[c]},\;[a][b])$ the multiplicity structure of $H//{\mathcal{S}}$. Therefore,
$[a][b]=\sum_{[c]\in H//{\mathcal{S}}}\overline{m}({[c]},\;[a][b])[c].$
We write shortly $[a]$ instead of $[a]_{\mathcal{S}}$ when no confusion is possible. Following [Har79], if ${(H,\mathcal{B})}$ is abelian, we can show that
(7.5) $\overline{m}({[c]},\;[a][b])=\sum_{w\in[c]}m(w,ab).$
It was proven in [Har79, Proposition 2.11] that if ${(H,\mathcal{B})}$ is an abelian dualizable probability group then,
$\begin{array}[]{cccc}\alpha:&{\mathcal{S}}^{\perp}&\to&\widehat{H//{\mathcal{S}}}\\ &\psi&\mapsto&\alpha(\psi)\end{array},$
with $\alpha(\psi)([a]_{\mathcal{S}}):=\psi(a)$, is an isomorphism of probability groups. It is clear that the same results remains true for dualizable ARN-hypergroups.
7.7. Applying Harrison’s results
In this subsection, we also use the short notation $H$ (or $\mathcal{B}$) for a hypergroup $(H,\mathcal{B})$, if $\mathcal{B}$ (or $H$) are implicitly understood. For any sub-hypergroup ${(L,\mathcal{S})}$ of a dualizable ANR-hypergroup ${(H,\mathcal{B})}$, note that [Har79, Proposition 2.11] implies that there is an isomorphism of hypergroups:
$\widehat{H}//\mathcal{S}^{\perp}\simeq\widehat{\mathcal{S}},\;\;[\mu]_{\mathcal{S}^{\perp}}\mapsto\mu\big{|}_{L}.$
In particular, for $\mathcal{S}=H_{ad}$,
(7.6) $\widehat{H}//G(\widehat{H})\simeq\widehat{H_{ad}}$
Dually, for ${\mathcal{S}}={(\widehat{H})_{ad}}$, we obtain that
(7.7) $H//G(H)\simeq\widehat{{(\widehat{H})_{ad}}}$
For $S=G(H)$, then $\widehat{H}//{(\widehat{H})_{ad}}\simeq\widehat{G(H)}$, and dually, for ${\mathcal{S}}={G(\widehat{H})}$, then $H//H_{ad}\simeq\widehat{{G(\widehat{H})}}.$
8. On lower and upper central series of hypergroups
Let $H$ be a RN-hypergroup. Following [GN08], let $H^{(0)}=H$, $H^{(1)}={H_{ad}}$, and $H^{(n)}=H^{(n-1)}_{ad}$, for all $n\geq 1$.
Definition 8.1.
The non-increasing sequence
$H=H^{(0)}\supseteq H^{(1)}\supseteq\dots\supseteq H^{(n)}\supseteq\dots$
will be called the upper central series.
Definition 8.2.
Let ${(L,\mathcal{S})}$ be a sub-hypergroup of an ARN-hypergroup ${(H,\mathcal{B})}$. Let ${\mathcal{S}}^{co}$ be the set of standard elements $x\in\mathcal{B}$ such that $xx^{*}\in\mathcal{{\mathcal{S}}}$. The commutator ${(L,\mathcal{S})}^{{co}}$ of ${(L,\mathcal{S})}$ in ${(H,\mathcal{B})}$ as the sub-hypergroup generated by ${\mathcal{S}}^{co}$.
As in [GN08, Remark 4.9], it follows that the linear span of ${\mathcal{S}}^{co}$ is already a sub-hypergroup of $H$. Moreover, [GN08, Lemma 4.15] works as well in the settings of ARN-hypergroups. Thus,
(8.1) $(\mathcal{S}^{{co}})_{ad}\subseteq\mathcal{S}\subseteq(\mathcal{S}_{ad})^{{co}}$
for any sub-hypergroup $\mathcal{S}$ of $\mathcal{B}$. By applying $\perp$ to (8.1), we obtain
(8.2) $\big{(}(\mathcal{S}^{{co}})_{ad}\big{)}^{\perp}\supseteq\mathcal{S}^{\perp}\supseteq\big{(}(\mathcal{S}_{ad})^{{co}}\big{)}^{\perp}.$
For an ARN-hypergroup ${(H,\mathcal{B})}$, define $H_{(0)}=\mathbb{C}$ and $H_{(n)}=(H_{(n-1)})^{{co}}$, for all $n\geq 1$. Then,
$H_{(1)}=\mathbb{C}^{{co}}=G{(H,\mathcal{B})}=H_{pt}.$
Definition 8.3.
Let $H$ be an ARN-hypergroup. The non-decreasing sequence
$\mathbb{C}=H_{(0)}\subseteq H_{(1)}\subseteq\dots\subseteq H_{(n)}\subseteq\dots$
will be called the lower central series of $H$.
Similarly to [GN08, Theorem 4.16], we can prove that $H^{(n)}=\mathbb{C}$ if and only if $H_{(n)}=H$, where $H$ is an ARN-hypergroup.
Definition 8.4.
An ARN-hypergroup is termed nilpotent if there exists some $n\in\mathbb{N}$ such that $H^{(n)}=\mathbb{C}$. The smallest such $n$ for which this condition is satisfied is referred to as the nilpotency class of $H$.
Lemma 8.5.
Let ${(H,\mathcal{B})}$ be a dualizable ARN-hypergroup. For any sub-hypergroup ${(L,\mathcal{S})}$,
(8.3) $(\mathcal{S}^{\perp})_{ad}\subseteq(\mathcal{S}^{co})^{\perp}.$
Proof.
It reduces to show that $\mu_{j}\star\mu_{j^{\#}}\in({\mathcal{S}}^{co})^{\perp}$, for all $\mu_{j}\in{\mathcal{S}}^{\perp}$.
Let $x\in{\mathcal{S}}^{co}$ be a standard element. It follows that $xx^{*}\in{\mathcal{S}}$, and therefore, $\mu_{j}(x){\mu_{j^{\#}}}(x)=\mu_{j}(xx^{*})=\mathrm{FPdim}(x)^{2}$, for all $\mu_{j}\in{\mathcal{S}}^{\perp}$. This implies
$[\mu_{j}\star\mu_{j^{\#}}]({\frac{x}{\mathrm{FPdim}(x)}})$ $=$ $\mu_{j}({\frac{x}{\mathrm{FPdim}(x)}})\mu_{j^{\#}}({\frac{x}{\mathrm{FPdim}(x)}})$
$=$ $\frac{1}{\mathrm{FPdim}(x)^{2}}\mu_{j}(x)\mu_{j^{\#}}(x)=1,$
which shows that $\mu_{j}\star{\mu_{j^{\#}}}\in({\mathcal{S}}^{co})^{\perp}$. Thus $({\mathcal{S}}^{\perp})_{ad}\subseteq(S^{co})^{\perp}$. ∎
Proposition 8.6.
Let ${(H,\mathcal{B})}$ be a dualizable ARN-hypergroup. With the above notations, for all $n\geq 0$,
(8.4) $\widehat{H}^{(n)}\subseteq(H_{(n)})^{\perp}$
and
(8.5) $H^{(n)}\subseteq(\widehat{H}_{(n)})^{\perp}.$
Proof.
We will prove the first inclusion by induction on $n$. If $n=0$, then
$\widehat{H}^{(0)}=\widehat{H}=\mathbb{C}^{\perp}=H_{(0)}^{\perp}.$
Now, suppose that ${\widehat{H}^{(n)}}\subseteq{H_{(n)}}^{\perp}$. Then
$\widehat{H}^{(n+1)}=({\widehat{H}^{(n)}})_{ad}\subseteq({H_{(n)}}^{\perp})_{ad}\subseteq({H_{(n)}}^{co})^{\perp}=H_{(n+1)}^{\perp}.$
By duality, we obtain the second inclusion. ∎
Observe that the equality holds for $n=1$ also, because
$\widehat{H}^{(1)}={(\widehat{H})_{ad}}=G(H)^{\perp}=H_{(1)}^{\perp}.$
8.1. Proof of Theorem 1.14
Proof.
If ${(H,\mathcal{B})}$ is nilpotent then $H_{(n)}=H$, for some $n$. Then, $\widehat{H}^{(n)}=\mathbb{C}$ by Equation (8.4), and therefore $^{\widehat{(H,\mathcal{B})}}$ is nilpotent. Idem for the converse with Equation (8.5).
8.2. Burnside property for nilpotent ARN-hypergroups
Let ${(H,\mathcal{B})}$ be an ARN-hypergroup. Recall that we call $x_{i}\in\mathcal{B}$ a vanishing element if there is $\mu_{j}\in{\widehat{\mathcal{B}}}$ such that $\mu_{j}(x_{i})=0$.
Lemma 8.7.
Let ${(H,\mathcal{B})}$ be an ARN-hypergroup and ${(L,\mathcal{S})}$ be a sub-hypergroup of ${(H,\mathcal{B})}$. Consider a standard element $x_{i}\in{\mathcal{B}}$ such that $[x_{i}]$ is a vanishing element in $H//{\mathcal{S}}$. Then, $x_{i}$ is also a vanishing element in ${(H,\mathcal{B})}$.
Proof.
Recall that $H//{\mathcal{S}}\simeq\lambda_{\mathcal{S}}H\lambda_{\mathcal{S}}$ as hypergroups, with $[x]\mapsto\lambda_{\mathcal{S}}x\lambda_{\mathcal{S}}$. If $\lambda_{\mathcal{S}}=\sum_{j\in\mathcal{I}_{\mathcal{S}}}F_{j}$ (Definition 6.3), then $\lambda_{\mathcal{S}}H\lambda_{\mathcal{S}}=\oplus_{j\in\mathcal{I}_{\mathcal{S}}}\mathbb{C}F_{j}$, as a subalgebra of $H$. Thus, the characters $\mu_{j}$, with $j\in\mathcal{I}_{\mathcal{S}}$, have distinct restriction to $\lambda_{\mathcal{S}}H\lambda_{\mathcal{S}}$ (since $\mu_{i}(F_{j})=\delta_{i,j}$), and so cover all the characters of $\lambda_{\mathcal{S}}H\lambda_{\mathcal{S}}$ (since $\dim_{\mathbb{C}}(\lambda_{\mathcal{S}}H\lambda_{\mathcal{S}})=|I_{\mathcal{S}}|$). If $[x_{i}]$ is a vanishing element in $H//{\mathcal{S}}$, then $\lambda_{\mathcal{S}}x_{i}\lambda_{\mathcal{S}}$ is a vanishing element in $\lambda_{\mathcal{S}}H\lambda_{\mathcal{S}}$, thus by above, $x_{i}$ is a vanishing element in $H$. ∎
Lemma 8.8.
Let ${(H,\mathcal{B})}$ be an ARN-hypergroup. A grouplike element $x_{i}\in{G{(H,\mathcal{B})}}$ is a constituent of $x_{j}x_{j^{*}}$ if and only if $x_{i}x_{j}=\mathrm{FPdim}(x_{i})x_{j}$. The set of all grouplike elements that are constituents of $x_{j}x_{j^{*}}$ form a subgroup of ${G{(H,\mathcal{B})}}$.
Proof.
By Frobenius reciprocity (2.18), $N_{j,j^{*}}^{i}=\frac{h_{j}}{h_{i}}N_{i,j}^{j}$, thus $x_{i}$ is a constituent of $x_{j}x_{j^{*}}$ if and only if $x_{j}$ is a constitutent of $x_{i}x_{j}$, if and only if (by Lemma 2.34) $\frac{x_{i}}{d_{i}}\frac{x_{j}}{d_{j}}=\frac{x_{j}}{d_{j}}$, meaning that $x_{i}x_{j}=\mathrm{FPdim}(x_{i})x_{j}$. The second assertion is a straightforward consequence of this. ∎
Proof of Theorem 1.15 By Theorem 1.14, it is enough to prove the following:
Proposition 8.9.
A nilpotent dualizable ARN-hypergroup ${(H,\mathcal{B})}$ is Burnside.
Proof.
We will proceed by induction on the nilpotency class of ${(H,\mathcal{B})}$.
If ${(H,\mathcal{B})}$ is of nilpotency class 111, then ${(H,\mathcal{B})}$ is pointed and we are done.
Now, suppose that ${(H,\mathcal{B})}$ is nilpotent of class $n$ and that the result holds for the nilpotency classes less than $n$. We will show that any standard element $x_{i}\in\mathcal{B}$ is either a vanishing or a grouplike element. Assume that $x_{i}\notin G(H)$. We will show that $x_{i}$ is a vanishing element.
Consider $[x_{i}]\in H//G(H)$. Since $H//G(H)\simeq\widehat{{(\widehat{H})_{ad}}}$ by (7.7), it is nilpotent, with nilpotency class $n-1$, by Theorem 1.14. By induction hypothesis, we know that $[x_{i}]$ is either a grouplike element or a vanishing element in $H//G(H)$.
If $[x_{i}]$ is a vanishing element, then Lemma 8.7 shows that $x_{i}$ is a vanishing element.
Now, if $[x_{i}]$ is a grouplike element of $H//G(H)$ and by abelian assumption,
$\lambda_{G(H)}=h_{[x_{i}]}\lambda_{G(H)}x_{i}\lambda_{G(H)}x_{i}^{*}\lambda_{G(H)}=\lambda_{G(H)}h_{[x_{i}]}x_{i}x_{i}^{*}\lambda_{G(H)},$
which implies that $h_{[x_{i}]}x_{i}x_{i}^{*}\in G(H)$. Thus, $h_{[x_{i}]}x_{i}x_{i}^{*}=\sum_{g\in G_{1}}g$, for a subgroup $G_{1}\subset G$ (by Lemma 8.8). Since $x_{i}$ is not a grouplike element in $H$, $G_{1}\neq\{1\}$. Then, there is $\mu_{j}$ such that $\mu_{j}(\Lambda_{G_{1}})=0$ where $\Lambda_{G_{1}}=\frac{1}{|G_{1}|}\big{(}\sum_{x\in G_{1}}x\big{)}$ is the integral of $G_{1}$. This implies that $\mu_{j}(x_{i}x_{i}^{*})=0$. Thus $\mu_{j}(x_{i})=0$, which means that $x_{i}$ is a vanishing element. ∎
Now, since $H$ and $\widehat{H}$ are simultaneously nilpotent, Theorem 1.15 follows.
9. Applications to fusion categories
Let $\mathcal{C}$ be a pivotal fusion category with a commutative Grothendieck ring. As in §1, we denote the set of isomorphism class representatives of simple objects of $\mathcal{C}$ by $\mathrm{Irr}(\mathcal{C}):=\{X_{1},\dots,X_{m}\}$ and let :=$\mathcal{I}:=\{1,\dots,m\}$. Let also $d_{i}:=\dim(X_{i})$ be the categorical dimension of $X_{i}$ for all $i\in\mathcal{I}$. For any simple object $X_{i}\in\mathrm{Irr}(\mathcal{C})$ denote by $x_{i}:=[X_{i}]$ the class of $X_{i}$ in the Grothendieck ring $K(\mathcal{C})$ of $\mathcal{C}$.
Let ${\mathcal{Z}(\mathcal{C})}$ be the Drinfeld center of the fusion category $\mathcal{C}$. The forgetful functor $F:{\mathcal{Z}(\mathcal{C})}\rightarrow\mathcal{C}$ admits a right adjoint functor $R:\mathcal{C}\rightarrow{\mathcal{Z}(\mathcal{C})}$. It is well-known that $A:=FR(\mathbf{1})$ has the structure of a central commutative algebra in $\mathcal{C}$ (meaning that $R(\mathbf{1})$ is a commutative algebra in ${\mathcal{Z}(\mathcal{C})}$), where $\mathbf{1}$ is the unit object of $\mathcal{C}$. The vector spaces
$\mathrm{{CE}(\mathcal{C})}:=\mathrm{{Hom}}_{\mathcal{C}}(\mathbf{1},A)\text{ and }{\mathrm{CF}(\mathcal{C})}:=\mathrm{{Hom}}_{\mathcal{C}}(A,\mathbf{1})$
are respectively called the space of central elements and the space of class functions of $\mathcal{C}$. Recall [Shi17] that ${\mathrm{CF}(\mathcal{C})}\simeq K(\mathcal{C})_{\mathbb{C}}$, the complexification of the Grothendieck ring $K(\mathcal{C})$ of $\mathcal{C}$.
For any simple object $X_{i}$ of $\mathcal{C}$ we denote by $\chi_{i}:=\mathrm{ch}(X_{i})\in{\mathrm{CF}(\mathcal{C})}$ its associated character [Shi17, Definition 3.9]. The central element space $\mathrm{{CE}(\mathcal{C})}$ has a basis of primitive orthogonal idempotents $E_{i}$ such that $<\chi_{i},E_{j}>=\delta_{i,j}d_{i}$ for all $1\leq i,j\leq m$.
As explained in [Shi17, Theorem 3.8] the adjunction between $F$ and $R$ gives a canonical isomorphism of algebras
(9.1) ${\mathrm{CF}(\mathcal{C})}\xrightarrow{\cong}\mathrm{End}_{{\mathcal{Z}(\mathcal{C})}}(R(\mathbf{1})),$
where ${\mathrm{CF}(\mathcal{C})}$ is the algebra of class functions on $\mathcal{C}$. Since ${\mathcal{Z}(\mathcal{C})}$ is also a fusion category we can write $R(\mathbf{1})=\bigoplus_{j\in\mathcal{I}}\mathcal{C}^{j}$ as a direct sum of simple objects in ${\mathcal{Z}(\mathcal{C})}$. Recall that $\mathcal{C}^{j}$ are called conjugacy classes for $\mathcal{C}$. The above isomorphism also gives a canonical bijection between the set of primitive central idempotents $\{F_{j}\}_{j\in\mathcal{I}}$ and the set of conjugacy classes $\{\mathcal{C}^{j}\}$.
For any fusion category $\mathcal{C}$, by abuse of notation, the RN-hypergroup ${(H,\mathcal{B},\mu_{1})}:=(K(\mathcal{C}),\mathrm{Irr}(\mathcal{C}),\mathrm{FPdim})$ is also denoted by $K(\mathcal{C})$. We denote by $^{\widehat{{K(\mathcal{C})}}}$ its dual hypergroup.
Recall also from [GN08] the notion of a universal group grading of a fusion category $\mathcal{C}$ and its adjoint subcategory ${\mathcal{C}_{\mathrm{ad}}}$.
Remark 9.1.
By [Bur20, Equation (4.8)], it follows that for any spherical fusion category, $\dim(\mathcal{C}^{j})=\frac{\dim(\mathcal{C})}{n_{j}}$. See also [Ost15, Theorem 2.13] for a related statement. Then Equation (2.11) implies that
(9.2) $\widehat{h}_{j}=\dim(\mathcal{C}^{j})$
in the dual hypergroup $^{\widehat{{K(\mathcal{C})}}}$. In particular, for a weakly-integral fusion category $\mathcal{C}$, by [ENO05, Proposition 8.27], $\dim(\mathcal{C}^{j})\in\mathbb{Z}$, and therefore $^{\widehat{{K(\mathcal{C})}}}$ is $h$-integral.
Corollary 9.2.
Let $\mathcal{C}$ be a a fusion category with a Burnside commutative Grothendieck ring and let $^{\widehat{{K(\mathcal{C})}}}$ be its dual. Then
(9.3) $\prod_{j=1}^{m}\mu_{j}=\sum_{X_{i}\in{\mathcal{C}_{\mathrm{pt}}}}{\mathrm{sgn}}(x_{i})\widehat{F}_{i}$
where ${\mathrm{sgn}}(x_{i})$ is the determinant of the permutation matrix $L_{{\frac{x_{i}}{d_{i}}}}$ on $K(\mathcal{C})$.
Proof.
Since $K(\mathcal{C})$ is a Burnside hypergroup the result follows from Corollary 4.4. ∎
Note that the dual hypergroup $^{\widehat{{K(\mathcal{C})}}}$ is denoted by $\widehat{{\mathrm{CF}(\mathcal{C})}}$ in [Bur21]. Theorem 3.4 from the same paper implies that for any pivotal fusion category there is a canonical isomorphism of $\mathbb{C}$-algebra
(9.4) $\alpha:\widehat{{\mathrm{CF}(\mathcal{C})}}\rightarrow\mathrm{{CE}(\mathcal{C})},\;\mu_{j}\mapsto\frac{C_{j}}{\dim(\mathcal{C}^{j})},$
where $C_{j}\in\mathrm{{CE}(\mathcal{C})}$ is the conjugacy class sum corresponding to the conjugacy class $\mathcal{C}^{j}$, defined for a pivotal fusion category by Shimizu as ${\mathcal{F}}^{-1}(F_{j})$. Here $\lambda\in{\mathrm{CF}(\mathcal{C})}$ is a cointegral of $\mathcal{C}$ such that $\langle\lambda,u\rangle=1$, see [Shi17, §5]. Recall also the Fourier transform of $\mathcal{C}$ associated to $\lambda$ is the linear map
(9.5) $\mathcal{F}_{\lambda}:\mathrm{{CE}(\mathcal{C})}\rightarrow{\mathrm{CF}(\mathcal{C})}\;\;\text{given by}\;\;a\mapsto\lambda\leftharpoonup\mathcal{S}(a)$
By [Bur20, Lemma 4.1 and (4.7)], for any spherical fusion category $\mathcal{C}$ over $\mathbb{C}$ with a commutative Grothendieck ring, we have
$\langle F_{i},C_{j}\rangle=\delta_{i,j}\dim(\mathcal{C}^{j}).$
This relation implies that the set $(F_{i})$ forms a dual basis to $(C_{j}/\dim(\mathcal{C}^{j}))$ with respect to the bilinear form $\langle$,$\rangle$ , ⟩. Consequently, this provides an easier way for defining the conjugacy class sums $(C_{j})$ in this case (see [Bur20, (4.11)] for a formula).
Corollary 9.3.
For any weakly-integral fusion category $\mathcal{C}$ with a commutative Grothendieck ring the following identity holds in $\mathrm{{CE}(\mathcal{C})}$:
(9.6) $\prod_{j=1}^{m}\frac{C_{j}}{\dim(\mathcal{C}^{j})}=\sum_{X_{i}\in{\mathcal{C}_{\mathrm{pt}}}}{\mathrm{sgn}}(x_{i})E_{i}$
where $E_{i}\in\mathrm{{CE}(\mathcal{C})}$ is the primitive central idempotent of $X_{i}\in{\mathcal{C}_{\mathrm{pt}}}$.
Proof.
By Corollary 4.21, we can apply Corollary 9.2. The result follows by applying the canonical isomorphism $\alpha$. ∎
For a spherical fusion category note the following:
Proposition 9.4.
Let $\mathcal{C}$ be a spherical fusion category with a commutative Grothendieck ring ${K(\mathcal{C})}$ and RN dual. Let $\mu_{j}\in{\widehat{{K(\mathcal{C})}}}$ be a linear character. Then $\mu_{j}$ is a grouplike element of $^{\widehat{{K(\mathcal{C})}}}$ if and only if $\dim(\mathcal{C}^{j})=1$.
Proof.
By Lemma 2.24 and the fact that $^{\widehat{{K(\mathcal{C})}}}$ is normalized, $\mu_{j}$ is a grouplike element if and only if ${\widehat{h}}_{j}=1$. This is also equivalent to $\dim(\mathcal{C}^{j})=1$, by Equation (9.2). ∎
Now, suppose that $\mathcal{C}$ is a pivotal fusion category and ${\mathcal{D}}\subseteq\mathcal{C}$ is a fusion subcategory. Following the notion of support defined in Definition 6.3, there exists a subset $\mathcal{I}_{\mathcal{D}}\subseteq\mathcal{I}$ such that
$\lambda_{{\mathcal{D}}}=\sum_{j\in\mathcal{I}_{\mathcal{D}}}F_{j},$
Please note that this set is called $\mathcal{L}_{\mathcal{D}}$ in [Bur20, §4.2]. Additionally, [Bur20, Lemma 4.6] implies
(9.7) ${\mathcal{I}}_{{}_{{\mathcal{C}_{\mathrm{ad}}}}}=\{j\in\mathcal{I}\;|\;\dim(\mathcal{C}^{j})=1\},$
for any fusion category with a commutative Grothendieck ring ${K(\mathcal{C})}$.
Proposition 9.5.
Let $\mathcal{C}$ be a fusion category with a commutative Grothendieck ring. Then $^{\widehat{{K(\mathcal{C})}}}$ is Burnside if and only if:
(9.8) $\prod_{i=1}^{m}{\frac{x_{i}}{d_{i}}}=\sum_{j\in{{\mathcal{I}}_{{}_{{\mathcal{C}_{\mathrm{ad}}}}}}}{\mathrm{sgn}}(\mu_{j})F_{j}$
Proof.
It follows from Corollary 4.8 and Theorem (6.5). ∎
Proof of Theorem 1.12
Proof.
It follows from Corollary 4.9 and Theorem (6.5), together with
$\sum_{j\in{{\mathcal{I}}_{{}_{{\mathcal{C}_{\mathrm{ad}}}}}}}F_{j}=\lambda_{{\mathcal{C}_{\mathrm{ad}}}}=\frac{1}{\dim({\mathcal{C}_{\mathrm{ad}}})}\big{(}\sum_{x_{i}\in{\mathcal{C}_{\mathrm{ad}}}}d_{i}x_{i}\big{)},$
where the last equality comes from Equation (5.2). ∎
Here is a reformulation of Theorem 1.12 for the case of a nilpotent finite group:
Theorem 9.6.
For any finite nilpotent group $G$, the dual $\widehat{\text{ch}(G)}$ is Burnside and therefore
(9.9) $\left(\prod_{x_{i}\in\mathrm{Irr}(G)}{\frac{x_{i}}{d_{i}}}\right)^{2}=\frac{|Z(G)|}{|G|}\left(\sum_{x_{i}\in\mathrm{Irr}(G/Z(G))}d_{i}x_{i}\right).$
Proof.
Recall that $K(\text{Rep}(G))$ is the character ring $\text{ch}(G)$. The result that $\widehat{\text{ch}(G)}$ is a RN hypergroup is established in [Bur21, Equation (3.12)]. Furthermore, [INW99, Theorem B] states that if $G$ is a nilpotent group, then $\text{ch}(G)$ is dual-Burnside. ∎
Theorem 1.15 leads to the following conclusion:
Corollary 9.7.
If $\mathcal{C}$ is a nilpotent fusion category with a commutative Grothendieck ring and RN dual, then both ${K(\mathcal{C})}$ and $^{\widehat{{K(\mathcal{C})}}}$ are Burnside.
Remark 9.8.
A nilpotent fusion category is weakly-integral [GN08], which means it is also pseudo-unitary and spherical [EGNO15]. It remains an open question whether every pseudo-unitary fusion category can be given a unitary structure [EGNO15, Remark 9.4.7]. However, a commutative fusion ring with a unitary categorification does have a RN dual [LPW21]. Therefore, we believe that the assumption of a RN dual in Corollary 9.7 could be omitted. Additionally, in the braided case, its Grothendieck ring possesses a RN dual according to [Bur21, Theorem 1.2].
10. Premodular categories
Recall that a premodular category is defined as a braided spherical fusion category. In the rest of this section, we will assume that $\mathcal{C}$ is a premodular category. By [Shi17, Example 6.14] there is $\mathbb{C}$-algebra map $\mathrm{f}_{Q}:{{\mathrm{CF}(\mathcal{C})}}\rightarrow\mathrm{{CE}(\mathcal{C})}$ given by the following formula:
(10.1) $\mathrm{f}_{Q}(\chi_{i})=\sum_{j\in\mathcal{I}}\frac{s_{ij}}{d_{j}}E_{j}.$
where $S=(s_{ij})$ is the $S$-matrix of $\mathcal{C}$ and $(E_{j})$ are the primitive central idempotents of $\mathrm{{CE}(\mathcal{C})}$ as defined in the previous section.
As described in [Bur22, §4], there is a map $M:\mathcal{I}\to\mathcal{I}$ (referred to as the braided partition function) such that if
$\mathrm{f}_{Q}(F_{j})=\sum_{i\in\mathcal{B}_{j}}E_{i},$
then $M(i)=j$. Let $\mathcal{I}_{2}\subseteq\mathcal{I}$ denote the set of all indices $j$ for which $\mathrm{f}_{Q}(F_{j})\neq 0$, meaning that $\mathcal{B}_{j}$ is not an empty set. Given that $\mathrm{f}_{Q}(1)=1$, we can thus partition the set of all isomorphism classes of simple objects $\text{Irr}(\mathcal{C})=\bigsqcup_{j\in\mathcal{I}_{2}}\mathcal{R}_{j}$, where $\mathcal{R}_{j}=\{[X_{i}]\;|\;i\in\mathcal{B}_{j}\}$. In other words, $\mathcal{R}_{j}=\{[X_{i}]\;|\;M(i)=j\}$. Thus, we obtain a unique function $M:\mathcal{I}\to\mathcal{I}_{2}$ with the property that $E_{i}\mathrm{f}_{Q}(F_{M(i)})\neq 0$ for all $i\in\mathcal{I}$.
The paper [BB15] introduces the notion of cosets of a fusion category $\mathcal{C}$ with respect to a fusion subcategory $\mathcal{D}$: two simple objects $X,Y$ in $\mathcal{C}$ are in the same (right) coset if and only if there is a simple object $S$ in $\mathcal{D}$ such that $X$ is a constituent of $Y\otimes S$. By [Bur22, Theorem 4.10] two simple objects $X_{i},X_{i^{\prime}}$ of a pseudo-unitary premodular fusion category $\mathcal{C}$ are in the same coset with respect to $\mathcal{C}^{\prime}:=\mathcal{Z}_{2}(\mathcal{C})$ if and only if $M(i)=M(i^{\prime})$. In other words, $(\mathcal{R}_{j})_{j\in\mathcal{I}_{2}}$ correspond exactly to these cosets. Denote also ${\mathrm{R}}_{j}:=\sum_{[X_{i}]\in\mathcal{R}_{j}}d_{i}\chi_{i}\in{\mathrm{CF}(\mathcal{C})}$ the regular part of their characters. By [Bur22, Equation (4.20)],
(10.2) $\dim({\mathrm{R}}_{j})=\dim({\mathcal{C}^{\prime}})\dim(\mathcal{C}^{j}),\;\text{for all}\;j\in\mathcal{I}_{2}.$
Proposition 10.1.
Suppose that $\mathcal{C}$ is a weakly-integral premodular category such that $\mathcal{C}^{\prime}\subseteq{\mathcal{C}_{\mathrm{pt}}}$ and acting freely on the set $\mathrm{Irr}(\mathcal{C})$. Then for any $j\in\mathcal{I}_{2}$ such that $\dim(\mathcal{C}^{j})>1$ there is an $x_{i}$ such that $\mu_{j}(x_{i})\neq 0$.
Proof.
First of all, ${K(\mathcal{C})}$ is $h$-integral because as a fusion ring, $h_{i}=1$ for all $i$. If $\mathcal{C}$ is weakly integral, then the dual $^{\widehat{{K(\mathcal{C})}}}$ is a rational RN hypergroup, as shown by [Bur21, Equation (4.3)]. Therefore, we can attempt to apply Theorem 4.22 for ${(H,\mathcal{B},\mu_{1})}={K(\mathcal{C})}$. To do this, we need to ensure that the additional condition (4.18) holds for all $i\in\mathcal{I}$, specifically that $\dim(\mathcal{C}^{j})\frac{|{\alpha_{{}_{ij}}}|^{2}}{d_{i}^{2}}\in\mathbb{A}$ for all $i\in\mathcal{I}$, where ${\alpha_{{}_{ij}}}:=\mu_{j}(x_{i})$. It was demonstrated in [Bur22, Lemma 4.2] that
(10.3) $\frac{\alpha_{{}_{iM(i^{\prime})}}}{d_{i}}=\frac{s_{{}_{ii^{\prime}}}}{d_{i}d_{i^{\prime}}}=\frac{\alpha_{{}_{i^{\prime}M(i)}}}{d_{i^{\prime}}}.$
for all $i,i^{\prime}\in\mathcal{I}$. Fix an index $i_{j}\in\mathcal{I}$ such that $M(i_{j})=j\in\mathcal{I}_{2}$, and take $i^{\prime}=i_{j}$. It follows that
(10.4) ${\alpha_{{}_{ij}}}=\frac{d_{i}}{d_{i_{j}}}\alpha_{i_{j}M(i)}.$
Therefore
[…]
But $\mathcal{C}$ is spherical, so $d_{i}=\overline{d_{i}}$ and therefore $|d_{i}|^{2}=d_{i}^{2}$. Thus
(10.6) $\dim(\mathcal{C}^{j})\frac{|{\alpha_{{}_{ij}}}|^{2}}{d_{i}^{2}}=\frac{\dim(\mathcal{C}^{j})}{d_{i_{j}}^{2}}\mid\alpha_{i_{j}M(i)}\mid^{2}$
But $\mathcal{C}^{\prime}\subseteq{\mathcal{C}_{\mathrm{pt}}}$, acting freely on $\mathrm{Irr}(\mathcal{C})$, thus $|\mathcal{R}_{j}|=\dim(\mathcal{C}^{\prime})$, and so by Equation (10.2), $\dim(\mathcal{C}^{j})=\frac{\dim(R_{j})}{\dim(\mathcal{C}^{\prime})}=d_{i_{j}}^{2}$, and the proof follows since $\alpha_{i_{j}M(i)}\in\mathbb{A}$. ∎
Roughly speaking, Proposition 10.1 can be intuitively understood as follows: a weakly integral premodular fusion category $\mathcal{C}$ is dual-Burnside with respect to $\mathcal{C}^{\prime}$.
10.1. Modular category case
In this section, let $\mathcal{C}$ denote a modular fusion category. We will utilize the results from the previous section to $K(\mathcal{C})$ and prove Theorems 1.16 and 1.19. Recall the definition of $\mathcal{I}_{{}_{\mathcal{D}}}$ as provided in Definition (6.3).
By composing Drinfeld’s map $\mathrm{f}_{Q}:{\mathrm{CF}(\mathcal{C})}\rightarrow\mathrm{{CE}(\mathcal{C})}$ with the inverse of the isomorphism $\alpha:\mathrm{{CE}(\mathcal{C})}\rightarrow\widehat{{\mathrm{CF}(\mathcal{C})}}$ from Equation (9.4), we obtain an algebra isomorphism $\widehat{\mathrm{f}}_{Q}:{\mathrm{CF}(\mathcal{C})}\rightarrow\widehat{{\mathrm{CF}(\mathcal{C})}}$. This can also be expressed as an algebra isomorphism $\widehat{\mathrm{f}}_{Q}:{K(\mathcal{C})}\rightarrow{\widehat{{K(\mathcal{C})}}}$. It was shown in [Bur21, Remark 4.2] that this is, in fact, an isomorphism of normalized hypergroups: $\widehat{\mathrm{f}}_{Q}:\overline{{K(\mathcal{C})}}\rightarrow{\widehat{{K(\mathcal{C})}}}$, where $\overline{{K(\mathcal{C})}}$ denotes the normalized version of ${K(\mathcal{C})}$, so that $\widehat{\mathrm{f}}_{Q}(\frac{x_{i}}{d_{i}})=\mu_{i}$.
Remark 10.2.
Drinfeld’s map $\mathrm{f}_{Q}$ is an algebra isomorphism since the S-matrix is invertible. It establishes a canonical bijection between the index sets of $\{F_{i}\}$ and $\{E_{i}\}$ which is given by $\mathrm{f}_{Q}(F_{i})=E_{i}$, for all $i\in\mathcal{I}$.
Corollary 10.3.
Let $\mathcal{C}$ be a modular fusion category. Then $K(\mathcal{C})$ is Burnside if and only if it is dual-Burnside.
Proof.
As mentioned above $\widehat{\mathrm{f}}_{Q}:\overline{{K(\mathcal{C})}}\rightarrow{\widehat{{K(\mathcal{C})}}}$ is an isomorphism of normalized hypergroup. Therefore $\overline{{K(\mathcal{C})}}$ (and thus ${K(\mathcal{C})}$) is Burnside if and only if $^{\widehat{{K(\mathcal{C})}}}$ is also Burnside.∎
Lemma 10.4.
Let $\mathcal{C}$ be a modular fusion category. Then with the above notations,
(10.7) $X_{i}\in\mathrm{Irr}({\mathcal{C}_{\mathrm{pt}}})\iff i\in{\mathcal{I}}_{\mathcal{C}_{\mathrm{ad}}}$
Proof.
Apply Theorem 6.5 with the isomorphism of normalized hypergroups $\widehat{\mathrm{f}}_{Q}$. ∎
Theorem 10.5.
In any modular fusion category $\mathcal{C}$ with ${K(\mathcal{C})}$ Burnside the following identity holds:
(10.8) $\prod_{i\in\mathcal{I}}\frac{x_{i}}{d_{i}}=\sum_{j\in{\mathcal{I}}_{\mathcal{C}_{\mathrm{ad}}}}{\mathrm{sgn}}(x_{j})F_{j}$
Proof.
Since $K(\mathcal{C})$ is Burnside, we can apply $\widehat{\mathrm{f}}_{Q}^{-1}$ to Equation (9.3), but $\widehat{\mathrm{f}}_{Q}({\frac{x_{i}}{d_{i}}})=\mu_{i}$, so we get
$\prod_{i\in\mathcal{I}}{\frac{x_{i}}{d_{i}}}=\sum_{X_{i}\in{\mathcal{C}_{\mathrm{pt}}}}{\mathrm{sgn}}(x_{i})F_{i}$
and Lemma 10.4 finishes the proof. ∎
Proof of Theorem 1.16
Proof.
Squaring Equation (10.8) we obtain that $\big{(}\prod_{i\in\mathcal{I}}\frac{x_{i}}{d_{i}}\big{)}^{2}=\sum_{j\in{\mathcal{I}}_{\mathcal{C}_{\mathrm{ad}}}}F_{j}=\lambda_{{\mathcal{C}_{\mathrm{ad}}}}$ by the definition of ${{\mathcal{I}}_{{}_{{\mathcal{C}_{\mathrm{ad}}}}}}$. The result follows by Equation (5.2).∎
Recall that the above theorem holds for weakly-integral modular categories since their Grothendieck rings are Burnside by [GNN09, Appendix] or [Bur23, Theorem 2].
Proof of Corollary 1.17
Proof.
In this case every invertible element $x_{j}$ has odd order since this order divides $|{G(\mathcal{C})}|$. Thus every ${\mathrm{sgn}}(x_{j})=1$ for any invertible object $X_{j}\in{\mathcal{C}_{\mathrm{pt}}}$. ∎
Proposition 10.6.
Let $\mathcal{C}$ be a fusion category such that ${K(\mathcal{C})}$ is commutative and the dual $^{\widehat{{K(\mathcal{C})}}}$ is Burnside. Then
$\frac{(\prod_{i\in\mathcal{I}}d_{i})^{2}}{\mathrm{FPdim}({\mathcal{C}_{\mathrm{ad}}})}\in\mathbb{Z}.$
Proof.
Let $P_{\mathcal{C}}$ be $\prod_{i\in\mathcal{I}}d_{i}$. Suppose that $\big{(}\prod_{i\in\mathcal{I}}x_{i}\big{)}^{2}=\sum_{i\in\mathcal{I}}N_{i}x_{i}$ with $N_{i}\in\mathbb{Z}_{\geq 0}$. Then equalizing coefficients in Equation (1.6), we obtain the following $\frac{N_{i}}{P_{\mathcal{C}}^{2}}=\frac{d_{i}}{\mathrm{FPdim}({\mathcal{C}_{\mathrm{ad}}})}$. Take $i=1$ then $N_{1}\mathrm{FPdim}({\mathcal{C}_{\mathrm{ad}}})=P_{\mathcal{C}}^{2}$ which proves the divisibility. ∎
Proof of Theorem 1.18.
Proof.
By [GN08, Theorem 3.10] $d_{i}^{2}\in\mathbb{Z}$ for all $i$ since $\mathcal{C}$ is weakly-integral. Note that since $^{\widehat{{K(\mathcal{C})}}}$ is Burnside the first divisibility result follows from Proposition 10.6. Thus ${\mathrm{FPdim}({\mathcal{C}_{\mathrm{ad}}})}$ is an integer (as a rational algebraic integer). Moreover, if $\mathcal{C}$ is nilpotent then [GN08, Corollary 5.3] implies $d_{i}^{2}\mid{\mathrm{FPdim}({\mathcal{C}_{\mathrm{ad}}})}$. Thus $\mathcal{V}({\mathcal{C}_{\mathrm{ad}}})=\bigcup_{i\in\mathcal{I}}\mathcal{V}(d_{i}^{2})$. ∎
Remark 10.7.
In the case of an integral fusion category Equation (1.10) can be written as
$\mathcal{V}({\mathcal{C}_{\mathrm{ad}}})=\bigcup_{i\in\mathcal{I}}\mathcal{V}(d_{i}).$
Recall that a weakly-integral fusion category with a commutative Grothendieck ring is Burnside by [Bur23, Theorem 2]. Moreover, a Burnside modular fusion category is dual-Burnside by Corollary 10.3. So a weakly-integral modular fusion category is dual-Burnside. Proof of Theorem 1.19
Proof.
The first divisibility follows from Proposition (10.6). Now, let $\mathcal{C}$ be a weakly-integral modular fusion category. By [ENO05, Proposition 8.27] we have that $\mathrm{FPdim}({\mathcal{C}_{\mathrm{ad}}})$ is an integer. By [GN08, Theorem 3.10] $d_{i}^{2}\in\mathbb{Z}$ for all $i$. The second item follows since for a modular category we have $\mathrm{FPdim}(\mathcal{C})=\mathrm{FPdim}({\mathcal{C}_{\mathrm{ad}}})\mathrm{FPdim}({\mathcal{C}_{\mathrm{pt}}})$, because the group associated to ${\mathcal{C}_{\mathrm{pt}}}$ is the universal grading group by [EGNO15, Lemma 8.22.5]. Thus $\mathcal{V}(\mathcal{C})=\mathcal{V}({\mathcal{C}_{\mathrm{ad}}})\cup\mathcal{V}({\mathcal{C}_{\mathrm{pt}}})$. On the other hand, from the first item, $\mathcal{V}({\mathcal{C}_{\mathrm{ad}}})\subseteq\bigcup_{i\in\mathcal{I}}\mathcal{V}(d_{i}^{{2}})$ which proves the inclusion $\mathcal{V}(\mathcal{C})\subseteq\mathcal{V}({\mathcal{C}_{\mathrm{pt}}})\cup\big{(}\bigcup_{i\in\mathcal{I}}\mathcal{V}(d_{i}^{{2}})\big{)}$.
Conversely, $\mathcal{V}({\mathcal{C}_{\mathrm{pt}}})\cup\big{(}\bigcup_{i\in\mathcal{I}}\mathcal{V}(d_{i}^{2})\big{)}\subseteq\mathcal{V}(\mathcal{C})$ since $\mathrm{FPdim}({\mathcal{C}_{\mathrm{pt}}})\mid\mathrm{FPdim}(\mathcal{C})$, and by [EGNO15, Proposition 8.14.6], $d_{i}^{{2}}\mid\mathrm{FPdim}(\mathcal{C})$ for all $i\in\mathcal{I}$. ∎
Proof of Theorem 1.21
Proof.
Since $\gcd(d_{i}^{2},d)=1$ and $d$ is square-free, it follows from Equation (1.13) that $d$ divides $\mathrm{FPdim}({\mathcal{C}_{\mathrm{pt}}})$. Now, consider a pointed fusion subcategory ${\mathcal{D}}\subseteq{\mathcal{C}_{\mathrm{pt}}}$ such that $\mathrm{FPdim}({\mathcal{D}})=d$. The existence of such a subcategory is straightforward to demonstrate because the associated groups are abelian, given that the category is braided. According to [EGNO15, Theorem 8.21.5], with $\mathcal{C}^{\prime}=\mathrm{Vec}$, we have $\mathrm{FPdim}({\mathcal{D}}^{\prime})=m$, but $\gcd(d,m)=1$. Consequently, by Lagrange’s theorem [EGNO15, Theorem 7.17.6], ${\mathcal{D}}\cap{\mathcal{D}}^{\prime}=\mathrm{Vec}$, which implies that ${\mathcal{D}}$ is non-degenerate, as stated in [EGNO15, Corollary 8.20.10]. Thus, by [EGNO15, Theorem 8.21.4], $\mathcal{C}\simeq{\mathcal{D}}\boxtimes{\mathcal{D}}^{\prime}$, and ${\mathcal{D}}^{\prime}$ is also non-degenerate. ∎
Remark 10.8.
Note that Theorem 1.21 improves [OY23, Theorem 4.5] since the authors also assume the existence of a Tannakian subcategory ${\mathcal{E}}=\mathrm{Rep}(G)\subseteq\mathcal{C}$ such that $\mathcal{C}^{0}_{G}\simeq\mathcal{C}(\mathbb{Z}_{d},q)\boxtimes\mathcal{A}$ for some non-degenerate braided fusion category $\mathcal{A}$.
Proof of Corollary 1.22:
Proof.
Immediate from Theorem 1.21, we just need to check that $gcd(d_{i}^{2},d)=1$, but it is clear since $d_{i}^{2}$ divides $\mathrm{FPdim}(\mathcal{C})$ and $d$ is a factor of the square-free part. ∎
Corollary 10.9.
An integral modular fusion category $\mathcal{C}$ can always be decomposed into ${\mathcal{D}}\boxtimes{\mathcal{D}}^{\prime}$, where both ${\mathcal{D}}$ and ${\mathcal{D}}^{\prime}$ are modular, ${\mathcal{D}}$ is pointed, and $\mathrm{FPdim}({\mathcal{D}})$ is the maximal square-free part of $\mathrm{FPdim}(\mathcal{C})$.
This means that the classification of integral modular fusion categories $\mathcal{C}$ simplifies to those where $\mathrm{FPdim}(\mathcal{C})$ lacks a square-free part, up to a Deligne tensor product with a pointed modular fusion category.
Corollary 10.10.
For an integral perfect modular fusion category $\mathcal{C}$, $\mathrm{FPdim}(\mathcal{C})$ has no square-free part. In other words, if a prime $p$ divides $\mathrm{FPdim}(\mathcal{C})$, then $p^{2}$ also divides $\mathrm{FPdim}(\mathcal{C})$.
Corollary 10.10 obviously extends to the weakly-integral case due to the following:
Proposition 10.11.
A weakly-integral perfect modular fusion category $\mathcal{C}$ is integral.
Proof.
Since ${\mathcal{C}_{\mathrm{pt}}}$ is trivial then ${\mathcal{C}_{\mathrm{ad}}}=\mathcal{C}$, because $({\mathcal{C}_{\mathrm{pt}}})^{\prime}={\mathcal{C}_{\mathrm{ad}}}$ by [EGNO15, Corollary 8.22.8], see also Remark 11.13. Moreover, [EGNO15, Exercise 9.6.12] asserts that the adjoint subcategory of a weakly-integral fusion category is integral. ∎
Corollary 10.12.
If $\mathcal{C}$ is a perfect, integral, even-dimensional modular fusion category, then 444 divides $\mathrm{FPdim}(\mathcal{C})$.
10.2. On Statements (1.25) and (1.26) and Conjecture (1.24)
In [LPR23, Theorem 5.6] it was shown that Statements (1.25) and (1.26) are equivalent. We will show that any of these statements implies Conjecture 1.24 formulated in [CP22].
Proposition 10.13.
The existence of a perfect odd dimensional modular fusion category implies the existence of a non-pointed simple integral modular fusion category.
Proof.
Let $\mathcal{C}$ be a perfect modular fusion category with an odd dimension. By Proposition 10.11, $\mathcal{C}$ must be integral.
[…]
However, due to the finite number of isomorphism classes of simple objects, $\mathcal{C}$ must contain at least one minimal subcategory. Let us call this minimal subcategory ${\mathcal{D}}$. If ${\mathcal{D}}$ is itself a modular fusion category, then ${\mathcal{D}}$ serves as an example of a non-pointed, simple, integral modular fusion category.
If ${\mathcal{D}}$ is not modular, then ${\mathcal{D}}^{\prime}\cap{\mathcal{D}}$ is non-trivial according to [EGNO15, Corollary 8.20.10], yet it is symmetric by design. But as discussed earlier, such symmetric subcategories cannot exist, leading to a contradiction. ∎
Recall that in [CP22] the authors have also shown that Conjecture 1.24 is equivalent to either of the following:
Conjecture 10.14.
Odd-dimensional fusion categories are solvable.
Conjecture 10.15.
Odd-dimensional modular fusion categories are solvable.
11. Applications and extra results
11.1. Perfect Drinfeld center
We will characterize the perfect fusion categories with a perfect Drinfeld center. That will provide a large class of perfect modular fusion categories.
Lemma 11.1.
Let $\mathcal{C}$ be a fusion category. Let $F:\mathcal{Z}(\mathcal{C})\to\mathcal{C}$ be the forgetful functor. Let $F_{1}:{\mathcal{Z}(\mathcal{C})}_{\mathrm{pt}}\rightarrow{\mathcal{C}_{\mathrm{pt}}}$ be the induced group homomorphism. Let $G$ be the universal grading group of $\mathcal{C}$. Then,
  1. (1)
    $\ker F_{1}$ is isomorphic to $\widehat{G_{ab}}$, with $G_{ab}:=G/G^{\prime}$ the abelianization of $G$,
  2. (2)
    $G$ is perfect if and only if $\ker F_{1}$ is trivial (i.e. $F(Z)=1$ implies $Z=1$).
Proof.
By definition of the center $\mathcal{Z}(\mathcal{C})$, see [EGNO15, Definition 7.13.1], $F(Z)=1$ if and only if $Z=(1,\gamma)$ with $\gamma\in\mathrm{Aut}_{\otimes}(\operatorname{id}_{\mathcal{C}})$. So $\mathrm{Aut}_{\otimes}(\operatorname{id}_{\mathcal{C}})$ is isomorphic to $\ker F_{1}$. Now, [GN08, Proposition 3.9] states that $\mathrm{Aut}_{\otimes}(\operatorname{id}_{\mathcal{C}})$ is isomorphic to $\widehat{G_{ab}}$. In particular, $\ker F_{1}$ is trivial if and only if $G^{\prime}=G$ (i.e. $G$ is perfect). ∎
Note that Lemma 11.1 covers [DNV15, Lemma 2.1].
Proposition 11.2.
Let $\mathcal{C}$ be a perfect fusion category. Then $\mathcal{Z}(\mathcal{C})$ is perfect if and only if the universal grading group of $\mathcal{C}$ is perfect.
Proof.
Let $Z$ be a simple object of $\mathcal{Z}(\mathcal{C})$ with $\mathrm{FPdim}(Z)=1$. The forgetful functor $F:\mathcal{Z}(\mathcal{C})\to\mathcal{C}$ is a tensor functor, so $\mathrm{FPdim}(F(Z))=1$, but $\mathcal{C}$ is perfect, so $F(Z)=1$. The result follows by Lemma 11.1. ∎
Corollary 11.3.
Let $\mathcal{C}$ be a braided fusion category. Then $\mathcal{Z}(\mathcal{C})$ is perfect if and only if $\mathcal{C}$ is perfect with a trivial universal grading group.
Proof.
If $\mathcal{Z}(\mathcal{C})$ is perfect, then $\mathcal{C}$ is also perfect because it can be embedded into $\mathcal{Z}(\mathcal{C})$ as a braided fusion category. According to Proposition 11.2, the universal grading group of $\mathcal{C}$ is perfect. However, since $\mathcal{C}$ is braided, its Grothendieck ring is commutative, which means its universal grading group must be abelian due to the faithful grading. But a perfect abelian group is trivial. Conversely, the result follows directly from Proposition 11.2, as the trivial group is perfect. ∎
Proposition 11.2 and Corollary 11.3 provide a complete characterization of when the Drinfeld center $\mathcal{Z}(\mathcal{C})$ is perfect assuming the fusion category $\mathcal{C}$ is perfect or braided. However, it is important to note that there are fusion categories that are neither perfect nor braided, yet their Drinfeld centers are perfect. An example of this is the fusion category $\mathcal{C}=\mathrm{Vec}(G)$, where $G$ is a non-abelian finite simple group. This fusion category is Morita equivalent to $\mathrm{Rep}(G)$, as illustrated in [EGNO15, Example 7.12.19]. As a result, their Drinfeld centers are braided equivalent, as shown in [EGNO15, Theorem 8.12.3]. Nonetheless, Corollary 11.7 below confirms that $\mathcal{Z}(\mathrm{Rep}(G))$ is perfect in this case.
Lemma 11.4.
The universal grading group of $\mathrm{Rep}(G)$ is isomorphic to the center $Z(G)$ of the finite group $G$.
Proof.
Immediate by Theorem 7.8, because a grouplike element $\mu_{j}$ in the dual hypergroup of the Grothendieck ring of $\mathrm{Rep}(G)$ corresponds to a column of the character table of $G$ with squared norm $n_{j}=|G|$, by Lemma 2.31, so to a conjugacy class of size $|G|/n_{j}=1$, thus to a central element. ∎
Corollary 11.5.
Let $G$ be a finite group. The Drinfeld center $\mathcal{Z}(\mathrm{Rep}(G))$ is perfect if and only if $G$ is perfect with a trivial center $Z(G)$.
Proof.
Immediate by Corollary 11.3 and Lemma 11.4, because $\mathrm{Rep}(G)$ is perfect if and only if $G$ is perfect. ∎
Remark 11.6.
[…]
Corollary 11.7.
Let $G$ be a non-abelian finite simple group. Then $\mathcal{Z}(\mathrm{Rep}(G))$ is perfect.
Proof.
Immediate by Corollary 11.5, because a non-abelian finite simple group is perfect with a trivial center. ∎
By Corollary 11.7, $\mathcal{Z}(\mathrm{Rep}(A_{5}))$ is a perfect integral modular fusion category of $\mathrm{FPdim}$ $60^{2}=3600$. Using Remark 11.6 and GAP, it is of rank 222222 and type
[[1,1],[3,2],[4,1],[5,1],[12,10],[15,4],[20,3]].[[1,1],[3,2],[4,1],[5,1],[12,10],[15,4],[20,3]].[ [ 1 , 1 ] , [ 3 , 2 ] , [ 4 , 1 ] , [ 5 , 1 ] , [ 12 , 10 ] , [ 15 , 4 ] , [ 20 , 3 ] ] .
Question 11.8.
Is there a non-trivial perfect integral modular fusion category of rank less than 222222, or $\mathrm{FPdim}$ of less than 360036003600?
Recall that a Hopf algebra $A$ is called perfect if $\mathrm{Rep}(A)$ is perfect, i.e. $G(A^{*})$ is a trivial group; that a finite dimensional semisimple Hopf algebra $A$ (over $\mathbb{C}$) is factorizable if and only if $\mathrm{Rep}(A)$ is modular; that $\mathcal{Z}(\mathrm{Rep}(A))=\mathrm{Rep}(D(A))$ where the Hopf algebra $D(A)$ is the Drinfeld double of $A$. So, for any non-abelian finite simple group $G$ then $D(G)$ is a finite dimensional semisimple factorizable perfect Hopf algebra (over $\mathbb{C}$) of dimension $|G|^{2}$. Let us clarify Question 11.8 for the case of Hopf algebras:
Question 11.9.
Is there a non-trivial finite-dimensional semisimple factorizable perfect Hopf algebra over $\mathbb{C}$ that has less than 222222 irreducible representations (up to isomorphism), or a dimension of less than 360036003600?
Lemma 11.10.
Let $A$ be a semisimple factorizable Hopf algebra. If $A^{*}$ is perfect then so is $A$.
Proof.
By [Sch01, Theorem 2.3(b)] there is an isomorphism $G(A^{*})\simeq G(A)\cap Z(A)$. If $G(A)=\{1\}$ then clearly $G(A^{*})=\{1\}$. ∎
[…]
11.2. Burnside and integrality properties
Remark 11.11.
According to Theorem 7.8, the dual hypergroup of a commutative fusion ring is perfect if and only if the universal grading group of the fusion ring is trivial.
Remark 11.12.
The Grothendieck ring of a modular fusion category is perfect if and only if its dual is perfect (since it is isomorphic to its dual as normalized hypergroup, see §10.1).
Remark 11.13.
By Remarks 11.11 and 11.12, a modular fusion category $\mathcal{C}$ is perfect if and only if its universal grading group is trivial (i.e. $\mathcal{C}_{ad}=\mathcal{C}$). Alternatively, recall that $({\mathcal{C}_{\mathrm{pt}}})^{\prime}={\mathcal{C}_{\mathrm{ad}}}$ by [EGNO15, Corollary 8.22.8].
Proof of Theorem 1.13
Proof.
We utilize the Galois action on $\widehat{\mathcal{B}}$ as described in Lemma 3.2. Let $\mu_{1}$ be the $\mathrm{FPdim}$ character. According to Proposition 3.5, any Galois conjugate of $\mu_{1}$, denoted by $\mu_{j}=\sigma\cdot\mu_{1}$, has a formal codegree given by $n_{j}=\sigma(\mathrm{FPdim}(H,\mathcal{B}))$. However, because $\mu_{j}=\sigma\cdot\mu_{1}$ is also a non-vanishing character, it must be a grouplike element due to the definition of a dual-Burnside hypergroup (Definition 4.7). Therefore, its formal codegree must also be $\mathrm{FPdim}(H,\mathcal{B})$, by Corollary 2.29. Thus, $\sigma(\mathrm{FPdim}(H,\mathcal{B}))=\mathrm{FPdim}(H,\mathcal{B})$ for every $\sigma$ in the Galois group. It follows that $\mathrm{FPdim}(H,\mathcal{B})$ is a rational number, completing the proof. ∎
As a consequence of Theorem 1.13, every commutative dual-Burnside fusion ring is weakly integral. However, the converse is not true, even for unitary integral fusion categories such as $\mathrm{Rep}(A_{7})$; refer to §12.2 for more examples. The following result is partially attributed to Andrew Schopieray; see [Sch23].
Corollary 11.14.
Let ${(H,\mathcal{B},\mu_{1})}$ be rational abelian normalizable hypergroup. If the dual $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is perfect and Burnside then ${(H,\mathcal{B},\mu_{1})}$ is integral, i.e $d_{i}\in\mathbb{Q}$.
Proof.
The dual $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is perfect, meaning that $\mu_{1}$ is the only grouplike linear character. As in the proof of Theorem 1.13, for any $\sigma\in\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$, $\sigma.\mu_{1}$ is also non-vanishing and thus grouplike by the dual-Burnside assumption. Consequently, $\sigma.\mu_{1}=\mu_{1}$, and $\sigma(d_{i})=d_{i}$ for all $\sigma$, which implies that $d_{i}\in\mathbb{Q}$. ∎
Corollary 11.15.
A commutative dual-Burnside fusion ring with a trivial universal grading group is integral.
Proof.
It follows from Corollary 11.14 and Remark 11.11. ∎
Recall Remark 6.17 about a trivial universal grading group.
Corollary 11.16.
A fusion category with a dual-Burnside commutative Grothendieck ring and a trivial universal grading group is integral.
Proof.
Immediate from Corollary 11.15. ∎
Proof of Theorem 1.27
Proof.
Recall that a modular fusion category is Burnside if and only if it is dual-Burnside (Corollary 10.3). Now, by Remark 11.13, the universal grading group is trivial, so by Corollary 11.16, it must be integral. The converse follows from [GNN09, Theorem 6.1]. ∎
Remark 11.17.
According to Theorem 1.27, a simple non-integral modular fusion category cannot be (dual-)Burnside. Many modular fusion categories of Lie type, known as Verlinde categories, belong to this class, as do the one described in [Sch22, Theorem 1] and the Drinfeld center of the Extended-Haagerup fusion categories. Consequently, none of these are (dual-)Burnside. Therefore, for the Grothendieck rings in this class, we can still infer the existence of a non-group-like and non-vanishing linear character (and basic element).
11.3. Near-group modular fusion categories
Let $G$ be a finite abelian group and $m$ be a non-negative integer. Let $K(G,m)$ be the fusion ring with basis $\mathcal{B}=G\cup\{\rho\}$ and fusion rules:
$\rho^{2}=\sum_{g\in G}g+m\rho\text{ and }g\rho=\rho g=\rho,$
for all $g\in G$. A fusion category whose Grothendieck ring is $K(G,m)$ is called near-group (or also Tambara-Yamagami when $m=0$), see for example [Izu17].
Proposition 11.18.
There is no modular fusion category of Grothendieck ring $K(G,m)$ if $G$ is non-trivial and $m>0$.
Proof.
Let $\mathcal{C}$ be a modular fusion category of Grothendieck ring $K(G,m)$. Given that $\rho^{2}=m\rho+\sum_{g\in G}g$, it follows that $\mathrm{FPdim}(\rho)^{2}=m\mathrm{FPdim}(\rho)+|G|$. Let $x_{+}>0$ and $x_{-}<0$ be the solutions to the equation $x^{2}-mx-|G|=0$. Thus, $\mathrm{FPdim}(\rho)=x_{+}$. Now, consider $\eta$, a linear character of $R$. For all $g\in G$, we have $\eta(\rho)=\eta(\rho g)=\eta(\rho)\eta(g)$. Therefore, $\eta(\rho)$ is nonzero if and only if $\eta(g)=1$ for all $g\in G$. This condition is equivalent to solving $\eta(\rho)^{2}-m\eta(\rho)-|G|=0$, which implies $\eta(\rho)=x_{\pm}$. We denote such characters as $\psi_{\pm}$. It follows that $\psi_{+}=\mathrm{FPdim}$ is the unit of $^{\widehat{\mathcal{B}}}$. Let $\epsilon$ be the trivial character of $G$. We identify any $\eta$ in $\widehat{G}\setminus\{\epsilon\}$ (non-empty since $G$ is non-trivial) with the corresponding element in $^{\widehat{\mathcal{B}}}$ that vanishes on $\rho$. Let $q=-\frac{x_{-}}{x_{+}}$, which is positive. The multiplication on $\widehat{R}$ is defined by:
$\eta\star\psi_{-}$ $=\eta,$
$\eta\star\eta^{\prime}$ $=(1-\delta_{\eta^{\prime},\eta^{*}})\eta\eta^{\prime}+\delta_{\eta^{\prime},\eta^{*}}(\frac{q}{1+q}\psi_{+}+\frac{1}{1+q}\psi_{-}),$
$\psi_{-}\star\psi_{-}$ $=q\psi_{+}+(1-q)\psi_{-}.$
[…]
Proposition 11.19.
Let $\mathcal{C}$ be a modular fusion category of Grothendieck ring $K(G,0)$. Then $G\simeq C_{1}$ or $C_{2}$.
Proof.
[…]
Remark 11.20.
Ising categories are the only modular fusion categories whose Grothendieck ring are $K(C_{2},0)$, see [DGNO10, Appendix B].
Corollary 11.21.
A modular fusion category is near-group if and only if its Grothendieck ring is $K(G,m)$ with $(G,m)=(C_{1},0),(C_{1},1),(C_{2},0)$.
Proof.
According to Proposition 11.18, either $G$ is trivial or $m=0$. If $G$ is trivial, the category has rank 2, and the conclusion follows from [Ost03b]. On the other hand, if $G$ is non-trivial and $m=0$, the result follows from Proposition 11.19 and Remark 11.20. ∎
Remark 11.22.
Note that modular (generalized) Tambara-Yamagami categories were classified before, see [Nat13, Lemma 5.3 and Theorem 5.4], [Tho11] and [Sie00].
12. Concrete examples
This section focuses on providing concrete examples of fusion rings and categories to which the main results of the paper apply as categorification criteria, along with counter-examples demonstrating the limits of these results.
12.1. Burnside property
Recall that a fusion ring is called Burnside if, for every basis element, its fusion matrix has a norm of 111 if and only if its determinant is nonzero. In other words, being grouplike is equivalent to being invertible (see Proposition 4.3). William Burnside originally proved that the Grothendieck ring of $\mathrm{Rep}(G)$ is Burnside for every finite group $G$. This result was expanded in [Bur23] to include all weakly integral fusion categories with a commutative Grothendieck ring, but first within the context of modular categories in [GNN09, Appendix]. In Theorem 1.2, it extends to every commutative fusion ring with an $h$-integral dual, and further to hypergroups in Theorem 4.17.
Remark 12.1.
By Theorems 1.8 and 7.9, along with the fact that $\mathrm{Rep}(G)_{pt}\simeq\mathrm{Rep}(G/G^{\prime})$, Burnside’s result can be restated as follows: for any finite group $G$,
$\left(\prod_{j\in\mathcal{I}}\frac{C_{j}}{|\mathcal{C}^{j}|}\right)^{2}=\frac{1}{|G^{\prime}|}\sum_{\mathcal{C}^{j}\subseteq G^{\prime}}C_{j},$
where $G^{\prime}$ is the commutator subgroup of $G$, $(\mathcal{C}^{j})$ are the conjugacy classes of $G$, and $C_{j}:=\sum_{g\in\mathcal{C}^{j}}g$ is the class sum associated with $\mathcal{C}^{j}$. It was already observed in [Har08].
A fusion ring $R$ is termed $\alpha$-Frobenius if, for every basis element $x$, the expression $\frac{\mathrm{FPdim}(R)^{\alpha}}{\mathrm{FPdim}(x)}$ is an algebraic integer [ENO25]. Kaplansky’s 6th conjecture posits that the Grothendieck ring of a complex fusion category is 111-Frobenius [ENO11, Question 1]. An exhaustive classification of all 111-Frobenius simple integral fusion rings, within certain specified limits, was provided in [LPW21]. These limits have been updated in [BP24] as follows:
$\begin{array}[]{c|cccccccc}\text{Rank}&\leq 5&6&7&8&9&10&11&12\\ \hline\cr\operatorname{\mathsf{FPdim}}\leq&10^{7}&10^{6}&10^{5}&20000&10000&5000&3000&1000\end{array}$
There are exactly 505505505 non-pointed examples (including 888 that are character rings of groups). Among them, we found only 444 non-Burnside ones. They have the type of $\mathrm{Rep}(A_{7})$, but different fusion data.
  • Rank: 999,
  • FPdim: $2520=2^{3}\times 3^{2}\times 5\times 7$,
  • Type: [[1,1],[6,1],[10,2],[14,2],[15,1],[21,1],[35,1]][[1,1],[6,1],[10,2],[14,2],[15,1],[21,1],[35,1]][ [ 1 , 1 ] , [ 6 , 1 ] , [ 10 , 2 ] , [ 14 , 2 ] , [ 15 , 1 ] , [ 21 , 1 ] , [ 35 , 1 ] ],
  • Fusion data 1:
    $\begin{smallmatrix}1&0&0&0&0&0&0&0&0\\ 0&1&0&0&0&0&0&0&0\\ 0&0&1&0&0&0&0&0&0\\ 0&0&0&1&0&0&0&0&0\\ 0&0&0&0&1&0&0&0&0\\ 0&0&0&0&0&1&0&0&0\\ 0&0&0&0&0&0&1&0&0\\ 0&0&0&0&0&0&0&1&0\\ 0&0&0&0&0&0&0&0&1\end{smallmatrix},\ \begin{smallmatrix}0&1&0&0&0&0&0&0&0\\ 1&0&0&0&0&1&0&1&0\\ 0&0&0&1&0&0&1&0&1\\ 0&0&1&0&0&0&1&0&1\\ 0&0&0&0&1&2&0&2&0\\ 0&1&0&0&2&0&1&0&1\\ 0&0&1&1&0&1&0&1&1\\ 0&1&0&0&2&0&1&2&1\\ 0&0&1&1&0&1&1&1&4\end{smallmatrix},\ \begin{smallmatrix}0&0&1&0&0&0&0&0&0\\ 0&0&0&1&0&0&1&0&1\\ 0&1&0&1&1&1&0&1&1\\ 1&0&0&0&1&1&1&1&1\\ 0&0&1&1&0&0&1&0&3\\ 0&0&1&1&0&1&1&1&2\\ 0&1&1&0&1&1&1&1&2\\ 0&0&1&1&0&1&1&1&4\\ 0&1&1&1&3&2&2&4&4\end{smallmatrix},\ \begin{smallmatrix}0&0&0&1&0&0&0&0&0\\ 0&0&1&0&0&0&1&0&1\\ 1&0&0&0&1&1&1&1&1\\ 0&1&1&0&1&1&0&1&1\\ 0&0&1&1&0&0&1&0&3\\ 0&0&1&1&0&1&1&1&2\\ 0&1&0&1&1&1&1&1&2\\ 0&0&1&1&0&1&1&1&4\\ 0&1&1&1&3&2&2&4&4\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&1&0&0&0&0\\ 0&0&0&0&1&2&0&2&0\\ 0&0&1&1&0&0&1&0&3\\ 0&0&1&1&0&0&1&0&3\\ 1&1&0&0&3&3&0&5&0\\ 0&2&0&0&3&0&2&2&2\\ 0&0&1&1&0&2&1&2&3\\ 0&2&0&0&5&2&2&4&2\\ 0&0&3&3&0&2&3&2&9\end{smallmatrix},$
    $\begin{smallmatrix}0&0&0&0&0&1&0&0&0\\ 0&1&0&0&2&0&1&0&1\\ 0&0&1&1&0&1&1&1&2\\ 0&0&1&1&0&1&1&1&2\\ 0&2&0&0&3&0&2&2&2\\ 1&0&1&1&0&3&0&3&2\\ 0&1&1&1&2&0&2&1&3\\ 0&0&1&1&2&3&1&4&3\\ 0&1&2&2&2&2&3&3&8\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&0&0&1&0&0\\ 0&0&1&1&0&1&0&1&1\\ 0&1&1&0&1&1&1&1&2\\ 0&1&0&1&1&1&1&1&2\\ 0&0&1&1&0&2&1&2&3\\ 0&1&1&1&2&0&2&1&3\\ 1&0&1&1&1&2&1&2&3\\ 0&1&1&1&2&1&2&2&5\\ 0&1&2&2&3&3&3&5&7\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&0&0&0&1&0\\ 0&1&0&0&2&0&1&2&1\\ 0&0&1&1&0&1&1&1&4\\ 0&0&1&1&0&1&1&1&4\\ 0&2&0&0&5&2&2&4&2\\ 0&0&1&1&2&3&1&4&3\\ 0&1&1&1&2&1&2&2&5\\ 1&2&1&1&4&4&2&6&4\\ 0&1&4&4&2&3&5&4&12\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&0&0&0&0&1\\ 0&0&1&1&0&1&1&1&4\\ 0&1&1&1&3&2&2&4&4\\ 0&1&1&1&3&2&2&4&4\\ 0&0&3&3&0&2&3&2&9\\ 0&1&2&2&2&2&3&3&8\\ 0&1&2&2&3&3&3&5&7\\ 0&1&4&4&2&3&5&4&12\\ 1&4&4&4&9&8&7&12&15\end{smallmatrix}$
  • Fusion data 2:
    $\begin{smallmatrix}1&0&0&0&0&0&0&0&0\\ 0&1&0&0&0&0&0&0&0\\ 0&0&1&0&0&0&0&0&0\\ 0&0&0&1&0&0&0&0&0\\ 0&0&0&0&1&0&0&0&0\\ 0&0&0&0&0&1&0&0&0\\ 0&0&0&0&0&0&1&0&0\\ 0&0&0&0&0&0&0&1&0\\ 0&0&0&0&0&0&0&0&1\end{smallmatrix},\ \begin{smallmatrix}0&1&0&0&0&0&0&0&0\\ 1&0&0&0&0&1&0&1&0\\ 0&0&0&1&0&0&1&0&1\\ 0&0&1&0&0&0&1&0&1\\ 0&0&0&0&1&2&0&2&0\\ 0&1&0&0&2&0&1&0&1\\ 0&0&1&1&0&1&0&1&1\\ 0&1&0&0&2&0&1&2&1\\ 0&0&1&1&0&1&1&1&4\end{smallmatrix},\ \begin{smallmatrix}0&0&1&0&0&0&0&0&0\\ 0&0&0&1&0&0&1&0&1\\ 0&1&1&2&1&1&1&1&0\\ 1&0&1&1&1&1&2&1&0\\ 0&0&1&1&0&0&1&0&3\\ 0&0&1&1&0&1&1&1&2\\ 0&1&2&1&1&1&2&1&1\\ 0&0&1&1&0&1&1&1&4\\ 0&1&0&0&3&2&1&4&5\end{smallmatrix},\ \begin{smallmatrix}0&0&0&1&0&0&0&0&0\\ 0&0&1&0&0&0&1&0&1\\ 1&0&1&1&1&1&2&1&0\\ 0&1&2&1&1&1&1&1&0\\ 0&0&1&1&0&0&1&0&3\\ 0&0&1&1&0&1&1&1&2\\ 0&1&1&2&1&1&2&1&1\\ 0&0&1&1&0&1&1&1&4\\ 0&1&0&0&3&2&1&4&5\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&1&0&0&0&0\\ 0&0&0&0&1&2&0&2&0\\ 0&0&1&1&0&0&1&0&3\\ 0&0&1&1&0&0&1&0&3\\ 1&1&0&0&3&3&0&5&0\\ 0&2&0&0&3&0&2&2&2\\ 0&0&1&1&0&2&1&2&3\\ 0&2&0&0&5&2&2&4&2\\ 0&0&3&3&0&2&3&2&9\end{smallmatrix},$
    $\begin{smallmatrix}0&0&0&0&0&1&0&0&0\\ 0&1&0&0&2&0&1&0&1\\ 0&0&1&1&0&1&1&1&2\\ 0&0&1&1&0&1&1&1&2\\ 0&2&0&0&3&0&2&2&2\\ 1&0&1&1&0&3&0&3&2\\ 0&1&1&1&2&0&2&1&3\\ 0&0&1&1&2&3&1&4&3\\ 0&1&2&2&2&2&3&3&8\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&0&0&1&0&0\\ 0&0&1&1&0&1&0&1&1\\ 0&1&2&1&1&1&2&1&1\\ 0&1&1&2&1&1&2&1&1\\ 0&0&1&1&0&2&1&2&3\\ 0&1&1&1&2&0&2&1&3\\ 1&0&2&2&1&2&2&2&2\\ 0&1&1&1&2&1&2&2&5\\ 0&1&1&1&3&3&2&5&8\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&0&0&0&1&0\\ 0&1&0&0&2&0&1&2&1\\ 0&0&1&1&0&1&1&1&4\\ 0&0&1&1&0&1&1&1&4\\ 0&2&0&0&5&2&2&4&2\\ 0&0&1&1&2&3&1&4&3\\ 0&1&1&1&2&1&2&2&5\\ 1&2&1&1&4&4&2&6&4\\ 0&1&4&4&2&3&5&4&12\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&0&0&0&0&1\\ 0&0&1&1&0&1&1&1&4\\ 0&1&0&0&3&2&1&4&5\\ 0&1&0&0&3&2&1&4&5\\ 0&0&3&3&0&2&3&2&9\\ 0&1&2&2&2&2&3&3&8\\ 0&1&1&1&3&3&2&5&8\\ 0&1&4&4&2&3&5&4&12\\ 1&4&5&5&9&8&8&12&14\end{smallmatrix}$
  • Fusion data 3:
    $\begin{smallmatrix}1&0&0&0&0&0&0&0&0\\ 0&1&0&0&0&0&0&0&0\\ 0&0&1&0&0&0&0&0&0\\ 0&0&0&1&0&0&0&0&0\\ 0&0&0&0&1&0&0&0&0\\ 0&0&0&0&0&1&0&0&0\\ 0&0&0&0&0&0&1&0&0\\ 0&0&0&0&0&0&0&1&0\\ 0&0&0&0&0&0&0&0&1\end{smallmatrix},\ \begin{smallmatrix}0&1&0&0&0&0&0&0&0\\ 1&1&0&0&0&1&1&0&0\\ 0&0&0&1&0&0&1&0&1\\ 0&0&1&0&0&0&1&0&1\\ 0&0&0&0&2&0&0&1&1\\ 0&1&0&0&0&3&1&1&0\\ 0&1&1&1&0&1&1&0&1\\ 0&0&0&0&1&1&0&3&1\\ 0&0&1&1&1&0&1&1&4\end{smallmatrix},\ \begin{smallmatrix}0&0&1&0&0&0&0&0&0\\ 0&0&0&1&0&0&1&0&1\\ 0&1&0&1&0&1&0&0&2\\ 1&0&0&0&0&1&1&0&2\\ 0&0&0&0&1&1&0&2&2\\ 0&0&1&1&1&0&1&1&2\\ 0&1&1&0&0&1&1&0&3\\ 0&0&0&0&2&1&0&3&3\\ 0&1&2&2&2&2&3&3&4\end{smallmatrix},\ \begin{smallmatrix}0&0&0&1&0&0&0&0&0\\ 0&0&1&0&0&0&1&0&1\\ 1&0&0&0&0&1&1&0&2\\ 0&1&1&0&0&1&0&0&2\\ 0&0&0&0&1&1&0&2&2\\ 0&0&1&1&1&0&1&1&2\\ 0&1&0&1&0&1&1&0&3\\ 0&0&0&0&2&1&0&3&3\\ 0&1&2&2&2&2&3&3&4\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&1&0&0&0&0\\ 0&0&0&0&2&0&0&1&1\\ 0&0&0&0&1&1&0&2&2\\ 0&0&0&0&1&1&0&2&2\\ 1&2&1&1&1&1&2&0&3\\ 0&0&1&1&1&0&1&2&3\\ 0&0&0&0&2&1&0&3&3\\ 0&1&2&2&0&2&3&0&5\\ 0&1&2&2&3&3&3&5&6\end{smallmatrix},$
    $\begin{smallmatrix}0&0&0&0&0&1&0&0&0\\ 0&1&0&0&0&3&1&1&0\\ 0&0&1&1&1&0&1&1&2\\ 0&0&1&1&1&0&1&1&2\\ 0&0&1&1&1&0&1&2&3\\ 1&3&0&0&0&6&2&3&0\\ 0&1&1&1&1&2&2&2&2\\ 0&1&1&1&2&3&2&3&3\\ 0&0&2&2&3&0&2&3&9\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&0&0&1&0&0\\ 0&1&1&1&0&1&1&0&1\\ 0&1&1&0&0&1&1&0&3\\ 0&1&0&1&0&1&1&0&3\\ 0&0&0&0&2&1&0&3&3\\ 0&1&1&1&1&2&2&2&2\\ 1&1&1&1&0&2&2&0&4\\ 0&0&0&0&3&2&0&5&4\\ 0&1&3&3&3&2&4&4&7\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&0&0&0&1&0\\ 0&0&0&0&1&1&0&3&1\\ 0&0&0&0&2&1&0&3&3\\ 0&0&0&0&2&1&0&3&3\\ 0&1&2&2&0&2&3&0&5\\ 0&1&1&1&2&3&2&3&3\\ 0&0&0&0&3&2&0&5&4\\ 1&3&3&3&0&3&5&0&7\\ 0&1&3&3&5&3&4&7&10\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&0&0&0&0&1\\ 0&0&1&1&1&0&1&1&4\\ 0&1&2&2&2&2&3&3&4\\ 0&1&2&2&2&2&3&3&4\\ 0&1&2&2&3&3&3&5&6\\ 0&0&2&2&3&0&2&3&9\\ 0&1&3&3&3&2&4&4&7\\ 0&1&3&3&5&3&4&7&10\\ 1&4&4&4&6&9&7&10&17\end{smallmatrix}$
  • Fusion data 4:
    $\begin{smallmatrix}1&0&0&0&0&0&0&0&0\\ 0&1&0&0&0&0&0&0&0\\ 0&0&1&0&0&0&0&0&0\\ 0&0&0&1&0&0&0&0&0\\ 0&0&0&0&1&0&0&0&0\\ 0&0&0&0&0&1&0&0&0\\ 0&0&0&0&0&0&1&0&0\\ 0&0&0&0&0&0&0&1&0\\ 0&0&0&0&0&0&0&0&1\end{smallmatrix},\ \begin{smallmatrix}0&1&0&0&0&0&0&0&0\\ 1&0&0&0&0&1&0&1&0\\ 0&0&0&1&0&0&1&0&1\\ 0&0&1&0&0&0&1&0&1\\ 0&0&0&0&1&2&0&2&0\\ 0&1&0&0&2&0&1&0&1\\ 0&0&1&1&0&1&0&1&1\\ 0&1&0&0&2&0&1&2&1\\ 0&0&1&1&0&1&1&1&4\end{smallmatrix},\ \begin{smallmatrix}0&0&1&0&0&0&0&0&0\\ 0&0&0&1&0&0&1&0&1\\ 1&0&0&2&1&1&2&1&0\\ 0&1&2&1&1&1&1&1&0\\ 0&0&1&1&0&0&1&0&3\\ 0&0&1&1&0&1&1&1&2\\ 0&1&2&1&1&1&2&1&1\\ 0&0&1&1&0&1&1&1&4\\ 0&1&0&0&3&2&1&4&5\end{smallmatrix},\ \begin{smallmatrix}0&0&0&1&0&0&0&0&0\\ 0&0&1&0&0&0&1&0&1\\ 0&1&2&1&1&1&1&1&0\\ 1&0&1&1&1&1&2&1&0\\ 0&0&1&1&0&0&1&0&3\\ 0&0&1&1&0&1&1&1&2\\ 0&1&1&2&1&1&2&1&1\\ 0&0&1&1&0&1&1&1&4\\ 0&1&0&0&3&2&1&4&5\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&1&0&0&0&0\\ 0&0&0&0&1&2&0&2&0\\ 0&0&1&1&0&0&1&0&3\\ 0&0&1&1&0&0&1&0&3\\ 1&1&0&0&3&3&0&5&0\\ 0&2&0&0&3&0&2&2&2\\ 0&0&1&1&0&2&1&2&3\\ 0&2&0&0&5&2&2&4&2\\ 0&0&3&3&0&2&3&2&9\end{smallmatrix},$
    $\begin{smallmatrix}0&0&0&0&0&1&0&0&0\\ 0&1&0&0&2&0&1&0&1\\ 0&0&1&1&0&1&1&1&2\\ 0&0&1&1&0&1&1&1&2\\ 0&2&0&0&3&0&2&2&2\\ 1&0&1&1&0&3&0&3&2\\ 0&1&1&1&2&0&2&1&3\\ 0&0&1&1&2&3&1&4&3\\ 0&1&2&2&2&2&3&3&8\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&0&0&1&0&0\\ 0&0&1&1&0&1&0&1&1\\ 0&1&2&1&1&1&2&1&1\\ 0&1&1&2&1&1&2&1&1\\ 0&0&1&1&0&2&1&2&3\\ 0&1&1&1&2&0&2&1&3\\ 1&0&2&2&1&2&2&2&2\\ 0&1&1&1&2&1&2&2&5\\ 0&1&1&1&3&3&2&5&8\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&0&0&0&1&0\\ 0&1&0&0&2&0&1&2&1\\ 0&0&1&1&0&1&1&1&4\\ 0&0&1&1&0&1&1&1&4\\ 0&2&0&0&5&2&2&4&2\\ 0&0&1&1&2&3&1&4&3\\ 0&1&1&1&2&1&2&2&5\\ 1&2&1&1&4&4&2&6&4\\ 0&1&4&4&2&3&5&4&12\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&0&0&0&0&1\\ 0&0&1&1&0&1&1&1&4\\ 0&1&0&0&3&2&1&4&5\\ 0&1&0&0&3&2&1&4&5\\ 0&0&3&3&0&2&3&2&9\\ 0&1&2&2&2&2&3&3&8\\ 0&1&1&1&3&3&2&5&8\\ 0&1&4&4&2&3&5&4&12\\ 1&4&5&5&9&8&8&12&14\end{smallmatrix}$
The determinant of the second fusion matrix for each fusion data set mentioned above is ±36$\pm$ 36± 36, while its norm is 666. This indicates that these integral fusion rings are non-Burnside and, consequently, cannot be categorified.
Similar classifications have been conducted for the non-1-Frobenius case. Interestingly, most—but not all—of the fusion rings identified are non-Burnside, and thus, not amenable to categorification (see the example below). This serves as supporting evidence for Kaplansky’s 6th conjecture.
  • Rank: 777,
  • FPdim: 798=2×3×7×19798=2$\times$ 3$\times$ 7$\times$ 19798 = 2 × 3 × 7 × 19,
  • Type: [[1,1],[7,1],[8,1],[9,3],[21,1]][[1,1],[7,1],[8,1],[9,3],[21,1]][ [ 1 , 1 ] , [ 7 , 1 ] , [ 8 , 1 ] , [ 9 , 3 ] , [ 21 , 1 ] ],
  • Fusion data:
    $\begin{smallmatrix}1&0&0&0&0&0&0\\ 0&1&0&0&0&0&0\\ 0&0&1&0&0&0&0\\ 0&0&0&1&0&0&0\\ 0&0&0&0&1&0&0\\ 0&0&0&0&0&1&0\\ 0&0&0&0&0&0&1\end{smallmatrix},\ \begin{smallmatrix}0&1&0&0&0&0&0\\ 1&0&0&1&1&1&1\\ 0&0&1&1&1&1&1\\ 0&1&1&1&1&1&1\\ 0&1&1&1&1&1&1\\ 0&1&1&1&1&1&1\\ 0&1&1&1&1&1&5\end{smallmatrix},\ \begin{smallmatrix}0&0&1&0&0&0&0\\ 0&0&1&1&1&1&1\\ 1&1&1&1&1&1&1\\ 0&1&1&2&1&1&1\\ 0&1&1&1&2&1&1\\ 0&1&1&1&1&2&1\\ 0&1&1&1&1&1&6\end{smallmatrix},\ \begin{smallmatrix}0&0&0&1&0&0&0\\ 0&1&1&1&1&1&1\\ 0&1&1&2&1&1&1\\ 1&1&2&1&1&2&1\\ 0&1&1&1&2&2&1\\ 0&1&1&2&2&1&1\\ 0&1&1&1&1&1&7\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&1&0&0\\ 0&1&1&1&1&1&1\\ 0&1&1&1&2&1&1\\ 0&1&1&1&2&2&1\\ 1&1&2&2&1&1&1\\ 0&1&1&2&1&2&1\\ 0&1&1&1&1&1&7\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&0&1&0\\ 0&1&1&1&1&1&1\\ 0&1&1&1&1&2&1\\ 0&1&1&2&2&1&1\\ 0&1&1&2&1&2&1\\ 1&1&2&1&2&1&1\\ 0&1&1&1&1&1&7\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&0&0&1\\ 0&1&1&1&1&1&5\\ 0&1&1&1&1&1&6\\ 0&1&1&1&1&1&7\\ 0&1&1&1&1&1&7\\ 0&1&1&1&1&1&7\\ 1&5&6&7&7&7&8\end{smallmatrix}$
The determinant of the third fusion matrix is 161616, while its norm is 888. This shows that the integral fusion ring is non-Burnside and therefore cannot be categorified.
12.2. Dual-Burnside property
We abbreviate “(A)RN” for “(abelian) real non-negative.” According to [LPW21, Corollary 8.5], the Grothendieck ring of a unitary fusion category, if commutative, has an RN dual (hypergroup).
Question 12.2.
Is the dual of the Grothendieck ring of a complex fusion category always RN?
Recall that a commutative fusion ring $F$ is called dual-Burnside if a column of its character table has a zero entry if and only if the squared norm of the column (the corresponding formal codegree) is strictly less than $\mathrm{FPdim}(F)$. When $F$ has an RN dual, it is equivalent to $P^{2}$ being idempotent, where
$P=\prod_{i\in\mathcal{I}}\frac{x_{i}}{\mathrm{FPdim}(x_{i})}$
and $\{x_{i}\}_{i\in I}$ is the set of basis elements (see Corollary 4.16). Note that for a non-abelian group $G$, the product of all its elements depends on the ordering; in fact, the set of all possible products forms a $G^{\prime}$-coset (see [DH82]).
According to [INW99, Theorem B], for any finite nilpotent group $G$, the Grothendieck ring of $\mathrm{Rep}(G)$ is dual-Burnside. This extends to every nilpotent dualizable ARN-hypergroup by Theorem 1.15. Thus, every commutative nilpotent fusion ring with an RN dual is both Burnside and dual-Burnside. All commutative nilpotent fusion rings we examined have RN duals.
Question 12.3.
Is there a commutative nilpotent fusion ring without an RN dual?
We define a finite group $G$ as dual-Burnside if $K(\mathrm{Rep}(G))$ is dual-Burnside. Therefore, a nilpotent finite group is dual-Burnside. There are exactly 144 finite groups of order less than 32, among which 30 are non-nilpotent. Of these, only $\mathrm{SL}(2,3)$ is dual-Burnside, while the other 29, such as $S_{3}$ and $D_{5}$, are not.
[…]
Remark 12.4.
By Theorem 1.8, a finite group $G$ is dual-Burnside if and only if
$\left(\prod_{\chi\in\mathrm{Irr}(G)}\frac{\chi}{\chi(1)}\right)^{2}=\frac{|Z(G)|}{|G|}\left(\sum_{\chi\in\mathrm{Irr}(G/Z(G))}\chi(1)\chi\right).$
12.3. Modular categories
Let us present some applications of Theorem 1.19. There are exactly 71 distinct types of half-Frobenius integral fusion rings with ranks up to 12 (see [ABPP]). Of these, 15 types (listed below) are excluded from modular categorification by Theorems 1.19 or 1.22.
[1,1,1,1,2,2],[1,1,1,1,2,2,2,2],[1,1,1,1,2,2,2,2,2],[1,1,1,1,2,2,2,4,4],[1,1,1,1,1,1,1,1,1,3],[1,1,1,1,2,2],[1,1,1,1,2,2,2,2],[1,1,1,1,2,2,2,2,2],[1,1,1,1,2,2,2,4,4],[1,1,1,1,1,1,1,1,1,3],[ 1 , 1 , 1 , 1 , 2 , 2 ] , [ 1 , 1 , 1 , 1 , 2 , 2 , 2 , 2 ] , [ 1 , 1 , 1 , 1 , 2 , 2 , 2 , 2 , 2 ] , [ 1 , 1 , 1 , 1 , 2 , 2 , 2 , 4 , 4 ] , [ 1 , 1 , 1 , 1 , 1 , 1 , 1 , 1 , 1 , 3 ] ,
[1,1,1,1,2,2,2,2,2,2],[1,1,1,1,2,2,2,4,4,4,4],[1,1,1,1,1,1,1,1,1,3,3,3],[1,1,1,1,1,1,1,1,2,2,2,2],[1,1,1,1,2,2,2,2,2,2],[1,1,1,1,2,2,2,4,4,4,4],[1,1,1,1,1,1,1,1,1,3,3,3],[1,1,1,1,1,1,1,1,2,2,2,2],[ 1 , 1 , 1 , 1 , 2 , 2 , 2 , 2 , 2 , 2 ] , [ 1 , 1 , 1 , 1 , 2 , 2 , 2 , 4 , 4 , 4 , 4 ] , [ 1 , 1 , 1 , 1 , 1 , 1 , 1 , 1 , 1 , 3 , 3 , 3 ] , [ 1 , 1 , 1 , 1 , 1 , 1 , 1 , 1 , 2 , 2 , 2 , 2 ] ,
[1,1,1,1,1,1,1,1,2,2,4,4],[1,1,1,1,2,2,2,2,2,2,2,2],[1,1,1,1,2,2,2,2,2,2,2,4],[1,1,1,1,2,2,2,4,4,4,4,4],[1,1,1,1,1,1,1,1,2,2,4,4],[1,1,1,1,2,2,2,2,2,2,2,2],[1,1,1,1,2,2,2,2,2,2,2,4],[1,1,1,1,2,2,2,4,4,4,4,4],[ 1 , 1 , 1 , 1 , 1 , 1 , 1 , 1 , 2 , 2 , 4 , 4 ] , [ 1 , 1 , 1 , 1 , 2 , 2 , 2 , 2 , 2 , 2 , 2 , 2 ] , [ 1 , 1 , 1 , 1 , 2 , 2 , 2 , 2 , 2 , 2 , 2 , 4 ] , [ 1 , 1 , 1 , 1 , 2 , 2 , 2 , 4 , 4 , 4 , 4 , 4 ] ,
[1,1,1,1,2,2,2,4,4,4,8,8],[1,1,2,2,2,2,3,3,6,6,6,6].[1,1,1,1,2,2,2,4,4,4,8,8],[1,1,2,2,2,2,3,3,6,6,6,6].[ 1 , 1 , 1 , 1 , 2 , 2 , 2 , 4 , 4 , 4 , 8 , 8 ] , [ 1 , 1 , 2 , 2 , 2 , 2 , 3 , 3 , 6 , 6 , 6 , 6 ] .
For example, let us provide all the fusion data for the first type above:
  • Rank: 666,
  • FPdim: $12=2^{2}\times 3$,
  • Type: [1,1,1,1,2,2][1,1,1,1,2,2][ 1 , 1 , 1 , 1 , 2 , 2 ],
  • Fusion data 1:
    $\begin{smallmatrix}1&0&0&0&0&0\\ 0&1&0&0&0&0\\ 0&0&1&0&0&0\\ 0&0&0&1&0&0\\ 0&0&0&0&1&0\\ 0&0&0&0&0&1\end{smallmatrix},\ \begin{smallmatrix}0&1&0&0&0&0\\ 1&0&0&0&0&0\\ 0&0&0&1&0&0\\ 0&0&1&0&0&0\\ 0&0&0&0&0&1\\ 0&0&0&0&1&0\end{smallmatrix},\ \begin{smallmatrix}0&0&1&0&0&0\\ 0&0&0&1&0&0\\ 1&0&0&0&0&0\\ 0&1&0&0&0&0\\ 0&0&0&0&0&1\\ 0&0&0&0&1&0\end{smallmatrix},\ \begin{smallmatrix}0&0&0&1&0&0\\ 0&0&1&0&0&0\\ 0&1&0&0&0&0\\ 1&0&0&0&0&0\\ 0&0&0&0&1&0\\ 0&0&0&0&0&1\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&1&0\\ 0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 0&0&0&0&1&0\\ 1&0&0&1&0&1\\ 0&1&1&0&1&0\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&0&1\\ 0&0&0&0&1&0\\ 0&0&0&0&1&0\\ 0&0&0&0&0&1\\ 0&1&1&0&1&0\\ 1&0&0&1&0&1\end{smallmatrix}$
  • Fusion data 2:
    $\begin{smallmatrix}1&0&0&0&0&0\\ 0&1&0&0&0&0\\ 0&0&1&0&0&0\\ 0&0&0&1&0&0\\ 0&0&0&0&1&0\\ 0&0&0&0&0&1\end{smallmatrix},\ \begin{smallmatrix}0&1&0&0&0&0\\ 1&0&0&0&0&0\\ 0&0&0&1&0&0\\ 0&0&1&0&0&0\\ 0&0&0&0&1&0\\ 0&0&0&0&0&1\end{smallmatrix},\ \begin{smallmatrix}0&0&1&0&0&0\\ 0&0&0&1&0&0\\ 0&1&0&0&0&0\\ 1&0&0&0&0&0\\ 0&0&0&0&0&1\\ 0&0&0&0&1&0\end{smallmatrix},\ \begin{smallmatrix}0&0&0&1&0&0\\ 0&0&1&0&0&0\\ 1&0&0&0&0&0\\ 0&1&0&0&0&0\\ 0&0&0&0&0&1\\ 0&0&0&0&1&0\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&1&0\\ 0&0&0&0&1&0\\ 0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 1&1&0&0&0&1\\ 0&0&1&1&1&0\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 0&0&0&0&1&0\\ 0&0&0&0&1&0\\ 0&0&1&1&1&0\\ 1&1&0&0&0&1\end{smallmatrix}$
  • Fusion data 3:
    $\begin{smallmatrix}1&0&0&0&0&0\\ 0&1&0&0&0&0\\ 0&0&1&0&0&0\\ 0&0&0&1&0&0\\ 0&0&0&0&1&0\\ 0&0&0&0&0&1\end{smallmatrix},\ \begin{smallmatrix}0&1&0&0&0&0\\ 1&0&0&0&0&0\\ 0&0&0&1&0&0\\ 0&0&1&0&0&0\\ 0&0&0&0&1&0\\ 0&0&0&0&0&1\end{smallmatrix},\ \begin{smallmatrix}0&0&1&0&0&0\\ 0&0&0&1&0&0\\ 1&0&0&0&0&0\\ 0&1&0&0&0&0\\ 0&0&0&0&1&0\\ 0&0&0&0&0&1\end{smallmatrix},\ \begin{smallmatrix}0&0&0&1&0&0\\ 0&0&1&0&0&0\\ 0&1&0&0&0&0\\ 1&0&0&0&0&0\\ 0&0&0&0&1&0\\ 0&0&0&0&0&1\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&1&0\\ 0&0&0&0&1&0\\ 0&0&0&0&1&0\\ 0&0&0&0&1&0\\ 0&0&0&0&0&2\\ 1&1&1&1&0&0\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 1&1&1&1&0&0\\ 0&0&0&0&2&0\end{smallmatrix}$
  • Fusion data 4:
    $\begin{smallmatrix}1&0&0&0&0&0\\ 0&1&0&0&0&0\\ 0&0&1&0&0&0\\ 0&0&0&1&0&0\\ 0&0&0&0&1&0\\ 0&0&0&0&0&1\end{smallmatrix},\ \begin{smallmatrix}0&1&0&0&0&0\\ 1&0&0&0&0&0\\ 0&0&0&1&0&0\\ 0&0&1&0&0&0\\ 0&0&0&0&1&0\\ 0&0&0&0&0&1\end{smallmatrix},\ \begin{smallmatrix}0&0&1&0&0&0\\ 0&0&0&1&0&0\\ 0&1&0&0&0&0\\ 1&0&0&0&0&0\\ 0&0&0&0&1&0\\ 0&0&0&0&0&1\end{smallmatrix},\ \begin{smallmatrix}0&0&0&1&0&0\\ 0&0&1&0&0&0\\ 1&0&0&0&0&0\\ 0&1&0&0&0&0\\ 0&0&0&0&1&0\\ 0&0&0&0&0&1\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&1&0\\ 0&0&0&0&1&0\\ 0&0&0&0&1&0\\ 0&0&0&0&1&0\\ 0&0&0&0&0&2\\ 1&1&1&1&0&0\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 1&1&1&1&0&0\\ 0&0&0&0&2&0\end{smallmatrix}$
They are excluded from modular categorification by Theorem 1.19 (1.14) because $\mathcal{V}(\mathcal{C})=\mathcal{V}(12)=\{2,3\}$, while $3\not\in\mathcal{V}(d_{i})$ for any $i$, and $\mathcal{V}(\mathcal{C}_{pt})=\{2\}$. Additionally, Theorem 1.22 rules them out since 3 is a (powerless) factor of $\mathrm{FPdim}(\mathcal{C})$, but does not divide $\mathrm{FPdim}(\mathcal{C}_{pt})=4$.
Note that the fusion data for cases 3 and 4 can be generalized. Consider a group $G$ with order $n^{2}$ and a group $K$ with order $m+1$. Then, analyze the following half-Frobenius integral fusion ring:
  • Rank: $n^{2}+m$,
  • FPdim: $n^{2}(m+1)$,
  • Type: $[[1,n^{2}],[n,m]]$,
  • Basis: =$\mathcal{B}=\{x_{g}\}_{g\in G}\cup\{\rho_{k}\}_{k\in K\setminus\{e\}}$,
  • Fusion data:
    • $\rho_{e}:=\frac{1}{n}\sum_{g\in G}x_{g}$,
    • $x_{g}x_{h}=x_{gh}$, for all $g,h\in G$,
    • $\rho_{k}\rho_{l}=n\rho_{kl}$, for all $k,l\in K$,
    • $x_{g}\rho_{k}=\rho_{k}x_{g}=\rho_{k}$, for all $(g,k)\in G\times K$,
According to Theorem 1.19 (1.14), it is excluded from modular categorification if $n>1$ and $m+1$ has a prime factor that does not divide $n$.
Note that the types [1,1,1,1,1,1,1,1,1,3,3,3][1,1,1,1,1,1,1,1,1,3,3,3][ 1 , 1 , 1 , 1 , 1 , 1 , 1 , 1 , 1 , 3 , 3 , 3 ] and [1,1,1,1,2,2,2,2,2,2,2,2][1,1,1,1,2,2,2,2,2,2,2,2][ 1 , 1 , 1 , 1 , 2 , 2 , 2 , 2 , 2 , 2 , 2 , 2 ] are excluded from modular categorification by Theorem 1.19, but not by Theorem 1.22, since their $\mathrm{FPdim}=2^{2}3^{2}$ have no powerless prime factor.
Availability of data and materials. Data for the computations in this paper are available on reasonable request from the authors.
Conflict of interest statement. On behalf of all authors, the corresponding author declares that there are no conflicts of interest.